A Complete Bibliography of Publications in Learned Publishing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.06

Title word cross-reference

10 [2157]. $100 [475]. $100.00 [353]. $12.50 [219]. $185.50 [314]. $25 [498]. $25.00 [637, 219, 139]. $28 [598]. 3 [955]. $35.00 [620]. $40 [52, 202]. $420 [240]. $46.00 [497]. $50.00 [400]. $60.00 [544]. $80.00 [117]. × [332, 288].

-D [955]. -year [2157].

/BL [324]. /Ph [181].

0 [94, 562, 399, 449, 541, 54, 119, 180, 543, 74, 137, 201, 448, 498, 473, 476, 524, 697, 352, 475, 117, 260, 421, 716, 75, 599, 140, 658, 76, 397, 598, 155, 157, 376, 497, 620, 114, 139, 400, 583, 618, 584, 286, 258, 138, 523, 677, 496, 521, 582, 617, 585, 657, 659]. 0-04-440237-6 [138]. 0-19-282669-7 [238]. 0-226-55014-1 [620]. 0-262-02275-3 [239]. 0-295-96651-3 [219]. 0-295-96652-1 [219]. 0-406-02094-9 [523]. 0-408-10716-2 [34].

0-442-01316-7 [617]. 0-470-21126-1 [200]. 0-521-35179-0 [94]. 0-521-35973-2 [94]. 0-521-47154-0 [697]. 0-521-55136-6 [716]. 0-521-55898-0 [716]. 0-566-05743-3 [155]. 0-566-07631-4 [677]. 0-670-83767-9 [618]. 0-7123-2045-8 [201]. 0-7123-3187-5 [312]. 0-7123-3227-8 [313]. 0-7123-3257-X [355]. 0-7185-1252-9 [75]. 0-7206-0720-5 [93]. 0-7206-7201-3 [93]. 0-7279-0822-7 [716]. 0-7494-0126-5 [599]. 0-7494-0900-2 [659]. 0-8103-1543-2 [117]. 0-8203-0889-7 [521]. 0-8242-0807-2 [377]. 0-8444-0554-X [114]. 0-847568-39-5 [137]. 0-85142-219-5 [140]. 0-85142-310-8 [582]. 0-85142-323-X [584]. 0-852-67274-8 [157]. 0-85386-146-3 [54]. 0-85386-64-1 [180]. 0-86291-620-8 [310]. 0-86656-616-3 [76]. 0-86656-654-6 [73]. 0-89391-273-5 [36]. 0-901737-05-4 [287]. 0-905958-78-0 [91]. 0-906148-09-X [74]. 0-906148-10-3 [657]. 0-907706-02-9 [330]. 0-907706-03-7 [330]. 0-914340-09-3 [543]. 0-930357-16-7 [139]. 0-930357-21-3 [222]. 0-935012-16-8 [598]. 0-936547-19-7 [620]. 0-94-2308-27-1 [286]. 0-942308-23-9 [119]. 0-942308-24-7 [220]. 0-942308-26-3 [314]. 0-942308-28-X [288]. 0-942308-30-1 [352]. 0-942308-31-X [353]. 0-942308-36-0 [475]. 0-942308-38-7 [498]. 0-942398-33-6 [400]. 0-947-56864-6 [658]. 0-947568-33-6 [55]. 0-947568-49-2 [397]. 0-947568-51 [354]. 0-947568-52-2 [448]. 0-947568-54-9 [476]. 0-947568-55-7 [473]. 0-947568-56-5 [497]. 0-947568-57-3 [524]. 0-948905-16-6 [92]. 0-948905-23-9 [95]. 0-948905-31-X [331]. 0-948905-35-2 [421]. 0-948905-38-7 [205]. 0-948905-39-5 [156]. 0-948905-47-6 [260]. 0-9516856-0-0 [376]. 0-9517527-0-7 [399]. 0-9517527-23 [496]. 0-9517527-3-1 [449]. 0-9521762-0-3 [583]. 0-9521762-1-1 [585]. 0-958-3866 [562]. 0-968905-33-6 [258]. 0044458061 [398]. 0074-4387 [222]. 01-373 [53]. 01-4 [259]. 020 [221]. 0265-9271 [285]. 03 [561]. 03-85403-4773-3 [561]. 0957-9566 [237]. 0EP [377].

1 [449, 541, 540, 180, 678, 352, 586, 563, 221, 472, 447, 204, 311, 286, 495, 585, 755]. 1-3 [202]. 1-55623-530-5 [619]. 1-55623-531-3 [619]. 1-85056-859-6 [311]. 1-850703-071-0 [541]. 1-85091-65-3 [204]. 1-85172-035-9 [495]. 1-85172-042-1 [522]. 1-85172-043-X [522]. 1-85172-049-9 [586]. 1-85431-023-2 [289]. 1-85433-107-8 [563]. 1-85739-084-9 [678]. 1-87069-14-4 [755]. 1-871224-020 [221]. 1-871224-11-X [447]. 1-871224-11-X/12-8 [447]. 1-87157745-2 [696]. 1-872757-00-6 [259]. 1-873521-00-6 [332]. 1-873836-01-5 [540]. 1-874728-00-3 [472]. 10.95 [93]. 10th [240, 435]. 11 [748, 709, 248, 279]. 11-year [2192]. 118pp [716]. 1198-9742 [637]. 11th [649]. 12 [250, 725]. 12-8 [447]. 12th [725, 232]. 13.95 [332]. 130pp [314]. 131pp [658]. 14 [249, 410, 278]. 14.95 [597]. 142pp [155]. 15 [277]. 15.00 [115, 355]. 15pp [53]. 15th [679]. 16 [726, 724, 705, 251]. 17 [275]. 17-18 [496]. 171pp [659]. 17th [956].

17th-century [956]. 18.00 [118, 120, 55]. 18.95 [289]. 19 [2158, 2167, 2165, 2188, 2122, 742, 2105, 2182, 2118]. 19.95/$35.95 [697]. 1980s [78]. 1986 [54]. 1986/87 [73]. 1987 [74]. 1988 [189, 96, 203, 202, 289]. 1989 [312]. 1990 [261, 310, 247, 354, 222, 240, 232]. 1990s [433, 367, 365]. 1991 [251, 324, 399, 448, 496]. 1993 [679, 562]. 1994 [638]. 1995 [698]. 1997 [756]. 1998 [1031]. 1EF [734]. 1HR [259]. 1st [73, 117]. 1XT [399, 496].

2 [94, 448, 117, 421, 397, 659]. 2-day [2192]. 2.0 [1972]. 2.95 [205]. 20 [706, 1489]. 20-year [1791]. 2000 [310]. 2003 [1155, 1163]. 2004 [1206, 1192]. 2005 [1255, 1250, 1241, 1251]. 2006 [1299]. 2007 [1350, 1336]. 200pp [73]. 2015 [1782]. 2017 [1907]. 2018 [1968]. 2019 [2037, 2181]. 206pp [156]. 20pp [449]. 20th [448, 203]. 21 [323, 707, 322]. 21st [1465, 1047]. 21st-century [1465]. 22 [230, 301]. 22-25 [73]. 22.50 [200]. 224pp [716, 677]. 22x [93]. 23 [280, 496]. 25 [92]. 25th [620]. 26-28 [448]. 27.00 [34]. 270pp [523]. 28-30 [312]. 29 [745, 708]. 29.95 [331, 95]. 2nd [331, 598, 583]. 2SG2 [447]. 2UN [237].

3 [54, 543, 524, 76, 561, 155, 581, 472, 204, 583, 657]. 3-527-29041-9 [581]. 3/4 [73]. 30.00 [203]. 36pp [157]. 37.95 [36]. 3866 [562]. 3HJ [180]. 3QH [401]. 3rd [156].

4 [677, 755]. 4.0 [2164]. 4.95 [182]. 45.00/$70.00 [638]. 48pp [311, 657]. 4JQ [332]. 4th [202, 523]. 4YP [449].

5 [540, 137, 599, 140, 497]. 5.00 [91]. 560 [872]. 5LB [472]. 5th [2054].

6 [260, 658, 311, 400, 258, 138]. 6.99 [621]. 66pp [51]. 6LE [290]. 6th [697].

7 [399, 498, 473, 716, 620, 139, 521, 617]. 7th [302].

8 [276, 201, 542, 563, 598, 157, 447, 582]. 82-00-07842-6 [52]. 86th [746]. ’87 [74]. ’88 [95]. 8879 [200]. 8AW [474]. 8DB [182]. 8UH [330, 447]. 8vo [258].

9 [688, 119, 476, 542, 678, 586, 75, 581, 618, 523, 495]. 9-918006-22-8 [542]. 9001 [1499]. ’90s [658, 352]. 92-825-8347-3 [181]. 98145-5096 [52]. 9JJ [696]. 9SB [53, 754]. 9th [584].

A&I [1409]. A&Is [1624]. A-Z [447]. A. [266]. A5 [266, 287]. Abate [117]. abbreviations [91]. Ablex [36]. abrupt [2121]. absence [800]. Abstracting [119, 544, 314, 353, 352, 475, 288, 400, 220, 244, 498]. Abstracts [251, 631, 2188, 104, 2228, 810]. abuses [387]. academia [2089, 2139, 362]. Academic [1776, 2216, 1320, 1604, 1203, 2162, 2152, 2153, 2012, 1638, 1357, 973, 461, 1165, 938, 1822, 118, 1222, 1629, 1795, 546, 2130, 964, 1979, 1828, 1752, 1778, 1961, 1942, 2041, 1146, 2225, 2186, 2105, 1987, 1366, 1276, 1913, 1543, 914, 1177, 568, 2217, 1639, 2191, 603, 667, 1657, 457, 824, 1132, 1267, 895, 1110, 1940, 2079, 1980, 1259, 2010, 1063, 2218, 2212, 2222, 2183, 722]. academics [1145, 2090, 1736, 297]. academies [2123, 459]. Academy [409]. accelerate [2053]. Acceptance [1697, 1640, 1901, 1295, 1483]. accepted [1007]. acces [938]. Access [1705, 1820, 1365, 1789, 420, 1611, 1314, 1836, 2166, 1466, 1578, 1247, 2135, 1268, 1176, 1343, 1652, 1379, 1482, 1601, 1822, 1889, 2026, 2142, 1222, 1966, 2018, 1371, 1481, 2078, 1575, 1331, 1096, 1111, 1166, 2092, 722, 2214, 1819, 1252, 1790, 974, 1210, 1332, 1912, 1992, 280, 1446, 1187, 1429, 1755, 1666, 1905, 1571, 1320, 1724, 1884, 2219, 1844, 1177, 1748, 620, 2064, 1936, 2107, 2123, 2146, 1242, 1849, 1612, 1731, 1886, 2215, 1965, 1434, 1815, 1982, 1220, 1693, 1303, 1719, 1763, 1751, 1330, 1283, 1467, 1599, 1687, 1991, 2184, 1128, 1500, 1046, 2069, 1794, 1727, 1335, 1055]. access [1239, 1123, 1514, 1954, 1169, 1444, 1189, 1211, 2145, 1325, 1129, 2174, 1409, 1910, 1911, 1089, 1573, 1223, 1260, 2203, 1909, 2053, 1973, 1576, 2086]. Accessibility [1919, 1922, 1916, 1928, 1921, 1925, 1918, 1930, 1917, 1972, 1924, 1893]. accessible [1923, 1926, 1916, 1450, 1920, 1927]. Accessing [1855]. accommodated [1735]. account [2148]. accountable [2227]. accounts [2216, 2218]. accuracy [1360, 1332]. accurate [1814]. acknowledge [1792]. Acknowledgement [1098]. ACORN [812]. acquisition [884, 759]. Acquisitions [2048, 553, 51]. across [1201, 2140, 522]. Act [189, 1031, 289, 842, 1031, 1390]. Action [772, 1923, 1682, 1742]. actionable [963]. Activating [296]. activities [42, 1056, 1828, 65, 107, 8, 25, 133]. activity [1792]. acts [2073]. actual [1673]. AD [872]. adapt [1435]. adaptation [2116]. adaptations [984]. add [2148, 530]. added [15, 130, 149, 1044, 36]. adding [1448]. address [7]. addressing [2061]. adherence [266]. administration [573]. administrative [2106, 2212]. Administrators [1994]. ado [1651]. adopters [1808]. Adopting [2174]. adoption [2175, 2139, 1991]. Advancement [1203, 1227]. advances [268]. Advantages [1234]. Adventures [1372, 1379, 1482]. advertising [455, 1309, 823, 1736, 1970]. advice [1929, 696, 1060, 583, 1354]. affairs [308, 326, 348, 394, 415, 443, 467, 516, 816, 373, 692, 633, 791, 865, 850, 880, 836, 756, 807, 821, 773]. affecting [2076]. affects [1688]. affiliations [2040, 2148]. affinity [1710]. affordable [1992]. afraid [611, 1031]. Africa [1575, 1763, 1347]. African [1981, 1108, 1046]. After [1199, 1829, 726, 687, 1241, 1624, 2035]. against [304, 2160, 2044, 1742]. AgBiotechNet^{T M} [990]. age [526, 1092, 1119, 35, 1813, 1596, 1750, 1199, 521, 1733, 1535]. agencies [954]. Agency [435]. Agenda [772]. agent [406]. Agents [919, 827, 401, 918, 942, 778]. aggregator [1524]. aggregators [1030]. AGM [230, 322]. agony [1093]. agreement [2055]. agreements [2142, 1151, 138, 523]. agricultural [1949]. agricultural/biological [1949]. ahead [620, 1285]. AI [2000]. aims [757]. al [542]. AL10 [474]. ALCS [734].

Aldershot [155]. Alexandria [598]. align [2053]. alignment [2147]. Alison [599, 659]. Allen [238]. Alliance [1742]. allocation [137]. alma [2230]. ALPSP [230, 419, 322, 725, 1336, 42, 689, 710, 727, 749, 791, 808, 767, 802, 865, 850, 876, 831, 845, 860, 875, 880, 997, 1155, 1127, 1140, 1206, 1181, 1255, 1299, 1350, 1782, 1846, 1907, 1968, 2037, 250, 323, 299, 302, 709, 724, 688, 248, 276, 279, 888, 1164, 1523, 65, 107, 192, 410, 705, 280, 906, 687, 836, 1387, 1205, 1609, 756, 784, 1207, 1275, 1300, 807, 8, 25, 84, 1033, 692, 821, 893, 909, 633, 773, 757, 640, 561]. ALPSP/Charlesworth [1155, 1206, 1255, 1299, 1350]. ALPSP/EASE [323]. ALPSP/PA [302]. ALPSP/Pira [250]. ALPSP/UKSG [687]. also [2182]. ALSP [133]. Alternative [1146, 1973, 1655, 1121, 814, 2181]. alternatives [461, 915]. Altmetric [1637]. altmetrics [1841, 1744]. America [1268, 436, 129, 609, 456, 1503, 704]. American [73, 203, 620, 1836, 1984, 1856, 2042, 1698, 1402, 144, 1858, 1249]. AMI [617]. among [2107, 2215, 2108, 1077]. amplification [2162]. Amsterdam [261, 183, 266]. Amy [240]. Analysing [2106, 1818]. Analysis [1832, 2071, 2165, 2164, 2171, 2135, 1842, 1381, 1575, 1151, 2182, 1936, 1680, 1786, 2117, 2029, 78, 2073, 1502, 1318, 882, 841]. Analytics [1720]. Anbyn [727, 749]. ancient [2123]. Anderson [155]. Andrew [542, 1066]. Andrews [75]. angle [268]. Animal [868, 1530]. Ann [544, 353, 498]. Anne [696]. anniversary [448, 435]. annotation [2016]. Annotum [1598]. Annual [746, 167, 137, 73, 285, 620, 673, 169, 275, 253, 203, 497, 232, 496]. anointed [1847]. anonymity [1319]. answer [1375]. answered [623]. answers [843]. Anthony [542, 52]. anthropause [2169]. anthropology [2156]. anti [22]. Anticipating [649]. antidotes [2149]. any [2001, 1363]. anybody [777]. anyway [527]. APC [1749, 1635]. APCs [1706]. apologies [1754]. appealing [1898]. appeals [1898]. appearing [2023]. appendices [182]. applications [1251, 365, 75]. appointed [1065]. approach [1926, 431, 1545, 2078, 1668, 898, 991, 1912, 1724, 1430, 1971, 2070, 1132]. approaches [1092, 1900, 1844, 2010]. appropriate [810, 1556]. approval [2190, 1688]. April [726, 410, 725, 1336]. aquatic [2180]. Arab [1953, 2036]. Arabia [2088]. Aransas [620]. arbiter [683]. arc [1059]. Archaeologist [292]. Archaeology [1080]. archival [1250]. Archive [1073, 1054, 669, 1196, 799]. archived [1332]. Archives [1343, 1071, 1417, 1238, 363, 1284]. Archiving [1040, 1122, 1266, 1586, 1136, 1668, 1724, 2219, 975, 1708, 1746]. Ardito [400]. Armageddon [668]. arrangement [20]. art [826]. Arthur [157, 288]. Article [1505, 2189, 1630, 1637, 2135, 367, 1021, 2143, 1640, 961, 1023, 1971, 1387, 1398, 1563, 1574, 1587, 1761, 1502, 1757]. article-level [1637]. articles [2012, 104, 1332, 1831, 1304, 1505, 1571, 1729, 1614, 2072, 1938, 1972, 1719, 2176, 1500, 1014, 1359, 1547, 1882, 2035, 1631]. artificial [1813, 1833]. ARVO [2146]. arXiv [2024]. ASA [918]. Asia [52, 874]. Asian [1699, 1895]. Ask [1282]. Aslib [285, 140, 584, 677, 582]. aspects [2224]. aspirations [2074]. assess [1522]. Assessing [745, 306, 1787, 1744]. Assessment [1165, 940, 1972]. assets [1178]. assignment [1994]. assistance [1104]. associated [1753]. Association [254, 746, 54, 180, 419, 137, 542, 52, 401, 673, 266, 313, 1176, 823, 258, 1352, 32, 50, 151, 385, 344, 247, 551, 134, 301, 821, 757]. Association/British [419]. associations [1641, 1486, 1354]. assumptions [1991, 1934]. assurance [1173]. astronomy [1304]. Athens [521, 952]. attention [1570]. Attitudes [2088, 1966, 2104, 186, 2077, 2083, 103]. attributes [2030]. attribution [1769]. Aubyn [689, 710, 764]. audiences [1821]. audit [1917, 339]. August [249, 448]. Australia [354, 2153, 1353, 2221, 1320, 2089]. Australian [2090, 2223, 1392, 1589, 1062]. authentic [1955]. Author [2161, 894, 1707, 1382, 1004, 1318, 734, 2148, 1273, 1955, 844, 1332, 2046, 927, 2194, 1430, 1952, 1708, 1722, 2140, 1279, 1611, 1986, 2227]. author-archived [1332]. author-funded [927]. Author-initiated [1707]. author-suggested [1952]. Authority [1709]. Authors [754, 53, 294, 1048, 1884, 948, 909, 1105, 1189, 2012, 2040, 1788, 683, 1686, 24, 697, 1699, 1644, 1908, 2173, 1640, 1320, 1939, 1952, 2149, 1544, 1660, 1768, 1237, 1280, 447, 1195, 345, 893, 1520, 2073, 926, 1779, 1642, 911, 2069, 180, 91, 355]. Authorship [1517, 2009, 1595, 1998, 2190, 1769, 1119, 1777, 1960, 2075, 1896]. autoethnographic [1971]. Automated [475]. availability [1162, 2173]. Available [754, 182, 696, 401, 496, 377, 734, 561]. avant [1228]. avant-garde [1228]. Avenue [472]. Avoiding [1535]. awakening [1572]. Award [343, 785, 2102, 2103, 2198, 575, 172, 64, 169, 170, 785, 253]. Awards [440, 464, 490, 513, 651, 691, 712, 729, 1155, 1206, 1255, 1299, 1350, 1782, 1907, 1968, 2037, 344, 1846]. awareness [381, 285, 1046]. away [1658, 1957].

B [35]. Back [1810, 668, 620, 1529]. background [1615]. Bacon [956]. bad [611, 1167]. badge [2175]. balance [1390, 913]. Banbury [312]. bane [323]. Bangladesh [1517]. bank [509, 1669]. Banks [1349]. Barnard [331, 260, 205, 156]. Baron [91]. Barrett [239]. barrier [2174]. Barriers [1723, 261, 385, 1742]. based [1519, 1983, 2178, 1405, 1872, 1831, 1370, 927, 1844, 2217, 2042, 1391, 2203, 986]. bases [1367]. Basham [597]. basic [2134, 330]. Bate [156]. Bath [376]. battle [2160, 1399]. battles [783]. Baumgarten [266]. Baverstock [599, 659]. BD7 [259]. be [1655, 2025, 1035, 1543, 2125, 1835, 1589, 916, 1642]. Beall [2160, 2111, 2215]. Beaulieu [524]. became [1330]. become [1790, 1718, 2176, 1335]. Bedford [180, 497]. been [1754]. before [553]. beginning [502]. behalf [561]. behaviour [2161, 1869, 1804, 1887, 1937, 1086, 2133]. behavioural [1897]. behaviours [2104, 2083]. behind [2155, 900]. Beijing [884]. being [1465, 1136, 1007, 1857]. beliefs [2075, 2074]. Bell [758]. Bella [203]. Benchmarking [1273]. beneath [698]. benefit [2209]. benefits [1314, 1925, 1571, 1762]. Benetech [1920]. bepress [1368]. Berg [1403]. Berks [377]. Bernard [2113]. Berne [266]. best [797, 1965, 1584]. beta [1329]. Bethesda [543]. better [2015, 1125, 1035, 1470, 2229, 1927]. Betty [119, 314, 352, 286, 220]. Between [1159, 2154, 2135, 1984, 2092, 870, 1556, 5, 2029, 1909, 1710]. Beyond [1586, 1769, 2208, 1925, 1672, 1019, 658, 2089, 1133]. Bi [183]. Bi-monthly [183]. bias [1697, 1579, 1603, 1483, 2183]. Bible [159]. BIBLINK [954]. bibliodiversity [1982]. bibliog [541, 36]. bibliographic [60, 954]. Bibliography [355, 313, 35]. bibliometric [2165, 2164, 1575, 1751, 1935, 2182]. big [611, 1822, 1472, 1159, 1378, 1125, 1246]. bigger [1974]. bilingual [1788]. Bill [618, 495, 926]. bills [1225]. biochemist [645]. Biography [788]. Bioline [988]. Biological [1225, 1949, 696, 91, 1952, 951]. Biology [543, 697, 1966, 1938]. biomedical [2161, 1898, 1394, 2213, 1104, 1722, 1236, 1751, 1296, 794, 1027, 1779, 1480]. BioOne [1121]. bioRxiv [2071]. Birmingham [292]. Birn [204]. BL [324]. Black [1683, 1889, 1912]. Blacklist [1956]. Blackstone [289, 289]. Blackwell [1358]. Blackwells [74]. Blaise [448]. Blake [696, 497]. bli [1861]. blind [1929, 1556, 1557]. Blockchain [2195, 2095, 1976]. blow [1650]. Blueprint [331, 260, 421, 95, 205, 258, 92, 156]. Blunden [638]. BMJ [716, 1791]. BNBRF [313]. boardroom [1823]. boards [2180, 2090, 1026]. boat [1331]. Bodies [675, 728, 750, 233, 246, 393, 414, 442, 466, 491, 515, 535, 558, 634, 652, 690, 711, 349, 371]. Bone [1485]. Book [1440, 1510, 1528, 1591, 94, 863, 290, 1455, 1476, 1525, 1526, 1567, 1662, 562, 1112, 238, 399, 411, 449, 511, 541, 540, 574, 637, 754, 707, 726, 748, 789, 51, 54, 119, 240, 200, 180, 237, 419, 543, 544, 1338, 1454, 1459, 638, 118, 120, 115, 1313, 1496, 74, 137, 201, 331, 314, 355, 353, 448, 498, 473, 476, 524, 542, 678, 697, 1703, 53, 73, 116, 181, 475, 1506, 1458, 1581, 1593, 1689, 1714, 52, 1607, 1619, 1647, 1474, 1511, 117, 260, 421, 1477, 716, 586, 1622, 563, 75, 332, 1605, 599, 95, 140, 221, 1413, 1050, 658, 1551]. Book [76, 397, 1410, 1663, 1701, 1712, 292, 1182, 1436, 597, 835, 1623, 561, 698, 598, 1298, 239, 696, 1491, 1497, 1676, 1424, 35, 91, 203, 398, 155, 157, 205, 312, 310, 497, 620, 175, 996, 1441, 1473, 1509, 194, 1772, 1726, 1737, 1494, 1495, 1507, 967, 401, 1457, 1412, 581, 1126, 1633, 1700, 1439, 219, 330, 259, 474, 472, 447, 202, 183, 266, 288, 287, 1492, 1592, 1594, 1274, 1323, 1373, 1384, 1437, 1786, 93, 204, 34, 114, 139, 222, 311, 400, 1527, 1791, 583, 1168, 1383, 1552, 1580, 1568, 1566, 1142, 1411, 1425, 619, 618, 1154, 1493, 1475, 584]. Book [55, 1606, 1648, 36, 313, 286, 1396, 1508, 1713, 258, 1397, 1426, 1438, 1539, 1540, 1569, 1634, 1661, 1715, 1716, 1423, 138, 289, 523, 677, 770, 819, 420, 1078, 1456, 1538, 1553, 1677, 1702, 1193, 496, 521, 1621, 495, 582, 92, 156, 220, 1664, 1678, 1646, 1036, 1094, 1243, 1739, 1770, 1771, 617, 1361, 522, 585, 621, 657, 755, 734, 848, 1287, 1022, 1554, 1620, 1738, 659, 283, 647, 231, 1656, 147, 1523, 1171, 1395, 1752, 1778, 1139, 1943, 1987, 1917, 1630, 273, 1212, 2222, 2087, 324, 590, 754, 313, 258]. booklet [184]. Books [17, 541, 2214, 330, 447, 583, 585, 1401, 1561, 250, 299, 449, 555, 890, 884, 1564, 1667, 82, 261, 2, 1926, 209, 2063, 1546, 1353, 1478, 1002, 1160, 61, 63, 1162, 1913, 1490, 96, 127, 661, 38, 1514, 1110, 163, 1377, 58, 659, 320, 52, 599, 620, 114]. Bookseller [562]. bookselling [1685, 621]. boon [323]. boosting [1162]. born [1920, 1404]. born-digital [1404]. Bosnia [430]. Boston [201, 355, 746]. bother [703, 1324]. bound [498, 220, 474]. bounds [221]. Bowker [678, 310]. Bowker-Saur [678, 310]. Box [449, 52, 330, 447, 183, 1919]. box-checking [1919]. boys [1149]. BPIF [630]. Bradford [259]. Bradley [584]. brand [2220, 2133]. Brazil [1572]. Brazilian [273]. Breaking [385, 1107]. Brian [638]. BRIC [1549, 1615, 1572, 1600]. bridges [870]. brief [804, 818, 834, 847, 862, 879, 891, 1808, 1870, 757, 874]. briefings [116]. Brighton [301]. Brindley [312]. Bring [1990, 1609]. Bringing [1406, 868, 1768, 92]. brink [1837]. British [291, 419, 118, 355, 678, 116, 563, 278, 561, 312, 310, 287, 679, 313, 2058, 2057, 30, 48, 71, 90, 112, 136, 154, 110, 177, 199, 216, 235, 283, 257, 327, 347, 372, 416, 403, 444, 468, 492, 517, 536, 409, 2, 147, 12, 127, 603, 163, 96]. Broadcast [1421]. Broader [1834]. Brobdingnag [870]. broken [1930]. Brookfield [155]. Bryn [73]. Bryson [618]. BS [96, 290]. BS4821 [291]. BSI [290, 96]. Buddhist [720]. budget [130, 102]. Building [1196, 1056, 2119, 2054, 58, 1871, 1869, 1121, 1775, 1796]. bulletin [285]. BUPA [576]. Bureau [339]. Business [388, 182, 619, 1798, 2156, 638, 1842, 1448, 1068, 1472, 990, 1008, 1260, 1985, 1065, 540, 449]. bust [1990]. Butler [420]. Butterworths [34, 523]. buyer [157]. buying [797]. buzzword [1019]. BV [183]. bytes [661].

C [620, 219, 52]. c&j [156]. C-50096 [52]. C. [400]. c/o [52]. CA [420, 1293, 1334]. Cabells [1956]. CAL [1926]. Calculating [1074]. calls [2137]. Calmer [1285]. Cambridge [94, 275, 697, 716, 239, 1869]. came [1733]. Campbell [419, 34]. Campus [259, 312]. Can [548, 1999, 2218, 2034, 426, 1998, 786, 593, 1706, 965, 1183, 586, 2025, 1533, 1219, 1543, 1640, 2125, 1835, 1844, 1957, 2045, 815, 1529, 1675, 361, 2053, 1616, 1874, 2095, 1815]. Canadian [281, 300]. cancellations [1284]. cancer [1819]. cannot [1471]. capable [1589]. Capital [1937]. care [2088, 1535]. career [2034, 1776, 2002, 2075, 2104, 2193, 1971, 2089, 1887, 1896, 1867, 1957, 2027, 2083, 2045, 1906, 2021, 1847]. careers [498, 621]. CARMA [1326]. Carol [472]. Cartoons [472]. case [1885, 2057, 2040, 1484, 1966, 2119, 2205, 2028, 1293, 2092, 1872, 1819, 2228, 1928, 2046, 1692, 83, 1134, 1808, 1584, 1974, 1959, 1927, 1708, 1965, 2120, 1899, 990, 2184, 1296, 2017, 1958, 2059, 872, 1941, 2038, 2079, 2022, 2014, 2183, 2087]. casebound [332]. Cased [419, 260, 421, 239, 36, 258, 377]. Casting [1548]. Catalogued [1488]. cataloguing [826]. catch [483]. Cause [783]. causes [1835, 2179]. caution [2188, 1079]. CBE [697]. CD [611, 1007, 226, 624, 753, 462, 554, 460, 584, 420]. CD-ROM [611, 226, 624, 753, 584, 420]. CD-ROMland [462]. CD-ROMs [1007, 460]. ceased [2223]. Ceci [1563]. celebration [448]. censorship [2068]. center [661]. Central [531, 1599]. Centre [324, 116, 157, 330, 757, 1157, 707, 726, 748, 722]. century [956, 1465, 1609, 1047]. Certifying [976]. c’est [1293]. chain [1221, 1015, 1913, 1930, 412, 1367, 1110]. Chairman [697, 689, 710, 727, 749, 25]. challenge [1074, 505, 1325]. Challenges [1248, 2196, 1204, 1102, 355, 1816, 964, 2051, 294, 2091, 1975, 1389, 936, 232, 1969, 2132, 1404]. challenging [1471]. chance [1640]. Change [707, 726, 748, 2100, 1983, 1293, 1334, 1912, 2186, 1913, 649, 1501, 505, 1887, 2027, 2126, 1296, 1158, 2082, 503]. changed [2084]. Changes [1752, 2051, 2228, 1214, 2091]. Changing [486, 1840, 1764, 532, 51, 1720, 406, 1961, 2084, 1017, 1883, 856, 1279, 1261]. chapter [1275]. chapters [1003]. character [243]. Characteristics [2158]. charge [2189, 401]. charges [2135, 1467, 1757]. Charging [839, 589, 1386]. charities [1687]. charity [269, 632]. Charlbury [401]. Charles [354, 138, 523, 1218]. Charleston [553]. Charlesworth [169, 785, 253, 64, 343, 785, 1155, 1206, 1255, 1299, 1350, 575]. Chatterton [73]. checking [1919]. checklists [2025]. cheers [356]. Cheltenham [332]. chemist [1366]. Chemistry [1225, 1579, 1319, 1175]. chemists [1345, 453]. Chicago [620]. Chichester [200]. Chief [2180, 1785]. Chile [147, 2189]. Chilean [2040]. China [2178, 958, 1643, 1695, 508, 1590, 1682, 1654, 1543, 1234, 1430, 2204, 1307, 1639, 1711, 1827, 1248, 1172, 1672, 1515, 126, 1308, 1600, 1084, 1388, 1516, 2079, 1406, 2085, 1259, 1271, 2010, 2226, 1063, 244, 1628, 2106, 2133, 2212, 2039]. Chinese [2155, 1371, 989, 1281, 1688, 1804, 1659, 1204, 1267, 1308, 1940, 1951, 2079, 2132, 2203, 1406, 1779, 1576, 1616, 2086, 1480, 1115, 2014, 2011]. choice [1803, 2094, 404]. choices [1074]. choose [1908, 1884, 1953]. Choosing [1354]. CHORUS [2092, 1679]. chose [1293]. Chris [563]. Christopher [266]. Ciba [279]. CIM [157]. CIM-TECH [157]. Cimtech [157]. circulation [2150]. Circulations [339]. cit [1075]. cit. [1217]. Citation [1655, 2213, 1950, 2203, 1829, 2154, 1743, 1399, 2225, 2143, 1643, 2043, 1895, 1883, 2149, 1938, 1364, 1765, 1480, 2141, 1466]. citations [2168, 2028, 2143, 1831, 963]. cite [2149]. cited [1814, 1959, 1734, 2022, 454]. CiteScore [2043]. Citing [1544, 290, 935]. City [203]. CLA [23, 478]. Clack [73]. clarification [757]. clarity [2137]. Clark [138, 523]. Clarke [696]. clash [380]. class [2085, 1143]. clear [1314, 795]. Clearance [801]. Close [399, 1558, 496, 433]. closed [1128]. closer [588, 1263]. Closing [1873]. Cloth [219]. club [1993]. cluster [1842]. cm [238, 399, 449, 541, 637, 239, 696, 93, 36, 313, 289]. CNIT [231]. Co [117, 131, 190, 5, 1668, 1304, 1202, 1742]. co-existence [1304]. Co-operation [5]. Co-operative [131, 190, 1202]. co-ordinated [1668, 1742]. coast [1856]. Code [1447, 1758, 1251]. coercive [2149]. cognition [2133]. cohort [2192]. Cole [114]. Colin [397]. Coling [658]. Collaborating [2092]. collaboration [2058, 2165, 1769, 1449, 1786, 881, 2059, 2082, 986]. Collaborative [642, 1489]. collaborator [1464]. Colleagues [1843, 1470]. Collecting [734].

Collection [1948, 1061, 755, 1164]. College [2158, 180, 73]. colleges [1069]. Colloquium [679]. Colombia [163]. Columbia [986]. Combining [1522, 1941]. Come [272, 1422, 1203, 1283]. comes [1210, 1596]. Coming [19, 37, 56, 77, 97, 121, 142, 158, 185, 206, 241, 263, 293, 315, 334, 357, 379, 402, 422, 451, 477, 501, 525, 564, 600, 622, 639, 660, 681, 699, 717, 735, 744, 526, 20, 852, 1414]. commended [171]. commentary [871, 1315, 1386]. commerce [943, 902]. commercial [268, 137, 322, 208, 915, 212]. Commission [902]. commissioning [1752]. Committee [772, 920, 697, 689, 710, 727, 749, 1205, 773, 1009]. common [497]. commons [1981, 1536, 1310]. Communicating [1562]. Communication [1729, 2137, 1885, 2180, 1232, 364, 73, 1179, 336, 1817, 1831, 1993, 1013, 1191, 949, 2104, 1905, 35, 1749, 1883, 1156, 1999, 866, 1804, 2074, 2083, 1041, 1172, 1195, 1261, 1071, 1794, 2036, 1529, 1149, 1233, 1837, 986, 542]. Communications [1709, 1336, 1992, 1862, 2185, 2117]. Communities [181, 775, 2095, 1270, 1685, 1710, 2133]. community [2058, 1247, 1319, 1001, 1993, 969, 2060, 100, 1999, 953, 1794, 2117, 2016, 1969, 1077, 2133, 449]. community-driven [1794]. community-owned [1999]. Company [363, 377]. comparative [1952, 1936]. compared [1578, 1800]. Comparing [1201, 1734, 2154]. Comparison [1710, 1909, 1249, 2218]. comparisons [2104]. Compass [1358]. competences [247, 1348]. competing [1240]. competition [1785, 929, 1446, 1729]. competitive [1683]. Compiled [421, 238, 157]. compiler [180]. compiling [210]. complete [94, 57, 2030]. complex [1913, 2093]. compliance [2161, 1923, 1135]. component [2176]. components [952]. comprehensive [1281]. compromise [1556]. Compton [237]. Computer [122, 597, 259, 679, 370, 1966, 500, 243, 668, 239, 474]. Computer-software [370]. Computerisation [118]. computerized [7]. Computers [301, 391, 721]. computing [521, 75]. concentration [1577]. concept [811, 1242, 6, 1781, 1173, 955, 1369]. conception [1135]. Conceptions [448]. conceptual [2024]. concern [1380, 783, 2138]. concerning [569]. concerns [893, 909]. Conference [746, 74, 73, 202, 1336, 943, 399, 448, 354, 312, 497, 496, 2181, 1671, 2212, 167, 553, 673, 275]. conferences [995, 2121, 19]. confident [1589]. confidentiality [684]. conflict [1517, 2226]. Congress [114]. connections [73]. consensus [737]. consent [2156]. consequences [1652]. Consideration [1994]. considerations [460]. Consolidation [942]. consortia [937, 1167]. Consortial [1107]. Consorting [1167]. Consortium [1203, 1463]. conspicuous [800]. constant [2030]. Consumer [2133, 1289]. consumers [1252]. consumption [2150]. Contemporary [1970]. Content [1608, 1120, 1637, 1519, 1069, 1809, 571, 1922, 1868, 898, 1250, 1821, 1585, 1542, 912, 1570, 1270, 1211, 1833, 1944, 1077]. Contents [1797, 1812, 1825, 1839, 1852, 1864, 1880, 1892, 1903, 1915, 1933, 1947, 1963, 1978, 1989, 2007, 2020, 2033]. contest [1608]. ConText [239, 1519, 1265, 2157, 663, 1608, 664]. Context-based [1519]. Continental [362]. Continued [1482]. continues [1311, 813]. continuing [1106, 1761]. continuity [667, 2100]. continuum [760]. contradiction [799]. Contributing [1938]. contribution [1769, 1392]. contributions [1998, 1783]. Contributor [1998]. contributors [335, 358, 893, 909]. contributorship [1998]. control [57, 627]. controlling [706]. convenience [2190, 1843]. Convention [266]. convert [1844]. converted [1166]. converting [2078]. cooperation [2119]. COPE [1009]. Copies [561]. Coping [503]. Copy [275, 1408, 721, 1152, 1324, 624, 1318, 399]. copy-edit [1324]. Copy-editing [1408, 721, 1318]. copying [408, 1062, 201]. Copyright [150, 465, 689, 710, 727, 749, 858, 843, 886, 1097, 438, 1117, 189, 454, 801, 1390, 949, 856, 766, 783, 765, 830, 800, 902, 14, 26, 43, 86, 108, 132, 173, 195, 214, 289, 274, 305, 325, 441, 514, 534, 557, 577, 594, 613, 653, 674, 623, 1317, 1961, 648, 1151, 481, 66, 408, 871, 1446, 961, 1023, 1185, 948, 288, 287, 1018, 1288, 1781, 872, 67, 225, 413, 1388, 1766, 1516, 783, 435, 677, 582]. Core [1272, 1940, 1869]. Corner [1097, 150, 465, 1374, 14, 26, 43, 86, 108, 132, 173, 195, 214, 274, 305, 325, 441, 514]. Corning [223, 545, 587]. Cornish [201]. Corp [36]. corporate [94, 1090]. corporation [473]. correct [2073]. correction [2087]. correlation [2154, 2218]. Correspondence [33, 680, 470, 450, 520, 538]. corresponding [2194]. Cost [3, 1040, 1315, 1749]. Cost-effective [3]. Costs [1571, 1314, 1673, 925, 1224, 292]. could [1655]. Council [180, 543, 697, 35, 287, 139, 222, 30, 48, 71, 90, 112, 136, 154, 177, 199, 216, 235, 283, 257, 327, 347, 372, 416, 403, 444, 468, 492, 517, 536, 873, 12, 1766]. count [1938]. COUNTER [1135, 1251]. Counterpoints [1763]. countrie [2157]. countries [2026, 1429, 1884, 1899, 625, 1055]. country [1669, 1123, 2174, 2183]. course [330]. courses [647]. Covenant [269]. cover [449, 266, 287]. coverage [1950]. COVID [2158, 2167, 2165, 2188, 2122, 2105, 2182, 2118]. COVID-19 [2158, 2167, 2165, 2188, 2122, 2105, 2182, 2118]. Craig [354]. Cranborne [474]. Creating [709, 1431, 1877, 1289, 2043]. creation [1188, 1860]. Creative [1310, 1536]. CRediT [1998, 1461, 1769]. Crisis [1013, 1194, 2188, 2047, 1899, 1653, 1149, 1516]. criteria [2190, 884, 2093, 1909]. Critical [1213, 1667, 316, 1854, 1789, 1899, 624, 854]. criticism [924]. CRM [1019]. Croatia [1692, 1407]. Croatian [1296, 1444, 1969]. Cronin [448]. cross [1829, 2167, 2140, 2227, 2226, 2105]. cross-disciplinary [2227]. cross-discipline [1829]. cross-publisher [2167]. cross-sectional [2140, 2226, 2105]. CrossCheck [1485, 1462, 1631]. CrossMark [1542]. CrossRef [1190, 1728, 1290]. crossroad [1822]. crossroads [533, 1290]. Crowcroft [698]. Crown [67]. Crowther [449]. Crummy [292]. Crunch [1427]. Crypto [2145]. cryptocurrencies [2145]. crystallography [1359]. Cues [2028]. cultural [2186]. culture [1091, 852, 1877]. Cummings [542]. Cunningham [544, 353, 498]. curated [2095]. Curation [1157, 1948]. curiosity [668]. currencies [39]. current [118, 381, 1992, 1943, 1666, 1801, 2189, 1161, 547, 679, 896, 285]. current-awareness [381]. customer [1479, 387, 1137, 286, 400]. customers [1644]. Czech [1784].

D [91, 581, 289, 955]. Daniel [581]. Dark [1535]. Darmstadt [249]. Data [1355, 1728, 175, 1740, 1741, 1792, 1186, 2206, 1742, 433, 1357, 509, 1956, 1579, 1743, 2205, 2200, 2208, 1479, 2201, 1221, 1453, 2173, 1975, 2175, 1974, 2204, 2207, 2209, 2077, 1965, 1544, 1806, 1746, 1745, 1535, 1562, 1359, 1372, 2203, 1521, 2022, 2141, 1935, 18, 1031]. Database [45, 392, 286, 1502, 2203]. databases [60, 387, 899, 1364, 939, 407, 2024]. datasets [1453]. date [252]. Dates [771, 790, 806, 820, 832, 849, 864, 878, 892, 908, 933, 946, 980, 998, 1011, 1024, 1038, 1052, 1067]. Datus [219]. David [221, 2107]. Davies [156]. Day [716, 2046, 1366, 708, 2192, 1416]. dBase [4]. DC [35]. de-selection [688]. dead [1795]. deal [1125, 1472, 1125, 1246]. dealing [665, 924, 1781]. Deals [1378]. Dear [33, 1795]. death [1754, 1286, 957, 1129]. débâcle [1183]. debate [2018]. deceive [1282]. December [709, 279, 278]. decide [688, 1996]. decision [1378, 723, 1133]. decision-making [1378]. decisions [2152, 1579, 1116, 704, 2131, 1603]. Declaration [2226]. decline [1452, 1799]. declining [2225]. deconstructed [895]. Decoupling [1987]. Deed [269]. defence [123]. Defense [231, 1504, 1079]. define [1399]. Defining [976, 2116, 1694]. delayed [1042]. Delhi [385]. delights [1888]. Delivering [1069, 1669]. Delivery [1070, 381, 1048, 429, 1305]. demand [1188, 1162, 857, 642, 809, 753]. demanding [604]. Demographics [1817]. demonstration [10]. Dempsey [376]. Denis [287]. Denley [180]. dépôt [160]. Department [448, 2092, 561, 1985, 278]. departments [118]. deposit [159, 318, 1124, 160]. depositing [1894]. Derek [75]. Descriptive [240]. deselection [704, 723]. Design [701, 324, 512, 571, 671, 264, 1872, 1873, 591, 286, 40]. Designs [189, 289]. desk [152, 196, 65, 107, 85, 133]. Desktop [57, 94, 95, 221, 83, 474, 156]. despatch [463]. Despatching [484]. details [2148, 2134]. detecting [1462, 1631]. determinants [2175]. detrimental [1997]. Detroit [240, 117]. Developing [898, 1345, 1400, 405, 1211, 2170, 1900, 1545, 950, 1669, 1899, 625, 1574, 1055, 1123]. Development [118, 116, 278, 994, 1034, 2012, 1005, 524, 1432, 2186, 1213, 1106, 1283, 2008, 1847]. Developments [773, 45, 113, 217, 1343, 965, 1311, 145, 228, 1223, 1063]. devices [2130]. diachronic [2227]. Diary [771, 790, 806, 820, 832, 849, 864, 878, 892, 908, 933, 946, 980, 998, 1011, 1024, 1038, 1052, 1067, 1081]. Dictionary [788, 260, 421, 618, 238, 117]. Did [2169, 1753, 2021]. difference [1800]. differences [2152, 1818, 2215, 1897, 2029]. different [1829, 2154, 1430, 2070, 1765, 659]. Digital [840, 801, 1389, 1613, 1709, 1516, 1940, 1062, 1268, 1092, 669, 1196, 1375, 911, 1448, 1136, 910, 738, 1116, 1119, 741, 1446, 1035, 1191, 2121, 1365, 739, 1303, 1597, 1680, 1750, 1791, 1195, 928, 1535, 986, 1404, 1710, 1157, 1790, 1535]. Digitisation [867]. digitised [839]. Digitizing [1417]. dilemma [2214, 1388]. dinosaurs [548]. Directions [486, 278]. Directive [1018, 902]. Directory [826]. disciplinary [362, 2227]. discipline [1829, 1735, 1841, 2203]. disciplines [1355]. disclosing [1178]. discomforts [1888]. discontinue [2223]. discount [638]. discounted [1575]. discourse [2162, 2117]. discoverability [1868, 1872, 1865, 1877, 2016]. discovering [1869]. Discovery [1876, 1813, 1870, 1868, 1869, 1875, 1321, 1762, 1871, 1874]. discussion [1637, 2028, 2224, 2045]. disk [157, 447]. diskette [619]. disks [243]. disputes [1595]. disruptor [2013]. disseminate [2202]. disseminated [2182]. Disseminating [52]. dissemination [461, 2046]. dissertations [291]. Dissimulate [2202]. Distinguishing [1542]. distributed [180, 447, 952]. Distribution [68, 346, 198, 11, 49, 352, 1001, 1008, 29]. Diversity [1893, 1998, 1949]. DIY [431]. Do [1575, 1512, 1244, 1993, 1930, 1994, 2167, 1686, 970, 1375, 1559, 1348, 1258, 586, 1385, 1996, 682, 1992, 2223, 899, 1107, 1939, 1544, 1346, 1324, 1613, 1200, 1053, 1195, 1582, 2021, 1352, 1799, 1767, 1874, 2142, 1339]. DOAJ [1263, 2023]. Document [429, 1114, 381, 157, 103]. document-delivery [381]. Documentation [1262]. documents [1136, 1209, 96, 155]. DOE [1727]. Does [919, 2002, 2074, 1882, 40, 2015, 2112, 1465, 546, 1803, 1150, 2124, 1049, 1356, 1986, 1800]. DOI [1732]. doing [1512, 604, 685, 1747]. dollars [1353]. Domesday [10]. Donation [531]. done [1465]. Donovan [2113]. don’ts [1060]. doomed [672]. door [433, 1220]. Doreen [696]. Dorner [355, 447, 734]. Dorothy [52, 155]. Dorrington [120]. Dorset [356]. Double [1319, 1557, 2085]. Down [1217, 1422]. download [1944, 1926]. download-and-go [1944]. downright [1888]. downturn [1487]. draw [1419]. drawer [1730]. Drayton [754, 53]. drift [2087]. drive [1997, 1913, 2209, 2059]. driven [1226, 1794]. drives [1282]. Driving [1158]. drop [1235]. drug [1970, 1583]. DSC [123]. DTP [500]. Ducklings [1683]. due [1461]. Dumpty [1322]. during [2169, 2188, 2126, 2082]. Dworkin [289]. dynamics [2155].

e-based [1405]. e-book [1656, 1917, 2222]. E-books [1401, 1110, 1564, 1546, 1478, 1160, 1490]. E-citations [963]. E-commerce [902]. e-delivery [1048]. e-journal [1370]. E-journals [1532, 1123, 1501]. E-learning [1106]. e-library [1160]. e-mail [602]. E-only [1235]. e-print [1064]. E-prints [1304, 1211]. e-products [1549]. e-publishing [1095]. each [1843]. EAJUS [1021]. Early [1808, 2075, 2193, 1971, 1887, 1896, 430, 2034, 2002, 2104, 2089, 1867, 1957, 2027, 2083, 2045, 1906, 2021]. Early-career [2193]. earth [2205]. EASE [323, 1520]. easier [1918, 582]. East [1699, 531]. Eastern [248, 126]. easy [597]. Ebook [1625, 1980]. ECDIN [184]. Ecological [2058]. economic [900, 1993, 1061, 1487]. Economics [592, 487, 181, 1134, 1013, 382, 504, 207, 566]. Economies [1848]. ecosystem [1668, 1980, 1944]. ecstasy [1093]. ECU [181]. ed [240, 331, 697, 117, 219, 156]. EDI [920]. EDICIC [2181]. Edinburgh [1985]. Edit [584, 1324]. Edited [449, 240, 544, 638, 314, 353, 448, 473, 524, 678, 352, 716, 658, 397, 312, 310, 497, 679, 286, 220, 755, 734, 120, 115, 74, 137, 73, 52, 117, 95, 239, 91, 203, 1901, 93, 114, 55, 138]. Editing [166, 592, 1108, 323, 670, 707, 364, 721, 682, 155, 330, 474, 1408, 521, 1318, 1843, 474]. Edition [523, 617, 53, 716, 91, 96, 311, 619, 138, 621]. Editor [523, 170, 321, 1670, 1764, 1752, 851, 620, 1430, 1952, 719, 337, 2144, 2140, 1948, 837, 921, 1795, 944, 968, 1153, 1264, 1297, 922, 1337, 1037, 979]. editor-selected [1952]. Editorial [224, 769, 934, 1098, 277, 341, 1026, 2096, 1301, 1, 1100, 788, 1617, 1747, 1969, 1101, 2180, 686, 1579, 57, 1898, 2090, 360, 2173, 242, 2138, 619, 1603, 1774, 503]. editorials [2162]. Editors [275, 399, 449, 543, 277, 496, 2034, 2180, 1788, 1753, 1785, 697, 1699, 124, 1152, 1276, 91, 2177, 2213, 1469, 1994, 2045, 1786, 2042, 1004, 1995, 306, 1969, 2038, 1642, 1816, 2224, 1362, 1442, 618]. edn [598, 583]. education [2156, 1923, 1001, 2230, 1211]. educational [2012, 1922, 1882, 2108]. Edward [697, 239]. EEI [598]. Effect [1628, 1419, 904, 1997, 2143, 1959, 1214, 940]. effective [3, 2207, 204, 585, 1462, 119]. Effectiveness [1618, 1163]. effects [1104, 1991, 2179, 266]. efficient [972, 900]. effort [2063]. efforts [1136]. Ege [2109]. eight [1486]. Eighteenth [551]. eJournals [1070]. electro [408]. electro-copying [408]. electro-manipulation [408]. electro-storage [408]. Electrocopying [193, 706]. Electronic [68, 13, 28, 44, 786, 976, 480, 645, 555, 1025, 268, 742, 930, 1021, 408, 725, 605, 342, 936, 666, 570, 915, 685, 799, 773, 952, 1077, 706, 737, 785, 865, 1357, 943, 826, 638, 1376, 964, 454, 1059, 838, 935, 898, 1040, 1001, 553, 35, 1028, 1043, 1007, 1006, 205, 312, 868, 759, 1060, 550, 1061, 1083, 765, 822, 814, 1124, 800, 813, 313, 1105, 1277, 1086, 1761, 987, 1122, 734, 831, 845, 738, 718, 986, 782, 590]. Electronics [607, 323, 592]. Eliminate [1901]. elite [2123]. Elizabeth [73]. Ellis [200]. ELP [261]. elsewhere [1257]. email [1070, 1292]. embargo [2219]. embrace [1129]. Embracing [1983, 1501]. Emerald [1432]. emergence [313]. Emerging [1760, 1720, 1841, 2084, 1110, 2038, 2174, 1950]. emotional [1793]. empirical [1187, 2217, 2133]. empowering [1790]. enabler [1789]. enabling [1808]. Encouraging [1823]. Encyclopedia [240]. end [524, 1733, 1778, 1090, 1944]. end-user [1090]. end-users [524, 1778]. endangered [1152, 755]. Enduring [2197]. Energy [2092]. engagement [1681, 1810, 2016]. engaging [2046, 1719]. Engineering [1406, 1257]. Engineers [298, 276]. England [2183]. English [2155, 1736, 1281, 1827, 1883, 1967, 1172, 1259, 2010, 1063, 2014]. English-language [2155, 1736, 1172, 1259, 2010, 1063]. enough [1296]. enriching [1637]. ensure [1998]. Enterprise [642]. enthusiasm [1059]. entities [2227]. environment [1069, 910, 1331, 1191, 982, 1657, 1597, 1195, 1909]. Environmental [797, 1406, 2205, 989]. environments [1504]. EPIC [986]. epidemiology [482]. Epsilon [51, 6]. Epubit [2133]. equally [1595]. equity [2101]. era [1376, 1966, 1992, 1854, 1365, 1006, 759, 1972]. erina [354]. Errata [2138]. Erratum [1170]. ESPERE [782, 718]. Essay [1141]. Essays [658, 540]. essential [1408]. established [1854, 989, 1786, 2029]. establishing [1500]. estimate [1505, 428]. eTextbook [2120]. Ethical [2156, 682, 1682, 1779, 1249]. Ethics [592, 999, 1541, 1537, 720, 1009, 543]. EUR [476, 181, 472, 523]. EURO [904]. Europaea [362]. Europe [449, 531, 299, 388, 129, 248, 126]. European [449, 261, 181, 2157, 364, 1228, 1982, 603, 902, 1766, 459, 562]. Europeana [1680]. EuroScience [1216]. Euston [318]. Evaluating [2154, 1720, 1917, 1550, 2101, 1610]. Evaluation [2093, 2178, 1841, 1832, 1364, 2070, 1734, 1937, 2195, 1144, 2014, 1000]. evaluations [1956]. Evans [563]. Event [1394]. events [19, 37, 56, 77, 97, 121, 142, 158, 185, 206, 223, 241, 263, 293, 315, 334, 357, 379, 402, 422, 451, 477, 501, 525, 545, 564, 587, 600, 622, 639, 660, 681, 699, 717, 735, 744, 768, 787, 803, 833, 846, 861, 877, 889, 907]. ever [2021]. every [1416]. everything [1265]. Evidence [1872, 1861, 2144, 1882, 1573, 1395, 1883]. Evidence-based [1872]. Evolution [2165, 1371, 1758, 686, 988, 1942, 2041, 1394, 1041, 1765, 1391, 1761, 1837]. evolve [548]. Evolving [925, 1156, 755, 1504, 1759, 1657, 1428, 1114]. exactly [1019]. exaggerated [1754]. Examination [2141, 1721, 854]. examining [1730]. Examples [2209]. Excellence [169, 343, 575, 253, 1416]. Excellent [1642]. exceptions [1390]. exchange [1792]. exchanging [1419]. Excited [124]. excluded [2023]. exclusivity [827]. exercise [2144, 1165, 940]. exercises [330]. exhibit [593, 629]. exhibiting [119, 601]. exhibition [401]. Exhibitions [19, 283, 606]. exhibitors [401]. existence [1452, 1304, 1505]. Exorcising [793]. Expanding [2220]. expectations [1042, 2136]. expenditure [2189]. expensive [1798]. experience [2161, 1190, 1705, 588, 1485, 1417, 161, 1124, 1793, 1906, 1499, 2181]. experienced [1922, 124]. Experiences [1788, 2169, 1971, 2131, 1746, 1929, 2012, 1807]. experiment [1070, 1849, 1514]. Experimental [2228]. experimentation [1329]. experiments [1873]. expert [354]. expertise [1871]. explained [1805]. exploding [1421]. exploiting [761]. exploration [2227]. exploratory [1517, 2040]. Exploring [2036]. Expressions [1380, 2138]. Extending [1369]. extensions [2220]. extent [1577]. Extenza [1163]. extraction [1675]. eye [1219, 1558]. eyes [2213].

f50p [399]. face [1418]. Facebook [2072]. faces [2194]. facilitators [1004]. facing [1816, 232]. fact [857]. factbases [1369]. factor [2076, 1470, 1959, 953, 1548]. Factors [2076, 2094, 2108, 1469, 1604, 2131, 1938, 1214, 1973]. facts [1990]. faculty [1586, 1420, 1501, 1144]. Fad [1834]. failed [1912]. failing [1927, 1826]. Fair [449, 924, 923, 1781, 231, 1780, 1178, 2205, 2031, 411, 511, 574]. fairness [581]. Fairs [19]. fake [2162, 1731]. fall [1794]. Fallacy [1421]. Falling [1953]. family [458]. family-style [458]. far [1925, 1219]. far-reaching [1925]. Fashion [1403]. fate [2202]. fault [1857]. favela [1982]. Fear [2162]. fears [480]. feature [780]. features [2132]. Feb [385]. febrile [1979].

February [299, 724, 742, 35]. federal [1679]. Federation [119, 419, 544, 314, 353, 498, 352, 475, 1832, 288, 400, 220]. feedback [2170]. Feeney [678, 310]. fees [345]. Feisty [1979]. Feldman [313]. few [1353]. fewer [890]. fiction [857, 332]. FID [161]. Field [292, 2077, 1843]. fields [1818, 2140]. Fifth [772, 129]. FIGARO [1132]. Fighting [1341]. FIGIT [687]. Figuring [1780]. file [7, 1730]. files [1868, 1732]. filtering [2068, 1144]. Final [181, 2035]. finally [1203]. finance [1706]. Financial [83, 1295, 2214, 1787, 1827]. find [1044, 2004, 1867, 1488]. Finding [360, 913, 1869]. findings [1523, 1340, 1630, 2122]. finished [484]. Finland [448]. Finnish [2012, 2184]. First [483, 1578, 275, 399, 1281, 962, 1595, 1690, 872, 2085, 1432, 2137]. first-class [2085]. fish [1306]. fit [1398, 1727]. five [2071, 1420, 389, 1773, 404, 757]. Flanders [1786]. flat [1518]. flow [261, 100]. Floyd [95]. focus [1340]. focused [1873]. Follett [546]. food [2077]. footprint [1600]. footsteps [2204]. force [626]. forces [984]. forcing [984]. foreign [1145, 39, 1675]. foreign-language [1675]. forgive [1582]. form [1028, 1430]. formal [2093]. formalization [1672]. format [697]. formats [454, 1625, 794]. forming [1071]. forms [1043]. Fortification [2231]. Forum [300, 630, 104, 251, 1216]. forward [232]. foster [2096]. Foundation [747, 279, 542, 280]. Four [2097, 2140]. Fourth [716, 91]. fowl [1306]. framework [1116, 1682]. Frameworks [2025]. France [1343]. Francis [458, 956, 1417]. Francisco [232]. Frank [117]. Fraud [719, 1639]. fraudulent [2216]. Free [53, 1059, 1083, 754, 1096, 1598, 401, 1080, 1008, 1514, 1089, 1488, 51]. freedom [949]. Freelance [275, 399, 449, 277, 496, 361]. freelancers [437, 277]. Freitas [287]. French [1629, 1966]. fresh [425]. freshwater [1938]. friendlier [2201]. frill [1408]. front [2013]. Frontiers [1406, 1683, 1745]. frontline [1746]. Frost [74]. FST [724]. FST/ [724]. Full [1875, 1706, 1087, 55, 1122, 1409, 938, 583]. full-text [1122]. function [1318]. functionality [1196, 1808]. Fund [355, 313]. funded [927, 1260]. funder [1636, 1861]. funders [1706, 2207, 1283]. funding [82, 1748, 1467, 1679]. funds [1307, 1599]. furies [1888]. further [1449, 1211]. fuss [2000]. Future [808, 278, 807, 831, 1342, 1638, 1923, 1561, 461, 552, 798, 1641, 1057, 1983, 82, 2200, 548, 827, 1353, 1942, 2041, 336, 1450, 1862, 2056, 1801, 1810, 1815, 1856, 547, 202, 457, 114, 567, 1356, 1958, 928, 796, 78, 1685, 1919, 1317, 1834, 590, 1271, 93]. futures [1146].

GA [521]. Gail [475]. gain [1686]. Gaining [1499]. Gale [240, 117]. game [1681]. gamification [1681]. gap [1873]. garbage [1732]. garde [1228]. Gardens [754, 53, 401]. gateways [952]. gathered [1955]. Gender [1897, 2152, 1823, 2174]. General [167, 105, 523, 1719]. Generalized [79]. generate [2169]. generation [2168, 1857, 672]. Geneva [266]. genre [2230]. Geometers [472]. George [716]. Georgia [521, 2044]. GeoScienceWorld [2059]. Gerald [289]. German [1857]. German-language [1857]. Germany [1136]. get [815, 1570, 1767]. Getting [994, 1044, 1965, 1626, 1059, 643, 472, 522]. ghostwriting [2039]. Gibson [696]. Gillian [34]. Giving [1377, 585]. GL54 [332]. Glasgow [679]. glimpse [958]. Global [2105, 1415, 2083, 1958, 1292, 2117, 1574, 1920, 1995, 288]. Globalization [1777]. globally [1742]. Glos [332]. Glossary [500]. GNSI [617]. go [593, 1559, 1658, 1944]. Gober [495]. Godalming [399, 496]. Going [1019, 1330, 2167, 1653, 584]. Gold [1889, 1652, 1912, 1693]. golden [1199]. goldrush [1199]. Goliaths [2107]. Good [894, 565, 279, 1125, 1162, 80, 733, 1167, 657, 1642]. goodbye [1090]. Google [1561, 1720, 1655, 1413, 1550, 2014]. Googled [1148]. Gordon [540]. Got [1570, 1432]. Government [815]. Gower [155, 677]. Graham [540, 118, 120, 115, 137, 201, 448, 473, 476, 524, 397, 354, 497, 55]. Grange [399, 496]. Granham [658]. grass [437]. Great [137, 1064, 1890, 1259]. greatly [1754]. Green [1693, 1889, 1573]. greener [437]. Greenhalgh [75]. grey [1773]. Grieves [678]. Group [343, 785, 324, 74, 137, 261, 53, 73, 716, 266, 575, 679, 657, 755, 2215, 1340, 708]. groups [585]. grown [588]. grows [1848]. Growth [1951, 424, 988, 1777, 2059]. GST [300]. GU7 [399, 496]. guard [1840]. Guardians [581]. Guest [1, 1100]. Guide [873, 156, 94, 541, 754, 200, 2055, 563, 801, 91, 155, 157, 205, 219, 474, 204, 619, 289, 287, 119, 498]. Guidelines [894, 24, 918, 2156, 180, 208, 1722, 2187, 971, 1520, 657, 2141, 865, 53, 139, 585, 734]. guides [1658]. guiding [1300]. gulfs [870]. Gurnsey [677]. Gutenberg [159, 1404]. Gutenberg-e [1404]. Gynecology [2035].

H [581, 377]. H.-D [581]. habits [1894]. HAL [1343]. half [182, 356, 1220, 1609]. Hall [716]. Halstead [200]. hamburger [1563]. Hampton [311]. Hancock [524]. Hancock-Beaulieu [524]. handbook [201, 401, 447, 733, 220]. Handing [850]. Handley [698]. handling [2149, 55]. hands [2181]. hands-on [2181]. Hans [540, 377]. Hanson [584]. Hants [155]. happen [1103]. happening [1543, 1254, 2016]. harbingers [2027, 2074]. hard [1918, 1107, 624, 1159]. Hardback [94, 638, 678, 697, 472, 93, 523, 447]. hardcover [221]. hare [702]. harmless [1849]. harmonised [800]. Harris [658, 397, 330]. harsh [2038]. Hartley [1534]. Harvesting [1071]. Hass [203]. Hatfield [157, 474, 447]. having [1997]. Haworth [73, 76]. Hazel [755]. hb [716, 522]. hdbk [581]. headers [641]. Headlines [1868, 563]. Heale [474]. health [2109, 1669, 2072]. healthcare [1309]. Heeks [118]. HEI [916]. held [1216, 449, 354, 312]. Héloïse’ [1668]. help [426, 1575, 1908, 1471, 2045, 815, 1675]. helped [1929]. helping [1998, 1044, 1843, 1821]. Helsinki [202]. Henry [377]. herbals [1005]. here [1559, 1625, 1128, 1665, 1003, 1767]. heroic [2073]. HERON [916]. Herts [51, 157, 330, 474, 447]. Herzegovina [430]. Hewson [221]. hidden [1314]. High [1695, 2156, 2229, 2182, 619]. High-impact [1695]. high-technology [619]. higher [1923, 1001, 2230, 1211]. highly [2182, 1959]. highway [608]. HighWire [1649, 1226]. hijacking [2196]. Hills [55].

Hindawi [1330]. hindered [1929]. HINTS [2022]. historical [1175, 2150]. histories [2230]. History [1073, 1225, 1870, 757]. hit [779]. Hitting [563]. hobby [432]. Hodge [475]. hold [1356]. Holistic [1877]. Holloway [182]. Home [868, 2183]. Homewood [619]. honesty [670]. Honorific [2190]. honour [658]. hope [546]. Horwood [200]. Horwood/Halstead [200]. hospital [2088]. host [1984]. hot [1209, 1483]. Hotbot [1005]. House [707, 726, 748, 330, 734, 1809, 551, 407, 92, 1327, 772]. Hua [1281]. Hub [2044, 2013]. human [249, 1941]. Humanities [809, 1486, 2155, 409, 1641, 1171, 1075, 35, 1724, 1786, 1953, 2036, 987, 2174]. Humpty [1322]. hunger [63]. hungry [1090]. hurt [2030]. Huth [697]. hybrid [1706, 1056]. Hyman [138]. Hyper [239]. hyperlinking [992]. Hyphenation [259, 259].

Ian [115, 658]. IASP [202]. Iberian [2181]. Ibero [1698]. iceberg [761]. ICEDIS [920]. ICMJE [2190, 1960]. ICSU [139, 222, 742]. idea [1111, 1203]. ideals [1556]. Ideas [1289]. Identification [1490]. identified [2025]. identifier [1940]. identifiers [1209, 963, 780, 781]. identify [1627]. Identifying [1755]. identity [2186]. iDs [1955]. IEE [1197]. IEPRC [638]. IEPRC/Pira [638]. If [2070, 1650, 1582]. IFLA [201]. ignorance [203]. II [1979, 705]. III [4]. Illus [541, 36]. illustration [576, 617]. image [157, 826]. images [826]. imbalanced [1725]. immunologist [702]. Impact [2186, 1548, 2154, 2156, 2122, 678, 15, 1221, 2225, 2143, 1470, 2173, 1695, 101, 1469, 1604, 1974, 2217, 1959, 1303, 953, 2185, 1514, 1877, 1214, 1866, 1223, 1284, 2118, 1973, 2011, 2218, 2076]. impacting [2108]. impacts [1008, 365]. impaired [1450]. impeding [2219]. imperative [1294]. Imperial [2158]. implementation [1578, 607, 1389]. Implementing [2205, 1705, 1531, 1807]. Implications [2144, 1250, 925, 1061, 547, 953, 570, 2038, 2176]. importance [1030, 2207, 2209]. important [1341, 628, 1576]. importation [163]. impressions [2137]. IMPRIMATUR [905]. imprint [739]. Improve [1901, 1684, 2218, 2231]. improved [1173]. Improving [2170, 1850, 1326, 1565, 1571, 1367, 1046, 1600, 987, 1416]. IMRD [2176]. IMRDi [2176]. in-house [407, 92, 1327]. incidence [2138]. inclusion [1522]. inclusive [1919]. inclusivity [1893]. income [1076, 1082]. inconsistently [2144]. increase [2225, 1292]. increases [1673]. increasing [1636, 1565]. incubator [1954]. independent [1636, 1764, 1753, 541, 1848, 2056, 2176, 2059]. Index [171, 637, 239, 495, 541, 2043, 522, 423, 1618, 210, 1950]. indexed [2178, 2215, 2144, 2108, 1488]. indexer [438]. indexer* [483]. Indexers [696, 203, 620, 620]. Indexes [240, 1395, 2177, 96, 1883, 36]. Indexing [754, 696, 529, 505, 760, 1005, 229, 390, 955, 210, 338, 244, 620, 498, 475, 696, 203, 620, 377]. India [1643, 1695, 2219, 1615, 1693]. Indian [1721]. indicators [2154, 1895]. Individual [237, 183, 1574]. industrial [137, 295, 569]. Industries [419]. Industry [276, 1311, 900, 1809, 647, 2122, 2116, 1021, 2002, 1451, 1820, 1240, 1590, 1958, 295, 1583, 1826, 506, 2164, 288]. industry-wide [1021]. inequality [2174]. infectious [482]. influence [1266, 1469, 1550, 1294]. influences [1949]. influencing [2094]. Information [27, 46, 69, 88, 109, 153, 174, 197, 215, 236, 172, 255, 284, 307, 328, 351, 395, 417, 445, 469, 493, 518, 537, 559, 578, 595, 614, 635, 654, 676, 693, 714, 731, 752, 772, 831, 1811, 1824, 1838, 1851, 1863, 1879, 1891, 1902, 1914, 1932, 1946, 1962, 1977, 1988, 2006, 2019, 2032, 2049, 2050, 2067, 2081, 2098, 2099, 2114, 2066, 2080, 2115, 2129, 2151, 2172, 2199, 2211, 780, 119, 240, 544, 184, 498, 678, 181, 352, 475, 1209, 165, 203, 434, 157, 1657, 472, 781, 400, 135, 679, 374]. Information [220, 2154, 1466, 1069, 1092, 461, 556, 608, 231, 938, 120, 268, 314, 355, 353, 448, 1900, 73, 116, 285, 1816, 1720, 740, 267, 294, 140, 146, 397, 280, 610, 1090, 102, 561, 239, 1015, 1044, 1420, 950, 100, 568, 1669, 155, 312, 412, 497, 620, 687, 2094, 912, 1867, 288, 815, 265, 661, 99, 1367, 1158, 955, 317, 383, 569, 36, 625, 1004, 566, 589, 1937, 297, 365, 882, 1890, 743, 2017, 678, 476, 524, 352, 310, 679, 55, 220]. information-hungry [1090]. information-seeking [1420]. informed [2156, 1956, 1869, 784]. infrastructure [2127, 1999, 1071]. infringement [872]. Ingwersen [476]. initial [1776]. initiated [1707]. initiative [2167, 1707, 1820, 2096, 930, 811, 1071, 1870, 986]. initiatives [605, 547, 1303, 1672]. Innovation [1358, 2058, 1184, 707, 726, 748, 1448, 1333, 1316, 1226, 2082]. innovation-oriented [1316]. innovations [2120]. innovative [1926, 1820, 1122]. input [1941]. inquiry [871]. insight [1573]. Insights [2166, 2089, 2054]. Instant [1042]. Institute [2057, 932, 291, 276, 292, 497, 917, 2017, 885, 901]. institutes [1923]. institution [99, 298]. Institutional [237, 1894, 1220, 1174, 2153, 1178, 2040, 2186, 2230, 914, 1981, 1262, 1500, 2147, 1923, 1467]. institutions [1984, 1405, 183]. instruction [598]. instructions [1779]. instrument [1802]. integrated [898, 1645]. Integrating [1584, 1359, 1479]. integrity [1743]. intellectual [943, 1068, 1221, 481]. intelligence [1813, 1833, 473]. Intelligent [1956, 473]. intentions [2133]. interaction [1430, 476]. Interactive [1173, 1029]. interdisciplinary [1400]. Interest [73, 859, 844, 887, 872, 2226]. interests [1240, 1377]. interface [249]. interlending [201]. interlibrary [793]. intermediaries [813, 786, 831]. Intermediary [805, 829]. International [249, 385, 302, 449, 51, 129, 638, 201, 1699, 52, 487, 725, 161, 649, 1827, 202, 266, 139, 1192, 1241, 1336, 743, 420, 572, 737, 540, 353, 448, 364, 1777, 2075, 989, 1895, 1660, 2139, 190, 1261, 1487, 1407, 319, 2108, 1115, 2011, 920, 98, 1203, 222]. international-indexed [2108]. Internationalization [1818, 2177, 2010, 1515]. Internationally [800]. internet [1992, 845, 1148, 612, 724, 852, 823, 988, 1070, 761, 705, 779, 529, 702, 766, 743, 1080, 1212, 544]. internet-era [1992]. interpret [428]. interpretation [1180]. Interview [1649, 1657, 2042]. interview-based [2042]. introduced [916]. Introducing [2173, 1093]. Introduction [1853, 489, 115, 221]. Investigating [2152, 2212, 1777]. investigation [2035]. invitation [640, 1799]. involvement [190]. IOP [2173]. IP [932, 1446, 917]. IPA [266]. IPA/STM [266]. IPR [1341]. Iranian [1814, 1967, 1843]. Irwin [619]. ISBN [94, 562, 238, 399, 449, 541, 540, 51, 54, 119, 200, 180, 543, 74, 137, 201, 331, 314, 355, 353, 448, 498, 473, 476, 524, 542, 678, 697, 73, 181, 352, 475, 52, 117, 260, 421, 716, 586, 563, 75, 332, 599, 95, 140, 221, 658, 76, 397, 561, 598, 239, 354, 696, 91, 398, 155, 157, 205, 312, 310, 376, 497, 620, 581, 219, 330, 259, 472, 447, 266, 288, 287, 93, 204, 34, 114, 139, 222, 311, 400, 583, 619, 618, 584, 55, 36, 313, 286, 258, 138, 289, 523, 677, 496, 521, 377, 495, 582, 92, 156]. ISBN [220, 617, 522, 585, 657, 755, 659]. ISI [1785, 738, 1364]. Isis [734]. island [1866]. ISO [200, 1499]. isolation [816]. ISSN [637, 237, 285, 222]. Issue [2050, 2067, 2081, 2098, 2099, 2114, 2115, 2129, 2151, 2172, 2199, 2211, 1795, 626]. issued [2137]. Issues [678, 486, 689, 710, 727, 749, 769, 858, 843, 886, 867, 1922, 116, 352, 1272, 665, 658, 76, 1162, 701, 1213, 936, 194, 288, 783, 765, 830, 869, 2137, 800, 902, 2185, 788, 684, 1267, 534, 557, 577, 594, 613, 653, 674, 896, 1779, 524]. iv [399, 354, 266, 845].

J [697, 52, 312]. J. [140]. Jack [34, 1890]. Jacket [1202]. Jan [385]. Jane [355, 447, 734]. January [697, 280, 251]. Japan [282, 2153, 2, 209, 828, 1643, 1695, 1007, 2089, 841]. Japanese [829]. JATS [1845]. JBJS [1802]. Jeffrey [2111]. Jill [1065]. JIME [1029]. JISC [1178, 1069, 1260]. JISC-funded [1260]. Joan [200]. job [1753]. jobbing [361]. John [754, 260, 421, 221, 354, 696, 310, 474, 472, 447, 114, 583, 258, 156, 585, 689, 710, 727, 749, 764]. join [640]. Joint [62, 1485]. jointly [266]. Jon [473, 698, 266]. Jonathan [586]. Jones [94]. Joseph [156]. jourals [630]. Journal [343, 1811, 1824, 1838, 1851, 1863, 1879, 1891, 1902, 1914, 1932, 1946, 1962, 1977, 1988, 2006, 2019, 2032, 2049, 2066, 2080, 894, 512, 424, 671, 237, 798, 456, 264, 487, 486, 1655, 1673, 2143, 927, 2177, 531, 1708, 21, 1387, 2196, 1602, 575, 359, 98, 2008, 1587, 210, 1950, 1784, 1327, 896, 1909, 2183, 2034, 1091, 984, 1329, 169, 321, 2161, 2169, 2154, 2152, 1247, 941, 367, 461, 533, 650, 1005, 1165, 54, 247, 279, 1956, 2026, 229, 384, 825, 331, 852, 1559, 1699, 2055, 2216, 549, 2028, 1021, 662, 1545, 1381, 1575, 484, 1577, 958]. journal [1452, 1945, 1031, 3, 2025, 1472, 1151, 2076, 168, 1134, 1831, 1993, 1370, 2186, 1187, 779, 1304, 1162, 1848, 1417, 1213, 2221, 2043, 1588, 1832, 1571, 1732, 720, 989, 2220, 914, 925, 1315, 1420, 2219, 2095, 2213, 1469, 2217, 1598, 1281, 1952, 1883, 1959, 1589, 489, 510, 1269, 1339, 1060, 1971, 591, 1061, 1161, 1224, 1263, 1346, 1398, 2187, 1972, 1237, 1719, 270, 382, 504, 1108, 2042, 187, 822, 2137, 1173, 1382, 2017, 1253, 1807, 1292, 1550, 1603, 1624, 1414, 1734, 1029, 1055, 1690, 895, 939, 1270, 225, 345, 1344, 1444, 1211, 1214, 1882, 987, 1120, 940, 2195, 1284, 460]. journal [1725, 1973, 1047, 1115, 1631, 2218, 786, 2162, 212, 1956, 1655, 1548, 720, 1994, 931, 637, 419, 2228, 1485, 868, 1281, 34, 1225, 1047]. Journal.fi [2184]. JournalGuide [1768]. journalism [1831]. Journals [903, 302, 129, 1164, 487, 725, 1095, 389, 649, 2215, 912, 1192, 1241, 1368, 1428, 1046, 684, 1347, 210, 404, 2023, 572, 1145, 984, 785, 2154, 2156, 2155, 2135, 1268, 973, 250, 724, 890, 51, 248, 1057, 1273, 1706, 1760, 1800, 1785, 271, 1816, 2178, 1371, 1272, 964, 1644, 209, 1842, 1546, 1059, 1381, 2078, 1655, 2090, 1111, 1503, 2214, 1908, 705, 1996, 1450, 1949, 1777, 1814, 76, 1692, 61, 1134, 810, 778, 2224, 1755, 857, 2173, 1643, 1240, 2223, 1276, 1729, 1522, 1093, 1691, 1386, 2093, 1007, 1736, 1718, 1939, 1844, 975, 1234, 1430, 1688, 1711, 1827, 1895, 80]. journals [39, 2146, 1967, 1901, 550, 1501, 188, 1731, 1708, 1886, 528, 1083, 866, 1351, 1815, 1768, 1722, 1289, 1303, 1532, 2144, 1751, 570, 971, 2176, 2184, 2140, 430, 733, 1128, 1296, 774, 1172, 2070, 2185, 794, 1698, 2069, 1407, 1286, 1897, 1295, 1970, 1953, 1204, 1267, 5, 897, 1123, 1881, 2029, 1954, 623, 1308, 78, 1675, 893, 909, 1084, 1086, 957, 1027, 1486, 1940, 1951, 2203, 1235, 506, 657, 1077, 1779, 1757, 1259, 1271, 1576, 1616, 2010, 2086, 1480, 1063, 1671, 2014, 2011, 785, 1379, 1358, 1499, 1954]. Journey [796, 1928]. Journeyman [586, 495, 522]. JSTOR [1919]. Judges [785]. July [748, 248, 354, 1603]. June [746, 324, 747, 247, 73, 181, 696, 497, 202, 232, 621]. JUSP [1545]. Just [987, 1652].

Kanban [1850]. Kansas [1363]. Kathy [92]. Kay [95]. keep [1028]. Keeping [1955, 1328, 784]. Keith [495]. Key [311, 2164, 1055, 1868, 1611, 311]. Keynes [290, 291]. keyword [2029]. killers [1064]. king [1750]. Kingdom [708, 2209]. Kirkman [583, 585]. Kirty [156]. kiss [1129]. Kitchen [2056]. KL [183]. knot [2047]. know [1512, 1470, 1339, 2191, 1346, 736, 1650, 1582]. Knowing [1502]. Knowledge [808, 2163, 1602, 1367, 807, 420, 249, 509, 2150, 1394, 1365, 203, 1669, 1981, 2191, 1763]. knowledgebases [1369]. known [220]. Kogan [599, 204, 621, 659]. Korea [282, 1643, 1695, 1905, 2017]. Korean [1832].

L [521, 1283]. label [1667]. labour [2191]. Lack [1899, 1733]. Landscape [2171, 1309, 1961, 2084, 1921, 1801, 1759, 1346, 1741, 1236, 1279, 1261]. Language [183, 2155, 1857, 884, 1736, 1827, 1296, 1172, 1675, 2203, 1259, 2010, 1063, 79, 183]. large [2130, 870, 2215, 585]. large-scale [2130]. larger [1821]. largest [532]. Laser [2079]. late [433]. latest [1957]. Latin [1268, 1503]. Launching [1344, 1598]. Laurence [117]. law [230, 929, 1901, 287, 1781, 1513, 1765]. Lawraine [137]. lawsuits [2044]. Lay [1719, 1614]. LC [532]. leading [1899, 1970]. lean [1850]. leared [630]. learn [1183, 553]. Learned [1878, 191, 1696, 302, 650, 129, 180, 212, 1164, 487, 838, 871, 725, 203, 649, 1185, 510, 866, 1434, 1192, 1241, 756, 251, 743, 1704, 98, 345, 572, 1260, 2167, 461, 724, 51, 247, 1641, 271, 15, 130, 168, 213, 610, 969, 242, 1787, 2052, 489, 1339, 1971, 208, 265, 567, 187, 227, 1286, 207, 228, 1033, 1239, 225, 623, 893, 909, 957, 1302, 504, 358, 981, 1039, 1099, 758, 1774, 1934]. Learning [808, 1589, 778, 1106, 1124, 1941, 807]. learnt [2206]. Leatherhead [250, 638]. led [1756, 991, 1815]. left [1856]. legacy [2160]. Legal [160, 318, 159, 1249, 160, 210]. legislative [1223]. legitimate [1755].

Leicester [75]. Leigh [73]. Lending [201]. Lending/UNESCO [201]. Leon [679]. less [2215]. lesson [2121]. Lessons [1917, 1257, 1791, 2206, 1918, 1971]. Let [1984]. Letchworth [51]. Letter [851, 1153, 1264, 1297, 922, 1362, 1337, 1037, 1442, 837]. Letters [921, 959, 944, 968, 979, 2144, 607, 1697]. level [1637, 2106, 2218]. Levine [288]. Lewis [619]. Liabilities [873]. liaison [62, 66]. libel [795]. Librarian [73, 1546, 76, 1401, 708, 603, 1762, 336]. Librarians [261, 452, 626, 412, 792, 369, 548, 588, 1017, 1532, 1562]. librarianship [397]. librarie [1377]. Libraries [746, 551, 687, 1638, 1268, 191, 1357, 688, 1222, 1629, 542, 546, 964, 1705, 1790, 1131, 914, 925, 1017, 568, 1185, 604, 550, 547, 1468, 1745, 1487, 317, 1710, 2222, 673, 137, 354, 376]. Library [344, 937, 1196, 118, 1376, 137, 678, 116, 738, 801, 278, 561, 553, 35, 312, 310, 376, 114, 985, 1954, 1876, 2154, 1561, 532, 1056, 82, 448, 1378, 181, 1756, 1816, 1448, 910, 658, 741, 76, 1160, 1748, 186, 2094, 401, 765, 144, 1328, 883, 1858, 1114, 755, 1874, 110, 254, 509, 884, 1403, 2048, 1345, 2120, 116]. library-led [1756]. Licence [947, 930, 865, 811, 662, 1536]. licences [1062]. Licensing [706, 251, 435, 734, 1519, 840, 552, 631, 1107, 856, 457, 252, 763]. lie [682]. life [608, 1366, 708, 667, 1866, 957, 1377, 687]. lifeblood [1068]. lifecycle [1232]. Light [1709, 1300]. like [2015, 995, 1603]. likely [2180]. Lilliput [870]. limitations [2044, 2073]. Limited [599, 974]. Linda [120]. line [2220, 774, 1679]. Lines [621]. Linford [290]. Linguistic [2228, 2029, 2130]. link [369, 1795, 316, 1090, 1131, 1930]. linked [982]. Linking [1087, 985, 1504, 1743, 954, 935, 1042, 1199]. links [1015]. LIS [278]. list [1670, 80, 2215, 58]. listened [2111]. listening [777]. Listings [959, 240]. Litchfield [844]. literacy [2170, 1092, 1920]. literary [1254]. literature [1175, 1996, 2084, 1610, 1315, 1386, 1537, 627, 453, 2022, 285]. little [1651, 1244, 588, 1416]. llbrarians [404]. loans [793]. Loanwords [117]. local [1785, 1292]. localization [1515]. Locknet [640]. logic [1981]. Logos [237]. London [230, 706, 323, 299, 324, 449, 540, 754, 707, 726, 748, 709, 724, 688, 747, 54, 180, 237, 182, 248, 247, 276, 279, 419, 118, 120, 115, 137, 331, 448, 473, 476, 524, 678, 53, 73, 116, 285, 260, 421, 322, 410, 586, 599, 140, 705, 658, 397, 300, 278, 277, 280, 725, 239, 354, 696, 91, 398, 205, 312, 310, 620, 708, 687, 219, 330, 287, 93, 204, 34, 1336, 251, 743, 618, 584, 313, 258, 138, 289, 523, 377, 495, 582, 92, 156, 734, 2158, 411, 511, 574]. Long [1136, 2192, 1342, 1900, 1682, 1938]. Long-term [1136, 2192, 1900, 1938]. longer [2001, 1673]. look [2015, 1263]. Looking [565, 620]. loop [1604]. Lorcan [376]. Lords [772]. love [1665]. Lowry [35]. loyalty [2133]. Ltd [298, 237, 260, 259, 311, 138, 377]. Lucas [240]. luck [1419, 1860]. Luxembourg [181]. Lynne [312]. Lyons [332].

M [200, 542, 475, 716]. même [1293]. MA [239]. machine [1375, 1941]. machines [1833]. Macintosh [619]. made [1136, 582]. Maeve [398]. Magazine [331]. magazines [1197]. mail [602]. Mailsort [106]. mainland [2085, 2212]. Maintaining [987]. major [1709]. make [1800, 2063, 2208, 2201, 1918, 420]. Makerere [2087]. makes [279, 2127]. Making [1923, 1916, 1450, 366, 610, 1862, 1378, 1926, 1131, 723, 1766, 1857, 205]. Malaysian [1829, 1785]. Malvern [137]. manage [646, 546, 1828, 665, 553]. Management [181, 1432, 445, 469, 493, 518, 537, 559, 578, 595, 614, 635, 654, 676, 693, 714, 731, 752, 573, 1465, 285, 1375, 1258, 3, 1250, 390, 1748, 1789, 1269, 1060, 2176, 1807, 1120, 76, 204, 755]. Manager [696, 1065]. Managing [706, 455, 1749, 1611]. mandate [1586, 1363]. mandates [1925]. Mandating [1746]. mandatory [1836]. manipulate [1604]. manipulation [408]. manual [697]. manuals [2141]. manuscript [323, 1955, 1888, 1640, 1392, 1807]. manuscripts [1898, 1332, 205, 1799]. many [182]. Mapping [1346, 1236, 1935, 955]. March [449, 301, 300]. Margot [638]. Marie [544, 353, 498]. Marine [420, 1938]. Mark [698, 1754, 1431, 598]. Market [125, 1659, 1188, 900, 937, 51, 2, 599, 991, 1002, 366, 778, 1549, 497, 176, 1615, 1524, 127, 21, 1289, 204, 311, 824, 164, 1158, 38, 78, 144, 1084, 1110, 1198, 1725, 841, 829, 479, 638]. market-led [991]. Market-oriented [1659]. Marketing [11, 29, 49, 299, 248, 1644, 209, 776, 914, 70, 111, 659, 320, 726, 825, 456, 828, 762, 1584, 407, 220, 1944, 495]. marketplace [1446]. markets [129, 745, 296, 405, 126]. marks [583]. Markup [79]. Marlborough [583]. marriage [1858]. Marriott [2120]. Martin [140]. Martyn [310]. Mary [678, 73, 310, 755]. Maryon [563]. mass [1831]. master [721]. masters [2003]. MasterVision [1479]. matching [1201]. mater [2230]. material [571, 290]. materials [1923, 454, 1001]. math [1845]. mathematics [1966]. Matilda [420]. matter [919, 1465, 899, 1049, 2137, 40]. matters [632, 2155, 1421, 1858]. mature [941]. Maureen [678]. Mawr [73]. maximise [1640]. Maximizing [2046]. Maxwell [1088]. May [230, 323, 119, 231, 322, 745, 203, 743, 1250]. McCarthy [696]. McDonald [354]. McIntosh [259]. MD [543]. me [1929, 1138, 904, 1595]. Meadows [34, 1890]. mean [2124]. meaning [1399]. means [1744, 1386]. measure [1227, 1514]. measured [1886]. measures [1734]. Measuring [1974, 2085, 1061, 1775, 1796, 1802, 2173]. Mechanical [298, 276]. mechanism [1001, 1604, 1128]. mechanisms [839]. Medals [254]. MedBiquitous [1270]. Medford [203]. media [1201, 2148, 1822, 911, 2216, 745, 1584, 1560, 1029, 365, 522, 563]. Medical [1091, 411, 511, 574, 229, 266, 1687, 530, 1583, 1329, 576, 2169, 1176, 53, 1949, 507, 2186, 696, 91, 1691, 2072, 1719, 1292, 1407, 1550, 1295, 1970, 1515, 2192, 1270, 1576, 2226, 1843, 2039]. Medicine [91, 1909, 985, 2228, 2183]. medium [359]. meet [1377, 459]. Meeting [167, 620, 383, 625, 2181, 203, 1208, 232, 105, 401]. Meetings [19, 62, 630]. meets [1833]. megajournal [2178]. megajournals [1800, 2070, 1910, 1911, 1721]. Mellon [542]. Member [2054, 211, 361]. members [754, 419, 638, 53, 285, 140, 157, 1501, 1434, 266, 584, 1352, 640, 755, 873, 53]. Memoriam [1349]. Mendeley [1882]. mentioned [2134, 2225]. Mentoring [1847]. MEP [298]. Merging [919]. Meritorious [170]. mess [1625]. message [522]. Metadata [1161, 1071, 1666, 912]. method [884, 2047]. methodology [1793]. methods [1923, 1760]. metrics [1637, 1466, 2178, 2043, 1522, 1294, 2203, 1655]. Mexico [127]. Meyer [266]. MI [240]. Michael [331, 260, 75, 205, 156]. Michèle [696]. Micheline [524]. Mickey [1171]. micro [2218]. micro-level [2218]. microbiologist [666]. Microsoft [1845]. Middx [311]. might [1913, 1335]. Mihill [563]. mill [472]. millennium [1025]. million [1505]. Millson [1230]. Milton [291, 290, 1171]. mind [608, 1470, 1003]. Mining [1728, 1587]. Minneapolis [745]. mirrored [988]. misconduct [1732]. Misrepresentation [2040]. missing [2018]. Mission [282, 143, 2147, 1855, 1860]. mix [1584]. MK14 [290]. mm [332, 288]. mobile [1680, 1626]. mobilization [2191]. model [1232, 384, 1405, 1472, 1790, 1993, 1202, 1121, 1464, 1368, 1400, 1531, 1794, 1008, 1347, 895, 1954, 1211, 2024, 2203, 1937, 201]. modelling [1563]. models [900, 1272, 1842, 1074, 1992, 2084, 1177, 990, 1118, 1784, 1260]. modern [159]. moment [1119]. money [463, 573, 610]. Monica [497]. Monograph [74, 2097, 1965, 1118]. Monographic [76]. monographs [936, 1404]. month [1765]. monthly [183]. moral [189, 765]. Morley [755]. Morris [1312]. Morrison [696]. Moscow [143]. most [366]. mostly [1282]. motivates [1729]. Motivations [1910, 893, 909]. motives [2131]. Mountain [420]. Mouse [1171]. move [1993]. movement [1822]. Moving [1342, 1679]. Much [1651, 1244, 1346]. multi [362, 2179]. multi-disciplinary [362]. multi-stakeholder [2179]. multidisciplinary [2070]. Multilingual [1520, 421]. multimedia [555, 1925, 734]. multiple [1718, 794]. Mulvany [620]. Munro [92]. must [1912, 1435]. mutant [984]. mutate [984]. my [598]. myth [2197, 2002, 779, 1080]. Mythbusting [2000, 1991]. myths [480, 1990, 897].

N [91]. n.a [222]. N7 [182]. name [1242, 1, 1529, 1635]. Nancy [620]. National [119, 544, 314, 355, 353, 498, 352, 475, 2104, 157, 288, 400, 788, 1407, 313, 985, 220, 952, 954, 1777, 2052, 2184, 1444, 1196, 930, 353]. natural [1675]. Nature [2220]. navigating [1971]. NBA [726]. near [1450]. Need [780, 2100, 669, 1375, 1348, 267, 1645, 1453, 1992, 1993, 2194, 2229, 2191, 625, 1521, 1642]. needed [1900, 2030]. needs [1034, 427, 302, 647, 1660, 1288, 295, 383, 1881, 297, 1114, 506]. neeh [401]. negative [2130, 2202]. negotiator [1035]. neither [1306]. NEJM [1329]. NESLI [881]. n’est [1563]. net [1548, 871]. Netherlands [1196, 320]. Network [485, 1117]. Networking [376, 382]. networks [1197, 342, 376]. Neuroscience [1463]. Never [1470, 1003, 380]. newer [1857]. News [804, 818, 834, 847, 862, 879, 891, 874, 30, 48, 71, 90, 112, 136, 154, 110, 151, 177, 199, 216, 235, 257, 347, 372, 416, 444, 468, 492, 517, 536, 134, 12, 997, 1181, 885]. next [2167, 2168, 275, 399, 1683, 1045, 1210, 389, 2206, 404, 1837]. NFAIS [544, 314, 475, 288, 286, 220]. NI [237]. nice [1111]. niche [736]. Nicola [330]. Nigeria [1145, 1236]. Nigerian [1751]. NIH [1363]. nimble [1418]. nineties [166]. Ninth [311, 487]. NISO [1387]. NISO/ALPSP [1387]. NJ [203, 36]. No [941, 74, 292, 157, 266, 288, 287, 251, 1866, 1836, 1697, 1058, 1673, 76, 696, 1608]. nomenclature [1387]. Non [1076, 1082, 2227, 2161, 2169, 965, 285, 2062, 75, 1121, 1017, 1883, 915, 584, 2059, 755, 1628, 2014, 53, 96, 332]. Non-author [2227]. non-China [1628]. non-commercial [915]. non-compliance [2161]. non-English [1883, 2014]. non-fiction [332]. non-medical [2169]. non-members [285, 584, 755, 53]. non-profit [2062, 1121, 2059]. non-scientific [75]. non-subscribers [96]. Non-subscription [1076, 1082]. non-technical [965]. non-virtual [1017]. nor [1306]. Nordic [2026, 1429]. norm [1718]. Normal [813, 2100]. Normative [2087, 2224]. North [182, 73, 2042, 436, 129, 609, 456, 1858]. Norwood [36]. Nostrand [617]. not-for-profit [1012, 1051]. not-so-harmless [1849]. note [311]. noted [262, 378]. Notes [335, 225, 358, 965]. nothing [1452]. notices [2138, 1292, 2073, 2231]. Nov [277]. November [706, 707, 276, 705, 278, 708]. number [1505]. numbers [1898, 1111, 2076]. numerator [1548]. NVQs [703, 277]. NW1 [696].

o [52]. OA [55, 1706, 1644, 1446, 2097, 1467]. OAI [1071]. OASIS [641, 663, 664]. Obituary [2113]. object [1940]. objectives [757]. Objectivity [2197]. objects [1209]. obligations [1923]. observations [1776, 1943]. obstacles [776]. Obstetrics [2035]. Obtainable [330]. Occasional [696]. Occupational [247]. occurs [1639]. O’Connor [398]. October [562, 250, 275, 637, 688, 638, 312, 251]. Odlyzko [1066]. OECD [1160]. offer [786, 965, 546]. Office [201, 181, 376, 686, 55]. Official [181, 815, 2218]. often [2134, 1884]. old [1242, 1368, 232]. Oldenbourg [1342]. Oliffe [332]. OMICS [2044]. On-demand [642, 809, 1162]. on-line [774]. One [2161, 356, 619, 1072, 1091, 385, 1652, 129, 761, 146, 588, 1558, 987, 2070]. Online [825, 1654, 1585, 1488, 1245, 855, 951, 984, 1201, 1099, 1268, 1209, 974, 1831, 969, 720, 975, 1598, 868, 103, 1791, 1672, 584, 1270, 1344, 1685, 2086, 2133, 1278, 822, 814, 1046, 1431, 120, 1225]. online-only [1344]. only [80, 1344, 1847, 1235]. Op [1075, 1217]. Open [1314, 1216, 1578, 1176, 1343, 2026, 1481, 1111, 1166, 1252, 1210, 1332, 1180, 1666, 1588, 1177, 1936, 1242, 1130, 1113, 1599, 1991, 1794, 1239, 1910, 1911, 1223, 1284, 1836, 2166, 1247, 2135, 1652, 938, 1379, 1482, 1601, 1822, 1889, 2142, 1222, 1966, 2018, 1371, 2078, 1331, 2214, 1819, 1790, 1912, 1992, 1446, 1187, 1429, 1755, 1905, 2125, 1320, 1724, 1884, 2175, 2219, 1844, 1748, 2064, 2204, 2123, 2146, 1849, 1999, 1731, 2215, 1965, 1434, 1815, 1220, 1693, 1303, 1719, 1763, 1751, 1602, 1518, 1330, 1283, 1467, 1687, 2184, 1128, 1558, 1672, 1500, 2069, 1335, 1514, 1954, 2016, 1169]. open [1444, 1189, 1211, 1372, 1325, 1129, 2174, 1409, 1573, 1260, 1909, 2053, 1973, 2086, 1935, 2205, 2078, 2107, 1982, 1602, 1071, 1611, 1870]. Open-access [1111]. Opening [1084, 1308]. Openness [1886, 2127, 2209]. OpenURL [1131]. operation [972, 5]. Operational [1911, 1061, 1416]. operative [131, 1202, 190]. opinion [1179]. opinions [1816]. Opportunities [2166, 2120, 897, 1723, 776, 662, 280, 1590, 1106, 1248, 567, 1404]. opportunity [2122, 2002, 429, 1177, 1259]. optical [243, 157]. Optimizing [823]. Opting [361]. options [1602, 92]. ORCID [1955, 1627, 1808]. order [2009, 644, 7]. ordering [68, 103]. ordinated [1668, 1742]. organisation [761]. organisers [401]. Organization [1147, 27, 46, 69, 88, 109, 153, 174, 197, 215, 236, 255, 284, 307, 328, 249, 351, 395, 417, 135, 374, 249, 481]. organizations [190]. Organizing [606, 1730]. oriented [2201, 1316, 1659]. origin [1628]. OSI [376]. other [149, 2046, 1386, 96, 2215, 1303, 971, 1734, 1310, 2024, 657, 1843]. otherwise [322, 419]. ounce [999]. our [1931, 2005, 2065, 2128, 2210, 2116, 1068, 1385, 1912, 203, 114, 1754, 1534, 1529, 1665]. out-of-print [1377]. outcome [384, 2035]. outcomes [1898, 2192]. outflow [1543]. output [1744, 2192]. outputs [2158, 2153, 2208]. outreach [1834]. outsource [1617, 1327]. outsourced [1095]. outsourcing [1471]. Over-promotion [2188]. Overburdening [2179]. overlay [939]. overlooked [1789]. Overseas [113, 217, 308, 326, 348, 394, 415, 443, 467, 516, 373]. oversee [2180]. overview [1922, 1272, 3, 1109, 1590, 311, 1260, 2010]. overwhelm [553]. Owen [93]. own [2180, 2224, 1237, 420]. owned [1999]. ownership [2150, 1749]. owns [1221]. OX7 [401]. Oxford [238, 74, 734, 1379, 1278, 1499, 1871]. Oxfordshire [657]. Oxon [401].

p [94, 754, 95, 91, 96, 330, 474, 202, 93]. p&p [399, 449]. P. [201]. p.a [183]. PA [324, 119, 475, 400, 220, 302, 300]. Pacheco [52]. Pacific [776, 778]. packing [447]. Page [599, 34, 621, 659]. PageMaker [597]. pages [240, 332, 698, 522]. pain [1555]. Paisley [420]. Pakistan [1885, 1899]. Panace [1247]. pandemic [2169, 2126, 2117, 2082]. Panel [785]. Paper [292, 398, 666, 219, 463, 716, 2004, 1028, 1043, 797, 1959, 472, 266, 287, 1295, 303, 716]. Paperback [399, 543, 544, 201, 314, 353, 448, 473, 476, 524, 542, 561, 698, 91, 311, 400, 289, 496, 495, 621, 94, 540, 259, 288, 93, 657]. Paperless [1269, 1222]. ’papers [1734, 696, 271, 1795, 2134, 507, 1695, 1543, 2182, 1883, 472, 2137, 1534, 1765, 2073, 1671]. paradigm [1014]. paradoxical [1252]. parasite [1238]. Paris [231, 742, 139, 222]. parity [1823]. Park [259]. Parker [1218]. Part [1910, 1911, 1014]. participation [1565]. Particle [1959]. partner [1433, 1354]. partnership [1545, 2092, 2058]. partnerships [970, 1827, 2120, 1279, 1858]. party [1524]. passion [1860]. past [1059, 1398, 1856, 1317, 1271]. Patents [189, 289]. path [624]. paths [1776, 2004]. Pathways [1865]. patient [2209]. patterns [1276, 1420, 1786, 1630, 2048]. pause [2169]. pave [2142]. pay [2214, 1160]. pay-per-view [1160]. Paying [1467, 1225]. payments [1749, 1611]. pays [549]. Pb [716, 563, 585, 522, 659]. pbk [238, 541, 586, 354, 696, 313, 582]. PC [619]. PDF [1873, 1972]. Peacock [260, 421, 258]. Peaks [1460]. Peer [1798, 2012, 1800, 1258, 1803, 1801, 1392, 1750, 1906, 1463, 1802, 640, 2015, 2170, 1809, 2197, 1760, 1667, 1484, 1319, 1326, 853, 2076, 1814, 1808, 871, 2075, 1522, 1810, 1736, 1835, 1464, 1598, 1565, 1556, 2149, 1806, 581, 1792, 2140, 1393, 1500, 1085, 1807, 1897, 2179, 2136, 1483, 1186, 1537, 1684, 1391, 718, 1027, 2024, 1245, 1103, 1799, 1767, 782, 855, 951, 1115, 1628, 854, 1000, 1573]. Peer-Review [1802, 1258, 2170, 2197, 1736, 1792, 1085, 1391, 1245, 1115, 1000]. peer-reviewed [1667, 1814, 1598, 1500, 1807]. Peggy [598]. PELICAN [1001]. Penetration [1751]. Penguin [618]. Pennsylvania [73]. People [233, 246, 393, 414, 442, 466, 491, 515, 535, 558, 634, 652, 675, 690, 711, 728, 750, 349, 371, 73]. Perceived [1547]. perception [1201]. Perceptions [1137, 1785, 2214, 1321, 1967, 1501, 2077, 1382, 2136, 951, 2086]. Perdue [619]. perform [2215]. Performance [2178, 1952, 1895, 2106]. performances [287]. Pergamon [1088]. perils [555]. period [1901]. periodical [1175]. periodicals [937, 82, 1319, 1064, 740, 96, 1515, 2145, 338, 1249]. periphery [1692, 1884, 2177, 2107]. perish [1231, 2002]. Personal [433, 637, 890, 966, 994, 1034, 942, 2030, 474, 1328]. perspective [2157, 1401, 1176, 1482, 1392, 1589, 1269, 1982, 567, 1087, 2179, 1587, 1291, 1910, 1911, 1940, 1858, 1980]. Perspectives [2153, 2058, 2150, 114, 2069, 1905]. persuasion [1399]. Pertti [448]. pervasive [1294]. Peter [180, 476, 563, 310, 93, 521, 1349]. Ph [181]. Pharmaceutical [1016, 1240, 1468]. PhD [420]. phenomenon [1088]. Phil [584]. Philadelphia [119, 544, 314, 353, 498, 352, 475, 288, 400, 286, 220]. Philip [292]. Philosopher [2214]. Phoenix [1118]. Photocopying [552, 252, 478, 763]. Phyllis [617]. Physical [1402, 1722, 2144, 1793, 992, 1697]. Physicists [1006]. physics [1228, 1603, 927, 1959]. Physiological [685]. pigs [1460]. Pilot [811, 1705, 1486]. pilots [2064]. pioneers [1415]. Pira [638, 250]. Piracy [22, 993, 1585, 1277, 1291]. Pirate [1912]. PIRUS2 [1574]. pixels [666]. place [1159]. Plagiarism [1513, 1291, 1341, 2088, 1461, 1694, 1462, 1631]. plagiarized [1995]. plan [1045, 1727, 2061, 2053]. planning [1958, 2192]. plant [1669]. Plantwise [1669]. platform [2184, 1444, 2096]. platforms [1917]. PLOS [2070]. Plus [1334, 1293, 4]. Plymouth [541]. podium [1823]. point [2018, 1447, 549, 829]. Points [859, 844, 887, 872]. policies [1836, 2161, 1266, 1898, 723, 2219, 186, 2187, 589, 2203, 2096, 353]. Policy [160, 543, 678, 1578, 2205, 397, 2173, 1921, 242, 2219]. political [2029]. Politics [2155]. Polytechnic [182]. pomposity [482]. poorly [2229]. Port [620]. portal [1545, 950]. portals [952]. Portrait [211]. Portuguese [176]. positive [2130]. possibilities [405]. possible [2197, 2145]. Post [1238, 1144, 51, 447, 1767]. Post-print [1238]. Post-publication [1144, 1767]. Potential [2195, 1026, 2196]. potentially [2149]. power [1636, 2155, 982]. powerful [182]. Pp [291, 399, 419, 543, 544, 421, 397, 398, 400, 420, 238, 541, 540, 637, 754, 54, 119, 200, 180, 182, 638, 118, 115, 137, 201, 261, 355, 353, 448, 498, 473, 476, 524, 542, 678, 697, 181, 352, 117, 260, 586, 563, 75, 599, 140, 221, 76, 597, 561, 598, 239, 354, 696, 203, 205, 376, 497, 620, 401, 581, 219, 259, 472, 447, 266, 288, 287, 204, 34, 114, 139, 222, 583, 619, 618, 584, 55, 36, 313, 258, 138, 289, 496, 521, 377, 495, 582, 92, 220, 617, 585, 621]. pp. [292]. ppl76 [475]. practical [541, 119, 2078, 563, 801, 330, 190, 1574, 460, 1975]. practicalities [591]. Practice [894, 2015, 1266, 1401, 1523, 598, 2229, 797, 733, 589, 657, 1447, 1584, 1758, 1251, 34, 419, 118, 120, 521]. Practices [1885, 1829, 1293, 1191, 1666, 2075, 1691, 1028, 1043, 1896]. Practising [1449]. Practitioner [2166, 507]. Pragmatic [935]. preach [1449]. preaching [1166]. precedents [138, 523]. precepts [956]. ’predatory [2162, 1652, 1881, 1757, 2160, 2163, 2090, 1945, 1939, 1849, 2131, 1763, 1751, 2069, 1953, 2029]. Predecessors [983]. predict [1057]. predicting [384]. predictions [2054]. predicts [1831]. preferences [974]. preferential [1055]. Preliminary [2144]. preparation [155]. Prepared [401]. preparing [96]. preprint [2161, 939, 983, 2024]. Preprints [2181, 2071, 2188, 1672]. presence [709, 700, 1791]. present [461, 672, 1856, 78, 987, 1317, 1271]. Presentation [291, 1139]. presentations [585]. preservation [1196, 1389, 1535]. Preserving [1743, 739, 928]. Press [94, 238, 637, 51, 200, 697, 73, 52, 1853, 1979, 716, 586, 75, 76, 698, 239, 620, 219, 139, 289, 521, 495, 522, 1002, 1854, 1614, 1860, 1072, 1858, 2222, 1649, 1983, 1965, 1226, 1292, 2079, 1871, 1985]. press/library [1858]. presses [1857, 1984, 128, 2230, 1854, 1248, 1855, 1981, 1965, 1856, 202, 1194, 2147, 628, 1325, 2118]. Pressing [1055, 1861]. pressures [2177]. prevailing [1991]. prevalence [2224, 2185]. prevalent [1995]. prevented [1835]. prevention [999, 1541]. Price [222, 220, 1673, 266, 288, 287, 93]. prices [1381]. Pricing [903, 1601, 486, 1386, 39, 1099, 1272, 1405, 1074, 1001, 1370, 186, 21, 1531, 38, 589]. Primary [146]. primer [1585]. Princeton [509, 1045]. Principles [894, 419, 34, 2015, 1601, 2064]. Print [1073, 1162, 1061, 1201, 275, 399, 1656, 1064, 1673, 1045, 1238, 1418, 1761, 1377, 1235, 586]. printed [367, 869, 657]. printers [426, 386]. Printing [419, 427, 260, 421, 139, 324]. prints [1304, 1211]. privacy [433, 684]. privatisation [473]. Prizes [576]. probably [1608]. problem [1031, 1994]. Problematizing [2131]. problems [2100, 662, 842, 405, 227]. Proceedings [399, 74, 137, 73, 203, 497, 620, 202, 679, 496, 448, 354, 312, 2181]. Process [1807, 1686, 2197, 1232, 2134, 1257, 301, 1888, 1588, 165, 1736, 1085, 503, 854]. processes [1850, 36]. processing [2135, 180, 2189, 157, 550, 1617, 1675]. Product [2220, 1058]. Production [1250, 2163, 650, 419, 1656, 3, 2150, 155, 489, 510, 1269, 359, 343, 258, 331]. productive [1304]. productivity [2157, 1227]. products [1610, 1549, 157]. Professional [180, 237, 1104, 130, 1146, 2186, 1197, 1095, 950, 1106, 1185, 231]. professionalization [816]. professionals [1817, 2091, 1407, 2192]. profile [828]. profiles [805]. Profit [1576, 250, 2062, 626, 1096, 61, 1121, 1012, 2059, 1089, 1051]. profitability [1471]. profits [1216, 1514]. programme [733, 1178]. programs [1575]. progress [881, 885]. Project [1431, 782, 826, 954, 1614, 809, 1404, 87, 738, 148, 812, 531, 916, 1596, 1602, 1132]. projects [1250, 1127, 1140]. promise [2200, 1767]. Promises [964]. promote [761, 2079]. Promoting [1092, 2130, 409, 1393]. Promotion [1614, 1145, 327, 403, 449, 2188, 147, 2218]. Proofreader [275]. Proofreaders [399, 449, 496, 277]. Proofreading [432, 598]. property [943, 1068, 1221, 481]. prophecy [1409]. prophylaxis [482]. proposal [1021, 1724, 733]. proposals [190, 902, 114]. proposed [1666, 1128, 2070]. prospect [1898, 1887]. prospects [555, 2195]. protecting [1068]. Protection [18, 1031, 481, 454, 1186, 370, 392, 413]. Protocol [1071, 84]. provide [1268, 1882]. provider [1042]. providers [1821]. providing [1222, 1096, 620, 1897, 1089]. province [2106]. province-level [2106]. provincial [2212]. provision [687]. Psychological [563]. psychologist [2175]. PsyclNFO [505]. pu [1861]. Public [643, 1691, 522, 646, 1484, 665, 2092, 1705, 1719, 1727, 2117, 1602]. Publication [1145, 157, 999, 1541, 1009, 1221, 292, 2109, 2182, 2094, 2123, 1971, 2030, 1316, 1786, 1467, 733, 1080, 794, 1369, 2192, 1294, 1611, 2132, 1144, 1767, 976, 1982, 1203, 543]. Publications [262, 378, 240, 181, 1227, 561, 311, 303, 68, 865, 1165, 1868, 1668, 2092, 1016, 1040, 213, 1936, 962, 1228, 1124, 1679, 2085, 657, 2212, 988, 841, 657]. Publicity [571, 522, 60]. Publish [1231, 541, 716, 2002, 2214, 1996, 1244, 1215, 1859, 1939, 1953, 1843, 947]. Published [562, 638, 697, 202, 266, 657, 159, 290, 1997, 2047, 1332, 1134, 1281, 1542, 1972, 183, 1765, 2073, 472]. Publisher [894, 910, 186, 736, 1279, 156, 1929, 2167, 1266, 1447, 911, 2028, 1331, 1875, 1151, 588, 245, 1755, 1848, 605, 1604, 2064, 1585, 208, 457, 265, 567, 1330, 1087, 1393, 1226, 1587, 1291, 1377, 1910, 1911, 2203, 1318, 1986, 212]. publisher-driven [1226]. Publishers [230, 385, 706, 324, 302, 54, 60, 180, 237, 247, 419, 369, 261, 52, 322, 316, 266, 251, 784, 313, 258, 1766, 2034, 426, 919, 2160, 1674, 488, 672, 1601, 1822, 1922, 970, 1375, 1064, 1348, 954, 1916, 1481, 124, 548, 662, 452, 2090, 870, 1031, 1818, 1790, 588, 277, 817, 1240, 829, 1095, 2229, 1121, 1918, 925, 950, 1177, 412, 687, 1060, 2191, 2215, 1161, 1012, 1957, 2045, 815, 1602, 1159, 5, 1826, 1535, 1562, 1270, 1762, 1630, 1120, 2038, 1102, 1051, 2053, 1874, 32, 50, 151, 840, 134, 697, 447, 495, 92]. Publishing [2058, 2057, 28, 44, 845, 894, 250, 612, 182, 331, 1448, 487, 260, 716, 95, 1073, 705, 991, 1160, 61, 2230, 508, 989, 205, 497, 2126, 266, 34, 644, 756, 1356, 251, 145, 1924, 642, 2069, 36, 1665, 258, 2082, 1277, 572, 986, 1249, 1929, 1188, 1091, 1636, 1216, 13, 169, 385, 737, 2162, 2163, 900, 1342, 1681, 1433, 191, 480, 1788, 645, 1686, 1723, 1735, 1764, 1696, 1652, 670, 647, 956, 180, 247, 798, 1025, 826, 638, 125, 2026, 1355, 1966, 1983, 2122, 742, 852, 970, 1756, 15, 1559, 57, 1776, 1371, 2055, 549, 1979, 2051]. publishing [2116, 1481, 1209, 1088, 1257, 1293, 1523, 1171, 421, 1229, 1577, 1752, 1778, 1961, 1942, 2041, 2062, 838, 958, 294, 935, 972, 1146, 1116, 2201, 924, 2002, 898, 1819, 1683, 1645, 1818, 1451, 682, 1817, 1453, 1820, 83, 81, 1429, 2224, 779, 2084, 2060, 2096, 725, 1202, 1213, 2124, 2221, 1590, 1654, 1095, 2056, 1121, 1075, 1823, 161, 1320, 1724, 605, 1315, 1028, 1043, 1464, 1104, 1177, 1006, 1748, 2052, 342, 485, 1307, 1827, 1959, 2107, 1615, 1967, 1927, 2193, 1849, 2003, 666, 1461, 2131, 1130, 1224, 1346, 1351, 1982, 2110, 603, 226, 1740, 1741, 2187, 2139, 1237, 1280, 1303, 1896, 1763]. publishing [474, 953, 202, 923, 270, 382, 1518, 504, 1400, 1200, 1653, 1173, 800, 813, 1659, 2017, 1958, 915, 1253, 1014, 1132, 685, 207, 228, 1033, 1414, 1133, 1583, 243, 895, 1514, 345, 1294, 1169, 1784, 1105, 1189, 809, 2101, 883, 1742, 2108, 2038, 2061, 940, 1516, 2079, 1174, 1327, 1354, 896, 460, 621, 657, 734, 1404, 1198, 1725, 1757, 1973, 2133, 2183, 2087, 845, 932, 1878, 2112, 745, 1302, 703, 2173, 917, 575, 885, 901, 1306, 637, 156, 94, 541, 355, 260, 221, 219, 139, 311, 358, 981, 1039, 1099, 758, 93, 232, 1774, 1934, 138, 523, 733]. Publishing-the [93]. Publons [2229, 2217]. PubMed [2144]. PubPeer [2138]. PubSCIENCE [992]. punctuation [583]. purchase [463]. purchasing [937, 82]. purpose [1398]. Purposes [2230, 2136]. pursuers [626]. pursuit [2003]. pushing [1745]. Putting [971, 1944, 168]. pvc [449].

qualitative [1272, 2186, 2136]. Quality [1722, 774, 1829, 1885, 2047, 2182, 1392, 1351, 1173, 2140, 1382, 1267, 1948, 1802, 1951, 1000]. Quarterly [237]. Quartet [87, 148]. queries [843]. quest [1148, 1503]. Question [1934, 1215, 1612]. questionable [2025, 1996]. questionnaire [2083, 2086, 1649]. Questionnaires [2225]. Questions [478, 623]. Quick [754].

R [238, 117, 1394, 551]. rôle [481]. Radiology [2057]. Ramsbury [583, 585]. Ramsdale [181]. range [868]. Rank [1655]. Ranking [1655, 2217]. ranks [2023]. Rapid [2167, 2225, 2182, 1041]. rapidly [1071]. raring [439]. rate [1734]. rates [1697, 1643]. rather [1651]. ratings [1484]. Raymond [582]. Rd [377]. Re [1257, 1810]. Re-engineering [1257]. reach [1753, 1821, 1524]. reaching [1925]. reactions [812]. read [1696]. ReadCube [1873]. reader [1929, 1385, 1873, 33]. reader-focused [1873]. Readers [1904, 973, 1385, 2046, 925, 243, 1675]. readership [1201]. Reading [16, 47, 72, 179, 218, 234, 256, 309, 329, 350, 396, 446, 471, 494, 519, 579, 615, 655, 694, 713, 730, 751, 31, 89, 1086, 375, 449, 1420, 1630, 1761, 1597]. ready [1533]. realities [870, 1240, 1017, 1911, 1077]. reality [1201, 900, 2197, 745, 2031, 725]. realizing [1026]. really [1468, 2070, 1008, 1352]. realm [1035]. reasons [2227]. reception [435]. recognition [1808]. Recognizing [1783, 1402]. Recollections [368]. Recommendations [290, 2015, 1923, 96, 1387, 1960, 2045]. Reconciling [1955]. record [1743, 2073]. recording [1147]. Records [1073]. recruitment [341]. recycling [2159, 2042]. Red [1202]. Reducing [292]. Redux [1979, 1853]. referee [1419]. refereed [1059, 1083]. Refereeing [162, 271]. referees [1402]. Reference [117, 722, 1504, 238, 2004, 1199]. references [1360, 1814]. referencing [290, 963]. reflect [1673]. reflected [2140]. Reflections [763, 2011, 1667]. reform [1659]. Reforming [2110]. refuse [1658]. regarding [1960]. regional [1949, 1974]. regions [2180, 2038, 2106, 2212]. Register [18]. registry [2095]. Reinhold [617]. reinvention [2116]. reinvigorated [1980]. reject [1244]. rejected [1898, 2035]. rejection [1555]. rejections [2030]. Related [816, 2164, 2208, 2217, 1018, 733]. relation [1764]. Relations [522, 646, 665, 1691, 643]. Relationship [894, 2138, 1428]. relationships [1984, 910, 1871]. releases [1292]. relevance [647, 381]. relevant [2001]. Reliability [1484, 581]. reliable [1802]. reliance [1704]. remains [1848]. remedies [2197]. removed [2023]. replace [1472, 361]. Reply [1534]. Report [785, 772, 544, 118, 355, 475, 312, 288, 876, 943, 888, 314, 181, 546, 905, 906, 2056, 1205, 311, 313, 893]. Reporting [1530, 995]. reports [9, 41, 808, 767, 831, 845, 860, 875, 2154, 2097, 784, 1754, 1802, 802, 1655]. repositories [1894, 1220, 1613, 1428, 1262, 1500, 1174, 2141]. repository [509, 1586]. represent [2002]. representing [1563]. reprinted [599]. Reproducibility [2047, 1717]. reproducible [1997]. Reprographic [1143]. Republic [1784]. Republication [1671]. repurposing [1707]. reputation [2076, 1775, 1796, 1830]. requests [2229]. required [1704]. Requirements [822, 1679]. Research [2164, 1165, 240, 118, 355, 542, 678, 116, 117, 1191, 312, 310, 814, 679, 313, 2203, 1998, 2158, 2169, 2156, 2159, 1723, 461, 125, 1706, 1190, 524, 1776, 2134, 2200, 2127, 2208, 2201, 2225, 1841, 278, 366, 1831, 35, 1610, 165, 2182, 2204, 2207, 2209, 1614, 2089, 866, 1806, 1899, 1957, 2074, 1719, 1236, 1530, 204, 99, 1400, 1283, 1687, 2176, 2140, 1958, 1158, 1560, 1679, 2192, 1308, 1118, 1359, 893, 909, 1730, 1406, 640, 2022, 1935, 2106, 1705, 1820, 2206, 1742, 940, 679, 397, 1976]. Research4Life [2171]. Researcher [1321, 2077, 1829, 1869, 2104, 1957, 1881, 2086, 1649]. researchers [2152, 1894, 267, 1828, 1575, 2201, 1996, 1627, 1240, 2075, 1729, 1975, 2088, 2193, 2089, 1657, 1532, 1613, 1804, 1887, 1896, 1867, 2027, 2074, 2083, 2045, 1786, 1340, 1906, 1261, 2021, 297, 2108, 2226]. ResearchGate [1894, 2044, 1830]. Researching [1531]. resilience [2119, 2116, 2121, 2124, 2123, 1704, 2126]. resilient [2127, 2125]. resolutions [2144]. resonance [1877]. resource [1394, 1762, 1847, 952, 137]. Resources [35, 1357, 2004]. responding [1694]. response [1066, 2105, 1205, 2117]. responses [1799]. Responsibilities [918, 682]. Responsibility [1451, 1960, 1948]. Responsible [2096]. Results [2206, 1846, 1578, 2040, 2202, 1385, 2086]. rethink [1530]. Rethinking [883]. retraction [2073, 2227, 2231]. retractions [2138, 2073, 2227]. retrieval [120, 115, 476, 1109, 679, 955]. Retrieving [740]. reusable [2208]. Reuse [2201, 2207, 2077]. Reuse-oriented [2201]. reveal [2215, 1790]. revenue [823, 1076, 1082]. Review [2171, 1697, 290, 562, 238, 399, 449, 541, 540, 637, 754, 51, 54, 119, 240, 200, 180, 419, 543, 544, 638, 118, 120, 115, 74, 137, 201, 331, 314, 355, 353, 448, 498, 473, 476, 524, 542, 678, 697, 53, 73, 116, 181, 475, 52, 117, 260, 421, 586, 563, 75, 332, 599, 95, 140, 221, 658, 76, 397, 292, 597, 561, 698, 598, 239, 696, 35, 91, 203, 398, 155, 157, 205, 312, 310, 497, 401, 581, 259, 474, 472, 447, 202, 183, 266, 288, 287, 93, 204, 1256, 34, 114, 139, 222, 311, 400, 1791, 583, 619]. Review [618, 584, 55, 1463, 36, 313, 286, 1141, 258, 138, 289, 523, 677, 420, 1802, 496, 521, 931, 495, 582, 92, 156, 220, 617, 522, 585, 621, 657, 755, 734, 659, 2015, 1798, 2156, 2170, 2163, 2012, 1809, 2197, 141, 1564, 1760, 1800, 1484, 1056, 1319, 285, 1326, 1689, 1175, 1258, 853, 924, 1996, 2076, 1841, 1808, 871, 2224, 2075, 1588, 1801, 1522, 1810, 1315, 1386, 1736, 1835, 1844, 1565, 1392, 1556, 1726, 2131, 1161, 1806, 2139, 1750, 1792, 2140, 1393, 1906, 1085, 1897, 2136, 1483, 1186, 1537, 1684, 1391, 718, 1557, 1027, 2024, 1245, 757, 640]. review [1103, 1799, 1767, 782, 855, 951, 1115, 1628, 1935, 854, 1000, 2167, 2048, 2084, 1976, 2031, 2097, 2150, 1975, 2004, 1373, 821, 1361, 581, 1959]. reviewed [17, 1667, 1814, 1598, 1500, 1807]. reviewer [2170, 1484, 1803, 2229, 1464, 2217, 1565, 1994, 2149, 1799, 1628, 1618]. reviewers [1931, 2005, 2065, 2128, 2210, 1949, 1952, 1589, 2149, 2179, 1805, 1799, 1878]. reviewing [1943]. Reviews [1510, 1528, 1591, 94, 768, 787, 803, 833, 846, 861, 877, 889, 907, 1476, 1525, 1526, 1112, 789, 959, 1338, 1313, 1496, 978, 1506, 1581, 1593, 1607, 1619, 1647, 1474, 1511, 1477, 716, 1622, 1605, 1551, 1410, 1182, 1623, 1298, 1491, 1497, 1424, 620, 1473, 1509, 1494, 1495, 1507, 967, 1412, 1126, 1633, 219, 330, 1492, 1592, 1594, 1274, 1323, 1384, 945, 1527, 1168, 1383, 1552, 1580, 1142, 1411, 1425, 1154, 1493, 1475, 1606, 1648, 1396, 1508, 1397, 1426, 1539, 1540, 1634, 1423, 770, 1538, 1553, 1193, 1621, 1243, 1287, 1010, 1554, 1620, 1440, 560, 580, 863, 1455, 1567]. reviews [1662, 1454, 1459, 1419, 539, 1703, 1458, 1714, 1395, 715, 1707, 1050, 1663, 1701, 1712, 1139, 1436, 835, 596, 1676, 695, 732, 996, 1441, 1772, 499, 1737, 1457, 636, 1700, 1439, 1437, 2140, 1568, 1566, 616, 753, 1713, 1438, 1569, 1661, 1715, 1716, 819, 1078, 1456, 1677, 1702, 656, 1664, 1678, 1646, 1036, 1094, 1739, 1770, 1771, 848, 1022, 1738, 1707]. Revise [1960]. Revised [53, 219]. revising [2135]. Revision [96, 2035]. revisions [599]. Revisited [462, 609, 2041]. revolution [1311, 1020, 1414, 1942, 2041]. Reward [1399]. Rewarding [1805]. rich [1445]. Richard [118, 289]. Riding [761]. right [2112, 913, 643, 815, 359, 1626, 1727]. Rights [759, 869, 706, 748, 1465, 1375, 189, 911, 1143, 439, 1221, 844, 1119, 1390, 2091, 1708, 1886, 287, 765, 1018, 926]. rigour [1994]. Rim [776, 778]. Rise [332, 1794, 1833]. rising [1074, 1118]. risks [1571, 2061]. Riv [354]. Riv-erina [354]. Riyadh [2088]. RMIT [1306]. Road [182, 696, 259, 474, 1416]. Robert [94, 541, 200, 419, 716, 34, 36, 495, 1230, 1088]. Robin [204]. Rochester [696]. rock [1159]. Rogerson [658]. Role [1998, 918, 1681, 840, 1764, 1753, 367, 1165, 1222, 1644, 1720, 406, 1331, 1752, 1875, 871, 1854, 1748, 2052, 568, 1883, 1524, 813, 1393, 1172, 883, 1102, 1944]. roles [1017, 1783, 1562]. ROM [611, 226, 624, 753, 584, 420]. Romanian [439]. RoMEO [1151]. ROMland [462]. ROMs [1007, 460]. Ronald [259]. roneoed [53]. room [2063]. Rounding [1745]. Roundtable [648, 2045]. Routledge [398]. RoweCom [1183]. Rowlands [115]. Royal [561, 91, 258, 1016]. RRO [977]. RSS [1644]. Ruben [679]. Rubriq [1684]. rules [2213, 55]. ruminations [368]. run [2090, 472]. Russia [1549]. Russian [1549].

S [36, 2061, 2053]. S&T [1895]. S022 [472]. Sage [844, 1927]. Saharan [1575]. sailing [1331]. sale [1583]. Sales [696, 455, 1162]. Sally [449, 1312]. same [1765]. sample [2190]. San [232]. sands [1787]. Sarajevo [667]. satisfaction [2140]. satisfactions [1686, 1753]. satisfy [2074]. Satisfying [506]. Saudi [2088]. Saur [678, 310]. save [463]. Saving [573]. savings [83]. saying [540]. scalable [1724]. scale [2130, 1802]. scams [1835]. scenario [1958]. scenarios [480]. scene [1228, 1161, 856, 1709]. Schipper [314, 353, 352, 286]. Scholar [1655, 1815, 1866, 1550, 2014]. scholar-led [1815]. Scholarly [894, 637, 852, 1559, 775, 1293, 1523, 1179, 745, 1692, 1429, 2221, 1905, 1420, 1307, 2187, 1709, 1336, 1253, 1195, 232, 1529, 1761, 1637, 1885, 1466, 1681, 1601, 1760, 1800, 1355, 542, 1272, 52, 1868, 2116, 1257, 1353, 2062, 935, 924, 2092, 336, 1683, 1817, 1992, 1134, 2230, 1862, 2096, 1013, 1191, 2124, 2104, 1590, 1654, 1505, 1095, 2056, 1823, 1610, 1320, 1744, 1386, 2095, 1821, 1549, 1844, 1177, 2052, 1749, 1883, 936, 1615, 1156, 1999, 1461, 1886, 2191, 1542, 2072, 1224, 1815, 1741, 1237, 1289, 1280, 1303, 1775, 1796, 1804, 1867, 2074, 2083, 202, 1041, 923, 815]. scholarly [963, 1108, 1563, 1653, 1659, 2017, 1261, 1071, 2117, 2036, 1587, 1270, 1954, 1537, 1149, 2145, 2101, 1547, 987, 1129, 2073, 1833, 940, 2195, 1174, 1233, 986, 2022, 2218, 2087, 212, 2056, 1887, 1937, 521]. scholars [1967, 570, 1953, 2036, 1630, 1866]. Scholarship [1278, 1448, 928, 35]. School [182]. SCI [1643, 2044, 2013]. Sci-Hub [2044, 2013]. SCIE [2178]. SCIENCE [1047, 772, 747, 280, 1832, 1205, 259, 2017, 2203, 1406, 1216, 737, 2154, 2155, 2180, 1670, 1268, 458, 1966, 448, 742, 2130, 1816, 2134, 1997, 1489, 810, 2177, 2094, 1147, 2077, 2110, 581, 2144, 1897, 1004, 1953, 2036, 1734, 1225, 1583, 2029, 1486, 1249, 976, 2157, 1883, 1698, 397, 398, 868]. ScienceDirect [1928]. sciences [1266, 2012, 2205, 1949, 1943, 1075, 2109, 1724, 1936, 1544, 1722, 1786, 1761, 2174, 992, 1909, 951, 1843, 696, 35]. Scientific [956, 543, 267, 280, 266, 139, 222, 265, 317, 1837, 2165, 2157, 556, 938, 1232, 1273, 268, 364, 1743, 1371, 740, 1088, 1449, 104, 1227, 716, 1668, 75, 1221, 1503, 2047, 1818, 1187, 2150, 1643, 949, 2182, 2220, 1028, 1043, 568, 1688, 2107, 2123, 1613, 661, 583, 1173, 1172, 1794, 1414, 1690, 243, 627, 1969, 2132, 2085, 1271, 1616, 1115, 1935, 2014, 697, 76, 617]. scientist [1227, 1119, 383, 569, 1890]. Scientists [497, 1252, 1562]. SCImago [2023, 1655]. scite [2168]. SCOPE [642]. Scoping [1198]. Scopus [2154, 2164, 2157, 2043, 1522, 2215, 1698]. Scoring [1802]. Scott [597]. Scottish [642, 1073, 227]. scrazzl [1610]. screening [1461, 1541]. search [626, 1869, 1727, 1413]. Searching [788]. Seasonal [1579, 1603, 1697, 1483].

Seattle [52, 219]. Second [617, 621, 1680, 496, 169]. secondary [146, 288, 1087, 314]. Secretary [401, 152, 196, 65, 107, 85, 133]. secrets [1927]. Section [1390]. sectional [2105, 2140, 2226]. sector [311]. security [1827]. see [1219]. Seeking [1390, 1720, 1420, 1761]. Seizing [1119]. Select [772]. selected [1381, 35, 1952, 1972]. selecting [1807]. selection [2152, 688, 884, 704, 76, 1803, 723, 1624, 1909, 2222]. Self [1266, 1643, 1480, 1686, 1674, 1764, 1586, 2028, 2143, 1803, 1708, 2061, 2087]. Self-archiving [1266, 1586, 1708]. Self-citation [1643, 1480, 2143]. self-citations [2028]. self-correction [2087]. self-publishers [1674]. self-publishing [1686, 1764, 2061]. self-selection [1803]. Selling [436, 276, 2091, 1549, 1615, 1572, 1600]. Semantic [1414, 1610, 1014, 1588]. semi [2177, 2107]. semi- [2107]. semi-periphery [2177]. Seminar [9, 41, 802, 302, 449, 51, 601, 767, 831, 845, 860, 875, 129, 888, 487, 906, 725, 1192, 1241, 784, 572, 649]. Seminars [630, 995, 98]. send [1055]. senior [1261]. sense [1857, 1805]. sensegiving [2117]. sensemaking [2117]. sensibility [1805]. separate [2194]. September [562, 687, 496]. September/October [562]. Serial [343, 688, 1064, 704, 723, 575, 126, 2001, 194, 657, 73, 657]. Serials [920, 74, 73, 76, 657, 755, 51, 253, 2084, 101, 412, 708, 78, 74]. serials-new [51]. Series [544, 475, 266, 288, 156, 314, 181, 239]. serious [1639]. serve [2060]. servers [983]. service [1466, 314, 822, 813, 1826, 1070]. Services [119, 314, 353, 498, 352, 475, 817, 829, 91, 1728, 288, 220, 2055, 1908, 146, 1160, 1321, 950, 400, 1684, 1920, 244, 544, 240, 286]. Serving [458, 1225, 1090]. sesame [1113]. Setting [962, 700, 1709, 1042]. settlement [1413]. several [1091]. SFEP [399, 449, 496, 816]. SFX [1032]. SG2 [330]. SGML [607, 200]. shades [1773]. shadow [1342]. shall [1282]. shame [1111]. shape [275, 399]. shaping [1638, 2193, 2131]. shared [2201]. shares [2072]. Sharing [2208, 2089, 1178, 2205, 2200, 1975, 2175, 2207, 2077, 1866, 1742]. Sheffield [496]. Shetlands [783]. Shiel [157]. shift [1998]. shifting [1787]. Shillingsburg [521]. shing [1861]. Shipton [332]. Shirley [474]. short [1682]. Shortcomings [1027]. Should [2148, 1469, 1718, 1952, 2149, 1582, 1129, 1874, 2063, 2111, 317, 1603, 530, 1534, 1535]. show [1089]. showcasing [1775, 1796]. shown [266, 288, 287]. shut [1220]. sic [53]. side [1188, 149]. significance [2112, 953]. Sigurdson [473]. simple [384]. Simplifying [1656, 1305]. since [1236]. Singapore [1143]. Single [1556, 1848, 479]. Single-blind [1556]. single-journal [1848]. Site [930, 811, 1329, 662]. sites [916]. situation [2171, 2189]. Six [2149]. Size [1049]. skewness [2028]. skill [432]. skills [1899, 2030]. SL5 [377]. slave [721]. SLG [562]. Small [324, 128, 1815, 1857, 488, 686, 1452, 870, 245, 1848, 202, 1602, 1159, 1292, 495, 1969, 585]. Smith [200, 598, 219, 495]. snapshot [1250]. Social [1560, 2155, 1266, 1268, 2148, 1822, 2134, 2216, 810, 1943, 1075, 35, 1584, 1724, 1936, 1544, 1786, 1953, 2036, 1761, 2174, 1486, 1909, 1249]. Societies [1435, 1433, 970, 2055, 2063, 130, 838, 1787, 1724, 2052, 208, 971, 227, 2054, 1704, 2059, 5, 1239, 1129, 1089, 1354, 1832]. Society [249, 808, 772, 275, 399, 449, 754, 180, 53, 745, 563, 277, 2060, 561, 696, 91, 203, 251, 679, 2054, 807, 232, 496, 734, 1342, 686, 724, 15, 2051, 1481, 2062, 213, 610, 1134, 969, 245, 242, 2121, 1095, 620, 1185, 2064, 962, 1434, 265, 567, 207, 228, 1033, 957, 2061, 1260, 1352, 2053, 2058, 706, 840, 212, 1016, 504, 685, 1402, 140]. society-published [1134]. Sociological [814]. Software [560, 580, 488, 141, 539, 715, 596, 695, 732, 499, 636, 616, 753, 656, 370, 1684]. solution [1481, 454, 1122]. solutions [867, 1666, 1095, 2196, 1267]. solve [1994]. Some [1943, 1040, 1755, 1276, 1060, 20, 623, 843, 1667, 2001]. song [321]. Source [100, 1076, 2229, 1999, 1602, 99, 1082]. Sources [1950, 155, 2149, 1544]. Soviet [66]. Spa [201, 355, 330, 447]. space [1680]. Spain [2214]. Spanish [1894, 1842, 1708, 1886, 164, 1500]. Special [780, 673, 1012, 2137, 2185, 1051, 746, 551]. Specialist [679, 1179, 1843]. specialized [2070]. specialty [1240]. species [1152, 602, 755]. specific [665, 2132, 2132]. specify [427]. spectre [793]. Spicer [541]. Spiral [498, 474, 221, 220]. spiralbound [447]. sponsors [971]. spot [1683, 1731]. Spread [224]. Spreading [1216, 738, 319]. spreadsheet [384]. Springtime [1208]. Square [180]. squeeze [1159]. SSP [1207]. St [689, 710, 727, 749, 764, 156]. stable [433]. staff [360, 361]. stage [757]. stakeholder [2179, 2136]. stand [1150]. Standard [96, 1021, 1574, 850, 79]. Standardizing [2159, 641]. Standards [291, 247, 1453, 1682, 2229, 2125, 340, 2187, 1792, 1924, 496]. Standing [340, 496]. Stankus [76]. Stanley [332]. start [425, 2111, 1626]. State [2044, 1943, 203]. Statement [160, 261]. States [1002]. Stationers [363]. Statistics [1629, 1378, 1756, 1021, 1546, 1545, 1305, 1251, 1574, 314]. statutory [1062]. stay [1984, 1665]. Staying [1418]. steamship [1331]. Steering [795]. STEM [1900]. stemmed [1543]. Step [597, 541]. step-by-step [541]. Stephanie [400]. Stephen [563]. steps [2206, 597]. Stevenage [330, 447]. Stevens [35]. still [1809, 1696, 1375, 1750, 1624, 757]. sting [1964]. STM [561, 266, 1188, 984, 160, 1247, 1213, 1161, 1014, 566, 2206, 882, 1077, 1198, 841, 2011]. Stockers [472]. Stone [1733]. stones [1733]. stop [1166, 1945]. storage [408, 157]. Stories [1756]. story [1143]. Strategic [821]. Strategies [81, 1967, 825, 1900, 950, 1500, 1204, 755, 1843, 845]. strategy [312, 2193]. Streamline [1707]. streams [1467]. Street [119, 734, 318]. structural [2176]. structured [810]. stuck [1398]. student [1901]. student-edited [1901]. students [1923, 2004]. Studies [448, 1723, 1735, 2119, 2047, 1151, 1808, 1254, 1937, 2112]. Study [1649, 137, 1517, 2057, 2040, 1484, 1579, 181, 2205, 2130, 2028, 2092, 1872, 1819, 1949, 1928, 974, 2046, 83, 1134, 2186, 1187, 2105, 561, 2075, 2223, 2109, 2175, 2217, 1952, 1959, 1967, 1927, 2088, 1965, 2120, 1657, 1709, 990, 1751, 2042, 2140, 1296, 2017, 1958, 2059, 2136, 2192, 2023, 893, 909, 1941, 1260, 1486, 2079, 1805, 1799, 2226, 2022, 2133, 2218, 2183, 1969]. Sturt [354]. Stutely [200]. Style [1658, 458, 956, 697, 337, 2141]. stylebook [619]. styles [2036]. sub [1575]. sub-Saharan [1575]. submission [1955, 2076, 1888, 718, 1245, 782]. submissions [1718, 1995]. submit [1533, 1603, 1483, 1534]. submitting [1273]. Subscribe [2078]. subscribers [96, 96]. Subscription [919, 401, 918, 573, 285, 406, 2078, 1673, 1076, 1831, 942, 778, 1093, 1844, 1749, 1082]. subscription-based [1831, 1844]. Subscriptions [269, 637, 54, 914, 183, 1328, 931]. subsidiary [1985]. success [2219]. Successful [779, 1247, 1820]. successfully [398, 2030]. such [941, 1058]. Sue [522]. suggest [1952]. suggested [1952, 2149]. suggestions [1755, 2149]. suitability [1832]. Suite [420]. summaries [2228, 1719]. Summarizing [2074]. summary [1179, 1614, 902]. summer [916]. Sunningdale [377]. supplement [76]. supplied [1357]. suppliers [1524]. supply [937, 1913, 1367, 317, 1110, 1114]. support [2034, 2153, 475, 2095, 1792, 1295]. supporting [1378, 1921, 2207, 1359]. surf [698]. Surgery [1485]. Surrey [496]. survey [2152, 2190, 1357, 1984, 118, 1629, 1523, 1898, 1163, 1385, 2202, 1736, 2083, 1261, 1487, 1295, 1123, 1995, 2086, 1631, 2039, 2166, 2054]. surveyed [964, 176]. surveys [1788, 1699, 178]. survival [384]. survive [1815, 1616]. Surviving [960, 667]. Susan [74, 55]. SUSHI [1305]. Sustainability [1819, 1184, 1787, 2059]. Sustainable [1636, 1954, 1586, 1121, 1724, 2064, 1335, 1919]. SW10 [754]. SWJO [53]. symbiont [1238]. symbols [91]. symposium [1216]. synthetic [1212]. system [556, 1627, 561, 1732, 1093, 568, 1688, 2074, 1316, 661, 794, 1807, 569, 566, 7, 1611, 2132, 882, 1115, 1343]. systematic [2163, 2224, 2131]. Systems [240, 1876, 1376, 524, 381, 1258, 239, 354, 157, 679, 1391, 1245, 1327, 1602, 157, 36].

Tågerud [473]. Table [1797, 1812, 1825, 1839, 1852, 1864, 1880, 1892, 1903, 1915, 1933, 1947, 1963, 1978, 1989, 2007, 2020, 2033]. tables [1453]. tabs [36]. tactics [585]. tail [1964]. tailoring [898]. take [1957]. taking [1822]. tale [702]. talking [1945, 585]. Tampere [448]. tapestry [1445]. targeted [2229, 1821]. task [626, 359]. tax [281, 15, 130, 149]. Taxonomy [1998]. Taylor [118, 120, 115, 137, 448, 473, 476, 524, 221, 658, 397, 354, 497, 474, 447, 55, 36, 458, 1417, 289]. team [1774]. teams [2173]. TECH [157]. Technical [292, 1682, 550, 266, 965, 104, 1643, 472, 583, 1271]. techniques [125, 447, 55, 585]. Technologies [381, 2164, 553, 1356, 1212, 1874]. Technology [178, 772, 116, 1832, 157, 183, 2017, 304, 1519, 369, 355, 965, 294, 1489, 610, 1993, 1921, 2125, 1610, 165, 101, 1147, 2077, 619, 763, 1105, 747, 280, 1205, 183]. Technology/ALPSP [280]. tel [53]. telemarketing [1163]. tell [1282, 1138]. Telling [995]. Ten [1564, 1246, 1613, 1226]. tender [971]. Tenth [572]. tenuous [1452]. Term [413, 1900, 1136, 2001, 1938, 2192]. terminology [2159]. terms [852, 500, 447, 799, 421]. Terrace [237]. Terrorism [1133]. Terry [584]. tertiary [2088]. test [916]. Testing [1370, 1991]. Texaco [628]. Texas [620]. Text [1109, 239, 390, 1728, 2042, 2159, 527, 330, 1087, 55, 1122, 1409, 115, 239]. textbook [1202, 1003]. texts [839, 760]. Thames [401]. thank [2125]. Thanking [1931, 2005, 2065, 2128, 2210]. theft [677]. their [2180, 1753, 1984, 647, 1828, 1026, 2224, 2075, 1237, 1280, 1896, 1380, 266, 227, 1284, 2053, 1757, 2231]. them [302, 2194, 1613, 472, 1582]. theory [521]. therapy [2144]. There [1058, 1836, 669, 798, 1125, 1336, 1128, 1516, 1767, 590]. TheScientificWorldJOURNAL [1184]. These [333, 1624]. theses [291, 2109]. thesis [2030]. thing [941, 1058]. think [1613]. Thinking [1333, 1899]. Third [51, 1524, 138, 131, 253, 190, 145]. third-party [1524]. Thirty [1638]. Thompson [377]. Thought [1276]. thoughts [2001, 1040, 1489]. Thousands [2090]. threat [1064, 2216, 429, 1177]. threats [897]. three [1883, 1614, 2138, 1295, 2010, 544]. three-year [1614]. thrive [1815]. Tier [1405, 1071]. Tier-based [1405]. tiered [1531]. tiger [1374, 1572]. Time [1945, 1696, 272, 1545, 1166, 1818, 1604, 1427, 1295, 1203, 1235]. time-loop [1604]. timeliness [1994]. times [1471]. Tipping [1443]. tips [1280, 447]. title [1214]. today [1348, 1471, 2004, 188, 1308]. together [1519, 1984, 59, 1449, 1479, 452, 168, 1749, 1768]. Token [2095]. Token-curated [2095]. Tolson [1065]. tomorrow [792, 1308]. tongues [1280]. Tony [76, 313]. too [433, 1244, 1595, 1642]. tool [762, 1832, 1847, 1462]. Tools [249, 1610, 187, 1684]. Toronto [637]. tortoise [702]. Total [240, 1749]. TR [434]. Tracing [2204]. trade [401, 187, 273, 659]. trading [911]. tradition [1107]. traditional [1764, 1800, 911]. traditionalism [1228]. Training [707, 726, 748, 431, 1229, 330, 270, 994, 1092, 277, 341, 733, 2192, 19, 400]. trans [362]. trans-Continental [362]. transcript [449]. transfer [1004, 1447, 1596, 1758]. transform [1999]. transformation [1992, 2121, 1233]. Transformative [2142]. transforming [1741, 1128]. Transition [1709, 960, 2228, 1013, 2064, 1248, 2146, 1680, 2053, 2093]. translation [1247]. translational [2153]. translators [1520]. transparency [2096, 2209, 2231]. transparent [1588]. transplanting [792]. trap [1953]. Trash [1711]. treatment [1541]. trenches [1409]. TrendMD [1821]. Trends [1381, 1364, 1638, 51, 1179, 2143, 1801, 78, 1063, 2071, 931, 54, 286]. tribute [764, 1890]. trough [1460]. true [1224]. trumping [1912]. Trust [1709, 1804, 1717]. Trustee [632]. trustworthiness [1829, 1885]. Truth [670, 853, 1077]. trying [1057]. Tsing [1281]. tumbling [1422]. tune [1650]. Turkey [2109]. Turning [2031, 1810]. Twain [1431, 1754]. Twenty [1302, 1773, 757, 275, 399]. Twenty-five [1773, 757]. twice [182]. twists [1368]. Two [1808, 330, 1089, 129, 1699, 146, 2194, 1556, 1991, 1897, 2029, 1721, 2119]. type [1732]. Typefaces [753]. typesetting [92]. Typographical [343, 253, 575]. Typography [528]. Tysoe [563].

UCL [698, 1965]. UCPPC [167]. Ugly [1683]. UK [74, 561, 310, 376, 259, 377, 657, 755, 772, 937, 556, 1165, 1056, 2122, 811, 189, 929, 722, 588, 1820, 1921, 1787, 1855, 936, 330, 1781, 824, 1599, 1211]. UK.AC.uni.ex [272]. UKSG [687]. ultimate [2013]. Unbundled [1057]. uncertainty [1418]. unconventional [296]. uncovered [1830]. Underlying [2015]. Understanding [1778, 1536, 1660, 1340, 1188, 1869, 1600]. UNESCO [201, 742]. Unesco/ [742]. Unions [139, 222]. uniquely [1627]. United [708, 1002, 2209]. units [1188, 91]. universal [1921, 1847]. Universities [961, 1023, 1836, 1069, 1735, 722, 1420, 1200, 1599, 2085, 1062]. University [94, 238, 509, 637, 647, 180, 448, 697, 52, 1979, 716, 75, 354, 620, 219, 1954, 2147, 521, 2118, 1871, 1985, 1047, 2222, 1857, 1984, 1983, 128, 658, 1002, 2230, 1854, 100, 1234, 1281, 1430, 1248, 1861, 1855, 1981, 1856, 1194, 1860, 1204, 1072, 1308, 628, 1325, 1986, 986, 1212, 1853, 2109, 2044, 2120, 2087, 542]. university-based [1983, 986]. Unpriced [35]. Unpublishable [1730]. unreasonable [2061]. Unruh [119, 314, 352, 286, 220]. Untangling [2047]. until [1175]. untrustworthy [2093]. Unwin [138]. update [1806, 901, 782]. Uploading [1732]. upper [1809]. Urdang [117]. US$30.50 [286]. US$40 [183]. US$50 [119]. US$70 [183]. USA [542, 52, 598, 155, 159, 420, 1980]. Usage [1378, 1466, 1357, 463, 1629, 1021, 1546, 1545, 1831, 1370, 1305, 1251, 1574, 1498, 1502, 2079, 1077]. usage-based [1370]. Use [102, 1720, 1209, 1546, 408, 810, 1321, 2004, 1303, 1613, 923, 204, 99, 1364, 1042, 1958, 1560, 243, 1780, 2145, 1761, 1940, 453, 2086, 1631]. used [333, 1601, 721]. useful [1054, 2014]. usefulness [2231]. useless [1282, 1498]. User [103, 1793, 200, 549, 974, 1993, 1090, 812, 400, 661, 295, 1077, 1710]. Userid [272]. usernames [2216]. users [524, 1778, 1258, 1512, 1044, 1669, 1303, 2184, 1488, 1053, 1502]. uses [387, 1380]. Using [1268, 1432, 762, 407, 2164, 2157, 1850, 1484, 1579, 1524, 1971, 187, 1550]. Utah [2120]. UTB [1202]. utopias [1489]. Utrecht [1954].

V [391]. VA [598]. VAAP [410, 84]. Vakkari [448]. valid [1802]. valuable [1624]. Value [15, 130, 149, 1188, 973, 678, 1448, 2062, 2230, 1044, 1547, 1882, 1502, 36]. Value-added [36]. Valuing [1068, 2153]. vanity [1169]. VAT [418, 747, 15, 479]. VCH [581]. Vendor [1357]. Vendor-supplied [1357]. vendors [548, 2222]. Ventures [221, 474, 447, 131, 190]. venue [2094]. verbal [482, 760]. version [1387, 1318]. versions [1972]. versus [571, 212, 1894, 2214, 857, 2147]. very [1237]. vi [180, 620, 259]. viable [1481]. Vickers [754, 310]. Vietnam [2164, 2038]. Vietnamese [2139, 2108]. View [420, 1829, 938, 1447, 811, 549, 149, 336, 704, 1160, 507, 829, 2093, 1861, 21, 603, 1856, 504, 38, 569, 25, 1762, 144, 319, 987, 2013, 966]. viewed [1982]. viewpoint [890, 1273, 942, 457, 295]. views [544, 1280, 265, 1906, 1407, 2042]. vii [696, 313]. viii [355]. Viking [618]. village [1415]. violation [1341]. Virology [1707]. Virtual [870, 1017, 1720, 1710, 1345]. virtually [1836]. Viruses [668, 122]. Visegrád [2157]. visibility [1998, 1148, 1421, 1503, 1640, 1292]. Vision [1860, 1561]. visions [2147]. visit [410]. visually [1450]. Vocabulary [1675]. voices [1179, 2116]. Vol [981, 1039, 251]. Volume [423, 240, 2229]. volumes [330]. Voluntary [1124]. vortex [1179]. vs [1332, 949, 1107, 1407, 1318, 1628].

W [542, 377]. WA [52]. Wade [472]. Wall [582]. Walls [1422]. Walnut [119]. Want [909, 1433, 970, 1258, 1385, 792, 1468, 1200, 1053, 1195, 893, 1105, 1352, 1986]. Wanted [1618, 890]. war [667]. Ward [522]. was [159, 540, 380, 2003]. Washington [52, 35, 219, 114]. waters [1285]. Way [420, 2142, 1865, 1330, 955, 1310, 1122]. ways [1775, 1796]. WC1A [734]. wC1B [180]. WCAG [1972]. weakest [1090]. Weathering [2051]. Web [698, 1908, 927, 1504, 2157, 709, 775, 1445, 898, 1489, 762, 701, 950, 1883, 1241, 700, 1698, 1014, 1391, 1766, 1521, 1588]. web-based [927, 1391]. Webb [495]. webinar [2056]. website [962, 1944]. WeChat [2079, 2218]. Weckert [354]. Weekend [137]. Weinberg [203]. Weinheim [581]. Wel [1283]. well [989, 1994, 2215, 2029]. well-established [989, 2029]. Wellisch [377]. Wendy [544, 314, 353, 498, 352, 286]. we’re [1363, 2070]. Western [884]. We’ve [1912]. Where [682, 1867, 1953, 1559, 629, 1339, 1461]. which [1576]. while [1225]. whistle [1650]. White [563, 2004]. Whitelist [1956]. Whitepaper [1975]. Whither [1478, 1418]. Who [1034, 1674, 1753, 611, 549, 1221, 1031, 2072, 1288, 2184, 1216, 1696, 302, 192, 1626, 2073, 1654]. Whose [1466, 527, 1203]. Whurr [237]. WIA [449]. Wicks [544, 498]. Wide [709, 775, 698, 701, 700, 1021, 63, 1548, 1558, 762]. Wiley [2054, 2166, 1358, 1805]. Will [553, 546, 904, 1488, 1103, 1028]. William [140, 420]. Wilson [377, 156]. Wilson-Davies [156]. Wiltshire [583]. Winchester [472]. Winter [105]. WIPO [943]. wish [1670]. within [1735, 1319, 1240, 2139, 144, 2023, 2133]. without [1045]. Witney [657]. wolves [1149]. Women [182]. Won’t [1138]. Wood [290, 137, 617]. Woodward [755]. Woodwark [472]. word [1216, 224, 180, 738, 869, 319, 1845]. wordprocessing [205]. Words [304, 2130, 598, 1237]. work [172, 23, 1992, 1803, 1866, 1766, 2222, 1874]. Workflow [1649, 1657, 1560]. workflows [1845, 1269]. Working [386, 59, 261, 208, 522, 1519, 1865, 1449, 452, 1920]. works [408, 413]. workshop [354]. World [486, 698, 743, 1636, 385, 159, 532, 1222, 1720, 63, 2177, 1518, 572, 1920, 2038, 1406, 709, 775, 762, 481, 131, 701, 190, 700, 145, 237]. world-wide [63, 762]. worlds [146]. Worshipful [363]. Worth [2063, 1395, 1500]. worthwhile [2003]. worthy [1530]. Wrangling [1845]. Write [1533, 1483, 716, 1729, 1603]. writer [1081, 238]. Writers [53, 439, 618, 734, 53, 205]. Writing [507, 576, 2159, 1850, 2130, 301, 391, 239, 1104, 474, 472, 530, 2192, 583, 332, 398, 447]. wrong [593, 1347].

X [331, 355, 353, 288, 522, 476, 352, 332, 203, 619, 289, 74, 114, 584, 2222]. X/12 [447]. xi [353, 398]. xii [638, 678, 155, 288, 495, 92]. xiii [36]. xiv [331, 620, 219]. XML [1845, 1020]. xv [697]. xvi [238]. xviii [117]. xxix [542]. xxv [239]. xxvi [377].

Yael [473]. Year [842, 1081, 2157, 116, 1045, 1614, 1791, 1765, 2192, 2206, 637, 116]. Yearbook [222]. years [2071, 275, 399, 1564, 1302, 389, 1246, 1613, 1226, 1624, 1747, 1954, 404, 757, 758]. Yesterday [188, 792]. York [73, 76, 397, 203, 377, 1208]. yourself [586]. you’ve [1432]. Yugoslavia [191].

Z [447, 377]. Zealand [1253]. Zeitlyn [586]. Zell [540]. zeroes [626]. Zhejiang [1047]. ’zine [987]. zone [667]. ZWOLLE [1150].

References

Paulus:1988:GEW

[1] Kurt Paulus. Guest editorial: What’s in a name? Learned Publishing, 1 (1):1, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Buchan:1988:MBB

[2] Amanda Buchan. The market for British books in Japan. Learned Publishing, 1(1):2–4, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Forbes:1988:CEJ

[3] Dennis Forbes. Cost-effective journal production: a management overview. Learned Publishing, 1(1):5–10, January 1988. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morley:1988:DIP

[4] Neil Morley. dBase III Plus. Learned Publishing, 1(1):11–14, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:1988:COB

[5] Alan Singleton. Co-operation between publishers and societies on journals. Learned Publishing, 1(1):15–18, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Muller:1988:EC

[6] Gervase Muller. The Epsilon concept. Learned Publishing, 1(1):19–22, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Summers:1988:AFC

[7] Lorna Summers. The address file: a computerized order system. Learned Publishing, 1(1):23–27, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1988:AA

[8] Edwin Shelock. ALPSP activities. Learned Publishing, 1(1):30–31, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:SRa

[9] Anonymous. Seminar reports. Learned Publishing, 1(1):32–34, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:DDa

[10] Anonymous. Domesday demonstration. Learned Publishing, 1(1):35–36, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:MDa

[11] Anonymous. Marketing and distribution. Learned Publishing, 1(1):37– 39, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McCarney:1988:BCN

[12] Flavia McCarney. British Council news. Learned Publishing, 1(1):40, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:EPa

[13] Anonymous. Electronic publishing. Learned Publishing, 1(1):41–43, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1988:CCa

[14] John St. Aubyn. Copyright corner. Learned Publishing, 1(1):44–45, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Caird-Daley:1988:VAT

[15] D. P. Caird-Daley. Value added tax: the impact of VAT on learned society publishing. Learned Publishing, 1(1):46–48, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:RAa

[16] Anonymous. Reading around. Learned Publishing, 1(1):48, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:BR

[17] Anonymous. Books reviewed. Learned Publishing, 1(1):49–51, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:DPR

[18] Anonymous. Data Protection Register. Learned Publishing, 1(1):52–53, January 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:CEC

[19] Anonymous. Coming events: Conferences, Meetings, Fairs, Exhibitions, Training. Learned Publishing, 1(1):54–56, January 1988. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Page:1988:CSA

[20] Gillian Page. Coming to some arrangement. Learned Publishing, 1(2): 57–60, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Merriman:1988:JPM

[21] John B. Merriman. Journal pricing — the market view. Learned Publishing, 1(2):61–66, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bradley:1988:PAP

[22] Clive Bradley. Piracy and anti–piracy. Learned Publishing, 1(2):66– 67, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clark:1988:CW

[23] Chairman Charles Clark. The CLA and its work. Learned Publishing, 1 (2):68–70, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1988:GA

[24] Journals Office Michele Benjamin. Guidelines for authors. Learned Publishing, 1(2):70–73, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1988:AAC

[25] Edwin Shelock. ALPSP activities: The Chairman’s view. Learned Publishing, 1(2):74–79, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1988:CCb

[26] John St. Aubyn. Copyright corner. Learned Publishing, 1(2):82–83, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:IOa

[27] Anonymous. Information organization. Learned Publishing, 1(2):84–85, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:EPb

[28] Anonymous. Electronic publishing. Learned Publishing, 1(2):86–88, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:MDb

[29] Anonymous. Marketing and Distribution. Learned Publishing, 1(2):89, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:BCNa

[30] Anonymous. British Council news. Learned Publishing, 1(2):90–92, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1988:RA

[31] A. B. Britton. Reading around. Learned Publishing, 1(2):93–95, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:PAa

[32] Anonymous. From the Publishers Association. Learned Publishing, 1(2): 96–98, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:CDR

[33] Anonymous. Correspondence: Dear Reader. Learned Publishing, 1(2): 99, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:1988:BRJ

[34] Kurt Paulus. Book review: Journal Publishing: Principles and Practice, by Gillian Page, Robert Campbell and Jack Meadows. Butterworths, London. 1987. 184 pp. ISBN 0-408-10716-2. £27.00. Learned Publishing, 1(2):100, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Katzen:1988:BRS

[35] May Katzen. Book review: Scholarship in the electronic age: a selected bibliography on research and communication in the humanities and social sciences, by A. Lowry and B. Stevens. Washington DC, Council on Library Resources Inc., February 1987. Unpriced. Learned Publishing, 1(2):100–101, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Seal:1988:BRV

[36] Alan Seal. Book review: Value-added processes in information systems, by Robert S. Taylor. Norwood, NJ: Ablex Publishing Corp., 1986. xiii, 257 pp. 24 cm. Illus, tabs, bibliog, indexes. ISBN 0-89391-273-5 (cased): £37.95. Learned Publishing, 1(2):101–102, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:CEa

[37] Anonymous. Coming events. Learned Publishing, 1(2):103–106, April 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ross:1988:PBM

[38] Neill Ross. The pricing of books: a market view. Learned Publishing, 1(3):105–115, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Long:1988:PJF

[39] Maurice Long. Pricing journals in foreign currencies. Learned Publishing, 1(3):116–118, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Trevitt:1988:DDM

[40] John Trevitt. Does design matter? Learned Publishing, 1(3):118–119, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:SRb

[41] Anonymous. Seminar reports. Learned Publishing, 1(3):120–123, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:AA

[42] Anonymous. ALPSP activities. Learned Publishing, 1(3):124–125, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1988:CCc

[43] John St. Aubyn. Copyright corner. Learned Publishing, 1(3):126–127, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:EPc

[44] Anonymous. Electronic publishing. Learned Publishing, 1(3):130–132, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:DDb

[45] Anonymous. Database developments. Learned Publishing, 1(3):133–134, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:IOb

[46] Anonymous. Information organization. Learned Publishing, 1(3):135, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:RAb

[47] Anonymous. Reading around. Learned Publishing, 1(3):136–137, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:BCNb

[48] Anonymous. British Council news. Learned Publishing, 1(3):138–139, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:MDc

[49] Anonymous. Marketing and distribution. Learned Publishing, 1(3):140– 142, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:PAb

[50] Anonymous. From the Publishers Association. Learned Publishing, 1(3): 143–144, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1988:BRTa

[51] Michele Benjamin. Book review: The Third International learned journals seminar: the changing market for serials-new trends in acquisitions. Epsilon Press, Letchworth, Herts. 1987. 66pp. ISBN 1850270171. £7.50, post free. Learned Publishing, 1(3):145, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clarke:1988:BRD

[52] Malcolm Clarke. Book review: Disseminating Asia’s scholarly books, edited by E. Pacheco: International Association of Scholarly Publishers (c/o Dorothy J. Anthony, University of Washington Press, PO Box C-50096, Seattle, WA 98145-5096, USA). 1987. ISBN 82-00-07842-6. $40. Learned Publishing, 1(3):146, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Burkhart:1988:BRG

[53] Sue Burkhart. Book review: Guidelines for medical writers. Revised edition 1985. Society of Authors, Medical Writers Group, 84 Drayton Gardens, London SWJO 9SB (tel: 01-373 6642). 15pp roneoed [sic]. Free to members; £1 non-members. Learned Publishing, 1(3):146–147, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1988:BRTb

[54] Michele Benjamin. Book review: Trends in journal subscriptions 1986. London: The Publishers Association. 1987. 31 pp. ISBN 0-85386-146-3. Learned Publishing, 1(3):147–148, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sakne:1988:BRI

[55] Allan Sakne. Book review: Information handling techniques for the office: full text rules OA?, edited by Susan Hills. Taylor Graham. 1987. 164 pp. ISBN 0-947568-33-6. £18.00. Learned Publishing, 1(3):148–149, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:CEb

[56] Anonymous. Coming events. Learned Publishing, 1(3):150–153, July 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Caplin:1988:DPC

[57] Stephen Caplin. Desktop publishing: complete editorial control. Learned Publishing, 1(4):153–155, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Welham:1988:BYL

[58] Robert Welham. Building up your list: books. Learned Publishing, 1 (4):156–158, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1988:WT

[59] Michele Benjamin and John Flower. Working together. Learned Publishing, 1(4):159–162, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1988:PPB

[60] Michele Benjamin. Publishers publicity and bibliographic databases. Learned Publishing, 1(4):162–163, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1988:PBJ

[61] H.K.B. Publishing books and journals for profit. Learned Publishing, 1(4):164, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:JLM

[62] Anonymous. Joint liaison meetings. Learned Publishing, 1(4):164, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1988:WWH

[63] H.K.B. The world-wide hunger for books. Learned Publishing, 1(4): 165–172, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:CA

[64] Anonymous. Charlesworth Award. Learned Publishing, 1(4):173, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1988:AAD

[65] Bernard Donovan. ALPSP activities: From the desk of the Secretary. Learned Publishing, 1(4):174–175, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1988:SCL

[66] H.K.B. Soviet copyright liaison. Learned Publishing, 1(4):176–178, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1988:CCd

[67] John St. Aubyn. Crown copyright. Learned Publishing, 1(4):178–179, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:DEO

[68] Anonymous. Distribution: Electronic ordering of publications. Learned Publishing, 1(4):182–183, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:IOc

[69] Anonymous. Information organization. Learned Publishing, 1(4):184– 185, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Long:1988:M

[70] Chairman Maurice Long. Marketing. Learned Publishing, 1(4):186–188, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:BCNc

[71] Anonymous. British Council news. Learned Publishing, 1(4):189–192, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:RAc

[72] Anonymous. Reading around. Learned Publishing, 1(4):193, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Burton:1988:BRS

[73] Paul F. Burton. Book review: Serial connections: people, information, communication. Proceedings of the North American Serials Interest Group 1st Annual Conference, 22-25 June, 1986, Bryn Mawr College, Bryn Mawr, Pennsylvania, edited by Leigh Chatterton and Mary Elizabeth Clack. New York, London: Haworth Press, 1987. 200pp. ISBN 0-86656-654-6 (The Serials Librarian 11 (3/4) 1986/87). Learned Publishing, 1(4):194–195, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1988:BRS

[74] A. B. Britton. Book review: Serials ’87: Proceedings of the UK Serials Group Conference 1987 (Serials Monograph No. 10), edited by Susan Frost. Oxford: Blackwells. £10.00. ISBN 0-906148-09-X. Learned Publishing, 1(4):195–196, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dunkley:1988:BRC

[75] Bernard Dunkley. Book review: Computing for non-scientific applications, by Derek Andrews and Michael Greenhalgh. Leicester: Leicester University Press, 1987. 346 pp. £15.95. ISBN 0-7185-1252-9. Learned Publishing, 1(4):197, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Green:1988:BRS

[76] P. R. Green. Book review: Scientific journals: issues in library selection and management. (Monographic supplement no. 3 to the Serials Librarian), by Tony Stankus. New York: Haworth Press, 1987. 218 pp. ISBN 0-86656-616-3. Learned Publishing, 1(4):197–198, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1988:CEd

[77] Anonymous. Coming events. Learned Publishing, 1(4):199–201, October 1988. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stevens:1989:MJS

[78] Allen Stevens. The market for journals and serials in the 1980s: an analysis of present and future trends. Learned Publishing, 2(1):1–7, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cave:1989:SGM

[79] Francis J. Cave. The Standard Generalized Markup Language. Learned Publishing, 2(1):8–14, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lomax:1989:YOG

[80] Robert Lomax. You are only as good as your list: journals. Learned Publishing, 2(1):15–19, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1989:SP

[81] H.K.B. Strategies for publishing. Learned Publishing, 2(1):20–23, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1989:FFL

[82] A. B. Britton. The future for the funding of library purchasing — books and periodicals. Learned Publishing, 2(1):24–25, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1989:FSC

[83] H.K.B. Financial savings: a case study in desktop publishing. Learned Publishing, 2(1):25–27, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1989:APV

[84] Edwin Shelock. The ALPSP protocol with VAAP. Learned Publishing, 2 (1):30, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1989:DS

[85] Bernard Donovan. From the desk of the Secretary. Learned Publishing, 2 (1):31, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1989:CCa

[86] John St. Aubyn. Copyright corner. Learned Publishing, 2(1):32–33, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:PQ

[87] Anonymous. Project Quartet. Learned Publishing, 2(1):34, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:IOa

[88] Anonymous. Information organization. Learned Publishing, 2(1):35–37, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1989:RA

[89] A. B. Britton. Reading around. Learned Publishing, 2(1):38–40, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:BCNa

[90] Anonymous. British Council news. Learned Publishing, 2(1):41–44, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Keen:1989:BRU

[91] Peter J. Keen. Book review: Units, symbols, and abbreviations: a guide for biological and medical editors and authors, (fourth edition); edited by D. N. Baron. London: Royal Society of Medicine Services. 1988. 64 p. Paperback. ISBN 0-905958-78-0. £5.00. Learned Publishing, 2(1):45–46, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:1989:BRB

[92] Mark Ware. Book review: Bringing typesetting in-house: the options for publishers, Kathy Munro. London: Blueprint, 132 + xii pp. ISBN 0-948905-16-6. £25. Learned Publishing, 2(1):46–48, January 1989. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Page:1989:BRP

[93] Gillian Page. Book review: Publishing-the future, edited by Peter Owen. London: Peter Owen. 1988. 128 p. 22x 14.3 cm. Hardback ISBN 0-7206-0720-5. £10.5; paperback 0-7206-7201-3 price £10.95. Learned Publishing, 2(1):48–49, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:BRC

[94] Anonymous. Book reviews: The complete guide to corporate desktop publishing, by Robert Jones. Cambridge University Press, 1988. 132 p. Hardback ISBN 0-521-35179-0; paperback ISBN 0-521-35973-2. Learned Publishing, 2(1):49, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Feeney:1989:BRD

[95] Mary Feeney. Book review: This is desktop ’88, edited by Kay Floyd. Blueprint Publishing, 1988. 160 p. ISBN 0-948905-23-9. £29.95. Learned Publishing, 2(1):49–51, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lavell:1989:BSR

[96] Cherry Lavell. British Standard recommendations for preparing indexes to books, periodicals and other documents BS 3700: 1988. (Revision of 1976 edition). 17 p. £10 to BSI subscribers, £20 to non-subscribers. Learned Publishing, 2(1):51–52, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:CEa

[97] Anonymous. Coming events. Learned Publishing, 2(1):53–56, January 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1989:ILJ

[98] Edwin Shelock. The International Learned Journal Seminars. Learned Publishing, 2(2):57–58, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pickering:1989:USI

[99] Roy Pickering. The use and source of information in a research institution. Learned Publishing, 2(2):59–67, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lamble:1989:SFI

[100] J. H. Lamble. Source and flow of information in a university community. Learned Publishing, 2(2):69–75, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kent:1989:ITS

[101] A. K. Kent. The impact of technology on serials. Learned Publishing, 2 (2):75–77, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Haygrath-Jackson:1989:UIB

[102] Angela R. Haygrath-Jackson. Use of the information budget. Learned Publishing, 2(2):78–86, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oakeshott:1989:UAO

[103] Priscilla Oakeshott. User attitudes to online document ordering. Learned Publishing, 2(2):87–89, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

D:1989:FAS

[104] B. T. D. A forum on abstracts of scientific or technical articles. Learned Publishing, 2(2):91, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

D:1989:WGM

[105] B. T. D. Winter General Meeting. Learned Publishing, 2(2):91–92, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

D:1989:M

[106] B. T. D. Mailsort. Learned Publishing, 2(2):92–93, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1989:AAD

[107] Bernard Donovan. ALPSP activities: From the desk of the Secretary. Learned Publishing, 2(2):94, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1989:CCb

[108] John St. Aubyn. Copyright corner. Learned Publishing, 2(2):95–96, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:IOb

[109] Anonymous. Information organization. Learned Publishing, 2(2):97–98, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:BLN

[110] Anonymous. British Library news. Learned Publishing, 2(2):98, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

S:1989:M

[111] E. A. S. Marketing. Learned Publishing, 2(2):99, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:BCNb

[112] Anonymous. British Council news. Learned Publishing, 2(2):100–103, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:OD

[113] Anonymous. Overseas developments. Learned Publishing, 2(2):103–104, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:1989:BRB

[114] Kurt Paulus. Book review: Books in our future: perspectives and proposals, edited by John Y. Cole. Washington: Library of Congress. 1987. 399 pp. ISBN 0-8444-0554-X. Learned Publishing, 2(2):105, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Blackmore:1989:BRT

[115] John Blackmore. Book review: Text retrieval — an introduction, edited by Rowlands Ian. London: Taylor Graham. 1987. 80 pp. £15.00. Learned Publishing, 2(2):105–106, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Burton:1989:BRL

[116] Paul Burton. Book review: Library and information briefings. London: British Library Research and Development and Library and Information Technology Centre, 1988. 10 issues per year. £50.00 per year. Learned Publishing, 2(2):106–107, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dauker:1989:BRL

[117] Hermi Däuker. Book review: Loanwords dictionary, 1st ed., edited by Laurence Urdang and Frank R. Abate. A Laurence Urdang Reference Book. Detroit: Gale Research Co. 1988. pp. xviii + 324. $80.00. ISBN 0-8103-1543-2. Learned Publishing, 2(2):107–108, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Blackmore:1989:BRC

[118] John Blackmore. Book review: Computerisation in academic departments — a survey of current practice, by Richard Heeks. British Library Research and Development Report 5939. London: Taylor Graham, 1987. 144 pp. £18.00. Learned Publishing, 2(2):108, April 1989. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1989:BRE

[119] Michele Benjamin. Book review: Effective exhibiting: a practical guide, by Betty Unruh. National Federation of Abstracting and Information Services, 1429 Walnut Street, Philadelphia, PA 19102. May 1988. 190 pp. US$50. ISBN 0-942308-23-9. Learned Publishing, 2(2):108–109, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Blackmore:1989:BRO

[120] John Blackmore. Book review: Online information retrieval in practice, edited by Linda Dorrington. London: Taylor Graham, 1987. £18.00. Learned Publishing, 2(2):109, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:CEb

[121] Anonymous. Coming events. Learned Publishing, 2(2):110–113, April 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

D:1989:CV

[122] B. T. D. Computer viruses. Learned Publishing, 2(3):113–118, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Carr:1989:DD

[123] Cindy Carr. The DSC’s defence. Learned Publishing, 2(3):119–120, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cooter:1989:EPE

[124] Margaret Cooter. Excited publishers — experienced editors. Learned Publishing, 2(3):121–123, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Birn:1989:MRT

[125] Robin J. Birn. Market research techniques in publishing. Learned Publishing, 2(3):125–128, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Skelley:1989:SMC

[126] Eva Skelley. Serial markets in China and Eastern Europe. Learned Publishing, 2(3):129–134, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McCarney:1989:MBB

[127] Flavia McCarney. The market for British books in Mexico. Learned Publishing, 2(3):134–137, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

D:1989:SUP

[128] B. T. D. Small university presses. Learned Publishing, 2(3):138, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1989:FIL

[129] Michele Benjamin. Fifth International Learned Journals Seminar: Europe and North America: two markets or one? Learned Publishing, 2(3):138–142, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

D:1989:VATa

[130] B. T. D. Value added tax and the 1989 budget — learned and professional societies. Learned Publishing, 2(3):142–144, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1989:COV

[131] H.K.B. Co-operative ventures with the Third World. Learned Publishing, 2(3):145–146, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1989:CCc

[132] John St. Aubyn. Copyright corner. Learned Publishing, 2(3):147–148, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

T:1989:AAD

[133] B. T. D. ALSP activities From the desk of the Secretary. Learned Publishing, 2(3):149–151, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

D:1989:PAN

[134] B. T. D. Publishers Association news. Learned Publishing, 2(3):151, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Raymer:1989:IO

[135] M. M. Raymer. Information organization. Learned Publishing, 2(3):152– 154, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:BCNc

[136] Anonymous. British Council news. Learned Publishing, 2(3):155–157, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1989:BRR

[137] A. B. Britton. Book review: Resource allocation in industrial and commercial libraries, edited by Lawraine Wood. Proceedings of the Library Association Industrial Group 1988 Annual Study Weekend, Great Malvern. London: Taylor Graham. 1988. 67 pp. ISBN 0-847568-39-5. Learned Publishing, 2(3):158–159, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1989:BRP

[138] John St. Aubyn. Book review: Publishing agreements — a book of precedents (third edition), edited by Charles Clark. London: Unwin Hyman Ltd. 1988. 238 pp. £20.00. ISBN 0-04-440237-6. Learned Publishing, 2(3): 159, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:1989:BRG

[139] Kurt Paulus. Book review: Guidelines for printing and publishing by International Council of Scientific Unions. Paris: ICSU Press, 1988. 52 pp. $25.00. ISBN 0-930357-16-7. Learned Publishing, 2(3):160, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Feeney:1989:BRI

[140] Mary Feeney. Book review: The information society, by Martin, William J.. London: Aslib, 1988. 174 pp. £24.00 (£18.00 to members of Aslib). ISBN 0-85142-219-5. Learned Publishing, 2(3):160–161, July 1989. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1989:SR

[141] Michele Benjamin. Software review. Learned Publishing, 2(3):163–165, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:CEc

[142] Anonymous. Coming events. Learned Publishing, 2(3):166–169, July 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1989:MM

[143] Bernard Donovan. Mission to Moscow. Learned Publishing, 2(4):169–172, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Strauch:1989:ALM

[144] Katina Strauch. The American library market: a view from within. Learned Publishing, 2(4):173–181, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Read:1989:PDT

[145] Tony Read. Publishing developments in the Third World. Learned Publishing, 2(4):183–190, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gannett:1989:PSI

[146] Elwood K. Gannett. Primary and secondary information services: two worlds or one? Learned Publishing, 2(4):191–205, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Campbell:1989:BBP

[147] William Campbell. British book promotion in Chile. Learned Publishing, 2(4):206–207, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1989:PQ

[148] H.K.B. Project Quartet. Learned Publishing, 2(4):208–210, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

D:1989:VATb

[149] B. T. D. Value added tax — the view from the other side. Learned Publishing, 2(4):211–212, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:CC

[150] Anonymous. Copyright corner. Learned Publishing, 2(4):213, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:PAN

[151] Anonymous. Publishers Association news. Learned Publishing, 2(4):214, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:DS

[152] Anonymous. From the desk of the Secretary. Learned Publishing, 2(4): 215, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:IOc

[153] Anonymous. Information organization. Learned Publishing, 2(4):216– 217, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:BCNd

[154] Anonymous. British Council news. Learned Publishing, 2(4):218–219, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1989:BRG

[155] David Lee. Book review: A guide to information sources for the preparation, editing and production of documents, by Anderson, Dorothy. Aldershot, Hants., Brookfield, USA: Gower. 1989. xii, 142pp. £19.50. ISBN 0-566-05743-3. Learned Publishing, 2(4):220, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:1989:BRD

[156] Mark Ware. Book review: Desktop Publishing, 3rd ed. By Wilson-Davies, Kirty, John Bate, Joseph St. and Barnard, Michael. London: Blueprint (Publisher’s Guide Series). 206pp. ISBN 0-948905-39-5. £19.95. (c&j). Learned Publishing, 2(4):220–222, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1989:BRO

[157] David Lee. Book review: Optical disk storage and document image processing systems: buyers’ guide to products and systems. (CIM-TECH Publication No. 29), compiled by Shiel, Arthur. Hatfield, Herts, Cimtech, National Centre for Information and Technology, 1989. 36pp. £18.00 (£15.00 to Cimtech members) ISBN 0-852-67274-8. Learned Publishing, 2(4):222, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1989:CEd

[158] Anonymous. Coming events. Learned Publishing, 2(4):223–225, October 1989. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Asser:1990:GBW

[159] Paul Nijhoff Asser. The Gutenberg Bible was published in the USA: legal deposit in the modern world. Learned Publishing, 3(1):1–7, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:SPS

[160] Anonymous. STM Policy Statement on Legal Deposit (‘dépôt légal’). Learned Publishing, 3(1):7–9, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Keenan:1990:IPF

[161] Stella Keenan. International publishing: the FID experience. Learned Publishing, 3(1):11–18, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smith:1990:R

[162] Jane Smith. Refereeing. Learned Publishing, 3(1):19–25, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Toney:1990:IBB

[163] Terence Toney. The importation of British books into Colombia. Learned Publishing, 3(1):25–27, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prout:1990:SM

[164] Peter Prout. The Spanish market. Learned Publishing, 3(1):28–31, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Keenan:1990:ITR

[165] Stella Keenan. Information technology and the research process. Learned Publishing, 3(1):32–39, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1990:EN

[166] Michele Benjamin. Editing into the nineties. Learned Publishing, 3(1): 40, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1990:UAG

[167] Hazel Bell. UCPPC Annual General Meeting and Conference. Learned Publishing, 3(1):41–46, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1990:LJP

[168] H.K.B. The learned journal: putting it together. Learned Publishing, 3(1):46–48, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:EJP

[169] Anonymous. Excellence in journal publishing: Second annual Charlesworth Award. Learned Publishing, 3(1):49–50, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:MAE

[170] Anonymous. Meritorious Award for editor. Learned Publishing, 3(1):51, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:IC

[171] Anonymous. Index commended. Learned Publishing, 3(1):51, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:IWA

[172] Anonymous. Information work award. Learned Publishing, 3(1):51, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1990:CCa

[173] John St. Aubyn. Copyright corner. Learned Publishing, 3(1):52–53, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:IOa

[174] Anonymous. Information organization. Learned Publishing, 3(1):54–55, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Long:1990:BD

[175] Maurice Long. Book data. Learned Publishing, 3(1):56, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Magalhaes:1990:PMS

[176] Rodrigo Magalhaes. Portuguese market surveyed. Learned Publishing, 3(1):57–58, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:BCNa

[177] Anonymous. British Council news. Learned Publishing, 3(1):59–61, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:TS

[178] Anonymous. Technology Surveys. Learned Publishing, 3(1):62, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:RAa

[179] Anonymous. Reading around. Learned Publishing, 3(1):63, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1990:BRW

[180] Michele Benjamin and Maeve O’Connor. Book review: From word processing to publishing: guidelines for authors. Denley, Peter (compiler). London: University College & Professional Publishers Council of The Publishers Association and the Association of Learned and Professional Society Publishers (distributed by the Publishers Association, 19 Bedford Square, London wC1B 3HJ). 1989. vi + 29 pp. ISBN 0-85386-64-1. Learned Publishing, 3(1):64, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Burton:1990:BRS

[181] Paul F. Burton. Book review: A study of library economics in the European Communities. Final report (June 1987). /Ph. Ramsdale. Luxembourg: Office for Official Publications of the European Communities, 1988. 149 pp. (Information Management series; EUR 11546 EN). ISBN 92-825-8347-3. 12.50 ECU. Learned Publishing, 3(1):65, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1990:WPT

[182] Michele Benjamin. Women in Publishing: twice as many, half as powerful? 1989. 66 pp + 4 appendices. Available from the Polytechnic of North London Business School, Holloway Road, London N7 8DB. £4.95. Learned Publishing, 3(1):65–66, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oakeshott:1990:BRL

[183] Priscilla Oakeshott. Book review: Language Technology. Bi-monthly; individual subscriptions US$40, institutions US$70, p.a.; published by Language Technology BV (po Box 70486, 1007 KL Amsterdam). Learned Publishing, 3(1):66–67, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bingham:1990:IBE

[184] A. Bingham. Information booklet on ECDIN. Learned Publishing, 3(1): 67, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:CEa

[185] Anonymous. Coming events. Learned Publishing, 3(1):70–72, January 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Law:1990:PPP

[186] Derek Law. Publishers’ pricing policies and library attitudes. Learned Publishing, 3(2):73–77, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Penfold:1990:LJU

[187] David W. Penfold. The learned journal: using the tools of the trade. Learned Publishing, 3(2):78–88, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Meadows:1990:YJT

[188] Jack Meadows. Yesterday’s journals and today’s. Learned Publishing, 3 (2):89–91, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clark:1990:UCD

[189] Charles Clark. The UK Copyright Designs and Patents Act 1988: moral rights. Learned Publishing, 3(2):92–94, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Parry:1990:COV

[190] Rex Parry. Co-operative ventures with the Third World: practical proposals and the involvement of international organizations. Learned Publishing, 3(2):95–101, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bakarsic:1990:LPL

[191] Kemal Bakarsic. Learned publishing and libraries in Yugoslavia. Learned Publishing, 3(2):102–110, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1990:AWW

[192] Bernard T. Donovan. ALPSP: who are we? Learned Publishing, 3(2): 111–114, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1990:E

[193] Michele Benjamin. Electrocopying. Learned Publishing, 3(2):115, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Marsh:1990:BSI

[194] Corril V. Marsh. Book and serial issues. Learned Publishing, 3(2):116– 117, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1990:CCb

[195] John St. Aubyn. Copyright corner. Learned Publishing, 3(2):117–118, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

D:1990:DS

[196] B. T. D. From the desk of the Secretary. Learned Publishing, 3(2): 119, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:IOb

[197] Anonymous. Information organization. Learned Publishing, 3(2):120– 123, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1990:D

[198] Edwin Shelock. Distribution. Learned Publishing, 3(2):124–125, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:BCNb

[199] Anonymous. British Council news. Learned Publishing, 3(2):126–127, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1990:BRS

[200] Michele Benjamin and Francis Cave. Book review: SGML — the user’s guide to ISO 8879, by Smith, Joan M. and Stutely, Robert. Chichester: Ellis Horwood/Halstead Press. 172 pp. ISBN 0-470-21126-1. £22.50. Learned Publishing, 3(2):128, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1990:BRM

[201] A. B. Britton. Book review: Model handbook for interlending and copying, by Cornish, Graham P.. Boston Spa: IFLA Office for International Lending/UNESCO, 1988. 80 pp. Paperback. ISBN 0-7123-2045-8. Learned Publishing, 3(2):128–129, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oakeshott:1990:BRF

[202] Priscilla Oakeshott. Book review: The future of small presses in scholarly publishing. Proceedings from the 4th International Conference on Scholarly Publishing, Helsinki, June 1-3, 1988. Published 1988 by the IASP, 159 p. $40. Learned Publishing, 3(2):129–130, April 1990. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lavell:1990:BRI

[203] Cherry Lavell. Book review: Indexing: the state of our knowledge and the state of our ignorance. Proceedings of the 20th annual meeting of the American Society of Indexers, New York City, 13 May 1988, edited by Weinberg, Bella Hass. Medford, NJ: Learned Information. 1989. x + 134 pp. £30.00. Learned Publishing, 3(2):130–131, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Page:1990:BRE

[204] Gillian Page. Book review: The effective use of market research: a guide for management. Birn, Robin. London: Kogan Page. 1988. 156 pp. ISBN 1-85091-65-3. Learned Publishing, 3(2):132, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1990:BRM

[205] David Lee. Book review: Making electronic manuscripts: a wordprocessing guide for writers. Barnard, Michael. London: Blueprint Publishing. 1989. (4), 28 pp. ISBN 0-948905-38-7. £2.95. Learned Publishing, 3(2): 132–133, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:CEb

[206] Anonymous. Coming events. Learned Publishing, 3(2):134–136, April 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1990:ELS

[207] Edwin Shelock. The economics of learned society publishing. Learned Publishing, 3(3):137–148, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:1990:WCP

[208] Sally Morris. Working with a commercial publisher: guidelines for learned societies. Learned Publishing, 3(3):149–154, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clarke:1990:MBJ

[209] Malcolm Clarke. Marketing books and journals in Japan. Learned Publishing, 3(3):155–161, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smith:1990:JIC

[210] Nigel Smith. Journal indexing: compiling the Legal Journals Index. Learned Publishing, 3(3):162–176, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:PM

[211] Anonymous. Portrait of a member. Learned Publishing, 3(3):177–178, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1990:SJL

[212] Michele Benjamin. The Scholarly Journal: The Learned Society versus the Commercial Publisher. Learned Publishing, 3(3):179–181, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1990:LSP

[213] H.K.B. The learned society and its publications. Learned Publishing, 3(3):181–182, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1990:CCc

[214] John St. Aubyn. Copyright corner. Learned Publishing, 3(3):183, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:IOc

[215] Anonymous. Information organization. Learned Publishing, 3(3):184– 186, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:BCNc

[216] Anonymous. British Council news. Learned Publishing, 3(3):187–189, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:OD

[217] Anonymous. Overseas developments. Learned Publishing, 3(3):190–191, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:RAb

[218] Anonymous. Reading around. Learned Publishing, 3(3):192, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

OConnor:1990:BRG

[219] Maeve O’Connor. Book reviews A guide to book publishing, revised ed. 1989. Smith, Datus C. Seattle & London: University of Washington Press. xiv + 268 pp. Cloth, ISBN 0-295-96651-3, $25.00; Paper, ISBN 0-295-96652-1, $12.50. Learned Publishing, 3(3):193, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:1990:BRI

[220] Mark Ware. Book review: Information marketing handbook. Edited by Betty Unruh. Philadelphia, PA: NFAIS (National Federation of Abstracting and Information Services). ISBN 0-942308-24-7. 156 pp, spiral bound. Price not known. Learned Publishing, 3(3):194, July 1990. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Feeney:1990:BRI

[221] Mary Feeney. Book review: Introduction to desktop publishing, by Hewson, David. John Taylor Book Ventures, 1989, 112 pp, hardcover or spiral bounds, ISBN 1-871224-020. Learned Publishing, 3(3):195, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:1990:BRI

[222] Kurt Paulus. Book review: International Council of Scientific Unions Yearbook 1990. Paris: ICSU. 1990. 402 pp. ISSN 0074-4387 ISBN 0-930357-21-3. Price n.a. Learned Publishing, 3(3):195–196, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:CEc

[223] Anonymous. Corning events. Learned Publishing, 3(3):198–200, July 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:ESW

[224] Anonymous. Editorial: Spread the word! Learned Publishing, 3(4):201, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1990:NLJ

[225] John D. St. Aubyn. Notes on learned journal copyright. Learned Publishing, 3(4):203–212, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Murphy:1990:CRP

[226] D. P. Murphy. CD-ROM publishing — is it for you? Learned Publishing, 3(4):213–216, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ritchie:1990:SLS

[227] Anthony Ritchie. The Scottish learned societies and their problems. Learned Publishing, 3(4):217–223, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1990:RDL

[228] Edwin Shelock. Recent developments in learned society publishing. Learned Publishing, 3(4):225–230, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Blake:1990:MJI

[229] Doreen Blake. Medical journal indexing. Learned Publishing, 3(4):231– 236, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:PLM

[230] Anonymous. Publishers and the law: 22 May 1990, London; ALPSP AGM. Learned Publishing, 3(4):237–240, October 1990. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1990:PIB

[231] Michele Benjamin. Professional information and book fair: 28–30 May 1990, La Defense, Paris, CNIT. Learned Publishing, 3(4):240, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1990:FFM

[232] Edwin Shelock. 1990 facing forward: meeting challenges old and new: 5–8 June 1990, San Francisco; 12th annual meeting of the Society for Scholarly Publishing. Learned Publishing, 3(4):241, October 1990. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:PB

[233] Anonymous. People and bodies. Learned Publishing, 3(4):242–243, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:RAc

[234] Anonymous. Reading around. Learned Publishing, 3(4):244–246, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:BCNd

[235] Anonymous. British Council news. Learned Publishing, 3(4):247–249, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:IOd

[236] Anonymous. Information organization. Learned Publishing, 3(4):249, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1990:LPJ

[237] Michele Benjamin. Logos — the Professional Journal for the Book World. Quarterly, ISSN 0957-9566. Individual, £30.00; Institutional, £45.00. Whurr Publishers Ltd, 1 Compton Terrace, London NI 2UN. Learned Publishing, 3(4):250, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1990:BRO

[238] Hazel Bell. Book review: The Oxford writers’ dictionary, compiled by Allen, R. E., Oxford: Oxford University Press, 1990. xvi, 448 pp. 20 cm. ISBN 0-19-282669-7 (pbk): £5.95. (Oxford reference). Learned Publishing, 3(4):250–251, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jones:1990:BRT

[239] Kevin Jones. Book review: Text, ConText, and Hyper Text: writing with and for the computer, edited by Edward Barrett. Cambridge, MA, and London: MIT Press, 1988. xxv, 368 pp. 24 cm. Index. ISBN 0-262-02275-3 (cased): £33.75. (MIT Press series in information systems.). Learned Publishing, 3(4):251–252, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1990:BRE

[240] Michele Benjamin. Book review: 1990 Encyclopedia of Information Systems and Services. Edited by Amy Lucas. Detroit, MI: Gale Research Publications Inc. 10th ed. Volume 1 Descriptive listings; volume 2 Indexes. Total pages 2200. $420. Learned Publishing, 3(4):252–253, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1990:CEd

[241] Anonymous. Coming events. Learned Publishing, 3(4):254–256, October 1990. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Huth:1991:EPL

[242] Edward J. Huth. The editorial policy of a learned society. Learned Publishing, 4(1):1–8, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Skaer:1991:UCD

[243] R. J. Skaer. The use of computer disks and optical character readers in scientific publishing. Learned Publishing, 4(1):9–15, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zeng:1991:AIS

[244] Lei Zeng. Abstracting and indexing services in China. Learned Publishing, 4(1):17–28, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hewitt:1991:SSP

[245] David Hewitt. The small society as publisher. Learned Publishing, 4 (1):29–34, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:PB

[246] Anonymous. People and bodies. Learned Publishing, 4(1):34, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1991:OSC

[247] Michele Benjamin. Occupational standards and competences in learned journal publishing: 27 June, 1990, London; Publishers Association.

Learned Publishing, 4(1):35–36, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1991:MJE

[248] Michele Benjamin. Marketing journals in Eastern Europe: 11 July, 1990, London; ALPSP. Learned Publishing, 4(1):36–38, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:TKO

[249] Anonymous. Tools for knowledge organization and the human interface: 14–17 August 1990, Darmstadt; International Society for Knowledge Organization. Learned Publishing, 4(1):38–39, January 1991. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Beare:1991:PBJ

[250] Geraldine Beare. Publishing books and journals for profit: 12 October 1990, Leatherhead; ALPSP/Pira. Learned Publishing, 4(1):39–41, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Phelan:1991:AFO

[251] Peter Phelan. Abstracts forum: 16 October, 1990, London; Publishers Licensing Society Learned Publishing Vol. 4 No. 1 January 1991. Learned Publishing, 4(1):41–43, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1991:PLD

[252] John St. Aubyn. Photocopying licensing up to date. Learned Publishing, 4(1):43–44, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1991:TES

[253] H.K.B. Typographical excellence in serials: Third annual Charlesworth Award. Learned Publishing, 4(1):45–48, January 1991. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:LAM

[254] Anonymous. The Library Association Medals. Learned Publishing, 4(1): 49, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:IOa

[255] Anonymous. Information organization. Learned Publishing, 4(1):50–53, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:RAa

[256] Anonymous. Reading around. Learned Publishing, 4(1):54–55, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:BCN

[257] Anonymous. British Council news. Learned Publishing, 4(1):55–56, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:1991:BRB

[258] Pippa Smart. Book review: Book Production, by Peacock, John. London: Blueprint Publishing in association with The Publishers Association. Royal 8vo. 494 pp. £29.95 (cased) ISBN 0-968905-33-6. Learned Publishing, 4(1):57–58, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

OConnor:1991:BRH

[259] Maeve O’Connor. Book review: Hyphenation, by McIntosh, Ronald. Bradford, UK: Computer Hyphenation Ltd (Bradford Science Park, 1 Campus Road, Bradford BD7 1HR). vi + 96 pp. £6.95 (£4.95 paperback). ISBN 1-872757-00-6 (and 01-4). Learned Publishing, 4(1):58–59, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dauker:1991:BRB

[260] Hermi Dáuker. Book review: The Blueprint dictionary of printing and publishing. By Peacock, John & Barnard, Michael. London: Blueprint Publishing Ltd. 1990. 280 pp. Cased, £19.95. ISBN 0-948905-47-6. Learned Publishing, 4(1):59, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1991:BFB

[261] A. B. Britton. Barriers to the flow of books: a statement by ELP, the Working Group of European Librarians and Publishers. 8 pp. Amsterdam, 1990. Learned Publishing, 4(1):59–61, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:PN

[262] Anonymous. Publications noted. Learned Publishing, 4(1):61–62, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:CEa

[263] Anonymous. Coming events. Learned Publishing, 4(1):63–64, January 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Charlesworth:1991:JD

[264] Neil Charlesworth and David Penfold. Journal design. Learned Publishing, 4(2):65–75, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pearce:1991:SIV

[265] Anthony Pearce. Scientific information: the views of a learned society and publisher. Learned Publishing, 4(2):77–81, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oakeshott:1991:BRE

[266] Priscilla Oakeshott. Book review: Effects of US adherence to the Berne Convention. Baumgarten, Jon A. and Meyer, Christopher A.. Published jointly by the International Publishers Association, Geneva and International Group of Scientific, Technical and Medical Publishers, Amsterdam, 1989, for their members. iv + 28 pp. A5, paper cover. No ISBN or price shown. (IPA/STM Publishing Series 2). Learned Publishing, 4(2):121– 122, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Coles:1991:SIW

[267] Bryan Coles. Scientific information: what researchers need. Learned Publishing, 4(2):82–84, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bradley:1991:EAS

[268] Cive Bradley. Electronic advances in scientific information: the commercial angle. Learned Publishing, 4(2):85–87, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hall:1991:SDC

[269] D. H. Hall. Subscriptions by Deed of Covenant: Is this charity? Learned Publishing, 4(2):88–89, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Page:1991:TJP

[270] Gillian Page. Training in journal publishing. Learned Publishing, 4(2): 90–91, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Briggs:1991:RPL

[271] C. Stephen Briggs. The refereeing of papers for learned journals. Learned Publishing, 4(2):93–95, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1991:CUA

[272] Michele Benjamin. Come in Userid 0231@UK.AC.uni.ex: your time is up. Learned Publishing, 4(2):98–101, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thompson:1991:BBT

[273] Michael Thompson. The Brazilian book trade. Learned Publishing, 4 (2):102–104, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1991:CCa

[274] John St. Aubyn. Copyright corner. Learned Publishing, 4(2):104–105, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1991:CPS

[275] Hazel K. Bell. Copy into print: the shape of the next twenty years: 17 October 1990, Cambridge; Society of Freelance Editors and Proofreaders’ first annual conference. Learned Publishing, 4(2):106, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1991:SIN

[276] Michele Benjamin. Selling to industry: 8 November 1990; Institute of Mechanical Engineers, London; ALPSP. Learned Publishing, 4(2):107– 108, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1991:NET

[277] H.K.B. NVQs: Editorial training for freelancers and publishers: 15 Nov. 1990, London; Society of Freelance Editors and Proofreaders. Learned Publishing, 4(2):108–109, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1991:FDL

[278] H.K.B. Future directions for LIS research: 14 November and 6 December 1990, London; British Library Research and Development Department. Learned Publishing, 4(2):110, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1991:WMG

[279] Michele Benjamin. What makes a good journal?: 11 December 1990, Ciba Foundation, London; ALPSP. Learned Publishing, 4(2):111, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1991:SIA

[280] H.K.B. Scientific information — access and opportunities: 23 January 1991, London; Foundation for Science and Technology/ALPSP. Learned Publishing, 4(2):112–114, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:NCT

[281] Anonymous. New Canadian tax. Learned Publishing, 4(2):115, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:MJK

[282] Anonymous. Mission to Japan and Korea. Learned Publishing, 4(2):115– 117, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:BCB

[283] Anonymous. British Council book exhibitions. Learned Publishing, 4(2): 117, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:IOb

[284] Anonymous. Information organization. Learned Publishing, 4(2):118– 119, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Burton:1991:CAB

[285] Paul F. Burton. Current awareness bulletin: a review of information management literature. London: Aslib. ISSN 0265-9271. Annual subscription £50.00 (Aslib members), £55.00 (non-members). Learned Publishing, 4(2):120, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shortridge:1991:BRT

[286] John Shortridge. Book review: Trends in database design and customer services. Edited by Wendy Schipper and Betty Unruh. NFAIS, 1990. Philadelphia US$30.50. ISBN 0-94-2308-27-1. Learned Publishing, 4(2): 121, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oakeshott:1991:BRL

[287] Priscilla Oakeshott. Book review: The law of copyright and rights in performances: The British Copyright Council’s Guide. Denis de Freitas. British Copyright Council, London. 76 pp. A5, paper cover, ISBN 0-901737-05-4. No price shown. Learned Publishing, 4(2):122, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oakeshott:1991:BRG

[288] Priscilla Oakeshott. Book review: Global copyright issues in the secondary information industry. Levine, Arthur. Philadelphia: National Federation of Abstracting and Information Services, xii + 96 pp. 214 × 273 mm, paperback, ISBN 0-942308-28-X. No price shown. (NFAIS Report Series 4). Learned Publishing, 4(2):122–123, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1991:BRB

[289] John St. Aubyn. Book review: Blackstone’s guide to the Copyright, Designs & Patents Act 1988. By Dworkin, Gerald & Taylor, Richard D, Blackstone Press, London, 1989 x, 454 pp. 22 cm. Paperback ISBN 1-85431-023-2 £18.95. Learned Publishing, 4(2):124, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

B:1991:BRB

[290] M. B. Book review: BS 5605: 1990 Recommendations for citing and referencing published material. £21.00. BSI Linford Wood, Milton Keynes MK14 6LE. Learned Publishing, 4(2):124–125, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

B:1991:BPT

[291] M. B. BS4821: 1990 Presentation of theses and dissertations. Pp. 28. £35.40. The British Standards Institute, Milton Keynes. Learned Publishing, 4(2):125, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1991:BRR

[292] H.K.B. Book review: Reducing publication costs, by Crummy, Philip. Birmingham: The Institute of Field Archaeologists’ Technical Paper No. 6, 1990. 16 pp. Learned Publishing, 4(2):125, April 1991. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:CEb

[293] Anonymous. Coming events. Learned Publishing, 4(2):127–128, April 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dorner:1991:AIT

[294] Jane Dorner. Authors and information technology: new challenges in publishing. Learned Publishing, 4(3):129–138, July 1991. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott:1991:WIN

[295] Elspeth Scott. What industry needs: the industrial user’s viewpoint. Learned Publishing, 4(3):139–144, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kogan:1991:AUM

[296] Philip Kogan. Activating unconventional markets. Learned Publishing, 4(3):145–148, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Swinnerton-Dyer:1991:INA

[297] Peter Swinnerton-Dyer. The information needs of academics and researchers. Learned Publishing, 4(3):149–151, July 1991. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:IME

[298] Anonymous. The Institution of Mechanical Engineers (MEP Ltd). Learned Publishing, 4(3):153–155, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1991:MBE

[299] Hazel K. Bell. Marketing books in Europe: 21 February 1991, London; ALPSP. Learned Publishing, 4(3):157–158, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1991:FCG

[300] H.K.B. Forum on Canadian GST: 12 March 1991, London; PA. Learned Publishing, 4(3):159–162, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dorner:1991:CWP

[301] Jane Dorner. Computers and the writing process: 22–23 March 1991, Brighton; Computers and Writing Association. Learned Publishing, 4(3): 162–164, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1991:PWN

[302] Hazel K. Bell. Publishers: who needs them?: 7th International Learned Journals Seminar, ALPSP/PA. Learned Publishing, 4(3):164–167, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thomson:1991:PP

[303] Ned Thomson. Publications on paper. Learned Publishing, 4(3):168– 169, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:WAT

[304] Anonymous. Words against technology. Learned Publishing, 4(3):170– 171, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1991:CCb

[305] John St. Aubyn. Copyright corner. Learned Publishing, 4(3):172–173, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stevenson:1991:AE

[306] Val Stevenson. Assessing editors. Learned Publishing, 4(3):173, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:IOc

[307] Anonymous. Information organization. Learned Publishing, 4(3):174– 177, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:OAa

[308] Anonymous. Overseas affairs. Learned Publishing, 4(3):178–180, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:RAb

[309] Anonymous. Reading around. Learned Publishing, 4(3):181–182, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1991:BRI

[310] David Lee. Book review: Information UK 2000. Edited by John Martyn, Peter Vickers and Mary Feeney. ISBN 0-86291-620-8. London: British Library Research; Bowker-Saur, 1990. Learned Publishing, 4(3):183–184, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:1991:BRK

[311] Kurt Paulus. Book review: Key note report — a market sector overview: book publishing. Ninth edition. Hampton, Middx: Key Note Publications Ltd. 1992. 48pp. Paperback. £169. ISBN 1-85056-859-6. Learned Publishing, 4(3):184–185, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1991:BRE

[312] David Lee. Book review: The electronic campus, an information strategy: proceedings of a conference held on 28-30 October 1988 at Banbury. Edited by Lynne J. Brindley. (Library and Information Research Report 73). ISBN 0-7123-3187-5. London: The British Library, 1989. Learned Publishing, 4(3):185–186, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Seal:1991:BRE

[313] Alan Seal. Book review: The emergence of the electronic book. Feldman, Tony. London: British National Bibliography Research Fund with the Publishers Association, 1990. vii, 74 pp. 30 cm. ISBN 0-7123-3227-8 (pbk): £30.00. (BNBRF report 46.). Learned Publishing, 4(3):186–187, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1991:BRS

[314] A. B. Britton. Book review: Secondary information service statistics. Edited by Wendy Schipper and Betty Unruh. Philadelphia: National Federation of Abstracting and Information Services, 1990. (NFAIS report series 2). 130pp. Paperback. ISBN 0-942308-26-3, $185.50. Learned Publishing, 4(3):187–188, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:CEc

[315] Anonymous. Coming events. Learned Publishing, 4(3):190–192, July 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dunn:1991:PCL

[316] Ronald G. Dunn and Peter T. Shepherd. Publishers: the critical link. Learned Publishing, 4(4):193–201, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Russon:1991:SIW

[317] David Russon. Scientific information: What libraries should supply. Learned Publishing, 4(4):203–205, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1991:LDE

[318] Hazel K. Bell. Legal deposit in Euston Street. Learned Publishing, 4 (4):207–214, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tolba:1991:SWI

[319] Hany Tolba. Spreading the word: an international view. Learned Publishing, 4(4):215–218, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

vanMuyen:1991:MBN

[320] Ad van Muyen. Marketing books in The Netherlands. Learned Publishing, 4(4):219–225, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:SJE

[321] Anonymous. The song of a journal editor. Learned Publishing, 4(4):226, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1991:PCO

[322] Bernard Donovan. Publishers: commercial and otherwise: 21 May 1991, London; ALPSP AGM. Learned Publishing, 4(4):227, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1991:EME

[323] Hazel K. Bell and Cherry Lavell. Electronics and manuscript editing: boon or bane?: 21 May 1991, London; ALPSP/EASE. Learned Publishing, 4(4):227–232, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1991:PDJ

[324] Hazel K. Bell. Printing by design: 19 June 1991, London; PA Small Publishers Group /BL Centre for the Book. Learned Publishing, 4(4): 232–235, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1991:CCc

[325] John St. Aubyn. Copyright corner. Learned Publishing, 4(4):236–237, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:OAb

[326] Anonymous. Overseas affairs. Learned Publishing, 4(4):238–241, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:BCP

[327] Anonymous. British Council promotion. Learned Publishing, 4(4):241– 242, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:IOd

[328] Anonymous. Information organization. Learned Publishing, 4(4):243– 244, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:RAc

[329] Anonymous. Reading around. Learned Publishing, 4(4):245–247, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

OConnor:1991:BRB

[330] Maeve O’Connor. Book reviews: Basic editing: a practical course. Harris, Nicola. London: Book House Training Centre. Two volumes. The text, 208 p, £20, ISBN 0-907706-02-9. The exercises, 160 p, £5, ISBN 0-907706-03-7. Obtainable from Spa Books, PO Box 47, Stevenage, Herts, SG2 8UH, UK. Learned Publishing, 4(4):248–249, October 1991. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1991:BRM

[331] A. B. Britton. Book review: Magazine and journal production. 2nd ed. Barnard, Michael. London: Blueprint Publishing, 1990. xiv, 378. ISBN 0-948905-31-X. £29.95. Learned Publishing, 4(4):249–250, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dunkley:1991:BRW

[332] Bernard Dunkley. Book review: Writing non-fiction. Lyons, Stanley. Stanley Lyons, 4 The Rise, Shipton Oliffe, Cheltenham, Glos. GL54 4JQ ISBN: 1-873521-00-6 180 × 115 mm ©1991 Stanley Lyons x + 193 pages, casebound £13.95. Learned Publishing, 4(4):250, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:U

[333] Anonymous. These I have used. Learned Publishing, 4(4):252–253, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:CEd

[334] Anonymous. Coming events. Learned Publishing, 4(4):254–256, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1991:NC

[335] Anonymous. Notes for contributors. Learned Publishing, 4(4):256, October 1991. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Friend:1992:LVF

[336] F. J. Friend. A Librarian’s view of the future of scholarly communication. Learned Publishing, 5(1):1–6, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Manley:1992:ES

[337] Don Manley. The editor’s style. Learned Publishing, 5(1):7–12, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wellisch:1992:IP

[338] Hans H. Wellisch. The indexing of periodicals. Learned Publishing, 5 (1):13–18, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1992:ABC

[339] H.K.B. Audit Bureau of Circulations. Learned Publishing, 5(1):19–20, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Minderhaud-Jones:1992:SS

[340] Marilyn Minderhaud-Jones and Sue Martin. Standing up for standards. Learned Publishing, 5(1):20–21, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1992:ETR

[341] H.K.B. Editorial training and recruitment. Learned Publishing, 5(1):21– 24, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1992:ENP

[342] David Lee. Electronic networks and publishing. Learned Publishing, 5 (1):25–27, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:CGA

[343] Anonymous. The Charlesworth Group Award for Typographical Excellence in Journal and Serial Production. Learned Publishing, 5(1):31–32, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:LAA

[344] Anonymous. Library Association Awards. Learned Publishing, 5(1):32– 33, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1992:LJP

[345] John St. Aubyn. Learned journal publishing and fees to authors. Learned Publishing, 5(1):34–35, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:D

[346] Anonymous. Distribution. Learned Publishing, 5(1):36, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:BCNa

[347] Anonymous. British Council news. Learned Publishing, 5(1):37–38, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:OAa

[348] Anonymous. Overseas affairs. Learned Publishing, 5(1):38–39, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lawrence:1992:PB

[349] T. E. Lawrence. People and bodies. Learned Publishing, 5(1):39–40, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:RAa

[350] Anonymous. Reading around. Learned Publishing, 5(1):41–42, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:IOa

[351] Anonymous. Information organization. Learned Publishing, 5(1):43–44, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Burton:1992:IDI

[352] Paul F. Burton. Information distribution issues for the 90s. Edited by Betty Unruh and Wendy Schipper. Philadelphia: National Federation of Abstracting and Information Services, 1991. x, 156 pp. ISBN 0-942308-30-1. Learned Publishing, 5(1):45–46, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1992:BRN

[353] A. B. Britton. Book review: National and international information policies. Edited by Wendy Schipper and Ann Marie Cunningham. Paperback, xi, 161 pp. Philadelphia: National Federation of Abstracting and Information Services, 1991. ISBN 0-942308-31-X. $100.00. Learned Publishing, 5(1):46–47, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jones:1992:LES

[354] Kevin P. Jones. Libraries and expert systems: proceedings of a conference and workshop held at Charles Sturt University, Riv-erina, Australia, July 1990. McDonald, Craig and Weckert, John. London: Taylor Graham, 1991. [iv], 232 pp. ISBN 0-947568-51 (pbk). Learned Publishing, 5(1):47– 48, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1992:BRA

[355] A. B. Britton. Book review: Authors and information technology: new challenges in publishing. Dorner, Jane. (British National Bibliography Research Fund Report 52). viii, 86 pp. Boston Spa, 1991. ISBN 0-7123-3257-X. £15.00. Learned Publishing, 5(1):48–49, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lavell:1992:OHC

[356] Cherry Lavell. One and a half cheers for Dorset. Learned Publishing, 5(1):49–50, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:CEa

[357] Anonymous. Coming events. Learned Publishing, 5(1):53–54, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:LPN

[358] Anonymous. Learned Publishing: notes for contributors. Learned Publishing, 5(1):55, January 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Penfold:1992:JPR

[359] David W. Penfold. Journal production: the right medium for the task. Learned Publishing, 5(2):57–63, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Flower:1992:FES

[360] Beverlie Flower. Finding editorial staff. Learned Publishing, 5(2):65– 71, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Whitcombe:1992:OJC

[361] Norma Whitcombe. Opting out and jobbing out: can the freelance replace the staff member? Learned Publishing, 5(2):73–76, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1992:AET

[362] Hazel K. Bell. Academia Europaea: trans-Continental, multidisciplinary. Learned Publishing, 5(2):77–80, April 1992. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Robertson:1992:WCS

[363] Michael Robertson. The Worshipful Company of Stationers and its archives. Learned Publishing, 5(2):81–85, April 1992. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Burrough:1992:EEI

[364] Joy Burrough and Jackie Senior. European editing for international scientific communication. Learned Publishing, 5(2):86–90, April 1992. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thorn:1992:NIM

[365] Sun Thorn. New information media: impacts and applications in the 1990s. Learned Publishing, 5(2):90–92, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1992:MMM

[366] H.K.B. Making the most of market research. Learned Publishing, 5(2): 93–94, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1992:RPJ

[367] Hazel K. Bell. The role of the printed journal article in the 1990s. Learned Publishing, 5(2):94–97, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

B:1992:RR

[368] H. K. B. and B. T. D. Recollections and ruminations. Learned Publishing, 5(2):97–98, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bhel:1992:PLT

[369] Hazhl K. Bhel. Publishers and librarians: the technology link. Learned Publishing, 5(2):99–102, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1992:CSP

[370] John St. Aubyn. Computer-software protection. Learned Publishing, 5 (2):103–104, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

P:1992:PB

[371] K. P. People and bodies. Learned Publishing, 5(2):105–107, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:BCNb

[372] Anonymous. British Council news. Learned Publishing, 5(2):108–109, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sayers:1992:OA

[373] Dorothy Sayers. Overseas affairs. Learned Publishing, 5(2):110–112, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Turpie:1992:IO

[374] Geraldine Turpie. Information organization. Learned Publishing, 5(2): 113, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thomson-Shore:1992:RA

[375] Thomson-Shore. Reading around. Learned Publishing, 5(2):114–115, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1992:LNO

[376] David Lee. Libraries, networks and OSI. Dempsey, Lorcan. Bath: UK Office for Library Networking, 1991, £28.00. [14], 232 pp. ISBN 0-9516856-0-0. Learned Publishing, 5(2):116–117, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Vickers:1992:IZH

[377] John A. Vickers. Indexing from A to Z, The H. W. Wilson Company. New York. Wellisch, Hans H., 1991, pp. xxvi, 461. ISBN 0-8242-0807-2 £31.00 (cased) Available in the UK from Thompson Henry Ltd, London Rd, Sunningdale, Berks SL5 0EP. Learned Publishing, 5(2):117–118, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:PN

[378] Anonymous. Publications noted. Learned Publishing, 5(2):118–119, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:CEb

[379] Anonymous. Coming events. Learned Publishing, 5(2):119–120, April 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1992:CNW

[380] Bernard Donovan. The clash that never was. Learned Publishing, 5(3): 121–122, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brown:1992:TRC

[381] D. J. Brown. Technologies of relevance to current-awareness and document-delivery systems. Learned Publishing, 5(3):123–135, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Page:1992:NEJ

[382] Gillian Page. Networking and the economics of journal publishing. Learned Publishing, 5(3):137–144, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott:1992:MSI

[383] Elspeth J. Scott. Meeting the scientist’s information needs. Learned Publishing, 5(3):145–152, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bodinham:1992:SJS

[384] Michael Bodinham and Robert Campbel. The survival of the journal: a simple spreadsheet model for predicting the outcome. Learned Publishing, 5(3):153–159, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:BBP

[385] Anonymous. Breaking the barriers; publishing for one world: New Delhi, 27 Jan.–2 Feb. 1992, International Publishers Association. Learned Publishing, 5(3):160, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

B:1992:WP

[386] H. K. B. Working with printers. Learned Publishing, 5(3):161–162, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1992:UAC

[387] H.K.B. The uses and abuses of customer databases. Learned Publishing, 5(3):162–163, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1992:BE

[388] Hazel K. Bell. Business in Europe. Learned Publishing, 5(3):163–164, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1992:JNF

[389] David Lee. Journals: the next five years. Learned Publishing, 5(3):164– 167, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jones:1992:TMI

[390] Kevin Jones. Text management and indexing. Learned Publishing, 5(3): 168–169, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dorner:1992:CWV

[391] Jane Dorner. Computers and writing V. Learned Publishing, 5(3):169– 171, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1992:DP

[392] John St. Aubyn. Database protection. Learned Publishing, 5(3):172– 173, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:PBa

[393] Anonymous. People and bodies. Learned Publishing, 5(3):174–175, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:OAb

[394] Anonymous. Overseas affairs. Learned Publishing, 5(3):176, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:IOb

[395] Anonymous. Information organization. Learned Publishing, 5(3):177– 179, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:RAb

[396] Anonymous. Reading around. Learned Publishing, 5(3):180–183, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Green:1992:BRR

[397] P. R. Green. Book review: Research policy in librarianship and information science. Edited by Colin Harris. Taylor Graham, London, New York, 1991 Pp 140. £24.00 ISBN 0-947568-49-2. Learned Publishing, 5(3):184– 185, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Leach:1992:BRW

[398] Chris Leach. Book review: Writing successfully in science. By O’Connor, Maeve. Routledge, London, 1991 Pp xi + 229. Paper, £8.95. ISBN 0044458061. Learned Publishing, 5(3):185–186, July 1992. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1992:BRC

[399] Hazel K. Bell. Book review: Copy into print: the shape of the next twenty years. Proceedings of the first SFEP conference. Society of Freelance Editors and Proofreaders. 1991. Pp iv, 60. 21 cm. Paperback, £4.00 (f50p p&p from SFEP, 10 Grange Close, Godalming, GU7 1XT). ISBN 0-9517527-0-7. Learned Publishing, 5(3):186, July 1992. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:1992:BRC

[400] Kurt Paulus. Book review: Customer services and user training. Ardito, Stephanie C.. Philadelphia, PA The National Federation of Abstracting and Information Services. 1991 Pp. 125. Paperback, US $50.00. ISBN: 0-942398-33-6. Learned Publishing, 5(3):186–187, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Muller:1992:BRM

[401] Gemase Müller. Book review: Meeting the neeh of library trade exhibitors. A handbook for exhibition organisers. 1991. 12 pp. Prepared by The Association of Subscription Agents. Available free of charge from The Secretary, Association of Subscription Agents, Thames Gardens, Charlbury. Oxon OX7 3QH. Learned Publishing, 5(3):187, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:CEc

[402] Anonymous. Coming events. Learned Publishing, 5(3):189–190, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:BCP

[403] Anonymous. British Council promotion. Learned Publishing, 5(3):191– 192, July 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stevenson:1992:JCL

[404] M. B. Stevenson. Journals: the choice for llbrarians through the next five years. Learned Publishing, 5(4):193–198, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Read:1992:DMP

[405] Tony Read. Developing markets: problems or possibilities? Learned Publishing, 5(4):199–204, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:1992:CRS

[406] John Cox. The changing role of the subscription agent. Learned Publishing, 5(4):205–209, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thorn:1992:UHD

[407] Sue Thorn. Using in-house databases for marketing. Learned Publishing, 5(4):211–219, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hadley:1992:EUC

[408] Colin P. Hadley. Electronic use of copyright works: electro-storage, electro-copying and electro-manipulation. Learned Publishing, 5(4):220– 224, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1992:BAP

[409] Hazel K. Bell and Bertrand Russell. The British Academy: promoting the humanities. Learned Publishing, 5(4):225–230, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1992:VVA

[410] Bernard Donovan. VAAP visit: 14 April 1992, London; ALPSP. Learned Publishing, 5(4):231–232, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1992:LMB

[411] Hazel K. Bell. London Medical Book Fair. Learned Publishing, 5(4): 232–233, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1992:LPS

[412] David Lee. Librarians and publishers: serials and the information chain. Learned Publishing, 5(4):234–236, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1992:TPC

[413] John St. Aubyn. Term of protection of copyright works. Learned Publishing, 5(4):236–238, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:PBb

[414] Anonymous. People and bodies. Learned Publishing, 5(4):239–241, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:OAc

[415] Anonymous. Overseas affairs. Learned Publishing, 5(4):242, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:BCNc

[416] Anonymous. British Council news. Learned Publishing, 5(4):243–244, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:IOc

[417] Anonymous. Information organization. Learned Publishing, 5(4):245– 247, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:V

[418] Anonymous. VAT. Learned Publishing, 5(4):247–248, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1992:BRJ

[419] Michele Benjamin. Book review: Journal production: Principles and practice. Campbell, Robert. London, Publishers Association/British Printing Industries Federation, Pp. 90. 1992. Cased. £7.50, ALPSP members; £12.50 otherwise. Learned Publishing, 5(4):249, October 1992. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Symondson:1992:BRH

[420] Brian Symondson. Book review: How to make your own CD-ROM. By Butler, Matilda PhD and Paisley, William PhD. Pp. 40. Knowledge Access International, Inc., 2685 Marine Way, Suite 1305, Mountain View, CA 94043, USA. Learned Publishing, 5(4):249–250, October 1992. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dauker:1992:BRM

[421] Hermi Däuker. Book review: Multilingual dictionary of printing and publishing terms. Compiled by Peacock, John. London: Blueprint. 1991. Pp. 252. Cased, £50.00. ISBN 0-948905-35-2. Learned Publishing, 5(4): 250, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:CEd

[422] Anonymous. Coming events. Learned Publishing, 5(4):253–254, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1992:VI

[423] Anonymous. Volume Index. Learned Publishing, 5(4):255–256, October 1992. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:JG

[424] Hazel K. Bell. Journal growth. Learned Publishing, 6(1):2–3, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1993:FS

[425] Bernard Donovan. A fresh start. Learned Publishing, 6(1):4–6, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Alden:1993:HCP

[426] William Alden. How can publishers help printers? Learned Publishing, 6(1):7–9, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Arrowsmith-Brown:1993:HSY

[427] Victoria Arrowsmith-Brown. How to specify your printing needs. Learned Publishing, 6(1):11–13, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott:1993:HIE

[428] Ian Scott. How to interpret an estimate. Learned Publishing, 6(1):15–20, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hunter:1993:DDT

[429] Karen Hunter. Document delivery: threat or opportunity? Learned Publishing, 6(1):21–24, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prastalo:1993:EJB

[430] Tatjana Praštalo and Neda Ćukac. Early journals in Bosnia and Herzegovina. Learned Publishing, 6(1):25–28, January 1993. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Corrigan:1993:TDA

[431] Adrian Corrigan. Training the DIY approach. Learned Publishing, 6 (1):29–33, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clarke:1993:PHS

[432] Gillian Clarke. Proofreading: hobby or skill? Learned Publishing, 6 (1):35–37, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:PDP

[433] Anonymous. Personal data privacy in the 1990s: too late to close the stable door? Learned Publishing, 6(1):38, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lavell:1993:TIN

[434] Cherry Lavell. TR 92: Information now. Learned Publishing, 6(1):38–39, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Whitford:1993:CLA

[435] Sir John Whitford. The Copyright Licensing Agency’s 10th anniversary reception. Learned Publishing, 6(1):42–43, January 1993. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:SNA

[436] Hazel K. Bell. Selling to North America. Learned Publishing, 6(1):44–45, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1993:GGF

[437] Michele Benjamin. Is the grass greener for freelancers? Learned Publishing, 6(1):45–46, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:CI

[438] Hazel K. Bell. Copyright and the indexer. Learned Publishing, 6(1):46– 47, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1993:RWR

[439] Bernard Donovan. Romanian writers raring for rights. Learned Publishing, 6(1):48–49, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:Aa

[440] Anonymous. Awards. Learned Publishing, 6(1):50–52, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1993:CCa

[441] John St. Aubyn. Copyright corner. Learned Publishing, 6(1):53–55, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:PBa

[442] Anonymous. People & bodies. Learned Publishing, 6(1):55–58, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:OAa

[443] Anonymous. Overseas affairs. Learned Publishing, 6(1):59–60, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:BCNa

[444] Anonymous. British Council news. Learned Publishing, 6(1):61–62, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:IMa

[445] Anonymous. Information management. Learned Publishing, 6(1):63–67, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:RAa

[446] Anonymous. Reading around. Learned Publishing, 6(1):68–71, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

OConnor:1993:BRW

[447] Maeve O’Connor. Book review: Writing on disk: an A-Z handbook of terms, tips and techniques for authors and publishers by Dorner, Jane, Herts, Hatfield: John Taylor Book Ventures (distributed by Spa Books 1992, PO Box 47, Stevenage, Herts 2SG2 8UH), 240 pp. £18.95 hardback £15.95 spiralbound + £1.50 post and packing. ISBN 1-871224-11-X/12-8. Learned Publishing, 6(1):73, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1993:BRC

[448] A. B. Britton. Book review: Conceptions of library and information science: proceedings of the international conference for the celebration of the 20th anniversary of the Department of Information Studies, University of Tampere, Finland, 26-28 August 1991, Edited by Vakkari Pertti and Blaise Cronin, London: Taylor Graham. 1992, Paperback. 314 pp., ISBN 0-947568-52-2. Learned Publishing, 6(1):74, January 1993. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:BRB

[449] Hazel K. Bell. Book review: Business in Europe: the promotion of books and reading in the European Community. Edited transcript of the seminar held at the London International Book Fair, 23 March 1992, Society of Freelance Editors and Proofreaders, 21 cm. 20pp (pvc cover). £5 (inc. p&p) from Sally Crowther, SFEP, PO Box 4YP, London WIA 4YP, ISBN 0-9517527-3-1. Learned Publishing, 6(1):75–77, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prior:1993:C

[450] Chair Albert Prior. Correspondence. Learned Publishing, 6(1):77, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:CEa

[451] Anonymous. Coming events. Learned Publishing, 6(1):78–79, January 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1993:LPW

[452] Bernard Donovan. Librarians and publishers: not working together. Learned Publishing, 6(2):3–6, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williams:1993:HCU

[453] Ivor A. Williams. How chemists use the literature. Learned Publishing, 6(2):7–14, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cornish:1993:CPM

[454] Graham P. Cornish. Copyright protection for materials in electronic formats: the CITED solution. Learned Publishing, 6(2):15–18, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Boden:1993:MAS

[455] Edward Boden. Managing advertising sales. Learned Publishing, 6(2): 19–21, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cameron:1993:JMN

[456] Jamie Cameron. Journal marketing in North America. Learned Publishing, 6(2):23–27, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Owen:1993:FLA

[457] Lynette Owen. The future for licensing: an academic publisher’s viewpoint. Learned Publishing, 6(2):29–35, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:TFS

[458] Hazel K. Bell. Taylor & Francis: Serving science, family-style. Learned Publishing, 6(2):37–39, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williams:1993:EAM

[459] Bernard J. S. Williams. European academies meet. Learned Publishing, 6 (2):41, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williams:1993:CRJ

[460] Bernard J. S. Williams. CD-ROMs and journal publishing: practical considerations. Learned Publishing, 6(2):41–44, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:DAR

[461] Hazel K. Bell. The dissemination of academic and research information through the learned journal and present and future alternatives. Learned Publishing, 6(2):44–45, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Symondson:1993:CRR

[462] Brian Symondson. CD-ROMland revisited. Learned Publishing, 6(2):45– 49, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:HSM

[463] Hazel K. Bell. How to save money on paper: purchase, usage and despatch. Learned Publishing, 6(2):50–51, April 1993. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:Ab

[464] Anonymous. Awards. Learned Publishing, 6(2):52–56, April 1993. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:CC

[465] Anonymous. Copyright corner. Learned Publishing, 6(2):57–58, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:PBb

[466] Anonymous. People and bodies. Learned Publishing, 6(2):58–61, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:OAb

[467] Anonymous. Overseas affairs. Learned Publishing, 6(2):62, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:BCNb

[468] Anonymous. British Council news. Learned Publishing, 6(2):63–64, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:IMb

[469] Anonymous. Information management. Learned Publishing, 6(2):65– 68, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Page:1993:C

[470] Gillian Page. Correspondence. Learned Publishing, 6(2):68, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:RAb

[471] Anonymous. Reading around. Learned Publishing, 6(2):69–70, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

OConnor:1993:BRH

[472] Maeve O’Connor. Book review: How to run a paper mill: writing technical papers and getting them published by John, Woodwark 1992. Cartoons by Carol, Wade Winchester: Information Geometers (47 Stockers Avenue, Winchester, S022 5LB) 128 pp. Hardback. EUR 9.50 ISBN 1-874728-00-3. Learned Publishing, 6(2):71, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1993:BRIa

[473] A. B. Britton. Book review: The intelligent corporation: the privatisation of intelligence. Edited by Sigurdson Jon and Tågerud Yael London: Taylor Graham, 1992 Paperback. 199 pp. ISBN 0-947568-55-7. Learned Publishing, 6(2):72, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

OConnor:1993:BRE

[474] Maeve O’Connor. Book review: Editing for desktop publishing: a guide to writing and editing on the personal computer. John, Taylor and Shirley, Heale 1992 Hatfield, Herts: John Taylor Book Ventures (7 Cranborne Road, Hatfield, Herts, AL10 8AW) 205 p. Spiral bound. Learned Publishing, 6(2):73, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Calvert:1993:BRA

[475] Drusilla Calvert. Book review: Automated support to indexing by Hodge, Gail M 1992 NFAIS Report Series. Philadelphia PA: National Federation of Abstracting & Information Services ppl76. $100 ISBN 0-942308-36-0. Learned Publishing, 6(2):73–74, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1993:BRIb

[476] A. B. Britton. Book review: Information retrieval interaction. Ingwersen, Peter. London: Taylor Graham, 1992 Paperback. x, 246 pp. EUR 30 ISBN 0-947568-54-9. Learned Publishing, 6(2):75, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:CEb

[477] Anonymous. Coming events. Learned Publishing, 6(2):78–80, April 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1993:QPC

[478] John St. Aubyn. Questions of photocopying and the CLA. Learned Publishing, 6(3):3–7, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Newmarch:1993:VSM

[479] Rob Newmarch. VAT and the Single Market. Learned Publishing, 6(3): 9–15, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Barden:1993:EPF

[480] Phil Barden. Electronic publishing: fears, myths, and scenarios. Learned Publishing, 6(3):17–20, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gervais:1993:PIP

[481] Daniel Gervais. Protection of intellectual property and copyright: the rôle of the World Intellectual Property Organization. Learned Publishing, 6(3):21–23, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gregory:1993:EPI

[482] Martin W. Gregory. The epidemiology and prophylaxis of infectious verbal pomposity. Learned Publishing, 6(3):25–29, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Adkins:1993:FCY

[483] Lesley Adkins and Roy Adkins. First catch your indexer*. Learned Publishing, 6(3):30–31, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Day:1993:DFJ

[484] Shirley Day. Despatching the finished journal. Learned Publishing, 6(3): 33–37, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1993:NP

[485] David Lee. Network publishing. Learned Publishing, 6(3):38–39, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Crane:1993:JPC

[486] Ian D. Crane. Journal pricing in a changing world: New directions, new issues. Learned Publishing, 6(3):40, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Crane:1993:EJP

[487] Ian D. Crane. The economics of journal publishing: Ninth International Learned Journals Seminar. Learned Publishing, 6(3):41–43, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:SSP

[488] Hazel K. Bell. Software for small publishers. Learned Publishing, 6(3): 43–45, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

M:1993:ILJ

[489] J. M. M. Introduction to learned journal production. Learned Publishing, 6(3):45–47, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:Ac

[490] Anonymous. Awards. Learned Publishing, 6(3):48–50, July 1993. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:PBc

[491] Anonymous. People and bodies. Learned Publishing, 6(3):50–51, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:BCNc

[492] Anonymous. British Council news. Learned Publishing, 6(3):52–53, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:IMc

[493] Anonymous. Information management. Learned Publishing, 6(3):54– 56, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:RAc

[494] Anonymous. Reading around. Learned Publishing, 6(3):57–58, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Vickers:1993:BRM

[495] John A. Vickers. Book review: Marketing for small publishers by Bill, Gober, Robert, Webb and Keith, Smith. London: Journeyman Press, 1992. xii, 128 pp. Index. Paperback. £8.99 ISBN 1-85172-035-9. Learned Publishing, 6(3):59, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Toop:1993:BRSa

[496] Sharon Toop. Book review: Standing up for standards: Proceedings of the second annual conference of the Society of Freelance Editors and Proofreaders. Sheffield, 17-18 September 1991. 94 pp. Paperback £5.60. Available from SFEP, 10 Grange Close, Godalming, Surrey GU7 1XT. ISBN 0-9517527-23. Learned Publishing, 6(3):60, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1993:BRC

[497] David Lee. Book review: The common market for information. Proceedings of the annual conference of the Institute of Information Scientists. June 1992, Bedford. Edited by Blake Monica Taylor Graham Publishing, 1992. £25.00 (US $46.00) 121 pp. ISBN 0-947568-56-5. Learned Publishing, 6(3):60–61, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1993:BRG

[498] A. B. Britton. Book review: Guide to careers in abstracting and indexing. By Ann Marie, Cunningham and Wendy, Wicks Philadelphia: National Federation of Abstracting and Information Services, 1992. 115 pp. Spiral bound. $25. ISBN 0-942308-38-7. Learned Publishing, 6(3):61–62, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McAllister:1993:SR

[499] Steve McAllister. Software reviews. Learned Publishing, 6(3):63–71, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McAllister:1993:GCD

[500] Steve McAllister. Glossary of computer and DTP terms. Learned Publishing, 6(3):71–79, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:CEc

[501] Anonymous. Coming events. Learned Publishing, 6(3):79–80, July 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Parker:1993:B

[502] Charles Parker and Robert Millson. In the beginning. Learned Publishing, 6(4):3–5, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ziemacki:1993:CCE

[503] Richard L. Ziemacki. Coping with change: the editorial process. Learned Publishing, 6(4):7–12, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pearce:1993:LSV

[504] Anthony Pearce. The Learned Society view of the economics of journal publishing. Learned Publishing, 6(4):13–16, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mulholland:1993:ICC

[505] Sarah Mulholland. Indexing and the challenge of change at PsyclNFO. Learned Publishing, 6(4):17–21, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Whittall:1993:SNI

[506] Jane Whittall. Satisfying the needs of industry for journals. Learned Publishing, 6(4):23–28, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hart:1993:WMP

[507] F. Dudley Hart. Writing medical papers: a practitioner’s view. Learned Publishing, 6(4):29–32, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hu:1993:PC

[508] Yeping Hu and Qiuping Li. Publishing in China. Learned Publishing, 6(4):33–36, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:PUL

[509] Hazel K. Bell. Princeton University Library: ‘Not a data bank but a repository of knowledge’. Learned Publishing, 6(4):37–40, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

M:1993:LJP

[510] J. M. M. Learned journal production. Learned Publishing, 6(4):41, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:LMB

[511] Hazel K. Bell. London Medical Book Fair. Learned Publishing, 6(4):41, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1993:JD

[512] Hazel K. Bell. Journal design. Learned Publishing, 6(4):41–43, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:Ad

[513] Anonymous. Awards. Learned Publishing, 6(4):45, October 1993. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1993:CCb

[514] John St. Aubyn. Copyright corner. Learned Publishing, 6(4):46–47, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:PBd

[515] Anonymous. People and bodies. Learned Publishing, 6(4):47–49, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:OAc

[516] Anonymous. Overseas affairs. Learned Publishing, 6(4):50–52, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:BCNd

[517] Anonymous. British Council news. Learned Publishing, 6(4):52–53, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:IMd

[518] Anonymous. Information management. Learned Publishing, 6(4):54–55, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:RAd

[519] Anonymous. Reading around. Learned Publishing, 6(4):56–58, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thorn:1993:C

[520] Sue Thorn. Correspondence. Learned Publishing, 6(4):58, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Toop:1993:BRSb

[521] Sharon Toop. Book review: Scholarly editing in the computing age: theory and practice. Peter, L. Shillingsburg Athens, GA: The University of Georgia Press. 178 pp. ISBN 0-8203-0889-7. Learned Publishing, 6(4): 59, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

White:1993:BRG

[522] Stephen White. Book review: Getting the message across Public Relations, Publicity and Working with the Media. Sue, Ward. Journeyman Press. 230 pages + index 1-85172-043-X pb. £12.95 pb. ISBN 1-85172-042-1 hb. Learned Publishing, 6(4):60, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1993:BRP

[523] John St. Aubyn. Book review: Publishing agreements: a book of precedents. 4th Edition General Editor: Clark Charles London: Butterworths 1993 Hardback 270pp EUR 37.50 ISBN 0-406-02094-9. Learned Publishing, 6(4):61–62, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1993:BRIc

[524] A. B. Britton. Book review: Information systems for end-users: research and development issues. Edited by Hancock-Beaulieu Micheline. London: Taylor Graham, 1992. 85 pp. Paperback. ISBN 0-947568-57-3. Learned Publishing, 6(4):62, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1993:CEd

[525] Anonymous. Coming events. Learned Publishing, 6(4):65–66, October 1993. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Askew:1994:CA

[526] Bill Askew. Coming of age. Learned Publishing, 7(1):3–4, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dorner:1994:WTI

[527] Jane Dorner. Whose text is it anyway? Learned Publishing, 7(1):5–9, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Miles:1994:TJ

[528] John Miles. Typography for journals. Learned Publishing, 7(1):11–16, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McFadden:1994:II

[529] T. G. McFadden. Indexing the Internet. Learned Publishing, 7(1):17–27, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott:1994:MWS

[530] J. T. Scott. Medical writing: should you add to it? Learned Publishing, 7(1):29–31, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mack:1994:ECE

[531] Arien Mack. The East and Central Europe Journal Donation Project. Learned Publishing, 7(1):33–38, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:LLL

[532] Hazel K. Bell. LC; the largest library in the changing world. Learned Publishing, 7(1):39–44, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:JC

[533] Hazel K. Bell. The journal at the crossroads. Learned Publishing, 7 (1):45–48, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1994:CIa

[534] John St. Aubyn. Copyright issues. Learned Publishing, 7(1):49–50, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:PBa

[535] Anonymous. People and bodies. Learned Publishing, 7(1):51–53, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:BCN

[536] Anonymous. British Council news. Learned Publishing, 7(1):54, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:IMa

[537] Anonymous. Information management. Learned Publishing, 7(1):54–58, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

deGlanville:1994:C

[538] Hugh de Glanville. Correspondence. Learned Publishing, 7(1):58, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brown:1994:SR

[539] David Brown. Software reviews. Learned Publishing, 7(1):59–63, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:BRW

[540] Hazel Bell. Book review: As I was saying: essays on the international book business by Gordon, Graham London: Hans Zell. 320 pp, 1993. paperback £16.95. ISBN 1-873836-01-5. Learned Publishing, 7(1):63–64, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:BRH

[541] Hazel K. Bell. Book review: How to publish a book: a practical stepby-step guide to independent publishing. Robert, Spicer Plymouth: How To Books, 1993. 139 pp. 22 cm. Illus, bibliog, index (pbk.): £8.99 ISBN 1-850703-071-0. Learned Publishing, 7(1):64, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1994:BRU

[542] A. B. Britton. Book review: University libraries and scholarly communication. By Anthony M., Cummings et al. [USA]: Association of Research Libraries for the Andrew W. Mellon Foundation, 1992. Paperback. xxix, 205 pp. ISBN 9-918006-22-8. Learned Publishing, 7(1):64–65, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1994:BRE

[543] Michele Benjamin. Book review: Ethics and Policy in Scientific Publication. Bethesda, MD: Council of Biology Editors. 1990. Pp. 290. Paperback. ISBN 0-914340-09-3. Learned Publishing, 7(1):65–66, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1994:BRT

[544] Michele Benjamin. Book review: Three views of the Internet. Edited by Marie Cunningham Ann and Wendy Wicks 1993 Paperback $60.00 NFAIS Report Series 3. Paperback. Pp. 106 Philadelphia: National Federation of Abstracting and Information Services. Learned Publishing, 7 (1):66, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:CEa

[545] Anonymous. Corning events. Learned Publishing, 7(1):71–72, January 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Campbell:1994:HWA

[546] Robert Campbell. How will academic libraries manage? What hope does the Follett report offer? Learned Publishing, 7(2):75–78, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Naylor:1994:ICF

[547] Bernard Naylor. The implications of current and future initiatives for libraries. Learned Publishing, 7(2):79–83, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:1994:CDE

[548] John Cox. Can dinosaurs evolve? The future of publishers, vendors and librarians. Learned Publishing, 7(2):85–88, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Coles:1994:WPJ

[549] Bryan R. Coles. Who pays for journal publishing? The user’s point of view. Learned Publishing, 7(2):89–91, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McMillan:1994:TPE

[550] Gail McMillan. Technical processing of electronic journals in libraries. Learned Publishing, 7(2):93–98, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cooper:1994:HER

[551] Alisa N. Cooper. The house at Eighteenth and R: Special Libraries Association. Learned Publishing, 7(2):99–103, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:PLN

[552] Hazel K. Bell. Photocopying licensing — now and in the future. Learned Publishing, 7(2):105–108, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kachmar:1994:WWL

[553] Diane Kachmar. Will we learn to manage the new electronic technologies before they overwhelm us? Charleston Library Acquisitions Conference. Learned Publishing, 7(2):108–111, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Symondson:1994:CC

[554] Brian Symondson. To CD or not to CD. Learned Publishing, 7(2):112– 115, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:EBM

[555] Hazel K. Bell. Electronic books and multimedia: prospects and perils. Learned Publishing, 7(2):116–117, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:SIS

[556] Hazel K. Bell. The scientific information system in the UK. Learned Publishing, 7(2):118–121, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1994:CIb

[557] John St. Aubyn. Copyright issues. Learned Publishing, 7(2):122–124, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:PBb

[558] Anonymous. People and bodies. Learned Publishing, 7(2):124–128, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:IMb

[559] Anonymous. Information management. Learned Publishing, 7(2):128– 131, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:SRa

[560] Anonymous. Software reviews. Learned Publishing, 7(2):131–134, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Holt:1994:BRS

[561] Harry Holt. Book review: The STM information system in the UK. A study on behalf of the Royal Society, the British Library and ALPSP. 218 pp. Paperback. £15.00. Copies available from the Publications Department, The Royal Society. ISBN 03-85403-4773-3. Learned Publishing, 7 (2):135–137, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Banks:1994:BRE

[562] Mary Banks. Book review: European Bookseller September/October 1993. Published by SLG ISBN 0-958-3866. Learned Publishing, 7(2): 137–138, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Duncan:1994:BRH

[563] Nigel Duncan. Book review: Hitting the headlines: a practical guide to media. Stephen, White, Peter, Evans, Chris, Mihill and Maryon, Tysoe British Psychological Society, 1993. 136 pp. Pb. ISBN 1-85433-107-8. Learned Publishing, 7(2):138, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:CEb

[564] Anonymous. Coming events. Learned Publishing, 7(2):141–142, April 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:LG

[565] Hazel Bell. Looking good. Learned Publishing, 7(3):147, July 1994. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1994:ESI

[566] Edwin Shelock. The economics of the STM information system. Learned Publishing, 7(3):149–151, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pearce:1994:LSP

[567] Anthony Pearce. A learned society publisher’s perspective of future opportunities. Learned Publishing, 7(3):153–157, July 1994. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Law:1994:RAL

[568] Derek Law. The role of academic libraries in the scientific information system. Learned Publishing, 7(3):159–162, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott:1994:VIS

[569] Elspeth J. Scott. The view of the industrial scientist concerning the information system. Learned Publishing, 7(3):163–166, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Olsen:1994:EJI

[570] Jan Olsen. Electronic journals: implications for scholars. Learned Publishing, 7(3):167–176, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:PMC

[571] Hazel K. Bell. Publicity material: content versus design. Learned Publishing, 7(3):177–178, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thorn:1994:PNW

[572] Sue Thorn. Publishing in the new world: The Tenth International Learned Journals Seminar. Learned Publishing, 7(3):178–185, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:SMA

[573] Hazel K. Bell. Saving money on administration and subscription management. Learned Publishing, 7(3):185–187, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:LMB

[574] Hazel K. Bell. London Medical Book Fair. Learned Publishing, 7(3):187– 188, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Page:1994:CGA

[575] Gillian Page. The Charlesworth Group Award for Typographical Excellence in Journal and Serial Publishing. Learned Publishing, 7(3):189–191, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:BPM

[576] Anonymous. BUPA Prizes for medical writing and illustration. Learned Publishing, 7(3):191–193, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1994:CIc

[577] John St. Aubyn. Copyright issues. Learned Publishing, 7(3):193–195, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:IMc

[578] Anonymous. Information management. Learned Publishing, 7(3):195– 199, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:RA

[579] Anonymous. Reading around. Learned Publishing, 7(3):199–202, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:SRb

[580] Anonymous. Software reviews. Learned Publishing, 7(3):203–206, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Neidle:1994:BRG

[581] Stephen Neidle. Book review: Guardians of science, fairness and reliability of peer review, by Daniel, H.-D. Weinheim: VCH 1993. 118 pp. hdbk. ISBN 3-527-29041-9. Learned Publishing, 7(3):207, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Waddingham:1994:BRC

[582] Anne Waddingham. Book review: Copyright made easier by Wall, Raymond A. London: Aslib, 1993. 393 pp. pbk. ISBN 0-85142-310-8. Learned Publishing, 7(3):207–208, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Peter:1994:BRF

[583] Sue Peter. Book review: Full marks: advice on punctuation for scientific and technical writing, 2nd edn. by Kirkman, John. Marlborough, Wiltshire: Ramsbury Books, 1993. 154 pp. ISBN 0-9521762-0-3. Learned Publishing, 7(3):208–209, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowlands:1994:BRG

[584] Jane Rowlands. Book review: Going online and CD-ROM by Bradley, Phil and Hanson, Terry. 9th Edit. London: Aslib 1993. 122 pp. £21.00 members, £26.00 non-members. ISBN 0-85142-323-X. Learned Publishing, 7(3):210, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:1994:BRG

[585] Alex Williamson. Book review: Guidelines for giving effective presentations: tactics and techniques for talking to large or small groups, by Kirkman, John Ramsbury Books, 1994. 184 pp. Pb. ISBN 0-9521762-1-1. Learned Publishing, 7(3):210–211, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1994:BRP

[586] Jane Donovan. Book review: Print: how you can do it yourself! by Zeitlyn, Jonathan London: Journeyman Press, 1992. 128 pp. pbk. £7.95. ISBN 1-85172-049-9. Learned Publishing, 7(3):211, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:CEc

[587] Anonymous. Corning events. Learned Publishing, 7(3):213, July 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Graham:1994:HUP

[588] Gordon Graham. How the UK’s publishers and librarians have grown (a little) closer: one publisher’s experience. Learned Publishing, 7(4): 219–222, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:1994:CIP

[589] Alan Singleton. Charging for information — pricing policies and practice. Learned Publishing, 7(4):223–232, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Worlock:1994:TFE

[590] Chairman David R. Worlock. Is there a future for the Electronic Book? Learned Publishing, 7(4):233–237, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mills:1994:PJD

[591] Jeremy Mills. The practicalities of journal design. Learned Publishing, 7 (4):239–243, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1994:EEE

[592] Michele Benjamin. Editing, ethics, electronics and economics. Learned Publishing, 7(4):245–251, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1994:HEW

[593] Hazel K. Bell. How to exhibit: what can go wrong. Learned Publishing, 7(4):251–252, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1994:CId

[594] Chariman John St. Aubyn. Copyright issues. Learned Publishing, 7(4): 253–254, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:IMd

[595] Anonymous. Information management. Learned Publishing, 7(4):255– 257, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hyman:1994:SR

[596] Kay Hyman. Software reviews. Learned Publishing, 7(4):259–269, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hayes:1994:BRP

[597] Jeff Hayes. Book review: PageMaker 5 in easy steps. Basham, Scott Computer Step, 1993. 223 pp £14.95. Learned Publishing, 7(4):270, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hughes:1994:BRM

[598] Glen Hughes. Book review: Mark my words — instruction and practice in proofreading, 2nd edn. by Smith, Peggy EEI: Alexandria, VA, USA. 480 pp. $28 ISBN 0-935012-16-8. Learned Publishing, 7(4):270–271, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Evans:1994:BRH

[599] Susan Evans. Book review: How to market books by Baverstock, Alison London: Kogan Page Limited 1990; reprinted with revisions 1993. 322 pp. £12.95 ISBN 0-7494-0126-5. Learned Publishing, 7(4):271–273, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1994:CEd

[600] Anonymous. Coming events. Learned Publishing, 7(4):274, October 1994. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1995:ES

[601] Michele Benjamin. Not the exhibiting seminar. Learned Publishing, 8 (1):3–6, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Secretary:1995:MS

[602] Executive Secretary and Linnean Society of London John Marsden. The e-mail of the species. Learned Publishing, 8(1):7–9, January 1995. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mowat:1995:EAP

[603] Ian R. M. Mowat. European academic publishing: the British academic librarian’s view. Learned Publishing, 8(1):11–17, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Matheson:1995:WUL

[604] Nina W. Matheson. What US libraries are demanding, and doing. Learned Publishing, 8(1):19–24, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Keyhani:1995:EPU

[605] Andrea Keyhani. Electronic publishing: US publishers’ initiatives. Learned Publishing, 8(1):25–32, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Teague:1995:OE

[606] Cheryl Teague. Organizing exhibitions. Learned Publishing, 8(1):33–37, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bedford:1995:ELS

[607] Jeremy Bedford. Electronics Letters — the SGML implementation. Learned Publishing, 8(1):39–44, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:LMI

[608] Hazel K. Bell. The life of the mind on the information highway. Learned Publishing, 8(1):45–46, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1995:NAR

[609] Michele Benjamin. North America revisited. Learned Publishing, 8(1): 47–48, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

HKB:1995:MMI

[610] H.K.B. Making money from information technology in a learned society. Learned Publishing, 8(1):48–49, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:WAB

[611] Hazel K. Bell. Who is afraid of the big bad CD-ROM? Learned Publishing, 8(1):49–51, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:PI

[612] Hazel K. Bell. Publishing and the Internet. Learned Publishing, 8(1):51– 52, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1995:CIa

[613] Chairman John St. Aubyn. Copyright issues. Learned Publishing, 8(1): 53–54, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:IMa

[614] Anonymous. Information management. Learned Publishing, 8(1):55–56, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:RAa

[615] Anonymous. Reading around. Learned Publishing, 8(1):57, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Robertson:1995:SR

[616] Michael Robertson. Software reviews. Learned Publishing, 8(1):59–63, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Webb:1995:BRS

[617] Ray Webb. Book review: Scientific Illustration (Second Edition) Phyllis Wood AMI GNSI Van Nostrand Reinhold 158 pp ISBN 0-442-01316-7.

Learned Publishing, 8(1):63–64, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Robertson:1995:BRP

[618] Michael Robertson. Book review: The Penguin dictionary for writers and editors, Bryson, Bill London: Viking, 1991. 403 pp. £12.99 ISBN 0-670-83767-9. Learned Publishing, 8(1):64–65, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Robertson:1995:BRH

[619] Michael Robertson. Book review: The high-technology editorial guide and stylebook. Perdue, Lewis Homewood, IL: Business One Irwin, 1991 x, 194 pp With diskette £25.99. [ISBN] 1-55623-531-3 (PC edition), 1-55623-530-5 (Macintosh edition). Learned Publishing, 8(1):66, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1995:BRI

[620] David Lee. Book reviews: Indexing books Mulvany, Nancy C. Chicago, London: University of Chicago Press, 1994 xiv, 320 pp. £23.95 ISBN 0-226-55014-1. Indexing, providing access to information: Looking back, looking ahead, the Proceedings of the 25th Annual Meeting of the American society of Indexers. Mulvany, Nancy C., editor. Port Aransas, Texas: American Society of Indexers, 1993. vi, 161 pp. $35.00 ISBN 0-936547-19-7. Learned Publishing, 8(1):66–67, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:1995:BRC

[621] Alex Williamson. Book review: Careers in publishing and bookselling. Second edition by Lines, June Kogan Page, 1994. 96 pp. Paperback, £6.99. Learned Publishing, 8(1):67–68, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:CEa

[622] Anonymous. Coming events. Learned Publishing, 8(1):69, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1995:CLJ

[623] John St. Aubyn. Copyright and learned journals: some questions answered. Learned Publishing, 8(2):75–86, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:1995:HCC

[624] David Powell. From hard copy to CD-ROM: a critical path. Learned Publishing, 8(2):87–91, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sharples:1995:MIN

[625] Carolyn Sharples. Meeting the information need in developing countries. Learned Publishing, 8(2):93–98, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1995:LSI

[626] Bernard T. Donovan. Librarians in search of an issue: task force zeroes in on profit pursuers. Learned Publishing, 8(2):99–102, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thorn:1995:YCY

[627] Sue Thorn. Are you in control of your scientific literature? Learned Publishing, 8(2):103–104, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thatcher:1995:WTI

[628] Sanford Thatcher. Why Texaco is important to university presses. Learned Publishing, 8(2):105–111, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davenport:1995:WWE

[629] Anne Davenport. Why and where to exhibit. Learned Publishing, 8(2): 113–117, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:SMB

[630] Hazel K. Bell. Seminars and meetings: BPIF leared jourals forum. Learned Publishing, 8(2):119, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:LA

[631] Hazel K. Bell. The licensing of abstracts. Learned Publishing, 8(2):119– 121, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:CTM

[632] Anonymous. Charity Trustee matters. Learned Publishing, 8(2):121, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thorn:1995:AA

[633] Sue Thorn. ALPSP affairs. Learned Publishing, 8(2):122, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:PBa

[634] Anonymous. People and bodies. Learned Publishing, 8(2):123–125, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:IMb

[635] Anonymous. Information management. Learned Publishing, 8(2):125– 128, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mulliner:1995:SR

[636] Claire Mulliner. Software reviews. Learned Publishing, 8(2):129, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:BRJ

[637] Hazel K. Bell. Book review: Journal of Scholarly Publishing. 26 (1) October 1994. Toronto: University of Toronto Press. 102 pp. 23 cm. Index. ISSN 1198-9742. Personal subscriptions $25.00 per year. Learned Publishing, 8(2):130–132, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bide:1995:BRE

[638] Mark Bide. Book review: The electronic publishing business and its market. Edited by Brian and Margot Blunden Leatherhead: IEPRC/ Pira International Hardback. xii + 628 pp. £45.00/$70.00 (10 discount for IEPRC and Pira members) Published October 1994. Learned Publishing, 8(2):132, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:CEb

[639] Anonymous. Coming events. Learned Publishing, 8(2):134–135, April 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:1995:PRR

[640] Alex Williamson. Peer review research: an invitation to ALPSP members to join Locknet. Learned Publishing, 8(3):139–140, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jarvis:1995:SHO

[641] John Jarvis. Standardizing ‘headers’: OASIS. Learned Publishing, 8(3): 141–143, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowlinson:1995:SSC

[642] Carolyn Rowlinson. SCOPE: The Scottish Collaborative On-demand Publishing Enterprise. Learned Publishing, 8(3):145–149, July 1995. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jackson:1995:PRG

[643] Head of Corporate Communication Marie Jackson. Public Relations — getting it right. Learned Publishing, 8(3):151–157, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Payne:1995:PO

[644] Alan Payne. Publishing to order. Learned Publishing, 8(3):159–162, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Barton:1995:EPB

[645] Geoffrey J. Barton. Electronic publishing and the biochemist. Learned Publishing, 8(3):163–167, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:HMP

[646] Hazel K. Bell. How to manage public relations. Learned Publishing, 8(3):169–170, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:UPC

[647] Hazel K. Bell. University publishing courses: their relevance to book industry needs. Learned Publishing, 8(3):170–173, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fowler:1995:RC

[648] John Fowler. Roundtable on copyright. Learned Publishing, 8(3):173– 176, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1995:ACI

[649] David Lee. Anticipating change: 11th International Learned Journals seminar. Learned Publishing, 8(3):176–178, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:LJP

[650] Hazel K. Bell. Learned journal production. Learned Publishing, 8(3):178– 180, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:A

[651] Anonymous. Awards. Learned Publishing, 8(3):181–184, July 1995. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:PBb

[652] Anonymous. People and bodies. Learned Publishing, 8(3):185–187, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1995:CIb

[653] John St. Aubyn. Copyright issues. Learned Publishing, 8(3):187–188, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:IMc

[654] Anonymous. Information management. Learned Publishing, 8(3):188– 191, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:RAb

[655] Anonymous. Reading around. Learned Publishing, 8(3):191–192, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Turner:1995:SR

[656] Robin Turner. Software reviews. Learned Publishing, 8(3):193–194, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:1995:BRS

[657] Alex Williamson. Book review: Serial publications: guidelines for good practice in publishing printed journals and other serial publications. Published by the UK Serials Group, Witney, Oxfordshire. 1994. 48pp. paperback. £16.00. ISBN 0-906148-10-3. Learned Publishing, 8(3):195–196, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Friend:1995:BRN

[658] Frederick J. Friend. Book review: The new university library: issues for the ’90s and beyond. Essays in honour of Ian Rogerson. Edited by Coling Harris. London, Taylor Granham, 1994. 131pp. ISBN 0-947-56864-6. Learned Publishing, 8(3):196–197, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Youens:1995:BRB

[659] Jane Youens. Book review: Are books different? Marketing in the book trade. Baverstock, Alison Kogan Page, 1993. 171pp. pb. ISBN 0-7494-0900-2. Learned Publishing, 8(3):197–198, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:CEc

[660] Anonymous. Coming events. Learned Publishing, 8(3):199, July 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pearce:1995:BBU

[661] Anthony Pearce. From books to bytes: the user at the center of the scientific information system. Learned Publishing, 8(4):203–208, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:1995:SLP

[662] John Cox. The site licence: problems and opportunities for journal publishers. Learned Publishing, 8(4):209–211, January 1995. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morgan:1995:OC

[663] Cliff Morgan. OASIS in context. Learned Publishing, 8(4):213–215, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

White:1995:OC

[664] Stephen White. OASIS in context. Learned Publishing, 8(4):217–220, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Duncan:1995:HMP

[665] Head of Communications Nigel Duncan. How to manage public relations: dealing with specific issues. Learned Publishing, 8(4):221–223, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mcgarva:1995:PPE

[666] Duncan Mcgarva and Dave Roberts. From paper to pixels: Electronic publishing and the microbiologist. Learned Publishing, 8(4):225–227, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mowat:1995:SSC

[667] Ian R. M. Mowat. Surviving in Sarajevo: the continuity of academic life in a war zone. Learned Publishing, 8(4):229–232, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ball:1995:VCC

[668] Stephen Ball. Viruses: from curiosity to Computer Armageddon, and back? Learned Publishing, 8(4):233–243, January 1995. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:TND

[669] Hazel K. Bell. Is there a need for a digital archive? Learned Publishing, 8(4):245, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:THE

[670] Hazel K. Bell. Truth and honesty in editing and publishing. Learned Publishing, 8(4):246–247, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:JD

[671] Hazel K. Bell. Journal design. Learned Publishing, 8(4):248–249, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1995:PGP

[672] Hazel K. Bell. ‘The present generation of publishers is doomed’. Learned Publishing, 8(4):249–251, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Muller:1995:SLA

[673] Gervase Muller. Special Libraries Association Annual Conference. Learned Publishing, 8(4):252, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1995:CIc

[674] Chairman John St. Aubyn. Copyright issues. Learned Publishing, 8(4): 253–254, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:PBc

[675] Anonymous. People and bodies. Learned Publishing, 8(4):254–257, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:IMd

[676] Anonymous. Information management. Learned Publishing, 8(4):257– 260, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1995:BRC

[677] John St. Aubyn. Book review: Copyright theft, by John Gurnsey Gower, Aslib 1995. 224pp £28.50 ISBN 0-566-07631-4. Learned Publishing, 8 (4):261–262, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Britton:1995:BRV

[678] A. B. Britton. Book review: The value and impact of information. Edited by Mary Feeney and Maureen Grieves (British Library Research: Information Policy Issues) London: Bowker-Saur, 1994. Hardback. xii, 282 pp. ISBN 1-85739-084-9. Learned Publishing, 8(4):262–263, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Roper:1995:IRN

[679] Vincent de P. Roper. Information retrieval: new systems and current research. Edited by Ruben Leon. Proceedings of the 15th Research Colloquium of the British Computer Society Information Retrieval Specialist Group, Glasgow, 1993. Learned Publishing, 8(4):263–265, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Henderson:1995:C

[680] Albert Henderson. Correspondence. Learned Publishing, 8(4):266, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1995:CEd

[681] Anonymous. Coming events. Learned Publishing, 8(4):267–268, January 1995. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Glen:1996:WDE

[682] John W. Glen. Where do ethical responsibilities in editing and publishing lie? Learned Publishing, 9(1):3–5, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Batchelor:1996:AA

[683] George Batchelor. An arbiter for authors? Learned Publishing, 9(1):7– 12, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sharp:1996:JIP

[684] David W. Sharp. Journals and issues of privacy and confidentiality. Learned Publishing, 9(1):13–15, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Semmens:1996:EPW

[685] Clive Semmens. Electronic publishing: what the Physiological Society is doing. Learned Publishing, 9(1):17–21, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1996:EEO

[686] Hazel K. Bell. The evolution of an editorial office in a small society. Learned Publishing, 9(1):23–27, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1996:LPI

[687] David Lee. Libraries, publishers and information provision: Life after FIGIT: 4 September 1995, London; ALPSP/UKSG. Learned Publishing, 9(1):29–31, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1996:SSS

[688] Hazel K. Bell. Serial selection and de-selection: how libraries decide: 9 October 1995, London; ALPSP. Learned Publishing, 9(1):31–32, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:CIJa

[689] Anonymous. Copyright issues: John St. Aubyn, Chairman, ALPSP Copyright Committee. Learned Publishing, 9(1):33–34, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:PBa

[690] Anonymous. People and bodies. Learned Publishing, 9(1):34–37, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:Aa

[691] Anonymous. Awards. Learned Publishing, 9(1):37–38, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Storrie:1996:AA

[692] Eileen Storrie and Sue Thorn. ALPSP affairs. Learned Publishing, 9 (1):39–43, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:IMa

[693] Anonymous. Information management. Learned Publishing, 9(1):44–46, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:RAa

[694] Anonymous. Reading around. Learned Publishing, 9(1):46–47, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kingston:1996:SRa

[695] Ian Kingston and Michael Robertson. Software reviews. Learned Publishing, 9(1):48–51, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jones:1996:BRI

[696] Richard M. Jones. Book review: Indexing the medical and biological sciences, by Blake, Doreen, Clarke, Michèle, McCarthy, Anne, Morrison, June with medical advice from Gibson, John (Society of Indexers Occasional Papers on Indexing no. 3) London, Society of Indexers, 1995 vii, 61 pp 21 cm ISBN 1-87157745-2 (pbk): £11.00 Available from Sales Manager, Society of Indexers, 38 Rochester Road, London NW1 9JJ. Learned Publishing, 9(1):52–53, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brooke:1996:BRS

[697] Oliver Brooke. Book review: Scientific style and format: The CBE manual for authors, editors, and publishers, 6th ed. Style Manual Committee, Council of Biology Editors (Huth, Edward J., Chairman). Cambridge: Cambridge University Press. Hardback, xv + 825 pp. £19.95/$35.95. Published January 1995. ISBN 0-521-47154-0. Learned Publishing, 9(1): 53–54, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Horrocks:1996:BRW

[698] Tony Horrocks. Book review: The World Wide Web — beneath the surf. Handley, Mark and Crowcroft, Jon UCL Press Paperback 198 pages 1995. Learned Publishing, 9(1):54–55, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:CEa

[699] Anonymous. Coming events. Learned Publishing, 9(1):56, January 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:1996:SPW

[700] Andrea Powell. Setting up a presence on the World Wide Web. Learned Publishing, 9(2):59–66, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Horrocks:1996:DIW

[701] Tony Horrocks. Design issues on the World Wide Web. Learned Publishing, 9(2):67–71, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morgan:1996:IIT

[702] Adrienne Morgan. The immunologist and the Internet — a tale of the tortoise and the hare? Learned Publishing, 9(2):73–77, April 1996. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Feest:1996:NPW

[703] Tim Feest. NVQs in Publishing: why bother? Learned Publishing, 9(2): 79–85, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Germain:1996:SSD

[704] J. Charles Germain. Serial selection and deselection decisions: the view from America. Learned Publishing, 9(2):87–93, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fowler:1996:PII

[705] John Fowler. Publishing and the Internet II: journals: 16 November, 1995, London; ALPSP. Learned Publishing, 9(2):95–97, April 1996. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:MER

[706] Anonymous. Managing electronic rights: controlling electrocopying: 20 November 1995, London; Publishers Licensing Society. Learned Publishing, 9(2):98–99, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1996:CIE

[707] Hazel K. Bell. Change and innovation in editing: 21 November 1995, London; Book House Training Centre. Learned Publishing, 9(2):100– 101, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1996:DLS

[708] David Lee. A day in the life of a serials librarian: 29 November 1995, London; United Kingdom Serials Group. Learned Publishing, 9(2):102– 103, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1996:CPW

[709] Hazel Bell. Creating a presence on the World–Wide Web: 11 December 1995, London; ALPSP. Learned Publishing, 9(2):104–105, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:CIJb

[710] Anonymous. Copyright issues: John St. Aubyn, Chairman, ALPSP Copyright Committee. Learned Publishing, 9(2):106–107, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:PBb

[711] Anonymous. People and bodies. Learned Publishing, 9(2):108–109, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:Ab

[712] Anonymous. Awards. Learned Publishing, 9(2):110–111, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:RAb

[713] Anonymous. Reading around. Learned Publishing, 9(2):112–114, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:IMb

[714] Anonymous. Information management. Learned Publishing, 9(2):115– 119, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dorner:1996:SR

[715] Jane Dorner. Software reviews. Learned Publishing, 9(2):120–123, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Deeley:1996:BRH

[716] Sue Deeley. Book reviews: How to write & publish a scientific paper. Fourth edition by Day, Robert A. Cambridge University Press, 1995. 224pp. ISBN 0-521-55898-0 Pb 0-521-55136-6 hb How to write a paper Edited by Hall, George M. BMJ Publishing Group 118pp. pb ISBN 0-7279-0822-7. Learned Publishing, 9(2):123–125, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:CEb

[717] Anonymous. Coming events. Learned Publishing, 9(2):127–128, April 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thorn:1996:EES

[718] Sue Thorn. eSPeRe: Electronic submission and peer review. Learned Publishing, 9(3):131–132, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lock:1996:FE

[719] Stephen Lock. Fraud and the editor. Learned Publishing, 9(3):133–140, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Keown:1996:JBE

[720] Damien Keown. The Journal of Buddhist Ethics: an online journal. Learned Publishing, 9(3):141–145, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Duckworth:1996:CUC

[721] Sharon Duckworth. Computers used in copy-editing: master or slave? Learned Publishing, 9(3):147–151, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Embers:1996:ARC

[722] Bill Embers. Academic Reference Centre access for UK universities. Learned Publishing, 9(3):153–155, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jenkins:1996:SSD

[723] Clare Jenkins. Serial selection and deselection: policies and decision– making. Learned Publishing, 9(3):157–162, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1996:LSJ

[724] Hazel K. Bell. The learned society, journals and the Internet: 16 February 1966, London; FST/ ALPSP. Learned Publishing, 9(3):163–166, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Horsfall:1996:EPR

[725] Amanda Horsfall. Electronic publishing: the reality: 12 April 1996, London; ALPSP 12th International Learned Journals Seminar. Learned Publishing, 9(3):166–171, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1996:CIM

[726] Hazel K. Bell. Change and innovation in marketing: marketing after the NBA: 16 April, London; Book House Training Centre. Learned Publishing, 9(3):171–173, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:CIJc

[727] Anonymous. Copyright issues: John St Anbyn, Chairman, ALPSP Copyright Committee. Learned Publishing, 9(3):174–175, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:PBc

[728] Anonymous. People and bodies. Learned Publishing, 9(3):175–177, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:Ac

[729] Anonymous. Awards. Learned Publishing, 9(3):177–179, July 1996. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:RAc

[730] Anonymous. Reading around. Learned Publishing, 9(3):179–181, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:IMc

[731] Anonymous. Information management. Learned Publishing, 9(3):181– 185, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kingston:1996:SRb

[732] Ian Kingston, Ian Kingston, Michael Robertson, and Ian Kingston. Software reviews. Learned Publishing, 9(3):186–194, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Priestley:1996:PPH

[733] Carol Priestley. A proposal for publication of A handbook of good practice in journals publishing and related training programme. Learned Publishing, 9(3):195–196, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:1996:BRG

[734] Alex Williamson. Book review: Guidelines for writers in electronic publishing and the multimedia. Edited by Dorner, Jane The Authors’ Licensing and Collecting Society, 1996 Available from ALCS, Isis House, 74 New Oxford Street, London WC1A 1EF. Learned Publishing, 9(3): 196–197, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:CEc

[735] Anonymous. Coming events. Learned Publishing, 9(3):199, July 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pearce:1996:PKY

[736] Anthony Pearce. Publisher, know your niche! Learned Publishing, 9 (4):203–205, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:TIC

[737] Anonymous. Towards an international consensus on electronic publishing in science. Learned Publishing, 9(4):207–209, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Czech:1996:SDW

[738] Isabel Czech. Spreading the digital word: ISI’s Electronic Library Project. Learned Publishing, 9(4):211–214, October 1996. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Meadows:1996:PDI

[739] Jack Meadows. Preserving the digital imprint. Learned Publishing, 9 (4):215–218, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clarke:1996:RIS

[740] Michèle Clarke. Retrieving information from scientific periodicals. Learned Publishing, 9(4):219–223, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Geleijnse:1996:DL

[741] Hans Geleijnse. The digital library. Learned Publishing, 9(4):225–232, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brown:1996:EPS

[742] David J. Brown. Electronic publishing in science: 19–23 February 1996, Paris; Unesco/ ICSU. Learned Publishing, 9(4):233–236, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:1996:IWI

[743] Andrea Powell. Internet World International 96: 21–23 May 1996, London; Learned Information. Learned Publishing, 9(4):237–240, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:CEd

[744] Anonymous. Coming events. Learned Publishing, 9(4):267–268, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1996:ARN

[745] Bernard Donovan. Assessing the reality of new markets and new media: 29–31 May, 1996, Minneapolis; Society for Scholarly Publishing. Learned Publishing, 9(4):240–241, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:SLA

[746] Anonymous. Special Libraries Association 86th Annual Conference: 9–12 June 1996, Boston. Learned Publishing, 9(4):242, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1996:VJL

[747] Hazel K. Bell. VAT 1996: 28 June 1996, London; Foundation for Science and Technology. Learned Publishing, 9(4):243–245, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1996:CIR

[748] Hazel K. Bell. Change and innovation in rights: 11 July 1996, London; Book House Training Centre. Learned Publishing, 9(4):245–247, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:CIJd

[749] Anonymous. Copyright issues: John St Anbyn, Chairman, ALPSP Copyright Committee. Learned Publishing, 9(4):248–249, October 1996. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1996:PB

[750] Michele Benjamin. People and bodies. Learned Publishing, 9(4):249–252, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:RAd

[751] Anonymous. Reading around. Learned Publishing, 9(4):253–254, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1996:IMd

[752] Anonymous. Information management. Learned Publishing, 9(4):255– 257, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Robertson:1996:SRT

[753] Michael Robertson. Software reviews: Typefaces on CD-ROM and on Demand. Learned Publishing, 9(4):257–265, October 1996. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1996:BRQ

[754] Hazel K. Bell. Book review: Quick guide 13: Indexing your book by Vickers, John The Society of Authors 1996. 12 pp £3 inc. p. and p. (free to members) Available from The Society of Authors, 84 Drayton Gardens, London SW10 9SB. Learned Publishing, 9(4):266, October 1996. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:1996:BRE

[755] Alex Williamson. Book review: Endangered species? Evolving strategies for library collection management. Edited by Woodward, Hazel and Morley, Mary UK Serials Group 1995 £25 (members); £27 (non-members) ISBN 1-87069-14-4. Learned Publishing, 9(4):266, October 1996. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pearce:1997:AAA

[756] Anthony Pearce. ALPSP Affairs: ALPSP and Learned Publishing, 1997. Learned Publishing, 10(1):2, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williams:1997:TFY

[757] Peter Williams. Twenty-five years — and still centre stage: a brief review of the history of ALPSP with a clarification of the aims and objectives of the Association. Learned Publishing, 10(1):3–6, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:1997:LPB

[758] David Lee. Learned Publishing: The Bell Years. Learned Publishing, 10(1):7–8, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McCracken:1997:RAE

[759] Richard McCracken. Rights acquisition in the electronic era. Learned Publishing, 10(1):9–13, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wellisch:1997:ICV

[760] Hans H. Wellisch. Indexing the continuum of verbal texts. Learned Publishing, 10(1):15–19, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dixon:1997:RIE

[761] Anne Dixon. Riding the iceberg: exploiting the Internet to promote one’s organisation. Learned Publishing, 10(1):21–24, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gedye:1997:UWW

[762] Richard Gedye. Using the World-Wide Web as a marketing tool. Learned Publishing, 10(1):25–32, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1997:RPL

[763] John St. Aubyn. Reflections on photocopying, licensing and new technology. Learned Publishing, 10(1):33–36, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1997:TJS

[764] Bernard Donovan. A tribute: John St. Aubyn. Learned Publishing, 10(1): 36, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:1997:CIM

[765] Charles Oppenheim. Copyright issues: Moral rights and the electronic library. Learned Publishing, 10(1):37–39, January 1997. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:1997:CI

[766] Charles Oppenheim. Copyright on Internet. Learned Publishing, 10(1): 39–41, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:ASRa

[767] Anonymous. ALPSP Seminar reports. Learned Publishing, 10(1):43–50, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:RREa

[768] Anonymous. Reviews of recent events. Learned Publishing, 10(1):51–59, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:EI

[769] Anonymous. Editorial issues. Learned Publishing, 10(1):60–66, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

StAubyn:1997:BR

[770] John St. Aubyn, Diana Rosenberg, Neville Goodman, and Alex Williamson. Book reviews. Learned Publishing, 10(1):67–70, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:DDa

[771] Anonymous. Diary dates. Learned Publishing, 10(1):71, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:HLF

[772] Anonymous. House of Lords: Fifth Report from the Select Committee on Science and Technology: Information Society: Agenda for Action in the UK. Learned Publishing, 10(1):73–96, January 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thorn:1997:AAE

[773] Sue Thorn. ALPSP Affairs: The Electronic Developments Committee. Learned Publishing, 10(2):98–100, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pullinger:1997:QLJ

[774] David Pullinger. Quality in on-line journals. Learned Publishing, 10(2): 101–108, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Charkin:1997:SCW

[775] Richard Charkin. Scholarly communities on the World Wide Web. Learned Publishing, 10(2):109–112, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clarke:1997:MPR

[776] Malcolm Clarke. Marketing to the Pacific Rim: opportunities and obstacles. Learned Publishing, 10(2):113–118, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davison:1997:AL

[777] Peter Davison. Is anybody listening? Learned Publishing, 10(2):119– 121, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Harwood:1997:MJP

[778] Paul Harwood. The market for journals in the Pacific Rim — what subscription agents are learning. Learned Publishing, 10(2):123–125, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hemingway:1997:SJP

[779] Clive Hemingway. Successful journal publishing on the Internet: hit or myth? Learned Publishing, 10(2):127–131, April 1997. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baron:1997:SFW

[780] Joel H. Baron. Special feature: Why We Need Information Identifiers. Learned Publishing, 10(2):132–134, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paskin:1997:II

[781] Norman Paskin. Information Identifiers. Learned Publishing, 10(2):135– 156, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wood:1997:PUE

[782] Dee Wood. Project update: Electronic submission and peer review — an update on the ESPERE project. Learned Publishing, 10(2):157–160, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:1997:CIC

[783] Charles Oppenheim. Copyright issues: Cause for concern? Copyright battles: The Shetlands. Learned Publishing, 10(2):161–164, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:1997:ASR

[784] Andrea Powell. ALPSP Seminar reports: Keeping Publishers Informed. Learned Publishing, 10(2):165–168, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:CGA

[785] Anonymous. The Charlesworth Group Award for electronic journals: Report of the Panel of Judges for the Charlesworth Group Award for Electronic Journals. Learned Publishing, 10(2):169–170, April 1997. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:EJI

[786] Anonymous. Electronic Journal Intermediaries: what can they offer you? Learned Publishing, 10(2):171–173, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:RREb

[787] Anonymous. Reviews of recent events. Learned Publishing, 10(2):175– 178, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Roberts:1997:EIS

[788] Rosemary Roberts. Editorial issues: Searching the New Dictionary of National Biography. Learned Publishing, 10(2):179–184, April 1997. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1997:BR

[789] Hazel K. Bell and Sue Deeley. Book reviews. Learned Publishing, 10 (2):185–187, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:DDb

[790] Anonymous. Diary dates. Learned Publishing, 10(2):189, April 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:AA

[791] Anonymous. ALPSP Affairs. Learned Publishing, 10(3):194–197, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Naylor:1997:WLW

[792] Bernard Naylor. What librarians want: transplanting yesterday into tomorrow. Learned Publishing, 10(3):199–204, July 1997. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brown:1997:ESI

[793] David J. Brown. Exorcising the spectre of interlibrary loans. Learned Publishing, 10(3):207–219, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Richardson:1997:SPB

[794] Martin J. Richardson. A system for the publication of biomedical journals in multiple formats. Learned Publishing, 10(3):221–225, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mathieson:1997:SCL

[795] Keith Mathieson. Steering clear of libel. Learned Publishing, 10(3):227– 230, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott:1997:JF

[796] Elspeth J. Scott. Journey to the future. Learned Publishing, 10(3):231– 235, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kuiper:1997:EPB

[797] Jan Kuiper. Environmental paper buying — best practice. Learned Publishing, 10(3):237–242, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1997:JPT

[798] Michele Benjamin. Journal publishing — is there a future? Learned Publishing, 10(3):243–245, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Semmens:1997:EAC

[799] Clive Semmens. Electronic archive — a contradiction in terms? Learned Publishing, 10(3):247–248, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prinsley:1997:CII

[800] Mark Prinsley. Copyright issues: Internationally harmonised copyright of electronic publishing: conspicuous by its absence. Learned Publishing, 10(3):249–254, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gadd:1997:CCD

[801] Elizabeth Gadd. Copyright clearance for the digital library: a practical guide. Learned Publishing, 10(3):255–259, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic). This paper was first published in Serials the journal of the United Kingdom Serials Group, Volume 10 No. 1, March 1997.

Anonymous:1997:ASRb

[802] Anonymous. ALPSP Seminar Reports. Learned Publishing, 10(3):261– 269, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:RREc

[803] Anonymous. Reviews of recent events. Learned Publishing, 10(3):271– 278, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:NBa

[804] Anonymous. News in brief. Learned Publishing, 10(3):279–280, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:IP

[805] Anonymous. Intermediary profiles. Learned Publishing, 10(3):281–284, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:DDc

[806] Anonymous. Diary dates. Learned Publishing, 10(3):286–287, July 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott:1997:AAF

[807] Peter Scott. ALPSP Affairs: The future of “Learning’ in a knowledge society. Learned Publishing, 10(4):290–299, October 1997. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:ARF

[808] Anonymous. ALPSP reports: The future of ‘learning’ in a knowledge society. Learned Publishing, 10(4):301–303, October 1997. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sykes:1997:DPH

[809] Phil Sykes. On-demand publishing in the humanities: project. Learned Publishing, 10(4):305–311, October 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hartley:1997:IAU

[810] James Hartley. Is it appropriate to use structured abstracts in social science journals? Learned Publishing, 10(4):313–317, October 1997. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Carter:1997:SLC

[811] Penny Carter. Site licence concept: a view of the UK Pilot Site Licence Initiative. Learned Publishing, 10(4):319–322, October 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kingston:1997:PAU

[812] Paula Kingston, Elizabeth Gadd, and Alan Poulter. Project ACORN: user reactions. Learned Publishing, 10(4):323–330, October 1997. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prior:1997:NSC

[813] Albert Prior. Normal service continues — the role of intermediaries in electronic publishing. Learned Publishing, 10(4):331–338, October 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Peters:1997:EAS

[814] Stuart Peters and Nigel Gilbert. The electronic alternative: Sociological Research Online. Learned Publishing, 10(4):339–343, October 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Owen:1997:OIS

[815] Tim Owen. Official information: scholarly publishers can help the Government get it right. Learned Publishing, 10(4):345–349, October 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1997:RAS

[816] Hazel K. Bell. Related affairs: SFEP: from isolation to professionalization. Learned Publishing, 10(4):351–353, October 1997. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Holden:1997:SP

[817] Bob Holden. Services to publishers. Learned Publishing, 10(4):355–357, October 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:NBb

[818] Anonymous. News in brief. Learned Publishing, 10(4):358–361, October 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Storrie:1997:BR

[819] Eileen Storrie and Hazel K. Bell. Book reviews. Learned Publishing, 10 (4):362–363, October 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1997:DDd

[820] Anonymous. Diary dates. Learned Publishing, 10(4):364, October 1997. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Storrie:1998:AAS

[821] Eileen Storrie. ALPSP Affairs: Strategic Review of the Association. Learned Publishing, 11(1):2–7, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Penfold:1998:ROE

[822] David Penfold. Requirements for an Online electronic journal service. Learned Publishing, 11(1):9–16, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Byrne:1998:OYA

[823] James A. Byrne. Optimizing your association’s revenue through Internet advertising. Learned Publishing, 11(1):17–22, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pearce:1998:AMU

[824] Anthony Pearce. The academic market in the UK. Learned Publishing, 11(1):23–31, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bol:1998:OJM

[825] Jennifer L. Bol. Online journal marketing strategies. Learned Publishing, 11(1):33–37, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bentley:1998:IDE

[826] J. Scott Bentley. The Image Directory: an electronic publishing project cataloguing art images. Learned Publishing, 11(1):39–41, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davenport:1998:AFE

[827] Anne Davenport. Agents — the future of exclusivity. Learned Publishing, 11(1):43–47, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davison:1998:JMP

[828] Peter Davison. Japan: a marketing profile. Learned Publishing, 11(1):49– 53, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ishibashi:1998:SPJ

[829] Mark Ishibashi. Services to publishers: The Japanese Market: an intermediary’s point of view. Learned Publishing, 11(1):54–56, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:1998:CI

[830] Charles Oppenheim. Copyright issues. Learned Publishing, 11(1):57–58, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:ASRa

[831] Anonymous. ALPSP Seminar reports: Electronic Information Intermediaries — the future. Learned Publishing, 11(1):59–65, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:DDa

[832] Anonymous. Diary dates. Learned Publishing, 11(1):66, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:RREa

[833] Anonymous. Reviews of recent events. Learned Publishing, 11(1):67–74, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:NBa

[834] Anonymous. News in brief. Learned Publishing, 11(1):75–76, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hirschmann:1998:BR

[835] P. N. Hirschmann and Mary Fox. Book reviews. Learned Publishing, 11 (1):77–80, January 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mellors-Bourne:1998:AA

[836] Robin Mellors-Bourne. ALPSP Affairs. Learned Publishing, 11(2):86– 87, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zielinski:1998:LE

[837] Christopher Zielinski. Letter to the editor. Learned Publishing, 11(2): 88–89, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1998:LSE

[838] Bernard Donovan. Learned societies and electronic publishing. Learned Publishing, 11(2):91–107, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bide:1998:CMD

[839] Mark Bide, Charles Oppenheim, and Anne Ramsden. Charging mechanisms for digitised texts. Learned Publishing, 11(2):109–118, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Balkwill:1998:DLR

[840] Richard Balkwill. Digital licensing — a role for the Publishers Licensing Society. Learned Publishing, 11(2):119–122, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yamakawa:1998:SPJ

[841] Takashi Yamakawa. STM Publications in Japan — a market analysis. Learned Publishing, 11(2):123–126, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kemp:1998:YPA

[842] Alasdair Kemp. Year 2000 problems: act now. Learned Publishing, 11 (2):127–131, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:CIS

[843] Anonymous. Copyright issues: Some answers to copyright queries. Learned Publishing, 11(2):133–135, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fowler:1998:PIS

[844] John Fowler. Points of interest: The Sage of Litchfield and authors’ rights. Learned Publishing, 11(2):136, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:ASRb

[845] Anonymous. ALPSP Seminar reports: Electronic Publishing Strategies — Publishing and the Internet IV. Learned Publishing, 11(2):137–144, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:RREb

[846] Anonymous. Reviews of recent events. Learned Publishing, 11(2):145– 151, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:NBb

[847] Anonymous. News in brief. Learned Publishing, 11(2):152–154, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:1998:BR

[848] Alex Williamson. Book reviews. Learned Publishing, 11(2):155–157, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:DDb

[849] Anonymous. Diary dates. Learned Publishing, 11(2):160–161, April 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:AAH

[850] Anonymous. ALPSP Affairs: Handing on the Standard. Learned Publishing, 11(3):166–168, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gadd:1998:LE

[851] Elizabeth Gadd. Letter to the editor. Learned Publishing, 11(3):169– 170, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brown:1998:SJP

[852] David J. Brown. Scholarly journal publishing: coming to terms with the Internet culture. Learned Publishing, 11(3):171–178, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:1998:TAP

[853] Bernard Donovan. The truth about peer review. Learned Publishing, 11(3):179–184, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

vanRooyen:1998:CEP

[854] Susan van Rooyen. A critical examination of the peer review process. Learned Publishing, 11(3):185–191, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wood:1998:OPR

[855] Dee Wood. Online peer review? Learned Publishing, 11(3):193–198, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:1998:CLC

[856] Sally Morris. Copyright and licensing — the changing scene. Learned Publishing, 11(3):199–204, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hodson:1998:DJF

[857] Richard Hodson. The demand for journals — fact versus fiction. Learned Publishing, 11(3):205–208, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:CI

[858] Anonymous. Copyright issues. Learned Publishing, 11(3):209–212, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:PI

[859] Anonymous. Points of interest. Learned Publishing, 11(3):213–215, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:ASRc

[860] Anonymous. ALPSP Seminar reports. Learned Publishing, 11(3):217– 227, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:RREc

[861] Anonymous. Reviews of recent events. Learned Publishing, 11(3):229– 234, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:NBc

[862] Anonymous. News in brief. Learned Publishing, 11(3):235–236, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:BR

[863] Anonymous. Book reviews. Learned Publishing, 11(3):238–239, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:DDc

[864] Anonymous. Diary dates. Learned Publishing, 11(3):242–243, July 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:AAG

[865] Anonymous. ALPSP Affairs: Guidelines for licence of electronic publications. Learned Publishing, 11(4):246–252, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:1998:LJC

[866] Sally Morris. Learned journals and the communication of research. Learned Publishing, 11(4):253–258, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Barrow:1998:DIS

[867] Edward Barrow. Digitisation: issues and solutions. Learned Publishing, 11(4):259–263, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lewis:1998:HER

[868] Gregory S. Lewis and John W. Edwards. Home on the electronic range: bringing the Journal of Animal Science online. Learned Publishing, 11 (4):265–270, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Owen:1998:RIP

[869] Lynette Owen. Rights issues and the printed word. Learned Publishing, 11(4):271–275, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Edmondson:1998:VRB

[870] John Edmondson. Virtual realities in Brobdingnag and Lilliput: gulfs and bridges between large and small publishers. Learned Publishing, 11 (4):277–282, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Harnad:1998:LIN

[871] Stevan Harnad. Learned inquiry and the Net: the role of peer review, peer commentary and copyright. Learned Publishing, 11(4):283–292, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smith:1998:PIF

[872] F. J. Smith. Points of interest: the first case of copyright infringement, AD 560. Learned Publishing, 11(4):293–296, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lawrey:1998:GCM

[873] Keith Lawrey. A guide to Council Members’ liabilities. Learned Publishing, 11(4):297–299, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davison:1998:NBA

[874] Peter Davison. News in Brief from Asia. Learned Publishing, 11(4):300, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:ASRd

[875] Anonymous. ALPSP Seminar reports. Learned Publishing, 11(4):301– 306, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:AR

[876] Anonymous. ALPSP report. Learned Publishing, 11(4):308–309, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:RREd

[877] Anonymous. Reviews of recent events. Learned Publishing, 11(4):311– 314, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:DDd

[878] Anonymous. Diary dates. Learned Publishing, 11(4):316–317, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1998:NBd

[879] Anonymous. News in brief. Learned Publishing, 11(4):318–320, October 1998. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:AA

[880] Anonymous. ALPSP Affairs. Learned Publishing, 12(1):2–4, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prior:1999:NPT

[881] Albert Prior. NESLI — progress through collaboration. Learned Publishing, 12(1):5–9, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:1999:SIS

[882] Anthony Watkinson. The STM information system: an analysis. Learned Publishing, 12(1):11–24, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Treloar:1999:RLR

[883] Andrew Treloar. Rethinking the library’s role in publishing. Learned Publishing, 12(1):25–31, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ben:1999:WLB

[884] Gu Ben. Western language books in the Beijing Library: criteria for acquisition and the method of selection. Learned Publishing, 12(1):33– 40, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thom:1999:IPN

[885] Rosie Thom. Institute of Publishing: News of progress. Learned Publishing, 12(1):41, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:CI

[886] Anonymous. Copyright issues. Learned Publishing, 12(1):43–49, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fowler:1999:PI

[887] John Fowler. Points of interest. Learned Publishing, 12(1):50, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:1999:ASR

[888] Michele Benjamin. ALPSP Seminar report. Learned Publishing, 12(1): 51–56, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:RREa

[889] Anonymous. Reviews of recent events. Learned Publishing, 12(1):59–62, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:1999:PVM

[890] Hazel K. Bell. Personal viewpoint: More books, fewer journals wanted? Learned Publishing, 12(1):63–64, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:NB

[891] Anonymous. News in brief. Learned Publishing, 12(1):65–69, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:DDa

[892] Anonymous. Diary dates. Learned Publishing, 12(1):70–71, January 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Swan:1999:WAWa

[893] Alma Swan and Sheridan Brown. ‘What authors want’: Report of the ALPSP research study on the motivations and concerns of contributors to learned journals. Learned Publishing, 12(2):74, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bammel:1999:GPA

[894] Jens Bammel. Guidelines: The publisher–author relationship: Principles of good practice in scholarly journal publishing. Learned Publishing, 12 (2):75–78, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smith:1999:DJN

[895] John W. T. Smith. The deconstructed journal — a new model for academic publishing. Learned Publishing, 12(2):79–91, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:1999:JPC

[896] Anthony Watkinson. Journal publishing: current issues. Learned Publishing, 12(2):93–96, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:1999:OTM

[897] Alan Singleton. Opportunities, threats and myths in journals. Learned Publishing, 12(2):97–105, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fletcher:1999:DIA

[898] Lloyd Alan Fletcher. Developing an integrated approach to electronic publishing: tailoring your content for the Web. Learned Publishing, 12 (2):107–117, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kaye:1999:WDD

[899] Laurie Kaye. What are databases and do they matter? Learned Publishing, 12(2):119–126, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Arjoon:1999:EMM

[900] Roger Arjoon. (In)Efficient market models: the reality behind economic models in the publishing industry. Learned Publishing, 12(2):127–133, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thom:1999:IPU

[901] Rosie Thom. Institute of Publishing: an update on the Institute of Publishing. Learned Publishing, 12(2):134, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prinsley:1999:CIC

[902] Mark Prinsley. Copyright issues: E-commerce — summary of European Commission proposals for a Directive. Learned Publishing, 12(2):135– 136, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Adler:1999:JP

[903] Kraig Adler and Wally Olsen. Journals pricing. Learned Publishing, 12 (2):137–139, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Campbell:1999:EWI

[904] Ruth E. Campbell. The EURO — will it effect me? Learned Publishing, 12(2):141–143, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dorner:1999:IR

[905] Jane Dorner. The IMPRIMATUR report. Learned Publishing, 12(2):145– 148, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Heald:1999:ASR

[906] Monique Heald. ALPSP Seminar report. Learned Publishing, 12(2):149– 151, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:RREb

[907] Anonymous. Reviews of recent events. Learned Publishing, 12(2):153– 161, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:DDb

[908] Anonymous. Diary dates. Learned Publishing, 12(2):164–165, April 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Swan:1999:WAWb

[909] Alma Swan. ‘What authors want’: the ALPSP research study on the motivations and concerns of contributors to learned journals. Learned Publishing, 12(3):170–172, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:1999:PLR

[910] John Cox. Publisher — library relationships in the digital environment. Learned Publishing, 12(3):173–178, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clark:1999:WWT

[911] Charles Clark. In what are we trading? Authors’ rights and publishers’ rights in traditional and digital media. Learned Publishing, 12(3):179– 190, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morgan:1999:JMI

[912] Cliff Morgan. Journals metadata: information about content. Learned Publishing, 12(3):191–195, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Grace:1999:FRB

[913] Mike Grace. Finding the right balance. Learned Publishing, 12(3):197– 200, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kidd:1999:MIJ

[914] Tony Kidd. Marketing institutional journal subscriptions to academic libraries. Learned Publishing, 12(3):201–207, July 1999. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowland:1999:EPN

[915] Fytton Rowland. Electronic publishing: non-commercial alternatives. Learned Publishing, 12(3):209–216, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mapp:1999:HPI

[916] Emma Mapp. The HERON Project to be introduced at HEI test sites this summer. Learned Publishing, 12(3):218, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lodge:1999:IIP

[917] Essex Lodge. IP Institute of Publishing. Learned Publishing, 12(3): 219, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Turner:1999:AGR

[918] Rollo Turner. ASA guidelines: The role and responsibilities of subscription agents. Learned Publishing, 12(3):222–224, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:MSA

[919] Anonymous. Merging subscription agents: Does it matter to publishers? Learned Publishing, 12(3):225, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:ICE

[920] Anonymous. International Committee on EDI for Serials (ICEDIS). Learned Publishing, 12(3):226–229, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Arunachalam:1999:LE

[921] Subbiah Arunachalam, John Davies, and Sally Morris. Letters to the Editor. Learned Publishing, 12(3):230–231, July 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:1999:LE

[922] Edwin Shelock. Letter to the Editor. Learned Publishing, 12(4):234, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Orlans:1999:FUU

[923] Harold Orlans. Fair use in US scholarly publishing. Learned Publishing, 12(4):235–244, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dunlop:1999:FDC

[924] Vivienne Dunlop. Fair dealing for criticism and review in scholarly publishing. Learned Publishing, 12(4):245–250, October 1999. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

King:1999:EJC

[925] Donald W. King and Carol Tenopir. Evolving journal costs: implications for publishers, libraries, and readers. Learned Publishing, 12(4):251–258, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wilson:1999:BRA

[926] David Gordon Wilson. A bill of rights for authors? Learned Publishing, 12(4):259–263, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Haynes:1999:NJP

[927] John Haynes. New Journal of Physics: a web-based and author-funded journal. Learned Publishing, 12(4):265–269, October 1999. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Russell:1999:PDS

[928] Kelly Russell, Ellis Weinberger, and Andy Stone. Preserving digital scholarship: the future is now. Learned Publishing, 12(4):271–280, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Desai:1999:UCL

[929] Kiran Desai and Mark Prinsley. UK competition law. Learned Publishing, 12(4):281–282, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Carr:1999:NES

[930] Reg Carr. National Electronic Site Licence Initiative. Learned Publishing, 12(4):283, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Turner:1999:RTJ

[931] Rollo Turner. Review of Trends in Journal Subscriptions — 1997. Learned Publishing, 12(4):285–286, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:IIP

[932] Anonymous. IP Institute of Publishing. Learned Publishing, 12(4):287, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:1999:DDc

[933] Anonymous. Diary dates. Learned Publishing, 12(4):290–292, October 1999. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:2000:E

[934] Michele Benjamin. Editorial. Learned Publishing, 13(1):2–3, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Doyle:2000:PCL

[935] Mark Doyle. Pragmatic citing and linking in electronic scholarly publishing. Learned Publishing, 13(1):5–14, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lonsdale:2000:ESM

[936] R. E. Lonsdale and C. J. Armstrong. Electronic scholarly monographs: issues and challenges for the UK. Learned Publishing, 13(1):15–24, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ball:2000:LPC

[937] David Ball and Jo Pye. Library purchasing consortia: the UK periodicals supply market. Learned Publishing, 13(1):25–35, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Berry:2000:FOA

[938] R. Stephen Berry. ‘Full and open access’ to scientific information: an academic’s view. Learned Publishing, 13(1):37–42, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smith:2000:JOP

[939] Arthur P. Smith. The journal as an overlay on preprint databases. Learned Publishing, 13(1):43–48, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Walford:2000:RAE

[940] Leo Walford. The Research Assessment Exercise: its effect on scholarly journal publishing. Learned Publishing, 13(1):49–52, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baggaley:2000:NST

[941] Richard Baggaley. No such thing as a mature journal. Learned Publishing, 13(1):53–57, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Harrison:2000:CSA

[942] Colin Harrison. Consolidation of subscription agents: a personal viewpoint. Learned Publishing, 13(1):60–62, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bammel:2000:WCE

[943] Jens Bammel. The WIPO conference on electronic commerce and intellectual property — report. Learned Publishing, 13(1):64–65, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Graham:2000:LE

[944] Peter S. Graham and John Cox. Letters to the Editor. Learned Publishing, 13(1):66–67, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Palmer:2000:R

[945] Richard Palmer and Adrian Corrigan. Reviews. Learned Publishing, 13 (1):68–69, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2000:DDa

[946] Anonymous. Diary dates. Learned Publishing, 13(1):70–71, January 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2000:LP

[947] Anonymous. Licence to Publish. Learned Publishing, 13(2):74, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2000:AC

[948] Sally Morris. Authors and copyright. Learned Publishing, 13(2):75– 76, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hugenholtz:2000:CVF

[949] P. Bernt Hugenholtz. Copyright vs. freedom of scientific communication. Learned Publishing, 13(2):77–81, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Klein:2000:WSP

[950] Donald C. Klein. Web strategies for professional publishers: developing an information services portal. Learned Publishing, 13(2):83–94, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wood:2000:OPR

[951] Dee Wood and Phil Hurst. Online peer review: perceptions in the biological sciences. Learned Publishing, 13(2):95–100, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wise:2000:AGP

[952] Alicia Wise, Chris Rusbridge, and Norman Wiseman. ATHENS, gateways, and portals: components of the Distributed National Electronic Resource. Learned Publishing, 13(2):101–104, April 2000. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

ONeill:2000:SIF

[953] Jill O’Neill. The significance of an impact factor: implications for the publishing community. Learned Publishing, 13(2):105–109, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clayphan:2000:BPL

[954] Robina Clayphan. The BIBLINK project: linking publishers and national bibliographic agencies. Learned Publishing, 13(2):111–117, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ross:2000:NWI

[955] Jan Ross. A new way of information retrieval: 3-D indexing and concept mapping. Learned Publishing, 13(2):119–123, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:2000:SPC

[956] Hazel Bell. Scientific publishing, 17th-century style: precepts of Francis Bacon. Learned Publishing, 13(2):124–126, April 2000. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thorn:2000:NLJ

[957] Sue Thorn. New life for journals or the death of the learned society? Learned Publishing, 13(2):127–128, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:2000:GJP

[958] Bernard Donovan. A glimpse of journal publishing in China. Learned Publishing, 13(2):128–129, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:2000:RLL

[959] Hazel Bell and Kathryn Toledano. Reviews, letters and listings. Learned Publishing, 13(2):130–136, April 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Doyle:2000:ST

[960] Mark Doyle. Surviving the transition. Learned Publishing, 13(3):138– 140, July 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Law:2000:UAC

[961] D. G. Law, R. L. Weedon, and M. R. Sheen. Universities and article copyright. Learned Publishing, 13(3):141–150, July 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McKinty:2000:SFW

[962] K. McKinty, S. Saadat, F. J. Smith, and B. T. Donovan. Setting up a first website for society publications. Learned Publishing, 13(3):151– 157, July 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paskin:2000:CAI

[963] Norman Paskin. E-citations: actionable identifiers and scholarly referencing. Learned Publishing, 13(3):159–166, July 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Chu:2000:PCE

[964] Heting Chu. Promises and challenges of electronic journals: academic libraries surveyed. Learned Publishing, 13(3):169–175, July 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brown:2000:WCT

[965] Iain D. Brown. What can technology offer? Notes on technical developments for the non-technical. Learned Publishing, 13(3):177–181, July 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williams:2000:PV

[966] Ronnie Williams. Personal View. Learned Publishing, 13(3):183–190, July 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morgan:2000:BR

[967] Cliff Morgan, Lesley Gray, Adam Marshall, Michael Mabe, and Susan Hezlet. Book reviews. Learned Publishing, 13(3):191–196, July 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jones:2000:LE

[968] Hugh Jones, Albert Henderson, and Hazel Bell. Letters to the Editor. Learned Publishing, 13(3):197–201, July 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Heward-Mills:2000:LSO

[969] Leon Heward-Mills. From learned society to online community. Learned Publishing, 13(4):202–203, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bull:2000:WDS

[970] David Bull and Susan Hezlet. What do societies and publishers want from publishing partnerships? Learned Publishing, 13(4):205–207, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Page:2000:PJT

[971] Gillian Page. Putting journals out to tender: guidelines for societies and other sponsors. Learned Publishing, 13(4):209–220, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dryburgh:2000:HEY

[972] Alastair Dryburgh. How efficient is your publishing operation? Learned Publishing, 13(4):221–227, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baldwin:2000:WRV

[973] Christine Baldwin and David Pullinger. What readers value in academic journals. Learned Publishing, 13(4):229–239, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Goldfinch:2000:LSO

[974] E. P. Goldfinch. A limited study of online access and user preferences. Learned Publishing, 13(4):241–245, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Leggate:2000:AOJ

[975] Peter Leggate and Mike Hannant. The archiving of online journals. Learned Publishing, 13(4):246–250, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2000:DCE

[976] Anonymous. Defining and Certifying Electronic Publication in Science. Learned Publishing, 13(4):251–258, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2000:WR

[977] Anonymous. What is an RRO? Learned Publishing, 13(4):259–260, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Braid:2000:R

[978] Andrew Braid, Stella Pilling, Alastair Dryburgh, and Judy Watkins. Reviews. Learned Publishing, 13(4):261–264, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williams:2000:LE

[979] Penny Williams, Albert Henderson, and Sinisa Maricic. Letters to the Editor. Learned Publishing, 13(4):265–266, October 2000. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2000:DDb

[980] Anonymous. Diary dates. Learned Publishing, 13(4):267–268, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2000:LPV

[981] Anonymous. Learned Publishing vol. 13, 2000. Learned Publishing, 13 (4):269–272, October 2000. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lynch:2001:PLE

[982] Clifford Lynch. The power of the linked environment. Learned Publishing, 14(1):3–5, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Till:2001:PPS

[983] James E. Till. Predecessors of preprint servers. Learned Publishing, 14 (1):7–13, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson:2001:MJH

[984] Kent Anderson. The mutant journal: how adaptations to online forces are forcing STM journals to mutate. Learned Publishing, 14(1):15–22, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smith:2001:LUN

[985] Kent A. Smith and Ed Sequeira. Linking at the US National Library of Medicine. Learned Publishing, 14(1):23–28, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wittenberg:2001:EPI

[986] Kate Wittenberg. The Electronic Publishing Initiative at Columbia (EPIC): a university-based collaboration in digital scholarly communication. Learned Publishing, 14(1):29–32, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tomlins:2001:JOM

[987] Christopher L. Tomlins. Just one more ’zine? Maintaining and improving the scholarly journal in the electronic present: a view from the humanities. Learned Publishing, 14(1):33–40, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Canhos:2001:BPH

[988] Vanderlei Canhos, Leslie Chan, and Barbara Kirsop. Bioline Publications: how its evolution has mirrored the growth of the Internet. Learned Publishing, 14(1):41–48, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kessler:2001:PWE

[989] Elisabeth Kessler and Zhang Keyu. Publishing a well-established international environmental journal in Chinese. Learned Publishing, 14(1): 49–53, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholson:2001:NBM

[990] David Nicholson. New business models: AgBiotechNet^{T M}, a case study. Learned Publishing, 14(1):55–59, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Grace:2001:PML

[991] Mike Grace. Publishing: a market-led approach. Learned Publishing, 14 (1):60–64, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Warnick:2001:PHP

[992] Walter L. Warnick. PubSCIENCE: hyperlinking to the physical sciences. Learned Publishing, 14(1):65–67, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Owen:2001:P

[993] Lynette Owen. Piracy. Learned Publishing, 14(1):67–70, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Albert:2001:GPA

[994] Tim Albert. Getting personal about training and development. Learned Publishing, 14(1):71–72, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:2001:TIL

[995] Hazel Bell. Telling it like it is: reporting conferences and seminars. Learned Publishing, 14(1):72–74, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mabe:2001:BR

[996] Michael Mabe. Book reviews. Learned Publishing, 14(1):75–76, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2001:AN

[997] Anonymous. ALPSP News. Learned Publishing, 14(1):77, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2001:DDa

[998] Anonymous. Diary dates. Learned Publishing, 14(1):78–79, January 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McLellan:2001:PEO

[999] Faith McLellan. Publication ethics: an ounce of prevention. Learned Publishing, 14(2):83–84, April 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

vanRooyen:2001:EPR

[1000] Susan van Rooyen. The evaluation of peer-review quality. Learned Publishing, 14(2):85–91, April 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hardy:2001:PPM

[1001] Rachel Hardy, Charles Oppenheim, and Iris Rubbert. PELICAN: a pricing mechanism for the electronic distribution of materials to the higher education community. Learned Publishing, 14(2):93–96, April 2001. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Greco:2001:MUP

[1002] Albert N. Greco. The market for university press books in the United States: 1985–1999. Learned Publishing, 14(2):97–105, April 2001. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Taylor:2001:NMT

[1003] David Taylor. Never mind the textbook — here’s the chapters. Learned Publishing, 14(2):107–112, April 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shashok:2001:AEF

[1004] Karen Shashok. Author’s editors: facilitators of science information transfer. Learned Publishing, 14(2):113–121, April 2001. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:2001:HHD

[1005] Hazel K. Bell. From herbals to Hotbot: the development of journal indexing. Learned Publishing, 14(2):123–130, April 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Langer:2001:PNE

[1006] James Langer. Physicists in the new era of electronic publishing. Learned Publishing, 14(2):131–137, April 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kubota:2001:HEJ

[1007] Teruzo Kubota. How are electronic journals and CD-ROMs being accepted in Japan? Learned Publishing, 14(2):139–143, April 2001. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Slowinski:2001:HFD

[1008] F. Hill Slowinski and Patrick Bernuth. How ‘free distribution’ impacts your business model: is it really free? Learned Publishing, 14(2):144– 148, April 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:2001:CCP

[1009] Alex Williamson. COPE: Committee on Publication Ethics. Learned Publishing, 14(2):149–150, April 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wood:2001:R

[1010] Dee Wood. Reviews. Learned Publishing, 14(2):151–158, April 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2001:DDb

[1011] Anonymous. Diary dates. Learned Publishing, 14(2):159–160, April 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2001:WSA

[1012] Sally Morris. What’s so special about not-for-profit publishers? Learned Publishing, 14(3):163–165, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Houghton:2001:CTE

[1013] John W. Houghton. Crisis and transition: the economics of scholarly communication. Learned Publishing, 14(3):167–176, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rzepa:2001:NPP

[1014] Henry S. Rzepa and Peter Murray-Rust. A new publishing paradigm: STM articles as part of the Semantic Web. Learned Publishing, 14(3): 177–182, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jones:2001:LIC

[1015] Simon Jones. The links in the information chain. Learned Publishing, 14(3):183–188, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fowler:2001:PRP

[1016] John Fowler. The publications of the Royal Pharmaceutical Society. Learned Publishing, 14(3):189–196, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Klugkist:2001:VNV

[1017] Alex C. Klugkist. Virtual and non-virtual realities: the changing roles of libraries and librarians. Learned Publishing, 14(3):197–204, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:2001:NDC

[1018] Charles Oppenheim. The new Directive on copyright and related rights. Learned Publishing, 14(3):205–212, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fletcher:2001:GBB

[1019] Lloyd Alan Fletcher. Going beyond the buzzword: what exactly is CRM? Learned Publishing, 14(3):213–222, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kasdorf:2001:XR

[1020] Bill Kasdorf. The XML revolution. Learned Publishing, 14(3):223–231, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cowhig:2001:EAJ

[1021] Jerry Cowhig. Electronic article and journal usage statistics (EAJUS): proposal for an industry-wide standard. Learned Publishing, 14(3):233– 236, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wood:2001:BR

[1022] Dee Wood and Michael Mabe. Book reviews. Learned Publishing, 14(3): 237–238, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Law:2001:UAC

[1023] D. G. Law, M. R. Sheen, and R. L. Weedon. Universities and article copyright. Learned Publishing, 14(3):239–240, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2001:DDc

[1024] Anonymous. Diary dates. Learned Publishing, 14(3):240–241, July 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bennett:2001:EPN

[1025] Liz Bennett. Electronic publishing in the new millennium. Learned Publishing, 14(4):243–246, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hames:2001:EBR

[1026] Irene Hames. Editorial boards: realizing their potential. Learned Publishing, 14(4):247–256, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wager:2001:SPR

[1027] Elizabeth Wager and Tom Jefferson. Shortcomings of peer review in biomedical journals. Learned Publishing, 14(4):257–263, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kircz:2001:NPE

[1028] Joost G. Kircz. New practices for electronic publishing 1: Will the scientific paper keep its form? Learned Publishing, 14(4):265–272, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shum:2001:JIJ

[1029] Simon Buckingham Shum and Tamara Sumner. JIME: an interactive journal for interactive media. Learned Publishing, 14(4):273–285, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Inger:2001:IA

[1030] Simon Inger. The importance of aggregators. Learned Publishing, 14 (4):287–290, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Evans:2001:WAD

[1031] Michael Evans. Who’s afraid of the Data Protection Act? Is the Data Protection Act 1998 a problem for journal publishers? Learned Publishing, 14(4):291–295, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Walker:2001:WS

[1032] Jenny Walker. What is SFX? Learned Publishing, 14(4):296–298, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shelock:2001:ALS

[1033] Edwin Shelock. ALPSP and learned society publishing. Learned Publishing, 14(4):299–302, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Albert:2001:WNP

[1034] Tim Albert. Who needs personal development? Learned Publishing, 14 (4):302–303, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Harris:2001:HBN

[1035] Lesley Ellen Harris. How to be a better negotiator in the digital realm. Learned Publishing, 14(4):303–306, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2001:BR

[1036] Anthony Watkinson, Judy Watkins, Alastair Dryburgh, Hazel Bell, and Susan Hezlet. Book reviews. Learned Publishing, 14(4):307–311, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2001:LE

[1037] Anthony Watkinson. Letter to the Editor. Learned Publishing, 14(4): 313–314, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2001:DDd

[1038] Anonymous. Diary dates. Learned Publishing, 14(4):315–316, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2001:LPV

[1039] Anonymous. Learned Publishing vol. 14, 2001. Learned Publishing, 14 (4):317–320, October 2001. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fox:2002:AEP

[1040] Peter Fox. Archiving of electronic publications — some thoughts on cost. Learned Publishing, 15(1):3–5, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Odlyzko:2002:RES

[1041] Andrew Odlyzko. The rapid evolution of scholarly communication. Learned Publishing, 15(1):7–19, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic). See response [1066].

Pullinger:2002:ILD

[1042] David Pullinger. Instant linking — delayed use: setting provider expectations. Learned Publishing, 15(1):21–25, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kircz:2002:NPE

[1043] Joost G. Kircz. New practices for electronic publishing 2: New forms of the scientific paper. Learned Publishing, 15(1):27–32, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kaser:2002:GIA

[1044] Richard T. Kaser. Getting it! The added value of helping users find information. Learned Publishing, 15(1):33–42, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Goodman:2002:YPP

[1045] David Goodman. A year without print at Princeton, and what we plan next. Learned Publishing, 15(1):43–50, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rosenberg:2002:AJO

[1046] Diana Rosenberg. African Journals Online: improving awareness and access. Learned Publishing, 15(1):51–57, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yuehong:2002:JZU

[1047] Zhang Yuehong (Helen). Journal of Zhejiang University (SCIENCE): a new journal for the 21st century. Learned Publishing, 15(1):59–62, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dorner:2002:AD

[1048] Jane Dorner. Authors and e-delivery. Learned Publishing, 15(1):63–68, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Naylor:2002:SDM

[1049] Bernard Naylor. Size does matter. Learned Publishing, 15(1):69–70, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fraser:2002:BR

[1050] Ron Fraser, Michael Mabe, and Anthony Watkinson. Book reviews. Learned Publishing, 15(1):71–74, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Welham:2002:WSA

[1051] Robert Welham and Sue Thorn. What’s so special about not-for-profit publishers? Learned Publishing, 15(1):75–77, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2002:DDa

[1052] Anonymous. Diary dates. Learned Publishing, 15(1):79–80, January 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowland:2002:WDU

[1053] Fytton Rowland. What do users want? Learned Publishing, 15(2):83– 84, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson:2002:UA

[1054] Kent Anderson. The useful archive. Learned Publishing, 15(2):85–89, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Silver:2002:PSK

[1055] Keith Silver. Pressing the ‘send’ key — preferential journal access in developing countries. Learned Publishing, 15(2):91–98, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Breaks:2002:BHL

[1056] Michael Breaks. Building the hybrid library: a review of UK activities. Learned Publishing, 15(2):99–107, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Berin:2002:UJT

[1057] Andrew Berin. Unbundled journals: trying to predict the future. Learned Publishing, 15(2):109–112, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dryburgh:2002:TNS

[1058] Alastair Dryburgh. There is no such thing as a product. Learned Publishing, 15(2):113–116, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Crawford:2002:FER

[1059] Walt Crawford. Free electronic refereed journals: getting past the arc of enthusiasm. Learned Publishing, 15(2):117–123, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McElroy:2002:DEJ

[1060] Emily McElroy. Dos and don’ts for electronic journal management: some advice to publishers. Learned Publishing, 15(2):125–128, April 2002. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Montgomery:2002:PEM

[1061] Carol Hansen Montgomery. Print to electronic: measuring the operational and economic implications of an electronic journal collection. Learned Publishing, 15(2):129–136, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wiseman:2002:DCS

[1062] Leanne Wiseman. Digital copying and the statutory licences in Australian universities. Learned Publishing, 15(2):137–147, April 2002. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yuehong:2002:ELA

[1063] Zhang Yuehong (Helen), Wang Sen, and Li Jiang. The English-language academic journals of China: trends and developments. Learned Publishing, 15(2):149–151, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Carim:2002:SKH

[1064] Lara Carim. Serial killers: how great is the e-print threat to periodicals publishers? Learned Publishing, 15(2):153–155, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tolson:2002:JTA

[1065] Jill Tolson. Jill Tolson appointed as Business Manager. Learned Publishing, 15(2):157, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hauptman:2002:RAO

[1066] Robert Hauptman. A response to Andrew Odlyzko. Learned Publishing, 15(2):158, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic). See [1041].

Anonymous:2002:DDb

[1067] Anonymous. Diary dates. Learned Publishing, 15(2):159–161, April 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:2002:VPO

[1068] John Cox. Valuing and protecting our intellectual property: the lifeblood of our business. Learned Publishing, 15(3):163–164, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baldwin:2002:DCU

[1069] Christine Baldwin. Delivering content to universities and colleges: JISC’s new information environment. Learned Publishing, 15(3):165–174, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cerdeira:2002:EDS

[1070] Hilda A. Cerdeira. eJournals Delivery Service: an email to Internet experiment. Learned Publishing, 15(3):175–178, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rusch-Feja:2002:OAI

[1071] Diann Rusch-Feja. The Open Archives Initiative and the OAI Protocol for Metadata Harvesting: rapidly forming a new tier in the scholarly communication infrastructure. Learned Publishing, 15(3):179–186, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sisler:2002:OUP

[1072] William P. Sisler. One university press. Learned Publishing, 15(3):187– 192, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Finkelstein:2002:SAP

[1073] David Finkelstein, Sarah Bromage, and Alistair McCleery. Scottish Archive of Print and Publishing History Records. Learned Publishing, 15(3):193–197, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ginn:2002:CPM

[1074] Claire Ginn. Calculating pricing models choices: rising to the challenge. Learned Publishing, 15(3):199–203, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

KHEC:2002:OCP

[1075] The Knight Higher Education Collaborative. Op. cit.: publishing in the humanities and social sciences. Learned Publishing, 15(3):205–216, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gedye:2002:NSR

[1076] Richard Gedye. Non-subscription revenue as a source of income. Learned Publishing, 15(3):217–220, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Worlock:2002:EJU

[1077] Kate Worlock. Electronic journals: user realities — the truth about content usage among the STM community. Learned Publishing, 15(3): 223–226, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tenopir:2002:BR

[1078] Carol Tenopir, Anthony Watkinson, Jill Membrey, and Andrea Powell. Book reviews. Learned Publishing, 15(3):227–230, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ewing:2002:DC

[1079] John H. Ewing. In defense of caution. Learned Publishing, 15(3):231– 233, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Richards:2002:IAM

[1080] Julian D. Richards. Internet Archaeology and the myth of free publication. Learned Publishing, 15(3):233–234, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Albert:2002:WDY

[1081] Tim Albert. A writer’s diary in the year 2015. Learned Publishing, 15(3):234–235, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:2002:NSR

[1082] Andrea Powell. Non-subscription revenue as a source of income. Learned Publishing, 15(3):238, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morgan:2002:FER

[1083] Cliff Morgan and Walt Crawford. Free electronic refereed journals. Learned Publishing, 15(3):238–239, July 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Taylor:2002:OJM

[1084] Ian Taylor. Opening the journals market in China. Learned Publishing, 15 (4):243–245, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowland:2002:PRP

[1085] Fytton Rowland. The peer-review process. Learned Publishing, 15(4): 247–258, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tenopir:2002:RBE

[1086] Carol Tenopir and Donald W. King. Reading behaviour and electronic journals. Learned Publishing, 15(4):259–265, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:2002:LFT

[1087] Andrea Powell. Linking to full text: the secondary publisher’s perspective. Learned Publishing, 15(4):267–272, October 2002. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:2002:PPR

[1088] Brian Cox. The Pergamon phenomenon 1951–1991: Robert Maxwell and scientific publishing. Learned Publishing, 15(4):273–278, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Walker:2002:TSS

[1089] Thomas J. Walker. Two societies show how to profit by providing free access. Learned Publishing, 15(4):279–284, October 2002. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Harwood:2002:YWL

[1090] Paul Harwood. You are the weakest link — goodbye: serving the information-hungry corporate end-user. Learned Publishing, 15(4):285– 290, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Albert:2002:MJP

[1091] Tim Albert and Alex Williamson. Medical journal publishing: one culture or several? Learned Publishing, 15(4):291–296, October 2002. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bawden:2002:PLD

[1092] David Bawden and Lyn Robinson. Promoting literacy in a digital age: approaches to training for information literacy. Learned Publishing, 15 (4):297–301, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kenneway:2002:INJ

[1093] Melinda Kenneway, Pam Sutherland, and Samuel C. Williamson. Introducing a new journals subscription system: the agony and the ecstasy. Learned Publishing, 15(4):302–306, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2002:BR

[1094] Anthony Watkinson, J. D. Lee, Anthony Watkinson, and Kurt Paulus. Book reviews. Learned Publishing, 15(4):307–312, October 2002. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jones:2002:JPO

[1095] Ruth Jones. Journals e-publishing: outsourced solutions for professional, scholarly and society publishers. Learned Publishing, 15(4):313–314, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Doyle:2002:HPP

[1096] Mark Doyle. How to profit by providing free access. Learned Publishing, 15(4):315, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2002:CC

[1097] Anonymous. Copyright corner. Learned Publishing, 15(4):317, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Benjamin:2002:EA

[1098] Michele Benjamin. Editorial acknowledgement. Learned Publishing, 15 (4):318, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2002:LPO

[1099] Anonymous. Learned Publishing online — pricing for 2003. Learned Publishing, 15(4):319, October 2002. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:2003:GE

[1100] Andrea Powell. Guest editorial. Learned Publishing, 16(1):3–4, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Welham:2003:E

[1101] Robert Welham. Editorial. Learned Publishing, 16(1):5, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Waltham:2003:CRP

[1102] Mary Waltham. Challenges to the role of publishers. Learned Publishing, 16(1):7–14, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:2003:WWH

[1103] Alex Williamson. What will happen to peer review? Learned Publishing, 16(1):15–20, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lagnado:2003:PWA

[1104] Max Lagnado. Professional writing assistance: effects on biomedical publishing. Learned Publishing, 16(1):21–27, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Swan:2003:AEP

[1105] Alma Swan and Sheridan Brown. Authors and electronic publishing: what authors want from the new technology. Learned Publishing, 16 (1):28–33, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Klein:2003:LNO

[1106] Donald Klein and Mark Ware. E-learning: new opportunities in continuing professional development. Learned Publishing, 16(1):34–46, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kohl:2003:CLV

[1107] David Kohl. Consortial licensing vs. tradition: breaking up is hard to do. Learned Publishing, 16(1):47–53, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pearce:2003:EAS

[1108] Carole Pearce. Editing an African scholarly journal. Learned Publishing, 16(1):54–60, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gilchrist:2003:TRO

[1109] Alan Gilchrist. Text retrieval: an overview. Learned Publishing, 16(1): 61–69, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Taylor:2003:BAM

[1110] David Taylor. E-books and the academic market: the emerging supply chain. Learned Publishing, 16(1):70–73, January 2003. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dryburgh:2003:OAJ

[1111] Alastair Dryburgh. Open-access journals — nice idea, shame about the numbers? Learned Publishing, 16(1):75–76, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Batchelor:2003:BR

[1112] Matthew Batchelor, Jane Moody, and Matthew Batchelor. Book reviews. Learned Publishing, 16(1):77–80, January 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2003:OS

[1113] Sally Morris. Open sesame? Learned Publishing, 16(2):83–84, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watts:2003:DSE

[1114] Louise Watts. Document supply: the evolving needs of the library. Learned Publishing, 16(2):85–90, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yuehong:2003:IPR

[1115] Zhang Yuehong (Helen), Yuan Yachun, and Jiang Yufei. An international peer-review system for a Chinese scientific journal. Learned Publishing, 16(2):91–94, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dryburgh:2003:NFD

[1116] Alastair Dryburgh. A new framework for digital publishing decisions. Learned Publishing, 16(2):95–101, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bide:2003:CN

[1117] Mark Bide. Copyright and the network. Learned Publishing, 16(2):103– 109, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Steele:2003:PRN

[1118] Colin Steele. Phoenix rising: new models for the research monograph? Learned Publishing, 16(2):111–122, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Frankel:2003:SMS

[1119] Mark S. Frankel. Seizing the moment: scientists’ authorship rights in the digital age. Learned Publishing, 16(2):123–128, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Upshall:2003:CMJ

[1120] Michael Upshall. Content management for journal publishers. Learned Publishing, 16(2):129–133, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Joseph:2003:BBS

[1121] Heather Dalterio Joseph. BioOne: building a sustainable alternative publishing model for non-profit publishers. Learned Publishing, 16(2): 134–138, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Turner:2003:AFT

[1122] Alan Turner and Pippa Steele. Archiving — the full-text solution: an innovative way to electronic archiving. Learned Publishing, 16(2):139– 142, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2003:JDC

[1123] Pippa Smart. E-journals: developing country access survey. Learned Publishing, 16(2):143–148, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:2003:VDE

[1124] David J. Powell. Voluntary deposit of electronic publications: a learning experience. Learned Publishing, 16(2):149–152, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Friend:2003:BDG

[1125] Frederick J. Friend. Big Deal — good deal? Or is there a better deal? Learned Publishing, 16(2):153–155, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Newton:2003:BR

[1126] Rachel Newton, Anthony Watkinson, and Dee Wood. Book reviews. Learned Publishing, 16(2):156–158, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2003:APa

[1127] Anonymous. ALPSP Projects. Learned Publishing, 16(2):159, April 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prosser:2003:HTP

[1128] David C. Prosser. From here to there: a proposed mechanism for transforming journals from closed to open access. Learned Publishing, 16(3): 163–166, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Velterop:2003:SSS

[1129] Jan Velterop. Should scholarly societies embrace open access (or is it the kiss of death)? Learned Publishing, 16(3):167–169, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2003:OP

[1130] Sally Morris. Open publishing. Learned Publishing, 16(3):171–176, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hellman:2003:OML

[1131] Eric Hellman. OpenURL: making the link to libraries. Learned Publishing, 16(3):177–181, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Savenije:2003:FPN

[1132] Bas Savenije. The FIGARO Project: a new approach towards academic publishing. Learned Publishing, 16(3):183–188, July 2003. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2003:TPD

[1133] Alan Singleton. Terrorism, the publishing decision and beyond. Learned Publishing, 16(3):189–192, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ha:2003:ESJ

[1134] Louisa Ha. The economics of scholarly journals: a case study on a societypublished journal. Learned Publishing, 16(3):193–199, July 2003. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shepherd:2003:CCC

[1135] Peter T. Shepherd. COUNTER: from conception to compliance. Learned Publishing, 16(3):201–205, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Connertz:2003:LTA

[1136] Thomas Connertz. Long-term archiving of digital documents: what efforts are being made in Germany? Learned Publishing, 16(3):207–211, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kahn:2003:PC

[1137] Deborah Kahn. Perceptions of the customer. Learned Publishing, 16(3): 213–217, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bell:2003:WYT

[1138] Hazel K. Bell. Won’t you tell me when? Learned Publishing, 16(3): 218, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hartley:2003:PBR

[1139] James Hartley. On the presentation of book reviews. Learned Publishing, 16(3):219–220, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2003:APb

[1140] Anonymous. ALPSP Projects. Learned Publishing, 16(3):221, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2003:ER

[1141] Alan Singleton. Essay review. Learned Publishing, 16(3):222–225, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Richardson:2003:BR

[1142] Peter Richardson, Anthony Watkinson, Louise Watts, Christine Reid, Chris Shelley, Nigel Lees, Andrea Powell, Martin Richardson, Anthony Watkinson, Anthony Watkinson, Sophie Leighton, and Andrea Powell. Book reviews. Learned Publishing, 16(3):226–238, July 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davison:2003:RRS

[1143] Peter Davison. Reprographic rights in Singapore: the story of CLASS. Learned Publishing, 16(4):243–247, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wets:2003:PPF

[1144] Kathleen Wets, Dave Weedon, and Jan Velterop. Post-publication filtering and evaluation: Faculty of 1000. Learned Publishing, 16(4):249–258, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Adomi:2003:PFJ

[1145] Esharenana E. Adomi and Chinedum Mordi. Publication in foreign journals and promotion of academics in Nigeria. Learned Publishing, 16(4): 259–263, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dryburgh:2003:AFA

[1146] Alastair Dryburgh. Alternative futures for academic and professional publishing. Learned Publishing, 16(4):265–270, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lindquist:2003:OTR

[1147] Mats G. Lindquist. Organization and technology for the recording of science. Learned Publishing, 16(4):271–276, October 2003. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Beckwith:2003:GQV

[1148] Kimberly Beckwith. Googled: the quest for visibility on the Internet. Learned Publishing, 16(4):277–283, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tananbaum:2003:WBS

[1149] Greg Tananbaum. Of wolves and and boys: the scholarly communication crisis. Learned Publishing, 16(4):285–289, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Harvey:2003:WDZ

[1150] Judith Harvey. What does ZWOLLE stand for? Learned Publishing, 16 (4):290–292, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gadd:2003:RSA

[1151] Elizabeth Gadd, Charles Oppenheim, and Steve Probets. RoMEO studies 4: an analysis of journal publishers’ copyright agreements. Learned Publishing, 16(4):293–308, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dunton:2003:CEE

[1152] Simon Dunton and Audrey D. Melkin. Are copy editors an endangered species? Learned Publishing, 16(4):309–315, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:2003:LE

[1153] Kurt Paulus. Letter to the Editor. Learned Publishing, 16(4):316, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Robson:2003:BR

[1154] Jodie Robson and Cathleen Blackburn. Book reviews. Learned Publishing, 16(4):317–318, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2003:ACA

[1155] Anonymous. ALPSP/Charlesworth and ALPSP Awards 2003. Learned Publishing, 16(4):319, October 2003. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Marks:2004:ESC

[1156] Jayne Marks and Timo Hannay. Evolving scholarly communication. Learned Publishing, 17(1):3–6, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Beagrie:2004:DCC

[1157] Neil Beagrie. The Digital Curation Centre. Learned Publishing, 17(1):7– 9, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Roosendaal:2004:DCR

[1158] Hans E. Roosendaal. Driving change in the research and HE information market. Learned Publishing, 17(1):11–16, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prosser:2004:BRH

[1159] David C. Prosser. Between a rock and a hard place: the big squeeze for small publishers. Learned Publishing, 17(1):17–22, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Green:2004:PBO

[1160] Toby Green. Publishing e-books: OECD’s pay-per-view and e-library services 1998–2003. Learned Publishing, 17(1):25–30, January 2004. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morgan:2004:MSJ

[1161] Cliff Morgan. Metadata for STM journal publishers: a review of the current scene. Learned Publishing, 17(1):31–37, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Higgins:2004:PDB

[1162] Suzanne Wilson Higgins. Print on-demand is boosting availability of books and journal issues — and that’s good for sales! Learned Publishing, 17(1):39–42, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Erzin:2004:ETE

[1163] Pinar Erzin. The 2003 Extenza telemarketing effectiveness survey. Learned Publishing, 17(1):43–45, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bonald:2004:ALJ

[1164] Laura Bonald and Sally Morris. The ALPSP Learned Journals Collection. Learned Publishing, 17(1):47–51, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bence:2004:RAJ

[1165] Valerie Bence and Charles Oppenheim. The role of academic journal publications in the UK Research Assessment Exercise. Learned Publishing, 17(1):53–68, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic). See erratum [1170] that supplies a missing table.

Dryburgh:2004:OAT

[1166] Alastair Dryburgh. Open access — time to stop preaching to the converted? Learned Publishing, 17(2):69–70, January 2004. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2004:CCG

[1167] Alan Singleton. Consorting with consortia: the good, the bad and the …. Learned Publishing, 17(2):71–75, January 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Peters:2004:BR

[1168] John Peters, Kimberly M. Beckwith, and Alastair Dryburgh. Book reviews. Learned Publishing, 17(2):76–78, January 2004. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stevenson:2004:OAN

[1169] Iain Stevenson. Is open access the new vanity publishing? Learned Publishing, 17(2):83–84, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2004:E

[1170] Anonymous. Erratum. Learned Publishing, 17(2):84, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic). See [1165].

Cronin:2004:MMM

[1171] Blaise Cronin and Kathryn La Barre. Mickey Mouse and Milton: book publishing in the humanities. Learned Publishing, 17(1):85–98, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ren:2004:RCE

[1172] Shengli Ren and Ronald Rousseau. The role of China’s English-language scientific journals in scientific communication. Learned Publishing, 17(1): 99–104, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Poschl:2004:IJC

[1173] Ulrich Pöschl. Interactive journal concept for improved scientific publishing and quality assurance. Learned Publishing, 17(1):105–113, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:2004:IRS

[1174] Mark Ware. Institutional repositories and scholarly publishing. Learned Publishing, 17(1):115–124, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cooke:2004:HRC

[1175] Helen Cooke. A historical review of the chemistry periodical literature until 1950. Learned Publishing, 17(1):125–134, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Banks:2004:OAM

[1176] Peter Banks. Open access: a medical association perspective. Learned Publishing, 17(1):135–142, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lamb:2004:OAP

[1177] Christine Lamb. Open access publishing models: opportunity or threat to scholarly and academic publishers? Learned Publishing, 17(1):143– 150, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Awre:2004:JFP

[1178] Chris Awre. The JISC’s FAIR Programme: disclosing and sharing institutional assets. Learned Publishing, 17(1):151–156, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davies:2004:SCT

[1179] J. Eric Davies and Helen Greenwood. Scholarly communication trends — voices from the vortex: a summary of specialist opinion. Learned Publishing, 17(1):157–167, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gregory:2004:OI

[1180] Peter Gregory. Open to interpretation. Learned Publishing, 17(1):169– 170, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2004:AN

[1181] Anonymous. ALPSP News. Learned Publishing, 17(1):171, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Halsall:2004:BR

[1182] Janet Halsall, Pippa Smart, and Anthony Watkinson. Book reviews. Learned Publishing, 17(1):172–174, April 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:2004:WCW

[1183] John Cox. What can we learn from the RoweCom débâcle? Learned Publishing, 17(3):179–182, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Allen:2004:TST

[1184] Anne Vindenes Allen and Graham Vaughan Lees. TheScientificWorld-JOURNAL: sustainability through innovation. Learned Publishing, 17 (3):183–187, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lees:2004:LPS

[1185] Nigel Lees. Learned and professional society libraries and copyright. Learned Publishing, 17(3):189–194, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2004:DPP

[1186] Alan Singleton. Data protection and peer review. Learned Publishing, 17(3):195–198, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hedlund:2004:OAS

[1187] Turid Hedlund, Tomas Gustafsson, and Bo-Christer Björk. The open access scientific journal: an empirical study. Learned Publishing, 17(3): 199–209, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Akie:2004:DSS

[1188] Ronald Akie, Joel Baron, and Gahl Berkooz. The demand side of STM publishing: understanding value creation in new market units. Learned Publishing, 17(3):211–218, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Swan:2004:AOA

[1189] Alma Swan and Sheridan Brown. Authors and open access publishing. Learned Publishing, 17(3):219–224, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brand:2004:CRE

[1190] Amy Brand. CrossRef and the research experience. Learned Publishing, 17(3):225–230, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Houghton:2004:RPS

[1191] John W. Houghton, Colin Steele, and Margaret Henty. Research practices and scholarly communication in the digital environment. Learned Publishing, 17(3):231–249, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:2004:ILJ

[1192] Kurt Paulus. 2004 International Learned Journals Seminar. Learned Publishing, 17(3):250–253, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Toledano:2004:BR

[1193] Kathryn Toledano and Anthony Watkinson. Book reviews. Learned Publishing, 17(3):254, July 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rix:2004:CWC

[1194] Tim Rix. Crisis? What crisis? The university presses. Learned Publishing, 17(4):259–260, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowlands:2004:SCD

[1195] Ian Rowlands, Dave Nicholas, and Paul Huntington. Scholarly communication in the digital environment: what do authors want? Learned Publishing, 17(4):261–273, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bellekom:2004:BPF

[1196] Chris Bellekom. Building preservation functionality in a digital archive: the National Library of The Netherlands. Learned Publishing, 17(4): 275–280, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jackson:2004:IPM

[1197] Paul Jackson. IEE professional magazines and networks. Learned Publishing, 17(4):281–289, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Worlock:2004:SSP

[1198] Kate Worlock. Scoping the STM publishing market. Learned Publishing, 17(4):291–298, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rapple:2004:AGG

[1199] Charlie Rapple. After the goldrush — the golden age of reference linking. Learned Publishing, 17(4):299–304, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pinfield:2004:WDU

[1200] Stephen Pinfield. What do universities want from publishing? Learned Publishing, 17(4):305–312, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson:2004:CPO

[1201] Kent R. Anderson. Comparing print and online readership: matching perception to reality across media. Learned Publishing, 17(4):313–315, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Huhn:2004:RJU

[1202] Volker Hühn and Michael Huter. The Red Jacket UTB: a model for co-operative textbook publishing. Learned Publishing, 17(4):316–318, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sosteric:2004:ICA

[1203] Michael Sosteric. The International Consortium for the Advancement of Academic Publication — an idea whose time has come (finally!). Learned Publishing, 17(4):319–325, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shuhua:2004:CSC

[1204] Wang Shuhua and Wang Hengjun. Challenges and strategies for Chinese university journals. Learned Publishing, 17(4):326–330, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2004:STC

[1205] Sally Morris. Science & Technology Committee report: the ALPSP response. Learned Publishing, 17(4):331, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2004:AAC

[1206] Anonymous. ALPSP and ALPSP/Charlesworth Awards 2004. Learned Publishing, 17(4):332–334, October 2004. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:2005:AS

[1207] Andrea Powell. ALPSP and SSP. Learned Publishing, 18(1):2, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Long:2005:SMN

[1208] Maurice Long. Springtime meeting in New York. Learned Publishing, 18 (1):3, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cousins:2005:IOH

[1209] Jill Cousins and Eamonn Neylon. Information objects are hot, documents are not: the use of identifiers in online publishing. Learned Publishing, 18(1):5–12, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Goodman:2005:OAW

[1210] David Goodman. Open access: what comes next? Learned Publishing, 18(1):13–24, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Swan:2005:DMP

[1211] Alma Swan, Paul Needham, Steve Probets, Adrienne Muir, Charles Oppenheim, Ann O’Brien, Rachel Hardy, Fytton Rowland, and Sheridan Brown. Developing a model for e-prints and open access journal content in UK further and higher education. Learned Publishing, 18(1):25–40, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yu:2005:STB

[1212] Pauline Yu. On synthetic technologies: the book, the university, the Internet. Learned Publishing, 18(1):41–50, January 2005. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hunter:2005:CID

[1213] Karen Hunter. Critical issues in the development of STM journal publishing. Learned Publishing, 18(1):51–55, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tempest:2005:EJT

[1214] David Tempest. The effect of journal title changes on impact factors. Learned Publishing, 18(1):57–62, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hemmings:2005:PPQ

[1215] Brian Hemmings, Peter Rushbrook, and Erica Smith. To publish or not to publish: that is the question? Learned Publishing, 18(1):63–65, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anguelov:2005:SWW

[1216] Simeon Anguelov, Pierre Baruch, and Françoise Praderie. Spreading the word: who profits from science publishing? A symposium held at the EuroScience Open Forum. Learned Publishing, 18(1):67–74, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hartley:2005:OC

[1217] James Hartley. Down with ‘op. cit.’. Learned Publishing, 18(1):75–77, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2005:CP

[1218] Anonymous. Charles Parker. Learned Publishing, 18(1):78, January 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Joseph:2005:FEC

[1219] Heather Joseph. As far as the eye can see. Learned Publishing, 18(2): 83–84, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mp:2005:IRO

[1220] Evan Harris Mp. Institutional repositories: is the open access door half open or half shut? Learned Publishing, 18(2):85–89, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Elliott:2005:WOS

[1221] Roger Elliott. Who owns scientific data? The impact of intellectual property rights on the scientific publication chain. Learned Publishing, 18(2):91–94, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bosc:2005:PWN

[1222] Hélène Bosc and Stevan Harnad. In a paperless world a new role for academic libraries: providing open access. Learned Publishing, 18(2):95– 99, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Waltham:2005:OAI

[1223] Mary Waltham. Open access — the impact of legislative developments. Learned Publishing, 18(2):101–114, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2005:TCS

[1224] Sally Morris. The true costs of scholarly journal publishing. Learned Publishing, 18(2):115–126, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Simoni:2005:SSW

[1225] Robert D. Simoni. Serving science while paying the bills: the history of the Journal of Biological Chemistry Online. Learned Publishing, 18 (2):127–130, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sack:2005:HPT

[1226] John Sack. HighWire Press: ten years of publisher-driven innovation. Learned Publishing, 18(2):131–142, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Daniel:2005:PMS

[1227] Hans-Dieter Daniel. Publications as a measure of scientific advancement and of scientists’ productivity. Learned Publishing, 18(2):143–148, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Montonen:2005:EPP

[1228] Claus Montonen. The European physics publications scene: avant-garde and traditionalism. Learned Publishing, 18(2):149–151, April 2005. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davies:2005:TP

[1229] Gill Davies. Training for publishing. Learned Publishing, 18(2):152– 156, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2005:RM

[1230] Anonymous. Robert Millson. Learned Publishing, 18(2):157–158, April 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Corbett:2005:PP

[1231] Sue Corbett. Publish or perish? Learned Publishing, 18(3):163–164, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bjork:2005:LMS

[1232] Bo-Christer Björk. A lifecycle model of the scientific communication process. Learned Publishing, 18(3):165–176, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Warner:2005:TSC

[1233] Simeon Warner. The transformation of scholarly communication. Learned Publishing, 18(3):177–185, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Li:2005:AUJ

[1234] Li Li. Advantages of university journals in China. Learned Publishing, 18(3):188–192, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:2005:OJI

[1235] Mark Ware. E-only journals: is it time to drop print? Learned Publishing, 18(3):193–199, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nwagwu:2005:MLB

[1236] Williams Nwagwu. Mapping the landscape of biomedical research in Nigeria since 1967. Learned Publishing, 18(3):200–211, July 2005. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2005:TVO

[1237] David Nicholas, Hamid R. Jamali M., Paul Huntington, and Ian Rowlands. In their very own words: authors and scholarly journal publishing. Learned Publishing, 18(3):212–220, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Richardson:2005:PPA

[1238] Martin Richardson. Post-print archives: parasite or symbiont? Learned Publishing, 18(3):221–223, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2005:OAL

[1239] Alan Singleton. Open access and learned societies. Learned Publishing, 18(3):223–228, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Irving:2005:PIR

[1240] Lewis Irving. The pharmaceutical industry, researchers and publishers: the realities of competing interests within specialty journals. Learned Publishing, 18(3):228–230, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:2005:ILJ

[1241] Kurt Paulus. 2005 International Learned Journals Seminar: The Web and after. Learned Publishing, 18(3):231–234, July 2005. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Marks:2005:OAN

[1242] Robert H. Marks. Open access: a new name for an old concept. Learned Publishing, 18(3):235–236, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2005:BR

[1243] Anthony Watkinson. Book reviews. Learned Publishing, 18(3):237–238, July 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Goldwyn:2005:DWR

[1244] Robert M. Goldwyn. Do we reject too little and publish too much? Learned Publishing, 18(4):243, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:2005:OSP

[1245] Mark Ware. Online submission and peer-review systems. Learned Publishing, 18(4):245–250, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2005:BDT

[1246] David Nicholas, Paul Huntington, Bill Russell, Anthony Watkinson, Hamid R. Jamali, and Carol Tenopir. The Big Deal — ten years on. Learned Publishing, 18(4):251–257, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Arroyo:2005:PSO

[1247] Cristina Márquez Arroyo, Laura Munoa, Fernando A. Navarro, María Verónica Saladrigas, and Karen Shashok. Panace@ — a successful open access journal from the STM translation community. Learned Publishing, 18(4):258–269, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Liping:2005:COC

[1248] Wang Liping. Challenges and opportunities: China’s university presses in transition. Learned Publishing, 18(4):271–274, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yacoubian:2005:PAL

[1249] George S. Yacoubian, Jr. Publishing in American legal and social science periodicals: an ethical comparison. Learned Publishing, 18(4):275–278, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Inger:2005:PCM

[1250] Simon Inger. Production and content management implications for archival projects: a snapshot in May 2005. Learned Publishing, 18(4): 279–285, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shepherd:2005:CNC

[1251] Peter T. Shepherd. COUNTER 2005: a new Code of Practice and new applications of COUNTER usage statistics. Learned Publishing, 18(4): 287–293, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gannon:2005:OAS

[1252] Frank Gannon. Open access: scientists as paradoxical consumers. Learned Publishing, 18(4):295–299, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowland:2005:SJP

[1253] Fytton Rowland. Scholarly journal publishing in New Zealand. Learned Publishing, 18(4):300–310, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scholes:2005:WHL

[1254] Robert Scholes. What is happening in literary studies? Learned Publishing, 18(4):311–315, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2005:AAC

[1255] Anonymous. ALPSP and ALPSP/Charlesworth Awards 2005. Learned Publishing, 18(4):316–317, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Parker:2005:R

[1256] Robert Parker. Review. Learned Publishing, 18(4):318, October 2005. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:2006:RES

[1257] John Cox. Re-engineering the scholarly publishing process — lessons from elsewhere. Learned Publishing, 19(1):3–4, January 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Domine:2006:PRM

[1258] Tom Domine. Peer-review management systems: what do users want? Learned Publishing, 19(1):5–7, January 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xian:2006:ELA

[1259] Jia Xian. English-language academic journals from China: a great opportunity. Learned Publishing, 19(1):9–13, January 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Waltham:2006:LSB

[1260] Mary Waltham. Learned society business models and open access: overview of a recent JISC-funded study. Learned Publishing, 19(1):15– 30, January 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowlands:2006:CSC

[1261] Ian Rowlands and Dave Nicholas. The changing scholarly communication landscape: an international survey of senior researchers. Learned Publishing, 19(1):31–55, January 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Probets:2006:DIR

[1262] Steve Probets and Celia Jenkins. Documentation for institutional repositories. Learned Publishing, 19(1):57–71, January 2006. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2006:WJJ

[1263] Sally Morris. When is a journal not a journal? A closer look at the DOAJ. Learned Publishing, 19(1):73–76, January 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Prosser:2006:LE

[1264] David C. Prosser. Letter to the Editor. Learned Publishing, 19(1):78, January 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Olivieri:2006:CE

[1265] René Olivieri. Context is everything. Learned Publishing, 19(2):83–84, April 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Antelman:2006:SAP

[1266] Kristin Antelman. Self-archiving practice and the influence of publisher policies in the social sciences. Learned Publishing, 19(2):85–95, April 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shuhua:2006:CAJ

[1267] Wang Shuhua and Paul R. Weldon. Chinese academic journals: quality, issues and solutions. Learned Publishing, 19(2):97–105, April 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Babini:2006:UDL

[1268] Dominique Babini and Pippa Smart. Using digital libraries to provide online access to social science journals in Latin America. Learned Publishing, 19(2):107–113, April 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Markwood:2006:PWJ

[1269] Priscilla Markwood. Paperless workflows in journal production: a management perspective. Learned Publishing, 19(2):115–124, April 2006. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smothers:2006:MJP

[1270] Valerie Smothers, Michael Clarke, and Craig Van Dyck. MedBiquitous and journal publishers: scholarly content and online medical communities. Learned Publishing, 19(2):125–132, April 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xian:2006:PPF

[1271] Jia Xian. The past, present and future of scientific and technical journals of China. Learned Publishing, 19(2):133–141, April 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Chesler:2006:CIP

[1272] Adam Chesler. Core issues in pricing models for scholarly journals: a qualitative overview. Learned Publishing, 19(2):142–146, April 2006. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bjork:2006:BSJ

[1273] Bo-Christer Björk and Jonas Holmström. Benchmarking scientific journals from the submitting author’s viewpoint. Learned Publishing, 19(2): 147–155, April 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:2006:BR

[1274] Charles Oppenheim and Charles Oppenheim. Book reviews. Learned Publishing, 19(2):156–158, April 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rockwood:2006:NCA

[1275] Irving Rockwood. A new chapter for ALPSP. Learned Publishing, 19(3): 163–164, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ju-fang:2006:TPS

[1276] Shao Ju-fang and Shen Hui-yun. Thought patterns of some editors of academic journals. Learned Publishing, 19(3):165–167, July 2006. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Taylor:2006:PEP

[1277] Alasdair Taylor. Publishing and electronic piracy. Learned Publishing, 19(3):168–174, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Goldsworthy:2006:OSO

[1278] Sophie Goldsworthy. Oxford Scholarship Online. Learned Publishing, 19(3):175–181, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Regazzi:2006:PAP

[1279] John J. Regazzi and Nicole A. Caliguiri. Publisher and author partnerships: a changing landscape. Learned Publishing, 19(3):183–192, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2006:TTT

[1280] David Nicholas, Hamid R. Jamali, and Ian Rowlands. On the tips of their tongues: authors and their views on scholarly publishing. Learned Publishing, 19(3):193–203, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Li:2006:FCC

[1281] Zhang Li, Yao Yuan, Zhang Fenglian, and Du Wentao. The first comprehensive Chinese university journal published in English: the Tsing Hua Journal. Learned Publishing, 19(3):204–208, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson:2006:ATS

[1282] Kent Anderson. Ask and they shall deceive: how what we ask drives what they tell us, and why it’s mostly useless. Learned Publishing, 19(3):209– 218, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pinfield:2006:WCD

[1283] Stephen Pinfield. A Wel(l)come development: research funders and open access. Learned Publishing, 19(3):219–225, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:2006:OAT

[1284] Mark Ware. Open archives and their impact on journal cancellations. Learned Publishing, 19(3):226–229, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:2006:CWA

[1285] Kurt Paulus. Calmer waters ahead? Learned Publishing, 19(3):230– 233, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sampson:2006:DLJ

[1286] Geoffrey Sampson. The death of learned journals. Learned Publishing, 19(3):234–235, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wilmot:2006:BR

[1287] Charlotte Wilmot, Charles Oppenheim, and Lesley Robinson. Book reviews. Learned Publishing, 19(3):236–238, July 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:2006:WNC

[1288] Charles Oppenheim. Who needs copyright? Learned Publishing, 19 (4):243–244, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2006:ICC

[1289] David Nicholas, Paul Huntington, Tom Dobrowolski, and Ian Rowlands. Ideas on creating a consumer market for scholarly journals. Learned Publishing, 19(4):245–249, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pentz:2006:CC

[1290] Ed Pentz. CrossRef at the crossroads. Learned Publishing, 19(4):250–258, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Taylor:2006:PPP

[1291] Kevin Taylor. Plagiarism and piracy: a publisher’s perspective. Learned Publishing, 19(4):259–266, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sambunjak:2006:PRE

[1292] Dario Sambunjak. Press releases and email notices increase local and global visibility of a small medical journal. Learned Publishing, 19(4): 267–271, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:2006:SPP

[1293] John Cox. Scholarly publishing practices: a case of plus ça change, plus c’est la même chose? Learned Publishing, 19(4):273–276, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Steele:2006:PIP

[1294] Colin Steele, Linda Butler, and Danny Kingsley. The publishing imperative: the pervasive influence of publication metrics. Learned Publishing, 19(4):277–290, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Schroter:2006:FST

[1295] Sara Schroter, Leanne Tite, and Ahmed Kassem. Financial support at the time of paper acceptance: a survey of three medical journals. Learned Publishing, 19(4):291–297, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Puliselic:2006:IEC

[1296] Lea Pulišelić and Jelka Petrak. Is it enough to change the language? A case study of Croatian biomedical journals. Learned Publishing, 19 (4):299–306, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowland:2006:LE

[1297] Fytton Rowland. Letter to the Editor. Learned Publishing, 19(4):307– 308, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Johnson:2006:BR

[1298] Ian M. Johnson, Paula Younger, Louise Ellis-Barrett, and Charles Oppenheim. Book reviews. Learned Publishing, 19(4):309–315, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2006:AAC

[1299] Anonymous. ALPSP and ALPSP/Charlesworth Awards 2006. Learned Publishing, 19(4):316–317, October 2006. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Russell:2007:AGL

[1300] Ian Russell. ALPSP: a guiding light. Learned Publishing, 20(1):3–4, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2007:E

[1301] Sally Morris and Priscilla Markwood. Editorial. Learned Publishing, 20 (1):5, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:2007:TYL

[1302] Bernard Donovan and Eileen Donovan. Twenty years of Learned Publishing. Learned Publishing, 20(1):7–10, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2007:IOA

[1303] David Nicholas, Paul Huntington, and Hamid R. Jamali. The impact of open access publishing (and other access initiatives) on use and users of digital scholarly journals. Learned Publishing, 20(1):11–15, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Henneken:2007:PJA

[1304] Edwin A. Henneken, Michael J. Kurtz, Guenther Eichhorn, Alberto Accomazzi, Carolyn S. Grant, Donna Thompson, Elizabeth Bohlen, Stephen S. Murray, Paul Ginsparg, and Simeon Warner. E-prints and journal articles in astronomy: a productive co-existence. Learned Publishing, 20(1):16–22, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pesch:2007:SSD

[1305] Oliver Pesch. SUSHI: simplifying the delivery of usage statistics. Learned Publishing, 20(1):23–29, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wright:2007:RPN

[1306] Janette Wright and Sara Hearn. RMIT Publishing: neither fish nor fowl. Learned Publishing, 20(1):31–35, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lifang:2007:SPF

[1307] Xu Lifang, Fang Qing, and Lian Xiaochuan. Scholarly publishing funds in China. Learned Publishing, 20(1):36–42, January 2007. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stanley:2007:COC

[1308] Adrian Stanley and Yan Shuai. China opening up: Chinese university journals and research — today and tomorrow. Learned Publishing, 20 (1):43–50, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brunelle:2007:HAL

[1309] Bette Brunelle and Chuck Richard. The healthcare advertising landscape. Learned Publishing, 20(1):51–57, January 2007. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Suthersanen:2007:CCO

[1310] Uma Suthersanen. Creative Commons — the other way? Learned Publishing, 20(1):59–68, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Haank:2007:IDR

[1311] Derk Haank. Industry developments: the revolution continues. Learned Publishing, 20(1):69–71, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williamson:2007:SM

[1312] Alex Williamson. Sally Morris. Learned Publishing, 20(1):73, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Blake:2007:BR

[1313] Monica Blake, Charles Oppenheim, Charles Oppenheim, Sue Lambert, and Charles Oppenheim. Book reviews. Learned Publishing, 20(1):74– 78, January 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson:2007:OAC

[1314] Rick Anderson. Open access — clear benefits, hidden costs. Learned Publishing, 20(2):83–84, April 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

King:2007:CJP

[1315] Donald W. King. The cost of journal publishing: a literature review and commentary. Learned Publishing, 20(2):85–106, April 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ohta:2007:IOP

[1316] Teruto Ohta. An innovation-oriented publication system. Learned Publishing, 20(2):107–111, April 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williams:2007:CPP

[1317] Alan Williams. Copyright: past, present and future. Learned Publishing, 20(2):113–119, April 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wates:2007:AVV

[1318] Edward Wates and Robert Campbell. Author’s version vs. publisher’s version: an analysis of the copy-editing function. Learned Publishing, 20(2):121–129, April 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brown:2007:DAP

[1319] Richard J. C. Brown. Double anonymity in peer review within the chemistry periodicals community. Learned Publishing, 20(2):131–137, April 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kennan:2007:AAS

[1320] Mary Anne Kennan. Academic authors, scholarly publishing, and open access in Australia. Learned Publishing, 20(2):138–146, April 2007. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jubb:2007:RUP

[1321] Michael Jubb, Hugh Look, and Sue Sparks. Researchers’ use and perceptions of discovery services. Learned Publishing, 20(2):147–153, April 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Walford:2007:HD

[1322] Leo Walford. Humpty Dumpty. Learned Publishing, 20(2):154–155, April 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:2007:BR

[1323] Charles Oppenheim, Huw Alexander, Charles Oppenheim, and Cliff Morgan. Book reviews. Learned Publishing, 20(2):156–158, April 2007. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2007:WDW

[1324] Sally Morris. Why do we bother to copy-edit? Learned Publishing, 20(3):163–164, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thatcher:2007:COA

[1325] Sanford G. Thatcher. The challenge of open access for university presses. Learned Publishing, 20(3):165–172, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Carmi:2007:IPR

[1326] Ran Carmi and Christof Koch. Improving peer review with CARMA. Learned Publishing, 20(3):173–176, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:2007:JPS

[1327] Mark Ware. Journal publishing systems: outsource or in-house? Learned Publishing, 20(3):177–181, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thommen:2007:KPS

[1328] Moritz Thommen and Carla Holmes. Keeping personal subscriptions out of the library. Learned Publishing, 20(3):182–186, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson:2007:MJB

[1329] Kent R. Anderson and Jean Nugent. A medical journal’s beta site: experimentation at NEJM. Learned Publishing, 20(3):187–190, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Peters:2007:GAW

[1330] Paul Peters. Going all the way: how Hindawi became an open access publisher. Learned Publishing, 20(3):191–195, July 2007. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

DeVries:2007:SBS

[1331] Saskia C. J. De Vries. From sailing boat to steamship: the role of the publisher in an open access environment. Learned Publishing, 20(3):196– 201, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Goodman:2007:OAA

[1332] David Goodman, Sarah Dowson, and Jean Yaremchuk. Open access and accuracy: author-archived manuscripts vs. published articles. Learned Publishing, 20(3):203–215, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Look:2007:TAI

[1333] Hugh Look. Thinking about innovation. Learned Publishing, 20(3):216– 219, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:2007:PCC

[1334] John Cox. Plus ça change …. Learned Publishing, 20(3):220–222, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shashok:2007:HMO

[1335] Karen Shashok. How might open access become sustainable? Learned Publishing, 20(3):223–224, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:2007:WTY

[1336] Kurt Paulus. Are we there yet? The ALPSP International Scholarly Communications Conference London, 13 April 2007. Learned Publishing, 20(3):225–229, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Suber:2007:LE

[1337] Peter Suber. Letter to the Editor. Learned Publishing, 20(3):230–231, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bennett:2007:BR

[1338] Fiona Bennett, Charles Oppenheim, Stuart Taylor, Jane Moody, and Jennifer Papin-Ramcharan. Book reviews. Learned Publishing, 20(3): 233–238, July 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Markwood:2007:DYK

[1339] Priscilla Markwood. It’s 2007. Do you know where your learned journal is? Learned Publishing, 20(4):243–244, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pitts:2007:URF

[1340] Sinae Pitts and Adrian Stanley. Understanding researchers: findings from a focus group. Learned Publishing, 20(4):245–251, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kulathuramaiyer:2007:FPI

[1341] Narayanan Kulathuramaiyer and Hermann Maurer. Fighting plagiarism and IPR violation: why is it so important? Learned Publishing, 20 (4):252–258, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Armbruster:2007:MOL

[1342] Chris Armbruster. Moving out of Oldenbourg’s long shadow: what is the future for society publishing? Learned Publishing, 20(4):259–266, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baruch:2007:OAD

[1343] Pierre Baruch. Open access developments in France: the HAL Open Archives System. Learned Publishing, 20(4):267–282, October 2007. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stalker:2007:LOO

[1344] Matt Stalker. Launching an online-only journal. Learned Publishing, 20 (4):283–289, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lees:2007:DVL

[1345] Nigel Lees. Developing a Virtual Library for 44,000 chemists. Learned Publishing, 20(4):290–297, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2007:MJP

[1346] Sally Morris. Mapping the journal publishing landscape: how much do we know? Learned Publishing, 20(4):299–310, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2007:JWM

[1347] Pippa Smart. Journals — the wrong model for Africa? Learned Publishing, 20(4):311–313, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cecchini:2007:WCD

[1348] Ivan Cecchini. What competences do today’s publishers need? Learned Publishing, 20(4):314–315, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ballard-Wood:2007:MPB

[1349] Aime M. Ballard-Wood. In Memoriam: Peter Banks. Learned Publishing, 20(4):316, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2007:ACA

[1350] Anonymous. ALPSP/Charlesworth Awards 2007. Learned Publishing, 20 (4):317–318, October 2007. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2008:WQJ

[1351] Sally Morris. What is quality in journals publishing? Learned Publishing, 21(1):4–6, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Waltham:2008:WDS

[1352] Mary Waltham. What do society and association members really want? Learned Publishing, 21(1):7–14, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Derricourt:2008:FDM

[1353] Robin Derricourt. For a few dollars more: a future for scholarly books in Australia? Learned Publishing, 21(1):15–21, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:2008:CPP

[1354] Mark Ware. Choosing a publishing partner: advice for societies and associations. Learned Publishing, 21(1):22–28, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Borgman:2008:DDS

[1355] Christine L. Borgman. Data, disciplines, and scholarly publishing. Learned Publishing, 21(1):29–38, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Penny:2008:PTW

[1356] Dan Penny. Publishing technologies: what does the future hold? Learned Publishing, 21(1):39–47, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baker:2008:VSU

[1357] Gayle Baker and Eleanor J. Read. Vendor-supplied usage data for electronic resources: a survey of academic libraries. Learned Publishing, 21 (1):48–57, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lafaye:2008:IWB

[1358] Vanessa Lafaye. Innovation and the Wiley — Blackwell Compass Journals. Learned Publishing, 21(1):58–62, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Strickland:2008:IRA

[1359] Peter R. Strickland, Brian Mcmahon, and John R. Helliwell. Integrating research articles and supporting data in crystallography. Learned Publishing, 21(1):63–72, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:2008:AR

[1360] Stephen K. Donovan. On accuracy in references. Learned Publishing, 21(1):74–75, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Welham:2008:BR

[1361] Robert Welham. Book Review. Learned Publishing, 21(1):77, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shepherd:2008:LE

[1362] Peter Shepherd. Letter to the Editors. Learned Publishing, 21(1):78, January 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Markwood:2008:NMW

[1363] Priscilla Markwood. The NIH mandate — we’re not in Kansas any more. Learned Publishing, 21(2):83–84, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pringle:2008:TUI

[1364] James Pringle. Trends in the use of ISI citation databases for evaluation. Learned Publishing, 21(2):85–91, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Koskinen-Olsson:2008:AKD

[1365] Tarja Koskinen-Olsson. Access to knowledge in the digital era. Learned Publishing, 21(2):93–102, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jabri:2008:DLA

[1366] Evelyn Jabri. A day in the life of an academic chemist. Learned Publishing, 21(2):105–109, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rapple:2008:KBI

[1367] Charlie Rapple. Knowledge bases: improving the information supply chain. Learned Publishing, 21(2):110–115, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Perciali:2008:JBN

[1368] Irene Perciali and Aaron Edlin. Journals at bepress: new twists on an old model. Learned Publishing, 21(2):116–122, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sandewall:2008:ECP

[1369] Erik Sandewall. Extending the concept of publication: factbases and knowledgebases. Learned Publishing, 21(2):123–131, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Harwood:2008:TUB

[1370] Paul Harwood and Albert Prior. Testing usage-based e-journal pricing. Learned Publishing, 21(2):133–139, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cheng:2008:EOA

[1371] Weihong Cheng and Shengli Ren. Evolution of open access publishing in Chinese scientific journals. Learned Publishing, 21(2):140–152, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tananbaum:2008:AOD

[1372] Greg Tananbaum. Adventures in open data. Learned Publishing, 21 (2):154–156, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:2008:BRa

[1373] Charles Oppenheim. Book Review. Learned Publishing, 21(2):158, April 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2008:TC

[1374] Sally Morris. The tiger in the corner. Learned Publishing, 21(3):163– 165, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Calow:2008:ASM

[1375] Duncan Calow and Rebecca Egan. Is the answer still in the machine: do publishers need digital rights management? Learned Publishing, 21(3): 167–175, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bley:2008:LSE

[1376] Robert Bley. Library systems in the electronic era. Learned Publishing, 21(3):176–186, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Vandooren:2008:GNL

[1377] Françoise Vandooren and Cécile Gass. Giving new life to out-of-print books: when publishers’ and libraries’ interests meet. Learned Publishing, 21(3):187–192, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bucknell:2008:USB

[1378] Terry Bucknell. Usage statistics for Big Deals: supporting library decision-making. Learned Publishing, 21(3):193–199, July 2008. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bird:2008:OJA

[1379] Claire Bird. Oxford Journals’ adventures in open access. Learned Publishing, 21(3):200–208, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Noonan:2008:ECT

[1380] Bridget M. Noonan and Debra Parrish. Expressions of concern and their uses. Learned Publishing, 21(3):209–213, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Creaser:2008:TJP

[1381] Claire Creaser and Sonya White. Trends in journal prices: an analysis of selected journals, 2000–2006. Learned Publishing, 21(3):214–224, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Regazzi:2008:APJ

[1382] John J. Regazzi and Selenay Aytac. Author perceptions of journal quality. Learned Publishing, 21(3):225–235, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Peters:2008:BR

[1383] Paul Peters. Book reviews. Learned Publishing, 21(3):237, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:2008:BRb

[1384] Charles Oppenheim. Book reviews. Learned Publishing, 21(3):238, July 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fisher:2008:WDO

[1385] Janet Fisher. What do our readers want? Results of the 2008 reader survey. Learned Publishing, 21(4):244–246, October 2008. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

King:2008:POM

[1386] Donald W. King and Frances M. Alvarado-Albertorio. Pricing and other means of charging for scholarly journals: a literature review and commentary. Learned Publishing, 21(4):248–272, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morgan:2008:JAV

[1387] Cliff Morgan. Journal article version nomenclature: the NISO/ALPSP recommendations. Learned Publishing, 21(4):273–277, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thatcher:2008:CCD

[1388] Sanford G. Thatcher. China’s copyright dilemma. Learned Publishing, 21 (4):278–284, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kirchhoff:2008:DPC

[1389] Amy J. Kirchhoff. Digital preservation: challenges and implementation. Learned Publishing, 21(4):285–294, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Givler:2008:SBR

[1390] Peter Givler. Seeking balance: rights and exceptions in Section 108 of the US Copyright Act. Learned Publishing, 21(4):295–299, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tananbaum:2008:EWB

[1391] Greg Tananbaum and Lyndon Holmes. The evolution of Web-based peerreview systems. Learned Publishing, 21(4):300–306, October 2008. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lu:2008:PRC

[1392] Yanping Lu. Peer review and its contribution to manuscript quality: an Australian perspective. Learned Publishing, 21(4):307–318, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Raphael:2008:PRP

[1393] Ellen Raphael. The publisher’s role in promoting peer review. Learned Publishing, 21(4):319–321, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Illig:2008:EKB

[1394] Joshua B. Illig and David Sampson. Event knowledge: a biomedical resource (r)evolution. Learned Publishing, 21(4):322–324, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dartnall:2008:IWI

[1395] Jean Dartnall. Are indexes worth it? Evidence from book reviews. Learned Publishing, 21(4):325–327, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2008:BR

[1396] Alan Singleton. Book reviews. Learned Publishing, 21(4):329–331, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2008:BR

[1397] Pippa Smart. Book reviews. Learned Publishing, 21(4):333, October 2008. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2009:JAF

[1398] Sally Morris. Is the journal article fit for purpose, or stuck in the past? Learned Publishing, 22(1):3–4, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davis:2009:RPB

[1399] Philip M. Davis. Reward or persuasion? The battle to define the meaning of a citation. Learned Publishing, 22(1):5–11, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pidgeon:2009:DNP

[1400] Sean Pidgeon. Developing a new publishing model for interdisciplinary research. Learned Publishing, 22(1):13–16, January 2009. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ball:2009:BPL

[1401] Rafael Ball. E-books in practice: the librarian’s perspective. Learned Publishing, 22(1):18–22, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sprouse:2009:RRA

[1402] Gene D. Sprouse. Recognizing referees at the American Physical Society. Learned Publishing, 22(1):23–26, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Earle:2009:BFL

[1403] Kathryn Earle. The Berg Fashion Library. Learned Publishing, 22(1):27– 35, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wittenberg:2009:GPO

[1404] Kate Wittenberg. The Gutenberg-e project: opportunities and challenges in publishing born-digital monographs. Learned Publishing, 22(1):36–41, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Chesler:2009:TBP

[1405] Adam Chesler and Susan King. Tier-based pricing for institutions: a new, e-based pricing model. Learned Publishing, 22(1):42–49, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wei:2009:BCR

[1406] Shi Wei and Nancy Benson. Bringing Chinese research to the world: Frontiers of environmental science & engineering in China. Learned Publishing, 22(1):50–56, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sambunjak:2009:NVI

[1407] Dario Sambunjak, Mirjana Huić, Darko Hren, Milica Katić, Ana Marušić, and Matko Marušić. National vs. international journals: views of medical professionals in Croatia. Learned Publishing, 22(1):57–70, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowland:2009:CEE

[1408] Fytton Rowland. Copy-editing — essential or frill? Learned Publishing, 22(1):71–72, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wagner:2009:AFT

[1409] A. Ben Wagner. A&I, full text, and open access: prophecy from the trenches. Learned Publishing, 22(1):73–74, January 2009. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Green:2009:BR

[1410] Toby Green. Book reviews. Learned Publishing, 22(1):75–77, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Richmond:2009:BRa

[1411] Carole Richmond. Book reviews. Learned Publishing, 22(1):77–78, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Murphy:2009:BR

[1412] Kevin Murphy. Book reviews. Learned Publishing, 22(1):78, January 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fisher:2009:GBS

[1413] Janet Fisher. The Google Book Search settlement. Learned Publishing, 22(2):82–83, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shotton:2009:SPC

[1414] David Shotton. Semantic publishing: the coming revolution in scientific journal publishing. Learned Publishing, 22(2):85–94, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ginsparg:2009:GVP

[1415] Paul Ginsparg. The global village pioneers. Learned Publishing, 22(2): 95–100, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

VanCampen:2009:ILL

[1416] Johan Van Campen and Eveline Hertzberger. ‘Improving little by little, every day’: the road to Operational Excellence. Learned Publishing, 22 (2):102–106, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Howells:2009:DJA

[1417] Matthew Howells, Ashleigh Bell, Nicholas Everitt, and Jennifer Mcmillan. Digitizing journal archives: the experience of Taylor & Francis. Learned Publishing, 22(2):109–112, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Seitter:2009:WPS

[1418] Keith L. Seitter and Kenneth F. Heideman. Whither print? Staying nimble in the face of uncertainty. Learned Publishing, 22(2):113–116, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bornmann:2009:LRD

[1419] Lutz Bornmann and Hans-Dieter Daniel. The luck of the referee draw: the effect of exchanging reviews. Learned Publishing, 22(2):117–125, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

King:2009:SJI

[1420] Donald W. King, Carol Tenopir, Songphan Choemprayong, and Lei Wu. Scholarly journal information-seeking and reading patterns of faculty at five US universities. Learned Publishing, 22(2):126–144, April 2009. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cookson:2009:WVM

[1421] Rod Cookson. Why visibility matters — exploding the Broadcast Fallacy. Learned Publishing, 22(2):146–152, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hannay:2009:WCT

[1422] Timo Hannay. Walls come tumbling down. Learned Publishing, 22(2): 153–154, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Spilsbury:2009:BR

[1423] Mark Spilsbury. Book reviews. Learned Publishing, 22(2):155, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kalk:2009:BR

[1424] Valentina Kalk. Book reviews. Learned Publishing, 22(2):155–157, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Richmond:2009:BRb

[1425] Carole Richmond. Book reviews. Learned Publishing, 22(2):157–158, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2009:BRa

[1426] Pippa Smart. Book reviews. Learned Publishing, 22(2):158, April 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2009:CT

[1427] Sally Morris. Crunch time. Learned Publishing, 22(3):163–164, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pinfield:2009:JRE

[1428] Stephen Pinfield. Journals and repositories: an evolving relationship? Learned Publishing, 22(3):165–175, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hedlund:2009:SPO

[1429] Turid Hedlund and Ingegerd Rabow. Scholarly publishing and open access in the Nordic countries. Learned Publishing, 22(3):177–186, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Li:2009:DFE

[1430] Li Li. A different form of editor — author interaction: the approach of university journals in China. Learned Publishing, 22(3):187–190, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Schiff:2009:CMT

[1431] Lisa Schiff. Creating the Mark Twain Project Online. Learned Publishing, 22(3):191–198, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Craven:2009:UWY

[1432] Alistair Craven and Graham Dallas. Using what you’ve got: the development of Emerald Management First. Learned Publishing, 22(3):199–205, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ashman:2009:WSW

[1433] Peter Ashman. What societies want from a publishing partner. Learned Publishing, 22(3):209–219, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morris:2009:LSM

[1434] Sally Morris and Sue Thorn. Learned society members and open access. Learned Publishing, 22(3):221–239, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ladds:2009:SMA

[1435] Andrew Ladds. Societies must adapt or die. Learned Publishing, 22(3): 240–242, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hayashi:2009:BR

[1436] Kazuhiro Hayashi. Book reviews. Learned Publishing, 22(3):243, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Oppenheim:2009:BR

[1437] Charles Oppenheim. Book reviews. Learned Publishing, 22(3):243–244, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2009:BRb

[1438] Pippa Smart. Book reviews. Learned Publishing, 22(3):244–245, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nix:2009:BR

[1439] Linda Nix. Book reviews. Learned Publishing, 22(3):245–248, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Adams:2009:BR

[1440] Clifford A. Adams. Book reviews. Learned Publishing, 22(3):248–249, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Marcovitch:2009:BR

[1441] Harvey Marcovitch. Book reviews. Learned Publishing, 22(3):249–252, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

White:2009:LE

[1442] Howard D. White. Letter to the Editors. Learned Publishing, 22(3):253– 254, July 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fisher:2009:T

[1443] Janet Fisher. Tipping over? Learned Publishing, 22(4):260–262, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stojanovski:2009:CNO

[1444] Jadranka Stojanovski, Jelka Petrak, and Bojan Macan. The Croatian national open access journal platform. Learned Publishing, 22(4):263– 273, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dodds:2009:WRT

[1445] Leigh Dodds. The Web’s rich tapestry. Learned Publishing, 22(4):275– 280, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Harper:2009:OIO

[1446] Georgia Harper. OA and IP: open access, digital copyright and marketplace competition. Learned Publishing, 22(4):283–288, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Campfens:2009:TCP

[1447] Yvonne Campfens and Ed Pentz. TRANSFER Code of Practice: the publisher’s point of view. Learned Publishing, 22(4):289–294, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Coleman:2009:PDL

[1448] Ross Coleman. Publishing and the digital library: adding value to scholarship and innovation to business. Learned Publishing, 22(4):297–303, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Csora:2009:PWW

[1449] Camelia Csora, Michiel Van Der Heyden, and Fabian Kersten. Practising what we preach: working together to further scientific collaboration. Learned Publishing, 22(4):304–311, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gardner:2009:MJA

[1450] John Gardner, Vladimir Bulatov, and Robert Kelly. Making journals accessible to the visually impaired: the future is near. Learned Publishing, 22(4):314–319, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gillingham:2009:RPI

[1451] Emily Gillingham and Sarah Bird. Responsibility and the publishing industry. Learned Publishing, 22(4):321–322, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:2009:DNT

[1452] Stephen K. Donovan. A decline to nothing? The tenuous existence of the small journal. Learned Publishing, 22(4):323–324, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Green:2009:WNP

[1453] Toby Green. We need publishing standards for datasets and data tables. Learned Publishing, 22(4):325–327, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bennett:2009:BRa

[1454] Linda Bennett. Book reviews. Learned Publishing, 22(4):328–329, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bailey:2009:BR

[1455] Rebecca Bailey. Book reviews. Learned Publishing, 22(4):329–330, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thatcher:2009:BR

[1456] Sanford G. Thatcher. Book reviews. Learned Publishing, 22(4):330–331, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Muller:2009:BR

[1457] Marius A. Müller. Book reviews. Learned Publishing, 22(4):331–332, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Carden:2009:BR

[1458] Mark Carden. Book reviews. Learned Publishing, 22(4):332–333, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bennett:2009:BRb

[1459] Linda Bennett. Book reviews. Learned Publishing, 22(4):333–334, October 2009. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2010:PTP

[1460] Alan Singleton. ‘Peaks and troughs’ or ‘pigs in troughs’? Learned Publishing, 23(1):3, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Meddings:2010:CWC

[1461] Kirsty Meddings. Credit where credit’s due: plagiarism screening in scholarly publishing. Learned Publishing, 23(1):5–8, January 2010. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2010:CET

[1462] Helen (Yuehong) Zhang. CrossCheck: an effective tool for detecting plagiarism. Learned Publishing, 23(1):9–14, January 2010. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Saper:2010:NPR

[1463] Clifford B. Saper and John H. R. Maunsell. The Neuroscience Peer Review Consortium. Learned Publishing, 23(1):15–16, January 2010. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kumar:2010:PRC

[1464] Malhar N. Kumar. The ‘peer reviewer as collaborator’ model for publishing. Learned Publishing, 23(1):17–22, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bide:2010:CRM

[1465] Mark Bide and Alicia Wise. 21st-century rights management: why does it matter and what is being done? Learned Publishing, 23(1):23–31, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Armbruster:2010:WMC

[1466] Chris Armbruster. Whose metrics? Citation, usage and access metrics as scholarly information service. Learned Publishing, 23(1):33–38, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pinfield:2010:POA

[1467] Stephen Pinfield. Paying for open access? Institutional funding streams and OA publication charges. Learned Publishing, 23(1):39–52, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nielsen:2010:WPL

[1468] Henning P. Nielsen and Sven-Olaf Vogt. What pharmaceutical libraries really want. Learned Publishing, 23(1):53–58, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Krell:2010:SEI

[1469] Frank-Thorsten Krell. Should editors influence journal impact factors? Learned Publishing, 23(1):59–62, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hartley:2010:NMI

[1470] James Hartley. Never mind the impact factor: colleagues know better! Learned Publishing, 23(1):63–65, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Glahn:2010:WOC

[1471] Peggy Glahn. Why outsourcing cannot help profitability in today’s challenging times. Learned Publishing, 23(1):67–68, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Friend:2010:JBM

[1472] Fred Friend. A journal business model to replace the big deal? Learned Publishing, 23(1):69–70, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Markwood:2010:BRa

[1473] Priscilla S. Markwood. Book reviews. Learned Publishing, 23(1):72–73, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:2010:BR

[1474] John Cox. Book reviews. Learned Publishing, 23(1):73–75, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowland:2010:BR

[1475] Fytton Rowland. Book reviews. Learned Publishing, 23(1):76, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Balkwill:2010:BRa

[1476] Richard Balkwill. Book reviews. Learned Publishing, 23(1):76–78, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Day:2010:BR

[1477] Sarah Day. Book reviews. Learned Publishing, 23(1):78, January 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fisher:2010:WB

[1478] Janet Fisher. Whither e-books? Learned Publishing, 23(2):83–84, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Culling:2010:ATN

[1479] James Culling. All together now: integrating customer data with Master-Vision. Learned Publishing, 23(2):85–92, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xue-Li:2010:SCC

[1480] Liu Xue-Li and Wang Mei-Ying. Self-citation in Chinese biomedical journals. Learned Publishing, 23(2):93–100, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cooney-Mcquat:2010:OAP

[1481] Sarah Cooney-Mcquat, Stefan Busch, and Deborah Kahn. Open access publishing: a viable solution for society publishers. Learned Publishing, 23(2):101–105, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bird:2010:CAO

[1482] Claire Bird. Continued adventures in open access: 2009 perspective. Learned Publishing, 23(2):107–116, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shalvi:2010:WWH

[1483] Shaul Shalvi, Matthijs Baas, Michel J. J. Handgraaf, and Carsten K. W. De Dreu. Write when hot — submit when not: seasonal bias in peer review or acceptance? Learned Publishing, 23(2):117–123, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bornmann:2010:RRR

[1484] L. Bornmann and H.-D. Daniel. Reliability of reviewers’ ratings when using public peer review: a case study. Learned Publishing, 23(2):124– 131, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Griffin:2010:JBJ

[1485] Cathy Griffin. The Journal of Bone & Joint Surgery’s CrossCheck experience. Learned Publishing, 23(2):132–135, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Waltham:2010:HSS

[1486] Mary Waltham. Humanities and social science journals: a pilot study of eight US associations. Learned Publishing, 23(2):136–143, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowlands:2010:EDL

[1487] Ian Rowlands and David Nicholas. The economic downturn and libraries: an international survey. Learned Publishing, 23(2):144–156, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rapple:2010:OIC

[1488] Charlie Rapple. Online. Indexed. Catalogued. Free. But will users find it? Learned Publishing, 23(2):157–161, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Freeman:2010:WTA

[1489] Ted Freeman. The Web@20: thoughts about utopias, technology, and collaborative science. Learned Publishing, 23(2):163–165, April 2010. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kilborn:2010:IB

[1490] Peter Kilborn. Identification of e-books. Learned Publishing, 23(2):166– 168, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jubb:2010:BRa

[1491] Michael Jubb. Book reviews. Learned Publishing, 23(2):169–170, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Okerson:2010:BR

[1492] Ann Okerson. Book reviews. Learned Publishing, 23(2):170–171, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Roos:2010:BR

[1493] Anna Marie Roos. Book reviews. Learned Publishing, 23(2):171–172, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mcilroy:2010:BRa

[1494] Thad Mcilroy. Book reviews. Learned Publishing, 23(2):172, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mcilroy:2010:BRb

[1495] Thad Mcilroy. Book reviews. Learned Publishing, 23(2):172–173, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bowes:2010:BR

[1496] William Bowes. Book reviews. Learned Publishing, 23(2):173–174, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jubb:2010:BRb

[1497] Michael Jubb. Book reviews. Learned Publishing, 23(2):174–175, April 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2010:WUU

[1498] Alan Singleton. Why usage is useless. Learned Publishing, 23(3):179– 184, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smith:2010:GIE

[1499] Maxine Smith. Gaining ISO 9001: the experience of Oxford Journals. Learned Publishing, 23(3):185–192, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rodriguez-Armentia:2010:IWE

[1500] Nerea Rodríguez-Armentia and Carlos B. Amat. Is it worth establishing institutional repositories? The strategies for open access to Spanish peerreviewed articles. Learned Publishing, 23(3):193–208, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mcclanahan:2010:ECP

[1501] Kitty Mcclanahan, Lei Wu, Carol Tenopir, and Donald W. King. Embracing change: perceptions of e-journals by faculty members. Learned Publishing, 23(3):209–223, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wan:2010:KYU

[1502] Gang (Gary) Wan and Zao Liu. Knowing your users: the value of article database usage analysis. Learned Publishing, 23(3):225–236, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Estrada-Mejia:2010:QVS

[1503] Catalina Estrada-Mejía and Clemente Forero-Pineda. The quest for visibility of scientific journals in Latin America. Learned Publishing, 23(3): 237–252, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson-Wilk:2010:DRM

[1504] Mark Anderson-Wilk. In defense of the reference: more than linking in evolving Web environments. Learned Publishing, 23(3):253–257, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jinha:2010:AME

[1505] Arif E. Jinha. Article 50 million: an estimate of the number of scholarly articles in existence. Learned Publishing, 23(3):258–263, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Campbell:2010:BR

[1506] Bob Campbell, John Willinsky, and Rick Anderson. Book reviews. Learned Publishing, 23(3):264–266, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mlambo-Thata:2010:BR

[1507] Buhle Mlambo-Thata. Book reviews. Learned Publishing, 23(3):266– 267, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2010:BR

[1508] Alan Singleton. Book reviews. Learned Publishing, 23(3):267–268, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Markwood:2010:BRb

[1509] Priscilla S. Markwood. Book reviews. Learned Publishing, 23(3):268– 269, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Albitz:2010:BR

[1510] Becky Albitz. Book reviews. Learned Publishing, 23(3):269–270, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Craig:2010:BR

[1511] Iain D. Craig. Book reviews. Learned Publishing, 23(3):270–271, July 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fisher:2010:DYK

[1512] Janet Fisher. Do you know what your users are doing? Learned Publishing, 23(4):277–278, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Saunders:2010:PL

[1513] Joss Saunders. Plagiarism and the law. Learned Publishing, 23(4):279– 292, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Snijder:2010:PFB

[1514] Ronald Snijder. The profits of free books: an experiment to measure the impact of open access publishing. Learned Publishing, 23(4):293–301, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shen:2010:CMP

[1515] Min Shen, Wan-Jie Jiang, and Shu sen Zheng. China’s medical periodicals: from localization to internationalization. Learned Publishing, 23 (4):303–311, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wang:2010:DPC

[1516] Zhigang Wang. Digital publishing in China: is there a copyright crisis? Learned Publishing, 23(4):313–318, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ahmed:2010:ACB

[1517] Hasan Shareef Ahmed, Abdullahel Hadi, and Nuzhat Choudhury. Authorship conflict in Bangladesh: an exploratory study. Learned Publishing, 23(4):319–325, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Park:2010:PWO

[1518] William Park. The publishing world is open, not flat. Learned Publishing, 23(4):326–328, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Armstrong:2010:CBL

[1519] Tracey Armstrong. Context-based licensing: technology and content working together. Learned Publishing, 23(4):329–330, October 2010. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ufnalska:2010:MEG

[1520] Sylwia B. Ufnalska. Multilingual EASE guidelines for authors and translators. Learned Publishing, 23(4):331–332, October 2010. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wilbanks:2010:WNW

[1521] John Wilbanks. We need a Web for data. Learned Publishing, 23(4): 333–335, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kahler:2010:CPR

[1522] Ove Kähler. Combining peer review and metrics to assess journals for inclusion in Scopus. Learned Publishing, 23(4):336–346, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:2010:SBP

[1523] Laura Cox. Scholarly book publishing practice: the ALPSP survey findings. Learned Publishing, 23(4):347–356, October 2010. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Marlow:2010:RAU

[1524] Martin Marlow. The role of the aggregator: using third-party suppliers to reach the market. Learned Publishing, 23(4):357–361, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Balkwill:2010:BRb

[1525] Richard Balkwill. Book reviews. Learned Publishing, 23(4):363–364, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Balkwill:2010:BRc

[1526] Richard Balkwill. Book reviews. Learned Publishing, 23(4):364–365, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Payne:2010:BR

[1527] David Payne. Book reviews. Learned Publishing, 23(4):365–366, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Alexander:2010:BR

[1528] Huw Alexander. Book reviews. Learned Publishing, 23(4):366–367, October 2010. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2011:SCC

[1529] Alan Singleton. Scholarly communication — can we have our name back? Learned Publishing, 24(1):3–4, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Osborne:2011:RAR

[1530] Nikki Osborne. Reporting animal research — worthy of a rethink? Learned Publishing, 24(1):5–7, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rapple:2011:RIN

[1531] Charlie Rapple. Researching and implementing a new tiered pricing model. Learned Publishing, 24(1):9–13, January 2011. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2011:JRN

[1532] David Nicholas, Ian Rowlands, and Peter Williams. E-journals, researchers — and the new librarians. Learned Publishing, 24(1):15–27, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hartley:2011:WWY

[1533] James Hartley. Write when you can and submit when you are ready! Learned Publishing, 24(1):29–32, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic). See reply [1534].

Shalvi:2011:WSW

[1534] Shaul Shalvi, Matthijs Baas, Michel J. J. Handgraaf, and Carsten K. W. De Dreu. When should we submit our papers? Reply to Hartley. Learned Publishing, 24(1):33–34, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic). See [1533].

Smit:2011:ADD

[1535] Eefke Smit, Jeffrey Van Der Hoeven, and David Giaretta. Avoiding a Digital Dark Age for data: why publishers should care about digital preservation. Learned Publishing, 24(1):35–49, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morgan:2011:UCC

[1536] Cliff Morgan. Understanding the Creative Commons licence. Learned Publishing, 24(1):51–53, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Souder:2011:ESP

[1537] Lawrence Souder. The ethics of scholarly peer review: a review of the literature. Learned Publishing, 24(1):55–72, January 2011. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thatcher:2011:BRa

[1538] Sanford G. Thatcher. Book reviews. Learned Publishing, 24(1):75–76, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2011:BRa

[1539] Pippa Smart. Book reviews. Learned Publishing, 24(1):76–77, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2011:BRb

[1540] Pippa Smart. Book reviews. Learned Publishing, 24(1):77–78, January 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott-Lichter:2011:PEP

[1541] Diane Scott-Lichter. Publication ethics: prevention, screening, and treatment. Learned Publishing, 24(2):84–85, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Meyer:2011:DPS

[1542] Carol Anne Meyer. Distinguishing published scholarly content with CrossMark. Learned Publishing, 24(2):87–93, April 2011. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jufang:2011:OAP

[1543] Shao Jufang and Shen Huiyun. The outflow of academic papers from China: why is it happening and can it be stemmed? Learned Publishing, 24(2):95–97, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mooney:2011:CDS

[1544] Hailey Mooney. Citing data sources in the social sciences: do authors do it? Learned Publishing, 24(2):99–108, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cradock:2011:JTP

[1545] Chris Cradock, Paul Meehan, and Paul Needham. JUSP in time: a partnership approach to developing a journal usage statistics portal. Learned Publishing, 24(2):109–114, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cox:2011:LUU

[1546] Laura Cox. Librarians’ use of usage statistics for journals and e-books. Learned Publishing, 24(2):115–121, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tenopir:2011:PVS

[1547] Carol Tenopir, Suzie Allard, Benjamin J. Bates, Kenneth J. Levine, Donald W. King, Ben Birch, Regina Mays, and Chris Caldwell. Perceived value of scholarly articles. Learned Publishing, 24(2):123–132, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hubbard:2011:CWN

[1548] Stephen C. Hubbard and Marie E. Mcveigh. Casting a wide net: the Journal Impact Factor numerator. Learned Publishing, 24(2):133–137, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kuznetsov:2011:SBR

[1549] Alexander Kuznetsov and Irina Razumova. Selling to the BRIC–Russia: scholarly e-products and the Russian market. Learned Publishing, 24 (2):139–144, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sanni:2011:EIM

[1550] S. A. Sanni and A. N. Zainab. Evaluating the influence of a medical journal using Google Scholar. Learned Publishing, 24(2):145–154, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fuller:2011:BR

[1551] Louise Fuller. Book reviews. Learned Publishing, 24(2):155–156, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Phillips:2011:BR

[1552] Angus Phillips. Book reviews. Learned Publishing, 24(2):156–157, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thatcher:2011:BRb

[1553] Sanford G. Thatcher. Book reviews. Learned Publishing, 24(2):157– 158, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Woodhead:2011:BR

[1554] Martin Woodhead. Book reviews. Learned Publishing, 24(2):158–159, April 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2011:PR

[1555] Alan Singleton. The pain of rejection. Learned Publishing, 24(3):162– 163, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Macri:2011:SBP

[1556] Erin M. Macri and Karim M. Khan. Single-blind peer review: an appropriate compromise between two ideals? Learned Publishing, 24(3): 164–165, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Vaux:2011:DBR

[1557] David L. Vaux. Double blind review. Learned Publishing, 24(3):165– 167, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pulverer:2011:COE

[1558] Bernd Pulverer. Close one eye and open wide! Learned Publishing, 24(3):167–169, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Campbell:2011:SJP

[1559] Robert Campbell and Alice Meadows. Scholarly journal publishing: where do we go from here? Learned Publishing, 24(3):171–181, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rowlands:2011:SMU

[1560] Ian Rowlands, David Nicholas, Bill Russell, Nicholas Canty, and Anthony Watkinson. Social media use in the research workflow. Learned Publishing, 24(3):183–195, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Barron:2011:LFG

[1561] Peter Barron. The library of the future: Google’s vision for books. Learned Publishing, 24(3):197–201, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smith:2011:CDN

[1562] Mackenzie Smith. Communicating with data: new roles for scientists, publishers and librarians. Learned Publishing, 24(3):203–205, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pettifer:2011:CEH

[1563] S. Pettifer, P. Mcdermott, J. Marsh, D. Thorne, A. Villeger, and T. K. Attwood. Ceci n’est pas un hamburger: modelling and representing the scholarly article. Learned Publishing, 24(3):207–220, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bennett:2011:TYB

[1564] Linda Bennett. Ten years of e-books: a review. Learned Publishing, 24(3):222–229, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ling:2011:IPR

[1565] Fay Ling. Improving peer review: increasing reviewer participation. Learned Publishing, 24(3):231–233, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rendell-Dunn:2011:BR

[1566] Alexis Rendell-Dunn. Book reviews. Learned Publishing, 24(3):235, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Balkwill:2011:BR

[1567] Richard Balkwill. Book reviews. Learned Publishing, 24(3):236–237, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rapple:2011:BR

[1568] Charlie Rapple. Book reviews. Learned Publishing, 24(3):237–238, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2011:BRc

[1569] Pippa Smart. Book reviews. Learned Publishing, 24(3):238–239, July 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott-Lichter:2011:GCG

[1570] Diane Scott-Lichter. Got content, get attention. Learned Publishing, 24 (4):245–246, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jubb:2011:CRB

[1571] Michael Jubb, Joel Cook, Daniel Hulls, David Jones, and Mark Ware. Costs, risks and benefits in improving access to journal articles. Learned Publishing, 24(4):247–260, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Myers:2011:SBB

[1572] David Myers. Selling to the BRIC: Brazil — an awakening tiger. Learned Publishing, 24(4):261–265, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wallace:2011:PGO

[1573] Julia M. Wallace. PEER: green open access — insight and evidence. Learned Publishing, 24(4):267–277, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shepherd:2011:PIA

[1574] Peter T. Shepherd. PIRUS2: individual article usage statistics — developing a practical, global standard. Learned Publishing, 24(4):279–286, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Davis:2011:DDJ

[1575] Philip M. Davis. Do discounted journal access programs help researchers in sub-Saharan Africa? A bibliometric analysis. Learned Publishing, 24 (4):287–298, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xiao-jun:2011:PAW

[1576] He Xiao-jun and Chen Zhen-ying. Profit or access: which is more important for Chinese medical journals? Learned Publishing, 24(4):299–302, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Didegah:2011:ECJ

[1577] Fereshteh Didegah and Ali Gazni. The extent of concentration in journal publishing. Learned Publishing, 24(4):303–310, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Armbruster:2011:OAP

[1578] Chris Armbruster. Open access policy implementation: first results compared. Learned Publishing, 24(4):311–324, October 2011. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bornmann:2011:SBE

[1579] Lutz Bornmann and Hans-Dieter Daniel. Seasonal bias in editorial decisions? A study using data from chemistry. Learned Publishing, 24 (4):325–328, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pratap:2011:BR

[1580] Anji Pratap. Book reviews. Learned Publishing, 24(4):331–332, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clark:2011:BR

[1581] Cindy Clark. Book reviews. Learned Publishing, 24(4):332–334, October 2011. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2012:SWF

[1582] Alan Singleton. Should we forgive them if they know not what they do? Learned Publishing, 25(1):3–5, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sismondo:2012:MPD

[1583] Sergio Sismondo. Medical publishing and the drug industry: is medical science for sale? Learned Publishing, 25(1):7–15, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kenneway:2012:ISM

[1584] Melinda Kenneway and Charlie Rapple. Integrating social media in the marketing mix: the case of Best Practice. Learned Publishing, 25(1):17– 20, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mccoyd:2012:OPP

[1585] Ed Mccoyd. Online piracy of publishers’ content: a primer. Learned Publishing, 25(1):21–28, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brand:2012:BMR

[1586] Amy Brand. Beyond mandate and repository, toward sustainable faculty self-archiving. Learned Publishing, 25(1):29–34, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smit:2012:JAM

[1587] Eefke Smit and Maurits Van Der Graaf. Journal article mining: the scholarly publishers’ perspective. Learned Publishing, 25(1):35–46, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Janowicz:2012:OTR

[1588] Krzysztof Janowicz and Pascal Hitzler. Open and transparent: the review process of the Semantic Web journal. Learned Publishing, 25(1):48–55, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lu:2012:LCC

[1589] Yanping Lu. Learning to be confident and capable journal reviewers: an Australian perspective. Learned Publishing, 25(1):56–61, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jei:2012:SPI

[1590] Xu Jei and Matthias Wahls. The scholarly publishing industry in China: overview and opportunities. Learned Publishing, 25(1):63–74, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Alexander:2012:BR

[1591] Huw Alexander. Book reviews. Learned Publishing, 25(1):75–76, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Okerson:2012:BRa

[1592] Ann Shumelda Okerson. Book reviews. Learned Publishing, 25(1):76–77, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clark:2012:BR

[1593] Cindy Clark. Book reviews. Learned Publishing, 25(1):78, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Okerson:2012:BRb

[1594] Ann Okerson. Book reviews. Learned Publishing, 25(1):78–79, January 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott-Lichter:2012:ADM

[1595] Diane Scott-Lichter. Authorship disputes: me first, me equally, me too, not me. Learned Publishing, 25(2):83–85, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mitchell:2012:PTC

[1596] Alison Mitchell and Elizabeth L. Winter. Project Transfer comes of age? Learned Publishing, 25(2):87–92, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2012:RDE

[1597] David Nicholas and David Clark. ‘Reading’ in the digital environment. Learned Publishing, 25(2):93–98, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Leubsdorf:2012:ALP

[1598] Carl Leubsdorf, Jr. Annotum: launching a peer-reviewed journal online for free. Learned Publishing, 25(2):99–106, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pinfield:2012:OAC

[1599] Stephen Pinfield and Christine Middleton. Open access central funds in UK universities. Learned Publishing, 25(2):107–117, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stanley:2012:SBC

[1600] Adrian Stanley. Selling to the BRIC: China — understanding and improving your footprint. Learned Publishing, 25(2):119–131, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bjork:2012:PPU

[1601] Bo-Christer Björk and David Solomon. Pricing principles used by scholarly open access publishers. Learned Publishing, 25(2):132–137, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Owen:2012:PKP

[1602] Brian Owen and Kevin Stranack. The Public Knowledge Project and Open Journal Systems: open source options for small publishers. Learned Publishing, 25(2):138–144, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Schreiber:2012:SBE

[1603] Michael Schreiber. Seasonal bias in editorial decisions for a physics journal: you should write when you like, but submit in July. Learned Publishing, 25(2):145–151, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Krell:2012:APT

[1604] Frank-Thorsten Krell. Academic publishers’ time-loop: another mechanism to manipulate impact factors? Learned Publishing, 25(2):153–154, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Esposito:2012:BR

[1605] Joseph J. Esposito. Book reviews. Learned Publishing, 25(2):155–156, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Saunders:2012:BR

[1606] Joss Saunders. Book reviews. Learned Publishing, 25(2):156–157, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clarke:2012:BRa

[1607] Anji Clarke. Book reviews. Learned Publishing, 25(2):157–158, April 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2012:CCN

[1608] Alan Singleton. Content, context? No contest … probably. Learned Publishing, 25(3):163–165, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paulus:2012:ABH

[1609] Kurt Paulus. ALPSP: bring on the half century! Learned Publishing, 25(3):167–174, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kavanagh:2012:SST

[1610] David J. Kavanagh and Paul Phillips. scrazzl: semantic technology for evaluating research tools and products in scholarly literature. Learned Publishing, 25(3):175–181, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thomson:2012:OAK

[1611] Simon Thomson and Robert Kurn. Open Access Key: a new system for managing author publication payments. Learned Publishing, 25(3):183– 188, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Meadows:2012:AQ

[1612] Alice Meadows, Robert Campbell, and Keith Webster. The access question. Learned Publishing, 25(3):189–194, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2012:DRT

[1613] David Nicholas, Ian Rowlands, Anthony Watkinson, David Brown, and Hamid R. Jamali. Digital repositories ten years on: what do scientific researchers think of them and how do they use them? Learned Publishing, 25(3):195–206, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mathelus:2012:PRA

[1614] Sharon Mathelus, Ginny Pittman, and Jill Yablonski-Crepeau. Promotion of research articles to the lay press: a summary of a three-year project. Learned Publishing, 25(3):207–212, July 2012. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Malik:2012:SBB

[1615] Rakesh Malik. Selling to the BRIC: the background to the scholarly publishing market in India. Learned Publishing, 25(3):213–218, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xiao-jun:2012:CSJ

[1616] He Xiao-jun, Chen Zhen-ying, and Shen Hui-yun. Chinese scientific journals: how they can survive. Learned Publishing, 25(3):219–224, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sellwood:2012:EPO

[1617] Sally Sellwood. Editorial processing: to outsource or not? Learned Publishing, 25(3):225–230, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baveye:2012:WRE

[1618] Philippe C. Baveye. Wanted: a ‘Reviewer Effectiveness Index’. Learned Publishing, 25(3):232–234, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clarke:2012:BRb

[1619] Anji Clarke. Book reviews. Learned Publishing, 25(3):235–236, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Woodhead:2012:BR

[1620] Martin Woodhead. Book reviews. Learned Publishing, 25(3):236, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ufnalska:2012:BR

[1621] Sylwia B. Ufnalska. Book reviews. Learned Publishing, 25(3):237–238, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Donovan:2012:BR

[1622] Stephen K. Donovan. Book reviews. Learned Publishing, 25(3):238, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hirschy:2012:BR

[1623] Norm Hirschy. Book reviews. Learned Publishing, 25(3):238–239, July 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott-Lichter:2012:NJS

[1624] Diane Scott-Lichter. New journal selection by A&Is — still valuable after all these years. Learned Publishing, 25(4):245–246, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mcilroy:2012:EFM

[1625] Thad Mcilroy. Ebook formats are a mess — here’s why. Learned Publishing, 25(4):247–250, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rapple:2012:GMR

[1626] Charlie Rapple. Getting mobile right: start with who and why, not what and how. Learned Publishing, 25(4):252–257, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Haak:2012:OSU

[1627] Laurel L. Haak, Martin Fenner, Laura Paglione, Ed Pentz, and Howard Ratner. ORCID: a system to uniquely identify researchers. Learned Publishing, 25(4):259–264, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2012:ERO

[1628] Xiangyi Zhang. Effect of reviewer’s origin on peer review: China vs. non-China. Learned Publishing, 25(4):265–270, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Boukacem-Zeghmouri:2012:SUF

[1629] Chérifa Boukacem-Zeghmouri and Joachim Schöpfel. Statistics usage by French academic libraries: a survey. Learned Publishing, 25(4):271–278, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tenopir:2012:ABR

[1630] Carol Tenopir, Rachel Volentine, and Donald W. King. Article and book reading patterns of scholars: findings for publishers. Learned Publishing, 25(4):279–291, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2012:SUC

[1631] Yuehong (Helen) Zhang and Xiaoyan Jia. A survey on the use of Cross-Check for detecting plagiarism in journal articles. Learned Publishing, 25 (4):292–307, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Berhane:2012:N

[1632] Daniel Berhane and David Payne. A new. Learned Publishing, 25(4): 308–315, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2012:BR

[1633] David Nicholas. Book reviews. Learned Publishing, 25(4):316–317, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2012:BR

[1634] Pippa Smart. Book reviews. Learned Publishing, 25(4):317–318, October 2012. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2013:AWN

[1635] Alan Singleton. APC — what’s in a name? Learned Publishing, 26(1):3– 4, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson:2013:SIP

[1636] Kent Anderson. Sustainable independent publishing in a world of increasing funder power. Learned Publishing, 26(1):5–10, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Adie:2013:AES

[1637] Euan Adie and William Roe. Altmetric: enriching scholarly content with article-level discussion and metrics. Learned Publishing, 26(1):11– 17, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Attis:2013:TTS

[1638] David Attis and Colin Koproske. Thirty trends shaping the future of academic libraries. Learned Publishing, 26(1):18–23, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lin:2013:WSA

[1639] Songqing Lin. Why serious academic fraud occurs in China. Learned Publishing, 26(1):24–27, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kalwij:2013:HAC

[1640] Jesse M. Kalwij and Christian Smit. How authors can maximise the chance of manuscript acceptance and article visibility. Learned Publishing, 26(1):28–31, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bennett:2013:FLA

[1641] Emma Bennett. The future of learned associations in the humanities. Learned Publishing, 26(1):32–41, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xiao-jun:2013:EEN

[1642] He Xiao-jun and Chen Zhen-ying. Excellent editors need to be good authors too. Learned Publishing, 26(1):42–44, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hongling:2013:SCR

[1643] Fang Hongling. Self-citation rates of scientific and technical journals in SCI from China, Japan, India, and Korea. Learned Publishing, 26(1):45– 49, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Chumbe:2013:MOJ

[1644] Santiago Chumbe and Roddy Macleod. Marketing OA journals now that authors are customers: the role of RSS. Learned Publishing, 26(1):51–56, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fund:2013:WNI

[1645] Sven Fund. We need integrated publishing. Learned Publishing, 26(1): 57–64, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wates:2013:BR

[1646] Edward Wates. Book reviews. Learned Publishing, 26(1):65–67, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Conrad:2013:BR

[1647] Lettie Conrad. Book reviews. Learned Publishing, 26(1):67–69, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Saunders:2013:BR

[1648] Joss Saunders. Book reviews. Learned Publishing, 26(1):69–70, January 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2013:HPR

[1649] Anonymous. HighWire Press: Researcher Workflow Interview Study Questionnaire. Learned Publishing, 26(2):1–4, April 2013. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott-Lichter:2013:IYB

[1650] Diane Scott-Lichter. If you blow a whistle, know the tune. Learned Publishing, 26(2):75, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Berry:2013:MAA

[1651] Michael Berry. Much ado about rather little. Learned Publishing, 26(2): 77, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Beall:2013:PPJ

[1652] Jeffrey Beall. Predatory publishing is just one of the consequences of gold open access. Learned Publishing, 26(2):79–84, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pinfield:2013:SPG

[1653] Stephen Pinfield. Is scholarly publishing going from crisis to crisis? Learned Publishing, 26(2):85–88, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jie:2013:OSP

[1654] Xu Jie and Yuan Xiaoqun. Online scholarly publishing in China: Who? What? How? Learned Publishing, 26(2):89–100, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Delgado-Lopez-Cozar:2013:RJC

[1655] Emilio Delgado-López-Cózar and Álvaro Cabezas-Clavijo. Ranking journals: could Google Scholar Metrics be an alternative to Journal Citation Reports and Scimago Journal Rank? Learned Publishing, 26(2):101– 114, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Borgstrom:2013:SBP

[1656] Liam Borgstrom. Simplifying e-book and print production. Learned Publishing, 26(2):115–122, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Newman:2013:IWA

[1657] Michael L. Newman and John Sack. Information workflow of academic researchers in the evolving information environment: an interview study. Learned Publishing, 26(2):123–131, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Joshi:2013:SGR

[1658] Yateendra Joshi. Style guides that refuse to go away. Learned Publishing, 26(2):133–134, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Qing:2013:MOR

[1659] Fang Qing and Xu Lifang. Market-oriented reform in Chinese scholarly publishing. Learned Publishing, 26(2):135–138, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mudrak:2013:UNI

[1660] Benjamin Mudrak. Understanding the needs of international authors. Learned Publishing, 26(2):139–147, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2013:BR

[1661] Pippa Smart. Book reviews. Learned Publishing, 26(2):148, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Balkwill:2013:BR

[1662] Richard Balkwill. Book reviews. Learned Publishing, 26(2):148–149, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gunter:2013:BR

[1663] Darrell W. Gunter. Book reviews. Learned Publishing, 26(2):149–150, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:2013:BRa

[1664] Mark Ware. Book reviews. Learned Publishing, 26(2):150–151, April 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2013:POL

[1665] Alan Singleton. Publishing — is our love here to stay …. Learned Publishing, 26(3):155–156, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hutchens:2013:OAM

[1666] Chad Hutchens. Open access metadata: current practices and proposed solutions. Learned Publishing, 26(3):159–165, July 2013. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Borghart:2013:LPR

[1667] Pieter Borghart. A label for peer-reviewed books? Some critical reflections. Learned Publishing, 26(3):167–171, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dillaerts:2013:HTC

[1668] Hans Dillaerts and Ghislaine Chartron. ‘Héloïse’: towards a co-ordinated ecosystem approach for the archiving of scientific publications? Learned Publishing, 26(3):173–179, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Leach:2013:PKB

[1669] Margo C. Leach and Shaun L. A. Hobbs. Plantwise knowledge bank: delivering plant health information to developing country users. Learned Publishing, 26(3):180–185, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Awati:2013:SEW

[1670] Mriganka Awati. A science editor’s wish list. Learned Publishing, 26(3): 187–188, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2013:RCP

[1671] Yuehong (Helen) Zhang and Xiaoyan Jia. Republication of conference papers in journals? Learned Publishing, 26(3):189–196, July 2013. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ren:2013:BOP

[1672] Xiang Ren. Beyond online preprints: formalization of open initiatives in China. Learned Publishing, 26(3):197–205, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gantz:2013:JPS

[1673] Paula Gantz. Journal print subscription price increases no longer reflect actual costs. Learned Publishing, 26(3):206–210, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baverstock:2013:WSP

[1674] Alison Baverstock and Jackie Steinitz. Who are the self-publishers? Learned Publishing, 26(3):211–223, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sun:2013:VEF

[1675] Jilan Sun. Vocabulary extraction in foreign-language journals: how natural language processing can help readers. Learned Publishing, 26(3): 224–227, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jubb:2013:BR

[1676] Michael Jubb. Book reviews. Learned Publishing, 26(3):228–229, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thatcher:2013:BR

[1677] Sanford G. Thatcher. Book reviews. Learned Publishing, 26(3):229– 230, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ware:2013:BRb

[1678] Mark Ware. Book reviews. Learned Publishing, 26(3):230–231, July 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott-Lichter:2013:CMR

[1679] Diane Scott-Lichter. CHORUS: moving research publications in line with US federal funding requirements. Learned Publishing, 26(4):236–238, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2013:SDT

[1680] David Nicholas and David Clark. The second digital transition: to the mobile space — an analysis of Europeana. Learned Publishing, 26(4): 240–252, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Armstrong:2013:NEG

[1681] David Armstrong. The new engagement game: the role of gamification in scholarly publishing. Learned Publishing, 26(4):253–256, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jian-ping:2013:TES

[1682] Lu Jian-ping. Technical and ethical standards in China: long on framework but short on action. Learned Publishing, 26(4):257–258, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Froehlich:2013:HSU

[1683] Peter Froehlich. How to spot Ugly Black Ducklings: the next competitive frontiers in scholarly publishing. Learned Publishing, 26(4):259–263, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stemmle:2013:RTS

[1684] Laura Stemmle and Keith Collier. Rubriq: tools, services, and software to improve peer review. Learned Publishing, 26(4):265–268, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tappuni:2013:OCF

[1685] Jane Tappuni. Are online communities the future of bookselling? Learned Publishing, 26(4):269–271, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baverstock:2013:WSD

[1686] Alison Baverstock and Jackie Steinitz. What satisfactions do selfpublishing authors gain from the process? Learned Publishing, 26(4): 272–282, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pinfield:2013:MRC

[1687] Stephen Pinfield. Medical research charities and open access. Learned Publishing, 26(4):285–302, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lin:2013:HAS

[1688] Songqing Lin. How the approval system affects Chinese scientific journals. Learned Publishing, 26(4):303–309, October 2013. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clark:2013:BR

[1689] Cindy Clark. Book review. Learned Publishing, 26(4):310, October 2013. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2014:FSJ

[1690] Alan Singleton. The first scientific journal. Learned Publishing, 27(1):2– 4, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kiernan:2014:PRP

[1691] Vincent Kiernan. Public relations practices at medical journals. Learned Publishing, 27(1):5–13, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Grgic:2014:SJP

[1692] Ivana Hebrang Grgić. Scholarly journals at the periphery: the case of Croatia. Learned Publishing, 27(1):15–20, January 2014. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mukherjee:2014:GGO

[1693] Bhaskar Mukherjee. Green and gold open access in India. Learned Publishing, 27(1):21–32, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wager:2014:DRP

[1694] Elizabeth Wager. Defining and responding to plagiarism. Learned Publishing, 27(1):33–42, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hongling:2014:HIP

[1695] Fang Hongling and Chang Hai-Min. High-impact papers from China, Japan, India, and Korea. Learned Publishing, 27(1):44–46, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baveye:2014:LPW

[1696] Philippe C. Baveye. Learned publishing: who still has time to read? Learned Publishing, 27(1):48–51, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Antonoyiannakis:2014:ARP

[1697] Manolis Antonoyiannakis. Acceptance rates in Physical Review Letters: no seasonal bias. Learned Publishing, 27(1):53–55, January 2014. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rodrigues:2014:IAJ

[1698] Rosângela Schwarz Rodrigues and Ernest Abadal. Ibero–American journals in Scopus and Web of Science. Learned Publishing, 27(1):56–62, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cerejo:2014:IJE

[1699] Clarinda Cerejo. International journal editors and East Asian authors: two surveys. Learned Publishing, 27(1):63–75, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2014:BR

[1700] David Nicholas. Book reviews. Learned Publishing, 27(1):76–77, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gunter:2014:BRa

[1701] Darrell W. Gunter. Book reviews. Learned Publishing, 27(1):77–78, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thatcher:2014:BR

[1702] Sanford G. Thatcher. Book reviews. Learned Publishing, 27(1):78, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bullock:2014:BR

[1703] Adrian Bullock. Book reviews. Learned Publishing, 27(1):78–79, January 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott-Lichter:2014:LSR

[1704] Diane Scott-Lichter. Learned societies: resilience not reliance required. Learned Publishing, 27(2):83–84, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Faulder:2014:ARE

[1705] Sarah Faulder and Shinwha Cha. Access to Research: the experience of implementing a pilot in public libraries. Learned Publishing, 27(2):85– 92, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bjork:2014:HRF

[1706] Bo-Christer Björk and David Solomon. How research funders can finance APCs in full OA and hybrid journals. Learned Publishing, 27(2):93– 103, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Emerman:2014:AIR

[1707] Michael Emerman. Author-initiated repurposing of reviews: the Streamline Reviews initiative at Virology. Learned Publishing, 27(2):104–106, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Melero:2014:JAR

[1708] R. Melero, J. M. Rodríguez-Gairín, F. Abad-García, and E. Abadal. Journal author rights and self-archiving: the case of Spanish journals. Learned Publishing, 27(2):107–120, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2014:TAS

[1709] David Nicholas, Anthony Watkinson, Rachel Volentine, Suzie Allard, Kenneth Levine, Carol Tenopir, and Eti Herman. Trust and authority in scholarly communications in the light of the digital transition: setting the scene for a major study. Learned Publishing, 27(2):121–134, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yan:2014:CBU

[1710] Yalan Yan and Xianjin Zha. Comparison between user affinity with digital libraries and virtual communities. Learned Publishing, 27(2):135– 143, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lin:2014:TJC

[1711] Songqing Lin and Lijuan Zhan. Trash journals in China. Learned Publishing, 27(2):145–154, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gunter:2014:BRb

[1712] Darrell W. Gunter. Book reviews. Learned Publishing, 27(2):155–156, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Slater:2014:BR

[1713] Eric S. Slater. Book reviews. Learned Publishing, 27(2):156–157, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clark:2014:BR

[1714] Cindy Clark. Book reviews. Learned Publishing, 27(2):157–158, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2014:BR

[1715] Pippa Smart. Book reviews. Learned Publishing, 27(2):158, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Souder:2014:BR

[1716] Lawrence Souder. Book reviews. Learned Publishing, 27(2):158–159, April 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Singleton:2014:RT

[1717] Alan Singleton. Reproducibility — and trust. Learned Publishing, 27(3): 163–166, July 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kumar:2014:SMS

[1718] Malhar N. Kumar. Should multiple submissions become the norm in journals? Learned Publishing, 27(3):167–172, July 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nunn:2014:LSO

[1719] Emily Nunn and Stephen Pinfield. Lay summaries of open access journal articles: engaging with the general public on medical research. Learned Publishing, 27(3):173–184, July 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clark:2014:EIS

[1720] D. J. Clark, David Nicholas, and Hamid R. Jamali. Evaluating information seeking and use in the changing virtual world: the emerging role of Google Analytics. Learned Publishing, 27(3):185–194, July 2014. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xia:2014:ETI

[1721] Jingfeng Xia. An examination of two Indian megajournals. Learned Publishing, 27(3):195–200, July 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nambiar:2014:QAG

[1722] Remya Nambiar, Priyanka Tilak, and Clarinda Cerejo. Quality of author guidelines of journals in the biomedical and physical sciences. Learned Publishing, 27(3):201–206, July 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baverstock:2014:BOR

[1723] Alison Baverstock and Jackie Steinitz. Barriers and opportunities for research in publishing studies. Learned Publishing, 27(3):207–221, July 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kennison:2014:TSS

[1724] Rebecca Kennison and Lisa Norberg. Toward a scalable and sustainable approach to open access publishing and archiving for humanities and social sciences societies: a proposal. Learned Publishing, 27(3):223– 235, July 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xia:2014:IJP

[1725] Jingfeng Xia. An imbalanced journal publishing market. Learned Publishing, 27(3):236–238, July 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McIlroy:2014:BRa

[1726] Thad McIlroy. Book review. Learned Publishing, 27(3):239, July 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott-Lichter:2014:UDP

[1727] Diane Scott-Lichter. US DOE public access plan: in search of the right fit. Learned Publishing, 27(4):242–244, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lammey:2014:CTD

[1728] Rachael Lammey. CrossRef’s text and data mining services. Learned Publishing, 27(4):245–250, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jubb:2014:CCW

[1729] Michael Jubb. Communication or competition: what motivates researchers to write articles for journals? Learned Publishing, 27(4):251– 252, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tsou:2014:URE

[1730] Andrew Tsou, Jutta Schickore, and Cassidy R. Sugimoto. Unpublishable research: examining and organizing the ‘file drawer’. Learned Publishing, 27(4):253–267, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mehrpour:2014:HSF

[1731] Saeed Mehrpour and Yaser Khajavi. How to spot fake open access journals. Learned Publishing, 27(4):269–274, October 2014. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jung:2014:UGF

[1732] Kyu Won Jung, Elizabeth Wagerxyv, and Sun Huh. Uploading garbage files to the DOI system of a journal: a new type of misconduct. Learned Publishing, 27(4):275–280, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Campbell:2014:SAC

[1733] Bob Campbell. ‘The Stone Age came to an end not for the lack of stones … ’. Learned Publishing, 27(4):281–282, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shuang-Shuang:2014:CPC

[1734] Gai Shuang-Shuang, Liu Xue-Li, Zhang Shi-Le, and Liu Rui-Yuan. Comparing ’papers cited rates’ with other measures for science journal evaluation. Learned Publishing, 27(4):283–290, October 2014. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baverstock:2014:HDP

[1735] Alison Baverstock and Jackie Steinitz. How is the discipline of publishing studies accommodated within universities? Learned Publishing, 27(4): 291–300, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kulemeka:2014:PRP

[1736] Owen Kulemeka. The peer-review process in English-language advertising journals: a survey of US academics. Learned Publishing, 27(4): 301–313, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McIlroy:2014:BRb

[1737] Thad McIlroy. Book reviews. Learned Publishing, 27(4):315, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Woodhead:2014:BR

[1738] Martin Woodhead. Book reviews. Learned Publishing, 27(4):315–317, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2014:BR

[1739] Anthony Watkinson. Book reviews. Learned Publishing, 27(4):317–318, October 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Murphy:2014:DP

[1740] Fiona Murphy and Alice Meadows. Data publishing. Learned Publishing, 27(5):S1–S2, September 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Murphy:2014:DSP

[1741] Fiona Murphy. Data and scholarly publishing: the transforming landscape. Learned Publishing, 27(5):S3–S7, September 2014. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Treloar:2014:RDA

[1742] Andrew Treloar. The Research Data Alliance: globally co-ordinated action against barriers to data publishing and sharing. Learned Publishing, 27(5):S9–S13, September 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Callaghan:2014:PIS

[1743] Sarah Callaghan. Preserving the integrity of the scientific record: data citation and linking. Learned Publishing, 27(5):S15–S24, September 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Khodiyar:2014:AMA

[1744] Varsha K. Khodiyar, Karen A. Rowlett, and Rebecca N. Lawrence. Altmetrics as a means of assessing scholarly output. Learned Publishing, 27(5):S25–S32, September 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Reilly:2014:RDL

[1745] Susan K. Reilly. Rounding up the data: libraries pushing new frontiers. Learned Publishing, 27(5):S33–S34, September 2014. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Norman:2014:MDA

[1746] Hazel Norman. Mandating data archiving: experiences from the frontline. Learned Publishing, 27(5):S35–S38, September 2014. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2015:EYH

[1747] Pippa Smart. Editorial: 350 years and how are we doing? Learned Publishing, 28(1):3, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lara:2015:LRM

[1748] Kate Lara. The library’s role in the management and funding of open access publishing. Learned Publishing, 28(1):4–8, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lawson:2015:TCO

[1749] Stuart Lawson. ‘Total cost of ownership’ of scholarly communication: managing subscription and APC payments together. Learned Publishing, 28(1):9–13, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2015:PRS

[1750] David Nicholas, Anthony Watkinson, Hamid R. Jamali, Eti Herman, Carol Tenopir, Rachel Volentine, Suzie Allard, and Kenneth Levine. Peer review: still king in the digital age. Learned Publishing, 28(1):15–21, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nwagwu:2015:PNP

[1751] Williams Ezinwa Nwagwu and Obinna Ojemeni. Penetration of Nigerian predatory biomedical open access journals 2007–2012: a bibliometric study. Learned Publishing, 28(1):23–34, January 2015. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dodds:2015:CRC

[1752] Francis Dodds. Changes in the role of the commissioning editor in academic book publishing. Learned Publishing, 28(1):35–42, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baverstock:2015:WIE

[1753] Alison Baverstock, Robert Blackburn, and Marfuga Iskandarova. Who are the independent editors, how did they reach their role and what are their associated job satisfactions? Learned Publishing, 28(1):43–53, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Powell:2015:ROD

[1754] Andrea Powell and Michael Pearson. The reports of our death have been greatly exaggerated (with apologies to Mark Twain). Learned Publishing, 28(1):55–57, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hill:2015:ILO

[1755] Tom Hill. Identifying legitimate open access journals: some suggestions from a publisher. Learned Publishing, 28(1):59–62, January 2015. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Busher:2015:SSL

[1756] Casey Busher and Irene Kamotsky. Stories and statistics from libraryled publishing. Learned Publishing, 28(1):64–68, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xia:2015:PJT

[1757] Jingfeng Xia. Predatory journals and their article publishing charges. Learned Publishing, 28(1):69–74, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Phillpotts:2015:ETC

[1758] James Phillpotts, Tim Devenport, and Alison Mitchell. Evolution of the Transfer Code of Practice. Learned Publishing, 28(1):75–79, January 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Luther:2015:EL

[1759] Judy Luther. The evolving landscape. Learned Publishing, 28(2):82– 83, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bjork:2015:ENM

[1760] Bo-Christer Björk and Turid Hedlund. Emerging new methods of peer review in scholarly journals. Learned Publishing, 28(2):85–91, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tenopir:2015:SAS

[1761] Carol Tenopir, Donald W. King, Lisa Christian, and Rachel Volentine. Scholarly article seeking, reading, and use: a continuing evolution from print to electronic in the sciences and social sciences. Learned Publishing, 28(2):93–105, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Stone:2015:BRD

[1762] Graham Stone. The benefits of resource discovery for publishers: a librarian’s view. Learned Publishing, 28(2):106–113, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nwagwu:2015:CAP

[1763] Williams Ezinwa Nwagwu. Counterpoints about predatory open access and knowledge publishing in Africa. Learned Publishing, 28(2):114–122, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baverstock:2015:HRI

[1764] Alison Baverstock, Robert Blackburn, and Marfuga Iskandarova. How the role of the independent editor is changing in relation to traditional and self-publishing. Learned Publishing, 28(2):123–131, April 2015. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shuang-Shuang:2015:CEL

[1765] Gai Shuang-Shuang, Liu Xue-Li, Zhang Shi-Le, and Liu Rui-Yuan. The citation evolution law of papers published in the same year but different month. Learned Publishing, 28(2):132–139, April 2015. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wade:2015:EPC

[1766] Angela Mills Wade. European Publishers Council: making copyright work on the Web. Learned Publishing, 28(2):140–142, April 2015. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Winker:2015:PPP

[1767] Margaret Winker. The promise of post-publication peer review: how do we get there from here? Learned Publishing, 28(2):143–145, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mudrak:2015:JBA

[1768] Benjamin Mudrak. JournalGuide: bringing authors and journals together. Learned Publishing, 28(2):147–149, April 2015. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brand:2015:BAA

[1769] Amy Brand, Liz Allen, Micah Altman, Marjorie Hlava, and Jo Scott. Beyond authorship: attribution, contribution, collaboration, and credit. Learned Publishing, 28(2):151–155, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2015:BRa

[1770] Anthony Watkinson. Book reviews. Learned Publishing, 28(2):156, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2015:BRb

[1771] Anthony Watkinson. Book reviews. Learned Publishing, 28(2):156–157, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mattingley:2015:BR

[1772] Sue Mattingley. Book reviews. Learned Publishing, 28(2):157–158, April 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2015:TFS

[1773] Pippa Smart. Twenty-five shades of grey. Learned Publishing, 28(3):163– 165, July 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2015:LPN

[1774] Pippa Smart. Learned Publishing: new editorial team. Learned Publishing, 28(3):167–168, July 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2015:NWBa

[1775] David Nicholas, Eti Herman, Hamid Jamali, Blanca Rodríguez-Bravo, Cherifa Boukacem-Zeghmouri, Tom Dobrowolski, and Stephanie Pouchot. New ways of building, showcasing, and measuring scholarly reputation. Learned Publishing, 28(3):169–183, July 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Carden:2015:APC

[1776] Mark Carden. Academic publishing career paths — initial research and observations. Learned Publishing, 28(4):184–194, July 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gazni:2015:GNJ

[1777] Ali Gazni. Globalization of national journals: investigating the growth of international authorship. Learned Publishing, 28(4):195–204, July 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dodds:2015:UEU

[1778] Francis Dodds. Understanding end-users in academic book publishing. Learned Publishing, 28(4):205–214, July 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wu:2015:EII

[1779] Yang Wu and Qiang Zou. The ethical issues in instructions for authors of Chinese biomedical journals. Learned Publishing, 28(4):216–222, July 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smith:2015:FFU

[1780] Kevin L. Smith. Figuring on fair use. Learned Publishing, 28(4):225– 227, July 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Owen:2015:FDC

[1781] Lynette Owen. Fair dealing: a concept in UK copyright law. Learned Publishing, 28(4):229–231, July 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2015:AA

[1782] Anonymous. ALPSP Awards 2015. Learned Publishing, 28(4):234–235, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Luther:2015:RRC

[1783] Judy Luther. Recognizing roles and contributions. Learned Publishing, 28 (4):236–237, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sucha:2015:JPM

[1784] Ladislava Zbiejczuk Suchá and Jela Steinerová. Journal publishing models in the Czech Republic. Learned Publishing, 28(4):239–249, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bodaghi:2015:CIP

[1785] Nahid Bayat Bodaghi, S. A. Sanni, and A. N. Zainab. In competition with ISI: the perceptions of chief editors of Malaysian local journals. Learned Publishing, 28(4):251–260, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ossenblok:2015:BES

[1786] Truyken L. B. Ossenblok, Raf Guns, and Mike Thelwall. Book editors in the social sciences and humanities: an analysis of publication and collaboration patterns of established researchers in Flanders. Learned Publishing, 28(4):261–273, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Johnson:2015:SSA

[1787] Rob Johnson and Mattia Fosci. On shifting sands: assessing the financial sustainability of UK learned societies. Learned Publishing, 28(4):274–281, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bartholomaus:2015:EBP

[1788] Elke Bartholomäus, Sandy Goldbeck-Wood, Meike Sewering, and Christopher Baethge. Experiences with bilingual publishing: surveys of authors and editors. Learned Publishing, 28(4):283–291, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lloyd:2015:AMO

[1789] Tim Lloyd. Access management: the overlooked but critical enabler. Learned Publishing, 28(4):292–297, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Glahn:2015:RDO

[1790] Peggy Glahn. Reveal Digital: an open access model empowering libraries to become publishers. Learned Publishing, 28(4):299–302, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Payne:2015:BRB

[1791] David Payne. Book review: The BMJ online: lessons from a 20-year digital presence. Learned Publishing, 28(4):304–308, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Paglione:2015:DES

[1792] Laura Dorival Paglione and Rebecca Naomi Lawrence. Data exchange standards to support and acknowledge peer-review activity. Learned Publishing, 28(4):309–316, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Priestley:2015:UEM

[1793] Owen Priestley. User experience methodology: from the physical to the emotional. Learned Publishing, 28(4):317–320, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Schopfel:2015:OAR

[1794] Joachim Schöpfel. Open access — the rise and fall of a community-driven model of scientific communication. Learned Publishing, 28(4):321–325, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cabanac:2015:DLI

[1795] Guillaume Cabanac. On the dead link issue in academic papers: Dear Editor. Learned Publishing, 28(4):326, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2015:NWBb

[1796] David Nicholas, Eti Herman, Hamid Jamali, Blanca Rodríguez-Bravo, Cherifa Boukacem-Zeghmouri, Tom Dobrowolski, and Stephanie Pouchot. New ways of building, showcasing, and measuring scholarly reputation. Learned Publishing, 28(4):327, October 2015. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:TCa

[1797] Anonymous. Table of contents. Learned Publishing, 29(1):1, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:PRE

[1798] Anonymous. Peer review: an expensive business. Learned Publishing, 29(1):3–4, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Willis:2016:WDP

[1799] Michael Willis. Why do peer reviewers decline to review manuscripts? A study of reviewer invitation responses. Learned Publishing, 29(1):5–7, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bjork:2016:PRM

[1800] Bo-Christer Björk and Paul Catani. Peer review in megajournals compared with traditional scholarly journals: Does it make a difference? Learned Publishing, 29(1):9–12, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jubb:2016:PRC

[1801] Michael Jubb. Peer review: the current landscape and future trends. Learned Publishing, 29(1):13–21, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thompson:2016:JPR

[1802] Stephen R. Thompson, Julie Agel, and Elena Losina. The JBJS Peer-Review Scoring Scale: a valid, reliable instrument for measuring the quality of peer review reports. Learned Publishing, 29(1):23–25, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hartley:2016:PCD

[1803] James Hartley, John Cowan, and Nick Rushby. Peer choice — does reviewer self-selection work? Learned Publishing, 29(1):27–29, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2016:CRS

[1804] David Nicholas, Jie Xu, Lifang Xu, Jing Su, and Anthony Watkinson. Chinese researchers, scholarly communication behaviour and trust. Learned Publishing, 29(1):31–38, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Warne:2016:RRS

[1805] Verity Warne. Rewarding reviewers — sense or sensibility? A Wiley study explained. Learned Publishing, 29(1):41–50, January 2016. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Murphy:2016:UPR

[1806] Fiona Murphy. An update on peer review and research data. Learned Publishing, 29(1):51–53, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Salem:2016:PSI

[1807] Ruwaida M. Salem, Natalie M. Culbertson, and Alison O’Connell. Process for selecting and implementing a manuscript management system: Experiences of a new peer-reviewed journal. Learned Publishing, 29(1): 55–59, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hanson:2016:EAO

[1808] Brooks Hanson, Rebecca Lawrence, Alice Meadows, and Laura Paglione. Early adopters of ORCID functionality enabling recognition of peer review: Two brief case studies. Learned Publishing, 29(1):60–63, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baverstock:2016:PRS

[1809] Alison Baverstock. Is peer review still the content industry’s upper house? Learned Publishing, 29(1):65–68, January 2016. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kennison:2016:BFR

[1810] Rebecca Kennison. Back to the future: (re)turning from peer review to peer engagement. Learned Publishing, 29(1):69–71, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:JIa

[1811] Anonymous. Journal information. Learned Publishing, 29(1):72, January 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:TCb

[1812] Anonymous. Table of contents. Learned Publishing, 29(2):73, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Luther:2016:DAA

[1813] Judy Luther. Discovery in an age of artificial intelligence. Learned Publishing, 29(2):75–76, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ghane:2016:HAC

[1814] Mohammad Reza Ghane. How accurate are cited references in Iranian peer-reviewed journals? Learned Publishing, 29(2):77–82, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Morrison:2016:SSL

[1815] Heather Morrison. Small scholar-led scholarly journals: Can they survive and thrive in an open access future? Learned Publishing, 29(2):83– 88, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Castella:2016:CFL

[1816] Candela Ollé Castellà, Alexandre López-Borrull, and Ernest Abadal. The challenges facing library and information science journals: Editors’ opinions. Learned Publishing, 29(2):89–94, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Greco:2016:DSP

[1817] Albert N. Greco, Robert M. Wharton, and Amy Brand. Demographics of scholarly publishing and communication professionals. Learned Publishing, 29(2):97–101, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gazni:2016:ISP

[1818] Ali Gazni and Zahra Ghaseminik. Internationalization of scientific publishing over time: Analysing publishers and fields differences. Learned Publishing, 29(2):103–111, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Foxall:2016:SOA

[1819] Katie Foxall and Audrey Nailor. Sustainability in open access publishing: the e cancer case study. Learned Publishing, 29(2):113–117, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Griffin:2016:ARI

[1820] Jonathan Griffin. Access to Research: an innovative and successful initiative by the UK publishing industry. Learned Publishing, 29(2):119–123, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kudlow:2016:THS

[1821] Paul Kudlow, Alan Rutledge, Aviv Shachak, Roger S. McIntyre, and Gunther Eysenbach. TrendMD: Helping scholarly content providers reach larger and more targeted audiences. Learned Publishing, 29(2):125–129, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bjork:2016:OAM

[1822] Bo-Christer Björk. The open access movement at a crossroad: Are the big publishers and academic social media taking over? Learned Publishing, 29(2):131–134, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kane:2016:PBE

[1823] Lauren Kane and Alice Meadows. From the podium to the boardroom: Encouraging gender parity in scholarly publishing. Learned Publishing, 29(2):135–138, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:JIb

[1824] Anonymous. Journal information. Learned Publishing, 29(2):140, April 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:TCc

[1825] Anonymous. Table of contents. Learned Publishing, 29(3):141, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2016:PFS

[1826] Pippa Smart. Are publishers failing as a service industry? Learned Publishing, 29(3):143–144, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lin:2016:IPP

[1827] Songqing Lin and Lijuan Zhan. International publishing partnerships for China’s English language journals and financial security. Learned Publishing, 29(3):145–153, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dagiene:2016:HRM

[1828] Eleonora Dagienė and Danutė Krapavickaitė. How researchers manage their academic activities. Learned Publishing, 29(3):155–163, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Abrizah:2016:DAA

[1829] Abdullah Abrizah, David Nicholas, Abdullah Noorhidawati, M. K. Yanti Idaya Aspura, and Fathiah Badawi. Not so different after all: Malaysian researchers’ cross-discipline view of quality and trustworthiness in citation practices. Learned Publishing, 29(3):165–172, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2016:RRU

[1830] David Nicholas, David Clark, and Eti Herman. ResearchGate: Reputation uncovered. Learned Publishing, 29(3):173–182, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ha:2016:HOU

[1831] Louisa Ha, Weiwei Jiang, Chang Bi, Ruonan Zhang, Tao Zhang, and Xiaoli Wen. How online usage of subscription-based journalism and mass communication research journal articles predicts citations. Learned Publishing, 29(3):183–192, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jeong:2016:ASK

[1832] Geum-Hee Jeong and Sun Huh. Analysis of the suitability of the Korean Federation of Science and Technology Societies journal evaluation tool. Learned Publishing, 29(3):193–198, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wade:2016:RMA

[1833] Alex D. Wade and Kuansan Wang. The rise of the machines: Artificial intelligence meets scholarly content. Learned Publishing, 29(3):201–205, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Williams:2016:BOF

[1834] Vicky Williams. Broader outreach: Fad or future? Learned Publishing, 29(3):207–209, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kulkarni:2016:WCP

[1835] Sneha Kulkarni. What causes peer review scams and how can they be prevented? Learned Publishing, 29(3):211–213, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson:2016:WTV

[1836] Rick Anderson. Why are there virtually no mandatory open access policies at American universities? Learned Publishing, 29(3):215–218, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Weilenmann:2016:SCB

[1837] Anne-Katharina Weilenmann. Scientific communication — on the brink of the next evolution? Learned Publishing, 29(3):219–221, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:JIc

[1838] Anonymous. Journal information. Learned Publishing, 29(3):224, July 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:TCd

[1839] Anonymous. Table of contents. Learned Publishing, 29(4):225, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Luther:2016:CG

[1840] Judy Luther. Changing of the guard. Learned Publishing, 29(4):227–228, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gonzalez-Valiente:2016:RAE

[1841] Carlos Luis González-Valiente, Josmel Pacheco-Mendoza, and Ricardo Arencibia-Jorge. A review of altmetrics as an emerging discipline for research evaluation. Learned Publishing, 29(4):229–238, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Claudio-Gonzalez:2016:CAB

[1842] Melba G. Claudio-González, Montserrat Martín-Baranera, and Anna Villarroya. A cluster analysis of the business models of Spanish journals. Learned Publishing, 29(4):239–248, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zeinolabedini:2016:CHE

[1843] Maryam Zeinolabedini and Javad Gholami. Colleagues helping each other to publish in medical sciences: Iranian field specialists’ convenience editing strategies. Learned Publishing, 29(4):249–258, October 2016. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Laakso:2016:HSB

[1844] Mikael Laakso, David Solomon, and Bo-Christer Björk. How subscription-based scholarly journals can convert to open access: a review of approaches. Learned Publishing, 29(4):259–269, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gebhard:2016:WMM

[1845] Caitlin Gebhard and Bruce Rosenblum. Wrangling math from Microsoft Word into JATS XML workflows. Learned Publishing, 29(4):271–279, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (elec-

tronic). URL https://jats.nlm.nih.gov/articleauthoring/tag-
library/1.1.

Anonymous:2016:AAR

[1846] Anonymous. ALPSP Awards results. Learned Publishing, 29(4):280–281, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thew:2016:MCD

[1847] David Thew. Mentoring as a career development tool: a universal resource or only for the ‘anointed’? Learned Publishing, 29(4):283–290, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Holmes:2016:ESH

[1848] Edith Holmes. Economies of small: How a single-journal publisher remains independent and grows. Learned Publishing, 29(4):293–299, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Martin:2016:HEP

[1849] Alexandre Martin and Tristan Martin. A not-so-harmless experiment in predatory open access publishing. Learned Publishing, 29(4):301–305, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bass:2016:IWP

[1850] Julian M. Bass. Improving writing processes using lean and Kanban. Learned Publishing, 29(4):307–310, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:JId

[1851] Anonymous. Journal information. Learned Publishing, 29(4):312, October 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:TCe

[1852] Anonymous. Table of contents. Learned Publishing, 29(S1):313, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cond:2016:UPR

[1853] Anthony Cond and Samantha Rayner. The University Press Redux: Introduction. Learned Publishing, 29(S1):314–315, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hill:2016:NEU

[1854] Mandy Hill. The new era of the university press: the critical role of established presses. Learned Publishing, 29(S1):316–319, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lockett:2016:NUP

[1855] Andrew Lockett and Lara Speicher. New university presses in the UK: Accessing a mission. Learned Publishing, 29(S1):320–329, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mudditt:2016:PPF

[1856] Alison Mudditt. The past, present, and future of American university presses: a view from the left coast. Learned Publishing, 29(S1):330– 334, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bargheer:2016:BSF

[1857] Margo Bargheer and Jutta Pabst. ‘Being small is not a fault’: Making sense of the newer generation of German-language university presses. Learned Publishing, 29(S1):335–341, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2016:WMM

[1858] Charles Watkinson. Why marriage matters: a North American perspective on press/library partnerships. Learned Publishing, 29(S1):342–347, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kember:2016:WP

[1859] Sarah Kember. Why publish? Learned Publishing, 29(S1):348–353, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shaw:2016:VMP

[1860] Alison Shaw. Vision, mission, passion, and luck: the creation of a university press. Learned Publishing, 29(S1):354–359, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Llewellyn:2016:UPE

[1861] Mark Llewellyn and Sumi David. On university pressing and evidence pu(bli)shing: the view from a funder. Learned Publishing, 29(S1):360– 365, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hill:2016:MFS

[1862] Steven A. Hill. Making the future of scholarly communications. Learned Publishing, 29(S1):366–370, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2016:JIe

[1863] Anonymous. Journal information. Learned Publishing, 29(S1):371, ???? 2016. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:TCa

[1864] Anonymous. Table of contents. Learned Publishing, 30(1):1, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Conrad:2017:PNW

[1865] Lettie Y. Conrad. Pathways to a new way of working: Discoverability. Learned Publishing, 30(1):3–4, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tenopir:2017:NSI

[1866] Carol Tenopir, Suzie Allard, Lisa Christian, Robert Anderson, Suzan Ali-Saleh, Dave Nicholas, Anthony Watkinson, and Hazel Woodward. No scholar is an island: the impact of sharing in the work life of scholars. Learned Publishing, 30(1):5–17, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2017:WHE

[1867] David Nicholas, Cherifa Boukacem-Zeghmouri, Blanca Rodríguez-Bravo, Jie Xu, Anthony Watkinson, A. Abrizah, Eti Herman, and Marzena Świgoń. Where and how early career researchers find scholarly information. Learned Publishing, 30(1):19–29, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Conrad:2017:HDF

[1868] Lettie Y. Conrad. Headlines from the discovery files: Key publications on scholarly content discoverability. Learned Publishing, 30(1):31–37, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Doshi:2017:FDH

[1869] Nisha Doshi. ‘Finding’ and ‘discovering’: How understanding researchers’ search and discovery behaviour informed the building of Cambridge Core. Learned Publishing, 30(1):39–44, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Varnum:2017:BHO

[1870] Kenneth J. Varnum. A brief history of the Open Discovery Initiative. Learned Publishing, 30(1):45–48, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wickes:2017:BDR

[1871] Abigail Wickes. Building discovery relationships and expertise at Oxford University Press. Learned Publishing, 30(1):49–53, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fallas:2017:EBD

[1872] Andrea Fallas and Suzanne Tatham. Evidence-based design for discoverability: a case study. Learned Publishing, 30(1):55–64, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hodgson:2017:CPG

[1873] Alex Hodgson and Lucas Schlager. Closing the PDF gap: ReadCube’s experiments in reader-focused design. Learned Publishing, 30(1):65–69, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhu:2017:SPW

[1874] Julie Zhu. Should publishers work with library discovery technologies and what can they do? Learned Publishing, 30(1):71–80, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dove:2017:FDW

[1875] John G. Dove. Full discovery: What is the publisher’s role? Learned Publishing, 30(1):81–86, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Akeroyd:2017:DST

[1876] John Akeroyd. Discovery systems: Are they now the library? Learned Publishing, 30(1):87–89, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sommer:2017:CIR

[1877] David Sommer. Creating impact resonance through a culture of Holistic Discoverability. Learned Publishing, 30(1):91–95, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:LPR

[1878] Anonymous. Learned Publishing Reviewers. Learned Publishing, 30(1): 97, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:JIa

[1879] Anonymous. Journal information. Learned Publishing, 30(1):100, January 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:TCb

[1880] Anonymous. Table of contents. Learned Publishing, 30(2):101, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2017:PJR

[1881] Pippa Smart. Predatory journals and researcher needs. Learned Publishing, 30(2):103–105, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Thelwall:2017:DMP

[1882] Mike Thelwall. Does Mendeley provide evidence of the educational value of journal articles? Learned Publishing, 30(2):107–113, April 2017. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Liu:2017:CRN

[1883] Weishu Liu. The changing role of non-English papers in scholarly communication: Evidence from Web of Science’s three journal citation indexes. Learned Publishing, 30(2):115–123, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kienc:2017:APC

[1884] Witold Kieńć. Authors from the periphery countries choose open access more often. Learned Publishing, 30(2):125–131, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ameen:2017:PQT

[1885] Kanwal Ameen. Practices of quality and trustworthiness in scholarly communication: a case from Pakistan. Learned Publishing, 30(2):133– 142, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Melero:2017:OSS

[1886] Remedios Melero, Mikael Laakso, and Miguel Navas-Fernández. Openness of Spanish scholarly journals as measured by access and rights. Learned Publishing, 30(2):143–155, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2017:ECRa

[1887] David Nicholas, Anthony Watkinson, Cherifa Boukacem-Zeghmouri, Blanca Rodríguez-Bravo, Jie Xu, Abdullah Abrizah, Marzena Świgoń, and Eti Herman. Early career researchers: Scholarly behaviour and the prospect of change. Learned Publishing, 30(2):157–166, April 2017. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hartley:2017:DDD

[1888] James Hartley and Guillaume Cabanac. The delights, discomforts, and downright furies of the manuscript submission process. Learned Publishing, 30(2):167–172, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bjork:2017:GGB

[1889] Bo-Christer Björk. Gold, green, and black open access. Learned Publishing, 30(2):173–175, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2017:JMT

[1890] Anthony Watkinson. Jack Meadows (1934–2016): a tribute to a great information scientist. Learned Publishing, 30(2):177–178, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:JIb

[1891] Anonymous. Journal information. Learned Publishing, 30(2):180, April 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:TCc

[1892] Anonymous. Table of contents. Learned Publishing, 30(3):181, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2017:DIA

[1893] Pippa Smart and Lettie Conrad. Diversity, inclusivity, and accessibility. Learned Publishing, 30(3):183–184, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Borrego:2017:IRV

[1894] Ángel Borrego. Institutional repositories versus ResearchGate: the depositing habits of Spanish researchers. Learned Publishing, 30(3):185– 192, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lin:2017:PAS

[1895] Wen-Yau Cathy Lin. The performance of Asian S&T journals in international citation indicators. Learned Publishing, 30(3):193–204, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2017:ECRb

[1896] David Nicholas, Blanca Rodríguez-Bravo, Anthony Watkinson, Cherifa Boukacem-Zeghmouri, Eti Herman, Jie Xu, Abdullah Abrizah, and Marzena Świgoń. Early career researchers and their publishing and authorship practices. Learned Publishing, 30(3):205–217, July 2017. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Schmaling:2017:GDP

[1897] Karen B. Schmaling and Arthur W. Blume. Gender differences in providing peer review to two behavioural science journals, 2006–2015. Learned Publishing, 30(3):221–225, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dambha-Miller:2017:APS

[1898] Hajira Dambha-Miller and Roger Jones. An appealing prospect? A survey into the numbers, outcomes, and editorial policies for appeals of rejected biomedical manuscripts. Learned Publishing, 30(3):227–231, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nauman:2017:LCT

[1899] Sarwat Nauman. Lack of critical thinking skills leading to research crisis in developing countries: a case of Pakistan. Learned Publishing, 30(3): 233–236, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brown:2017:NAN

[1900] David Brown. Are new approaches needed for developing long-term strategies for STEM information? Learned Publishing, 30(3):237–243, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Manley:2017:ISE

[1901] Stewart Manley. Improve student-edited law journals: Eliminate the acceptance period. Learned Publishing, 30(3):245–250, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:JIc

[1902] Anonymous. Journal information. Learned Publishing, 30(3):252, July 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:TCd

[1903] Anonymous. Table of contents. Learned Publishing, 30(4):253, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2017:RWR

[1904] Pippa Smart. Readers — what readers? Learned Publishing, 30(4): 255–257, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Joung:2017:SCO

[1905] Kyoung Hee Joung and Jennifer Rowley. Scholarly communication and open access: Perspectives from Korea. Learned Publishing, 30(4):259– 267, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rodriguez-Bravo:2017:PRE

[1906] Blanca Rodríguez-Bravo, David Nicholas, Eti Herman, Chérifa Boukacem-Zeghmouri, Anthony Watkinson, Jie Xu, Abdullah Abrizah, and Marzena Świgoń. Peer review: the experience and views of early career researchers. Learned Publishing, 30(4):269–277, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:AA

[1907] Anonymous. ALPSP Awards 2017. Learned Publishing, 30(4):278–279, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Forrester:2017:NWS

[1908] Amy Forrester, Bo-Christer Björk, and Carol Tenopir. New web services that help authors choose journals. Learned Publishing, 30(4):281–287, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wijewickrema:2017:JSC

[1909] Manjula Wijewickrema and Vivien Petras. Journal selection criteria in an open access environment: a comparison between the medicine and social sciences. Learned Publishing, 30(4):289–300, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wakeling:2017:OAMa

[1910] Simon Wakeling, Valérie Spezi, Jenny Fry, Claire Creaser, Stephen Pinfield, and Peter Willett. Open access megajournals: the publisher perspective (part 1: Motivations). Learned Publishing, 30(4):301–311, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wakeling:2017:OAMb

[1911] Simon Wakeling, Valérie Spezi, Claire Creaser, Jenny Fry, Stephen Pinfield, and Peter Willett. Open access megajournals: the publisher perspective (part 2: Operational realities). Learned Publishing, 30(4):313– 322, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Green:2017:WFP

[1912] Toby Green. We’ve failed: Pirate black open access is trumping green and gold and we must change our approach. Learned Publishing, 30 (4):325–329, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jubb:2017:SCA

[1913] Michael Jubb and Richard Fisher. The supply chain for academic books: Why is it so complex and how might we drive change? Learned Publishing, 30(4):331–335, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2017:JId

[1914] Anonymous. Journal information. Learned Publishing, 30(4):336, October 2017. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:TCa

[1915] Anonymous. Table of contents. Learned Publishing, 31(1):1, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Conrad:2018:MAM

[1916] Lettie Y. Conrad and Bill Kasdorf. Making accessibility more accessible to publishers. Learned Publishing, 31(1):3–4, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McNaught:2018:EBP

[1917] Alistair McNaught, Ruth MacMullen, Sue Smith, and Vicky Dobson. Evaluating e-book platforms: Lessons from the e-book accessibility audit. Learned Publishing, 31(1):5–10, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kasdorf:2018:WAH

[1918] Bill Kasdorf. Why accessibility is hard and how to make it easier: Lessons from publishers. Learned Publishing, 31(1):11–18, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Trimble:2018:AJB

[1919] Lauren Trimble. Accessibility at JSTOR: From box-checking to a more inclusive and sustainable future. Learned Publishing, 31(1):21–24, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Turner:2018:BGL

[1920] Brad Turner. Benetech global literacy services: Working towards a ‘born accessible’ world. Learned Publishing, 31(1):25–29, January 2018. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

House:2018:TUA

[1921] Emma House, Richard Orme, and Mark Bide. Towards universal accessibility: The UK policy landscape and supporting technology. Learned Publishing, 31(1):31–34, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bowes:2018:OCA

[1922] Frederick Bowes III. An overview of content accessibility issues experienced by educational publishers. Learned Publishing, 31(1):35–38, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Axelrod:2018:MMA

[1923] Jamie Axelrod. Making materials accessible to students in higher education institutes: Institutional obligations, methods of compliance, and recommendations for future action. Learned Publishing, 31(1):39–44, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rothberg:2018:PAS

[1924] Madeleine Rothberg. Publishing with accessibility standards from the inside out. Learned Publishing, 31(1):45–47, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Iglesias:2018:BMF

[1925] Violaine Iglesias. Beyond the mandates: the far-reaching benefits of multimedia accessibility. Learned Publishing, 31(1):49–54, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Burges:2018:CDI

[1926] Jo Burges. CAL Download — an innovative approach to making books more accessible. Learned Publishing, 31(1):57–62, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Manis:2018:SFB

[1927] Caroline Manis and Huw Alexander. The secrets of failing better: Accessible publishing at SAGE. A case study. Learned Publishing, 31(1): 63–68, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gies:2018:SAJ

[1928] Ted Gies. The ScienceDirect accessibility journey: a case study. Learned Publishing, 31(1):69–76, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Abbott:2018:HPH

[1929] George Abbott. How publishing has helped and hindered me: Experiences and advice from a blind reader and publisher. Learned Publishing, 31(1):79–82, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kerscher:2018:DYB

[1930] George Kerscher. Do you have a broken link in your accessibility chain? Learned Publishing, 31(1):83–85, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:TOR

[1931] Anonymous. Thanking our reviewers. Learned Publishing, 31(1):87, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:JIa

[1932] Anonymous. Journal information. Learned Publishing, 31(1):88, January 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:TCb

[1933] Anonymous. Table of contents. Learned Publishing, 31(2):89, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2018:QAA

[1934] Pippa Smart. Question all assumptions: Learned Publishing at 30. Learned Publishing, 31(2):91–94, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2018:MSR

[1935] Yun Zhang, Weina Hua, and Shunbo Yuan. Mapping the scientific research on open data: a bibliometric review. Learned Publishing, 31(2): 95–106, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Liu:2018:OAP

[1936] Weishu Liu and Yanchao Li. Open access publications in sciences and social sciences: a comparative analysis. Learned Publishing, 31(2):107– 119, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Soheili:2018:EIB

[1937] Faramarz Soheili, Ali Akbar Khasseh, Afshin Mousavi Chelak, and Mohammad Tavakolizadeh-Ravari. An evaluation of information behaviour studies through the Scholarly Capital Model. Learned Publishing, 31 (2):121–129, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nash:2018:CFL

[1938] Joshua R. Nash, Rafael J. Araújo, and Geoffrey S. Shideler. Contributing factors to long-term citation count in marine and freshwater biology articles. Learned Publishing, 31(2):131–139, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kurt:2018:WDA

[1939] Serhat Kurt. Why do authors publish in predatory journals? Learned Publishing, 31(2):141–147, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wang:2018:DOI

[1940] Weilang Wang, Lvxiang Deng, Bin You, Ping Zhang, and Yifeng Chen. Digital object identifier and its use in core Chinese academic journals: a Chinese perspective. Learned Publishing, 31(2):149–154, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Upshall:2018:CHI

[1941] Michael Upshall. Combining human input with machine learning: a case study. Learned Publishing, 31(2):157–160, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dodds:2018:FAP

[1942] Francis Dodds. The future of academic publishing: Revolution or evolution? Learned Publishing, 31(2):163–168, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hartley:2018:SOC

[1943] James Hartley. Some observations on the current state of book reviewing in the social sciences. Learned Publishing, 31(2):169–171, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wilcock:2018:PED

[1944] Jacob Wilcock. Putting an end to download-and-go: the website’s role in a content marketing ecosystem. Learned Publishing, 31(2):173–179, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Eriksson:2018:TST

[1945] Stefan Eriksson and Gert Helgesson. Time to stop talking about ‘predatory journals’. Learned Publishing, 31(2):181–183, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:JIb

[1946] Anonymous. Journal information. Learned Publishing, 31(2):184, April 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:TCc

[1947] Anonymous. Table of contents. Learned Publishing, 31(3):185, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2018:CCQ

[1948] Pippa Smart. Collection, curation, and quality: the editor’s responsibility. Learned Publishing, 31(3):186–188, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gaston:2018:WIR

[1949] Thomas Gaston and Pippa Smart. What influences the regional diversity of reviewers: a study of medical and agricultural/biological sciences journals. Learned Publishing, 31(3):189–197, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Somoza-Fernandez:2018:JCE

[1950] Marta Somoza-Fernández, Josep-Manuel Rodríguez-Gairín, and Cristóbal Urbano. Journal coverage of the Emerging Sources Citation Index. Learned Publishing, 31(3):199–204, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wang:2018:GQC

[1951] Yuandi Wang, Yu Chen, Meijun Liu, and Ruifeng Hu. Growth and quality of Chinese journals from 1949 to 2013. Learned Publishing, 31 (3):205–214, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Liang:2018:SAS

[1952] Yan Liang. Should authors suggest reviewers? A comparative study of the performance of author-suggested and editor-selected reviewers at a biological journal. Learned Publishing, 31(3):216–221, July 2018. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shehata:2018:WAS

[1953] Ahmed Maher Khafaga Shehata and Mohammed Fathy Mahmoud Elgllab. Where Arab social science and humanities scholars choose to publish: Falling in the predatory journals trap. Learned Publishing, 31(3): 222–229, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sondervan:2018:SOA

[1954] Jeroen Sondervan and Fleur Stigter. Sustainable open access for scholarly journals in 6 years — the incubator model at Utrecht University Library Open Access Journals. Learned Publishing, 31(3):230–234, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dunford:2018:KIA

[1955] Robin Dunford and Bruce Rosenblum. Keeping it authentic: Reconciling ORCID iDs gathered at submission with the author manuscript. Learned Publishing, 31(3):236–240, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bisaccio:2018:CJW

[1956] Michael Bisaccio. Cabells’ Journal Whitelist and Blacklist: Intelligent data for informed journal evaluations. Learned Publishing, 31(3):243– 248, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2018:WPC

[1957] David Nicholas, Anthony Watkinson, Abdullah Abrizah, Cherifa Boukacem-Zeghmouri, Jie Xu, Blanca Rodríguez-Bravo, Marzena Świgoń, and Eti Herman. What publishers can take away from the latest early career researcher research. Learned Publishing, 31(3):249– 253, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rhoods:2018:FGR

[1958] Samira Rhoods and Anca Babor. The future of global research: a case study on the use of scenario planning in the publishing industry. Learned Publishing, 31(3):254–260, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Liu:2018:EPH

[1959] Weishu Liu, Fang Liu, Chao Zuo, and Junwen Zhu. The effect of publishing a highly cited paper on a journal’s impact factor: a case study of the Review of Particle Physics. Learned Publishing, 31(3):261–266, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Helgesson:2018:RIR

[1960] Gert Helgesson and Stefan Eriksson. Revise the ICMJE Recommendations regarding authorship responsibility! Learned Publishing, 31(3):

267–269, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dodds:2018:CCL

[1961] Francis Dodds. The changing copyright landscape in academic publishing. Learned Publishing, 31(3):270–275, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:JIc

[1962] Anonymous. Journal information. Learned Publishing, 31(3):276, July 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:TCd

[1963] Anonymous. Table of contents. Learned Publishing, 31(4):329, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2018:ST

[1964] Pippa Smart. A sting in the tail? Learned Publishing, 31(4):331–333, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Montgomery:2018:GBD

[1965] Lucy Montgomery, Cameron Neylon, Alkim Ozaygen, and Tama Leaver. Getting the best out of data for open access monograph presses: a case study of UCL Press. Learned Publishing, 31(4):335–344, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Boukacem-Zeghmouri:2018:FPA

[1966] Chérifa Boukacem-Zeghmouri, Hans Dillaerts, Thierry Lafouge, Pascal Bador, and Aude Sauer-Avargues. French publishing attitudes in the open access era: the case of mathematics, biology, and computer science. Learned Publishing, 31(4):345–354, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Maniati:2018:SPE

[1967] Mahmood Maniati and Alireza Jalilifar. Strategies for publishing in English journals: a study of the perceptions of Iranian scholars. Learned Publishing, 31(4):355–365, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:AA

[1968] Anonymous. ALPSP Awards 2018. Learned Publishing, 31(4):366–367, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Vrana:2018:ECS

[1969] Radovan Vrana. Editorial challenges in a small scientific community: Study of Croatian editors. Learned Publishing, 31(4):369–374, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Scott:2018:CDA

[1970] James R. Scott, Mark Gibson, and Rebecca S. Benner. Contemporary drug advertising in leading US medical journals. Learned Publishing, 31 (4):375–380, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Merga:2018:ECE

[1971] Margaret K. Merga, Shannon Mason, and Julia Morris. Early career experiences of navigating journal article publication: Lessons learned using an autoethnographic approach. Learned Publishing, 31(4):381–389, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nganji:2018:AAP

[1972] Julius T. Nganji. An assessment of the accessibility of PDF versions of selected journal articles published in a WCAG 2.0 era (2014–2018). Learned Publishing, 31(4):391–401, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xia:2018:AJI

[1973] Jingfeng Xia and Megan P. Smith. Alternative journal impact factors in open access publishing. Learned Publishing, 31(4):403–411, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lamanna:2018:MRI

[1974] Camillo Lamanna and Stevan Bruijns. Measuring regional impact: the case for bigger data. Learned Publishing, 31(4):413–416, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kessler:2018:WPC

[1975] Rachel Kessler. Whitepaper: Practical challenges for researchers in data sharing: Review. Learned Publishing, 31(4):417–419, October 2018. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hill:2018:BRR

[1976] Tom Hill. Blockchain for research: Review. Learned Publishing, 31(4): 421–422, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic). Refers to an article by Joris van Rossum in Digital Science (doi:10.6084/m9.figshare.5607778.v1).

Anonymous:2018:JId

[1977] Anonymous. Journal information. Learned Publishing, 31(4):424, October 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:TCe

[1978] Anonymous. Table of contents. Learned Publishing, 31(S1):277, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cond:2018:FFA

[1979] Anthony Cond and Lara Speicher. ‘Feisty and febrile academic publishing’: The University Press Redux II. Learned Publishing, 31(S1):278–279, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2018:AEE

[1980] Charles Watkinson. The academic eBook ecosystem reinvigorated: a perspective from the USA. Learned Publishing, 31(S1):280–287, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Luescher:2018:AUP

[1981] Thierry M. Luescher and François van Schalkwyk. African university presses and the institutional logic of the knowledge commons. Learned Publishing, 31(S1):288–298, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mounier:2018:PFB

[1982] Pierre Mounier. ‘Publication favela’ or bibliodiversity? Open access publishing viewed from a European perspective. Learned Publishing, 31(S1):299–305, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brand:2018:ECP

[1983] Amy Brand. Embracing change: The MIT Press and the future of university-based publishing. Learned Publishing, 31(S1):306–310, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bayer:2018:LST

[1984] Lisa Bayer. Let’s stay together: a survey of relationships between American university presses and their host institutions. Learned Publishing, 31(S1):311–314, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wright:2018:EUP

[1985] Timothy E. Wright. Edinburgh University Press: From a department to a subsidiary business. Learned Publishing, 31(S1):315–317, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Winters:2018:WDA

[1986] Jane Winters. What does an author want from a university publisher? Learned Publishing, 31(S1):318–322, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hill:2018:DAB

[1987] Steven A. Hill. Decoupling the academic book. Learned Publishing, 31(S1):323–327, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2018:JIe

[1988] Anonymous. Journal information. Learned Publishing, 31(S1):328, September 2018. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:TCa

[1989] Anonymous. Table of contents. Learned Publishing, 32(1):1, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Plume:2019:BFB

[1990] Andrew Plume. Bring the facts, bust the myths. Learned Publishing, 32(1):3–5, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pollock:2019:OAM

[1991] Dan Pollock and Ann Michael. Open access mythbusting: Testing two prevailing assumptions about the effects of open access adoption. Learned Publishing, 32(1):7–12, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Green:2019:OAA

[1992] Toby Green. Is open access affordable? Why current models do not work and why we need internet-era transformation of scholarly communications. Learned Publishing, 32(1):13–25, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hartley:2019:DWN

[1993] John Hartley, Jason Potts, Lucy Montgomery, Ellie Rennie, and Cameron Neylon. Do we need to move from communication technology to user community? A new economic model of the journal as a club. Learned Publishing, 32(1):27–35, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McGlinchey:2019:DJA

[1994] Noel McGlinchey, Tom Hunter, Jack Bromley, Ruth Fisher, Anna Debiec-Waszak, and Thomas Gaston. Do Journal Administrators solve the reviewer assignment problem as well as editors? Consideration of reviewer rigour and timeliness. Learned Publishing, 32(1):37–46, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2019:HPP

[1995] Pippa Smart and Thomas Gaston. How prevalent are plagiarized submissions? Global survey of editors. Learned Publishing, 32(1):47–56, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Frandsen:2019:WDR

[1996] Tove Faber Frandsen. Why do researchers decide to publish in questionable journals? A review of the literature. Learned Publishing, 32 (1):57–62, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Drummond:2019:DRS

[1997] Chris Drummond. Is the drive for reproducible science having a detrimental effect on what is published? Learned Publishing, 32(1):63–69, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Allen:2019:HCW

[1998] Liz Allen, Alison O’Connell, and Veronique Kiermer. How can we ensure visibility and diversity in research contributions? How the Contributor Role Taxonomy (CRediT) is helping the shift from authorship to contributorship. Learned Publishing, 32(1):71–74, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McGonagle-OConnell:2019:CWT

[1999] Alison McGonagle-O’Connell and Kristen Ratan. Can we transform scholarly communication with open source and community-owned infrastructure? Learned Publishing, 32(1):75–78, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Upshall:2019:MAW

[2000] Michael Upshall. Mythbusting AI: What is all the fuss about? Learned Publishing, 32(1):79–83, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cook:2019:TSR

[2001] Eleanor I. Cook and Regina Romano Reynolds. Is the term ‘serials’ relevant any longer? Some thoughts on the matter. … Learned Publishing, 32(1):85–89, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dyke:2019:DEC

[2002] Gareth Dyke. Does the early career ‘publish or perish’ myth represent an opportunity for the publishing industry? Learned Publishing, 32(1):90– 94, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mauti:2019:PPW

[2003] Charlotte Mauti. The pursuit of publishing: Was a masters in publishing worthwhile? Learned Publishing, 32(1):95–96, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kessler:2019:WPP

[2004] Rachel Kessler. White paper: Paths to reference — how today’s students find and use reference resources: Review. Learned Publishing, 32(1):97– 98, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:TOR

[2005] Anonymous. Thanking our reviewers. Learned Publishing, 32(1):99, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:JIa

[2006] Anonymous. Journal information. Learned Publishing, 32(1):100, January 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:TCb

[2007] Anonymous. Table of contents. Learned Publishing, 32(2):101, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2019:JDW

[2008] Pippa Smart. Journal development: Why and how? Learned Publishing, 32(2):103–105, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Helgesson:2019:AO

[2009] Gert Helgesson and Stefan Eriksson. Authorship order. Learned Publishing, 32(2):106–112, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xu:2019:ICE

[2010] Jie Xu, Jia Yun Wang, Lihong Zhou, and Fen Liu. Internationalization of China’s English-language academic journals: an overview and three approaches. Learned Publishing, 32(2):113–125, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2019:RII

[2011] Yuehong (Helen) Zhang, Fang Bao, Jian Wu, and Hanfeng Lin. Reflections on the international impact of Chinese STM journals. Learned Publishing, 32(2):126–136, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Atjonen:2019:PRD

[2012] Päivi Atjonen. Peer review in the development of academic articles: Experiences of Finnish authors in the educational sciences. Learned Publishing, 32(2):137–146, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2019:SHN

[2013] David Nicholas, Chérifa Boukacem-Zeghmouri, Jie Xu, Eti Herman, David Clark, Abdullah Abrizah, Blanca Rodríguez-Bravo, and Marzena Świgoń. Sci-Hub: the new and ultimate disruptor? View from the front. Learned Publishing, 32(2):147–153, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2019:GSU

[2014] Yang Zhang and Huilian Lun. Is Google Scholar useful for the evaluation of non-English scientific journals? The case of Chinese journals. Learned Publishing, 32(2):154–162, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Allen:2019:WDB

[2015] Heidi Allen, Alexandra Cury, Thomas Gaston, Chris Graf, Hannah Wakley, and Michael Willis. What does better peer review look like? Underlying principles and recommendations for better practice. Learned Publishing, 32(2):163–175, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Staines:2019:WHO

[2016] Heather Staines. What’s happening with open annotation? Discoverability, engagement, community. Learned Publishing, 32(2):177–180, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Rhee:2019:KIS

[2017] Hea Lim Rhee. Korea Institute of Science and Technology Information’s scholarly journal publishing: a case study. Learned Publishing, 32(2): 181–187, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Buck:2019:DOA

[2018] Stuart Buck. The debate over open access is missing the point. Learned Publishing, 32(2):188–190, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:JIb

[2019] Anonymous. Journal information. Learned Publishing, 32(2):191, April 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:TCc

[2020] Anonymous. Table of contents. Learned Publishing, 32(3):193, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2019:WDE

[2021] Pippa Smart. What did early career researchers ever do for us? Learned Publishing, 32(3):195–197, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yoon:2019:HRD

[2022] JungWon Yoon, EunKyung Chung, Jae Yun Lee, and Jihyun Kim. How research data is cited in scholarly literature: a case study of HINTS. Learned Publishing, 32(3):199–206, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Sun:2019:JRD

[2023] Li Sun. Journals removed from DOAJ appearing within SCImago’s ranks: a study of excluded journals. Learned Publishing, 32(3):207–211, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wang:2019:CPR

[2024] LingFeng Wang and YaQing Zhan. A conceptual peer review model for arXiv and other preprint databases. Learned Publishing, 32(3):213– 219, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Frandsen:2019:HCQ

[2025] Tove Faber Frandsen. How can a questionable journal be identified: Frameworks and checklists. Learned Publishing, 32(3):221–226, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bjork:2019:OAJ

[2026] Bo-Christer Björk. Open access journal publishing in the Nordic countries. Learned Publishing, 32(3):227–236, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2019:ECR

[2027] David Nicholas, Anthony Watkinson, Cherifa Boukacem-Zeghmouri, Blanca Rodríguez-Bravo, Jie Xu, Abdullah Abrizah, Marzena Świgoń, David Clark, and Eti Herman. So, are early career researchers the harbingers of change? Learned Publishing, 32(3):237–247, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Copiello:2019:SJS

[2028] Sergio Copiello. On the skewness of journal self-citations and publisher self-citations: Cues for discussion from a case study. Learned Publishing, 32(3):249–258, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Soler:2019:LDB

[2029] Josep Soler and Ying Wang. Linguistic differences between wellestablished and predatory journals: a keyword analysis of two journals in political science. Learned Publishing, 32(3):259–269, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Merga:2019:CRH

[2030] Margaret K. Merga, Shannon Mason, and Julia E. Morris. ‘The constant rejections hurt’: Skills and personal attributes needed to successfully complete a thesis by publication. Learned Publishing, 32(3):271–281, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hill:2019:TFR

[2031] Tom Hill. Turning FAIR into reality: Review. Learned Publishing, 32(3): 283–286, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:JIc

[2032] Anonymous. Journal information. Learned Publishing, 32(3):288, July 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:TCd

[2033] Anonymous. Table of contents. Learned Publishing, 32(4):289, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ainsworth:2019:HCP

[2034] Megan Ainsworth. How can publishers support early career journal editors? Learned Publishing, 32(4):291–292, October 2019. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Venkatesh:2019:FOA

[2035] Kartik K. Venkatesh, Rebecca S. Benner, Thomas W. Riggs, and Nancy C. Chescheir. Final outcome of articles rejected after revision at Obstetrics & Gynecology: an investigation. Learned Publishing, 32 (4):295–303, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shehata:2019:ESC

[2036] Ahmed Maher Khafaga Shehata. Exploring the scholarly communication styles of Arab social science and humanities scholars. Learned Publishing, 32(4):304–311, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:AA

[2037] Anonymous. ALPSP Awards 2019. Learned Publishing, 32(4):312–313, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Vuong:2019:HWP

[2038] Quan-Hoang Vuong. The harsh world of publishing in emerging regions and implications for editors and publishers: the case of Vietnam. Learned Publishing, 32(4):314–324, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zou:2019:SMG

[2039] Qiang Zou, Jingqiu Ma, and Xiaoyang Sheng. A survey of medical ghostwriting in China. Learned Publishing, 32(4):325–334, October 2019. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bachelet:2019:MIA

[2040] Vivienne C. Bachelet, Francisco A. Uribe, Rubén A. Díaz, Alonso F. Vergara, Fabiana Bravo-Córdova, Víctor A. Carrasco, Francisca J. Lizana, Nicolás Meza-Ducaud, and María S. Navarrete. Misrepresentation of institutional affiliations: the results from an exploratory case study of Chilean authors. Learned Publishing, 32(4):335–344, October 2019. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dodds:2019:FAP

[2041] Francis Dodds. The future of academic publishing: Revolution or evolution revisited. Learned Publishing, 32(4):345–354, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pemberton:2019:TRV

[2042] Michael Pemberton, Susanne Hall, Cary Moskovitz, and Chris M. Anson. Text recycling: Views of North American journal editors from an interview-based study. Learned Publishing, 32(4):355–366, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

James:2019:CMC

[2043] Chris James, Lisa Colledge, Wim Meester, Norman Azoulay, and Andrew Plume. CiteScore metrics: Creating journal metrics from the Scopus citation index. Learned Publishing, 32(4):367–374, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Manley:2019:LRL

[2044] Stewart Manley. On the limitations of recent lawsuits against Sci-Hub, OMICS, ResearchGate, and Georgia State University. Learned Publishing, 32(4):375–381, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

OBrien:2019:HPE

[2045] Anna O’Brien, Chris Graf, and Kate McKellar. How publishers and editors can help early career researchers: Recommendations from a roundtable discussion. Learned Publishing, 32(4):383–393, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Green:2019:MDE

[2046] Toby Green. Maximizing dissemination and engaging readers: the other 50% of an author’s day: a case study. Learned Publishing, 32(4):395–405, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Franca:2019:RCS

[2047] Thiago F. A. França and José Maria Monserrat. Reproducibility crisis, the scientific method, and the quality of published studies: Untangling the knot. Learned Publishing, 32(4):406–408, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Galbraith:2019:LAP

[2048] James Galbraith. Library Acquisitions Patterns: Review. Learned Publishing, 32(4):409–411, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2019:JId

[2049] Anonymous. Journal information. Learned Publishing, 32(4):412, October 2019. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:IIa

[2050] Anonymous. Issue information. Learned Publishing, 33(1):1, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Conrad:2020:WCC

[2051] Lettie Y. Conrad, Peter Richardson, and Rebecca Rinehart. Weathering changes and challenges to society publishing. Learned Publishing, 33(1): 2–3, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Late:2020:RLS

[2052] Elina Late, Laura Korkeamäki, Janne Pölönen, and Sami Syrjämäki. The role of learned societies in national scholarly publishing. Learned Publishing, 33(1):5–13, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wise:2020:HSP

[2053] Alicia Wise and Lorraine Estelle. How society publishers can accelerate their transition to open access and align with Plan S. Learned Publishing, 33(1):14–27, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Roscoe:2020:BNS

[2054] Jonathan Roscoe. Building new societies: Insights and predictions from the 5th Wiley Society Member Survey. Learned Publishing, 33(1):29–36, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Clarke:2020:JPS

[2055] Michael T. Clarke. The journal publishing services agreement: a guide for societies. Learned Publishing, 33(1):37–41, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jones:2020:RFI

[2056] Phill Jones. A report on the future of independent and scholarly publishing webinar by the Scholarly Kitchen. Learned Publishing, 33(1):43–46, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderton:2020:PBI

[2057] Sophia Anderton. Publishing at the British Institute of Radiology: a case study. Learned Publishing, 33(1):48–51, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Aime:2020:PPB

[2058] Emilie Aimé, Erika L. Newton, and Catherine Hill. Publishing perspectives from the British Ecological Society: Partnership, community, collaboration, and innovation. Learned Publishing, 33(1):52–56, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Segers:2020:GCS

[2059] Marc Segers, Brett Rubinstein, Laura FitzGerald, and Alix Vance. Geo-ScienceWorld: a case study in non-profit collaboration to drive sustainability and growth for independent societies. Learned Publishing, 33(1): 57–60, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Holmes:2020:SPA

[2060] Edith Holmes. Society publishing: It’s all about the community you serve. Learned Publishing, 33(1):61–63, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Vuong:2020:PSP

[2061] Quan-Hoang Vuong. Plan S, self-publishing, and addressing unreasonable risks of society publishing. Learned Publishing, 33(1):64–68, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dodenhoff:2020:WVN

[2062] Richard Dodenhoff. Why I value non-profit society scholarly publishing. Learned Publishing, 33(1):69–70, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Compton:2020:WEW

[2063] Dana Compton and Angela Cochran. Worth the effort: Why societies should make room for books. Learned Publishing, 33(1):71–74, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Legge:2020:TSO

[2064] Malavika Legge. Towards sustainable open access: a society publisher’s principles and pilots for transition. Learned Publishing, 33(1):76–82, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:TOR

[2065] Anonymous. Thanking our reviewers, 2019. Learned Publishing, 33(1): 83, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:JIa

[2066] Anonymous. Journal information. Learned Publishing, 33(1):84, January 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:IIb

[2067] Anonymous. Issue information. Learned Publishing, 33(2):85, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2020:FC

[2068] Pippa Smart. Is filtering censorship? Learned Publishing, 33(2):86– 87, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Salehi:2020:PPO

[2069] Mohammad Salehi, Mohammad Soltani, Hadis Tamleh, and Shohreh Teimournezhad. Publishing in predatory open access journals: Authors’ perspectives. Learned Publishing, 33(2):89–95, April 2020. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Repiso:2020:IPO

[2070] Rafael Repiso, Alicia Moreno-Delgado, and Daniel Torres-Salinas. If PLOS ONE were really 101 different specialized journals: a proposed approach to the evaluation of multidisciplinary megajournals. Learned Publishing, 33(2):96–103, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anderson:2020:BTA

[2071] Kent R. Anderson. bioRxiv: Trends and analysis of five years of preprints. Learned Publishing, 33(2):104–109, April 2020. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mohammadi:2020:WSH

[2072] Ehsan Mohammadi, Nilofar Barahmand, and Mike Thelwall. Who shares health and medical scholarly articles on Facebook? Learned Publishing, 33(2):111–118, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Vuong:2020:LRN

[2073] Quan-Hoang Vuong. The limitations of retraction notices and the heroic acts of authors who correct the scholarly record: an analysis of retractions of papers published from 1975 to 2019. Learned Publishing, 33(2):119– 130, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2020:DSC

[2074] David Nicholas, Anthony Watkinson, Abdullah Abrizah, Blanca Rodríguez-Bravo, Cherifa Boukacem-Zeghmouri, Jie Xu, Marzena Świgoń, and Eti Herman. Does the scholarly communication system satisfy the beliefs and aspirations of new researchers? Summarizing the Harbingers research. Learned Publishing, 33(2):132–141, April 2020. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jamali:2020:ECR

[2075] Hamid R. Jamali, David Nicholas, Anthony Watkinson, Abdullah Abrizah, Blanca Rodríguez-Bravo, Cherifa Boukacem-Zeghmouri, Jie Xu, Tatiana Polezhaeva, Eti Herman, and Marzena Świgon. Early career researchers and their authorship and peer review beliefs and practices: an international study. Learned Publishing, 33(2):142–152, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gaston:2020:FAJ

[2076] Thomas E. Gaston, Francesca Ounsworth, Tessa Senders, Sarah Ritchie, and Emma Jones. Factors affecting journal submission numbers: Impact factor and peer review reputation. Learned Publishing, 33(2):154–162, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Melero:2020:RAP

[2077] Remedios Melero and Carolina Navarro-Molina. Researchers’ attitudes and perceptions towards data sharing and data reuse in the field of food science and technology. Learned Publishing, 33(2):163–179, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Crow:2020:SOP

[2078] Raym Crow, Richard Gallagher, and Kamran Naim. Subscribe to Open: a practical approach for converting subscription journals to open access.

Learned Publishing, 33(2):181–185, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wang:2020:UWP

[2079] Xiaofeng Wang, Ying Deng, Xuan Lü, Xinlei Zhang, and Lei Yang. The usage of WeChat to promote academic publishing in China: a case study on Chinese Laser Press. Learned Publishing, 33(2):187–191, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:JIb

[2080] Anonymous. Journal information. Learned Publishing, 33(2):192, April 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:IIc

[2081] Anonymous. Issue information. Learned Publishing, 33(3):193, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smart:2020:PDP

[2082] Pippa Smart. Publishing during pandemic: Innovation, collaboration, and change. Learned Publishing, 33(3):194–197, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nicholas:2020:GQS

[2083] David Nicholas, Hamid R. Jamali, Eti Herman, Anthony Watkinson, Abdullah Abrizah, Blanca Rodríguez-Bravo, Cherifa Boukacem-Zeghmouri, Jie Xu, Marzena Świgoń, and Tatiana Polezhaeva. A global questionnaire survey of the scholarly communication attitudes and behaviours of early career researchers. Learned Publishing, 33(3):198–211, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Herman:2020:CCL

[2084] Eti Herman, John Akeroyd, Gaelle Bequet, David Nicholas, and Anthony Watkinson. The changed — and changing — landscape of serials publishing: Review of the literature on emerging models. Learned Publishing, 33(3):213–229, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wei:2020:MSP

[2085] Fangfang Wei and Guijie Zhang. Measuring the scientific publications of double first-class universities from mainland China. Learned Publishing, 33(3):230–244, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xu:2020:CRP

[2086] Jie Xu, Chen He, Jing Su, Yuanxiang Zeng, Zixian Wang, Fei Fang, and Wenhui Tang. Chinese researchers’ perceptions and use of open access journals: Results of an online questionnaire survey. Learned Publishing, 33(3):246–258, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

vanSchalkwyk:2020:NDS

[2087] François van Schalkwyk. Normative drift and self-correction in scholarly book publishing: the case of Makerere University. Learned Publishing, 33(3):259–268, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Marar:2020:ART

[2088] Sumayyia D. Marar and Muaawia A. Hamza. Attitudes of researchers towards plagiarism: a study on a tertiary care hospital in Riyadh, Saudi Arabia. Learned Publishing, 33(3):270–276, July 2020. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Merga:2020:SRA

[2089] Margaret Merga and Shannon Mason. Sharing research with academia and beyond: Insights from early career researchers in Australia and Japan. Learned Publishing, 33(3):277–286, July 2020. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Downes:2020:TAA

[2090] Mike Downes. Thousands of Australian academics on the editorial boards of journals run by predatory publishers. Learned Publishing, 33(3):287– 295, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hodder:2020:SRC

[2091] Clare Hodder. Selling rights: Changes and challenges for rights professionals. Learned Publishing, 33(3):296–298, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Dylla:2020:CPA

[2092] H. Frederick Dylla and Jeffrey Salmon. Collaborating for public access to scholarly publications: a case study of the partnership between the US Department of Energy and CHORUS. Learned Publishing, 33(3):300– 306, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kratochvil:2020:EUJ

[2093] Jiří Kratochvíl, Lukáš Plch, Martin Sebera, and Eva Koriťáková. Evaluation of untrustworthy journals: Transition from formal criteria to a complex view. Learned Publishing, 33(3):308–322, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lee:2020:FIC

[2094] Jongwook Lee, Kiduk Yang, and Dong-Geun Oh. Factors influencing the choice of a publication venue in library and information science. Learned Publishing, 33(3):323–332, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kosmarski:2020:TCR

[2095] Artyom Kosmarski and Nikolay Gordiychuk. Token-curated registry in a scholarly journal: Can blockchain support journal communities? Learned Publishing, 33(3):333–339, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Horbach:2020:PRE

[2096] Serge P. J. M. Horbach, Wytske M. Hepkema, and Willem Halffman. The Platform for Responsible Editorial Policies: an initiative to foster editorial transparency in scholarly publishing. Learned Publishing, 33(3): 340–344, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hill:2020:FRO

[2097] Tom Hill. Four reports on the OA monograph: Review. Learned Publishing, 33(3):345–347, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:IId

[2098] Anonymous. Issue information. Learned Publishing, 33(3):348, July 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:IIe

[2099] Anonymous. Issue information. Learned Publishing, 33(4):349, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ainsworth:2020:NCP

[2100] Megan Ainsworth. The need for change and the problems of change: Continuity and the ‘new normal’. Learned Publishing, 33(4):350–351, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Taylor:2020:EES

[2101] Simone Taylor, Susan Spilka, Kristen Monahan, Isabel Mulhern, and Jeri Wachter. Evaluating equity in scholarly publishing. Learned Publishing, 33(4):353–367, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:Aa

[2102] Anonymous. Award. Learned Publishing, 33(4):368, October 2020. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:Ab

[2103] Anonymous. Award. Learned Publishing, 33(4):369, October 2020. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jamali:2020:NCE

[2104] Hamid R. Jamali, David Nicholas, Eti Herman, Cherifa Boukacem-Zeghmouri, Abdullah Abrizah, Blanca Rodríguez-Bravo, Jie Xu, Marzena Świgon’, Tatiana Polezhaeva, and Anthony Watkinson. National comparisons of early career researchers’ scholarly communication attitudes and behaviours. Learned Publishing, 33(4):370–384, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Helliwell:2020:GAR

[2105] Jack A. Helliwell, William S. Bolton, Joshua R. Burke, Jim P. Tiernan, David G. Jayne, and Stephen J. Chapman. Global academic response to COVID-19: Cross-sectional study. Learned Publishing, 33(4):385–393, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2020:ARP

[2106] Guijie Zhang and Fangfang Wei. Analysing the research performance of province-level administrative regions in China. Learned Publishing, 33 (4):395–409, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Macan:2020:DAG

[2107] Bojan Macan, Lea Škorić, and Jelka Petrak. David among Goliaths: Open access publishing in scientific (semi-)periphery. Learned Publishing, 33(4):410–417, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Trinh:2020:FII

[2108] Thao-Phuong-Thi Trinh, Trung Tran, Hien-Thu-Thi Le, Tien-Trung Nguyen, and Hiep-Hung Pham. Factors impacting international-indexed publishing among Vietnamese educational researchers. Learned Publishing, 33(4):419–429, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Karaca:2020:PHS

[2109] Ozan Karaca, S. Ayhan Çalışkan, and Halil İbrahim Durak. The publication of health sciences theses in Turkey: a study of Ege University. Learned Publishing, 33(4):431–436, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Moustafa:2020:RSP

[2110] Khaled Moustafa. Reforming science publishing. Learned Publishing, 33 (4):437–440, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Downes:2020:WWS

[2111] Mike Downes. Why we should have listened to Jeffrey Beall from the start. Learned Publishing, 33(4):442–448, October 2020. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baverstock:2020:WSD

[2112] Alison Baverstock. What significance does Publishing Studies have right now? Learned Publishing, 33(4):449–452, October 2020. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Frsb:2020:OBD

[2113] Sue Thorn Frsb. Obituary: Bernard Donovan, 1927–2020. Learned Publishing, 33(4):454–455, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2020:IIf

[2114] Anonymous. Issue information. Learned Publishing, 33(4):456, October 2020. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:IIa

[2115] Anonymous. Issue information. Learned Publishing, 34(1):1–3, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Conrad:2021:DRO

[2116] Lettie Y. Conrad. Defining resilience for our industry: Voices of adaptation and reinvention in scholarly publishing. Learned Publishing, 34(1): 4–5, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Shaw:2021:SSD

[2117] Lisa Cuevas Shaw. From sensemaking to sensegiving: a discourse analysis of the scholarly communications community’s public response to the global pandemic. Learned Publishing, 34(1):6–16, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Watkinson:2021:UPI

[2118] Charles Watkinson. University presses and the impact of COVID-19. Learned Publishing, 34(1):17–24, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brown:2021:BRT

[2119] Josh Brown, Phill Jones, Alice Meadows, and Fiona Murphy. Building resilience through cooperation: Two case studies. Learned Publishing, 34(1):25–29, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mower:2021:OEI

[2120] Allyson Mower. Opportunities for eTextbook innovations and partnerships: a case study of the Marriott Library at the University of Utah. Learned Publishing, 34(1):30–34, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Iglesias:2021:LRA

[2121] Violaine Iglesias, Steven Rathgeb Smith, and Dan Gibson. A lesson in resilience: the abrupt digital transformation of society conferences in 2020. Learned Publishing, 34(1):35–42, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Brinton:2021:ICU

[2122] Jeremy Brinton. The impact of COVID-19 on the UK publishing industry: Findings and opportunity. Learned Publishing, 34(1):43–48, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mallett:2021:RSP

[2123] Jonathan J. Mallett. The resilience of scientific publication: From elite ancient academies to open access. Learned Publishing, 34(1):49–56, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Irfanullah:2021:WDR

[2124] Haseeb Md. Irfanullah. So, what does resilience mean for scholarly publishing? Learned Publishing, 34(1):57–63, January 2021. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kasdorf:2021:WWC

[2125] Bill Kasdorf. Why we can be resilient: We have technology and open standards to thank. Learned Publishing, 34(1):64–67, January 2021. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

OConnor:2021:PDP

[2126] Niamh O’Connor. Publishing during a pandemic: Resilience and change. Learned Publishing, 34(1):68–70, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cousijn:2021:WOM

[2127] Helena Cousijn, Ginny Hendricks, and Alice Meadows. Why openness makes research infrastructure resilient. Learned Publishing, 34(1):71–75, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:TOR

[2128] Anonymous. Thanking our reviewers 2020. Learned Publishing, 34(1):76– 77, January 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:IIb

[2129] Anonymous. Issue information. Learned Publishing, 34(2):79–81, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cao:2021:PSL

[2130] Xiaoke Cao, Lei Lei, and Ju Wen. Promoting science with linguistic devices: a large-scale study of positive and negative words in academic writing. Learned Publishing, 34(2):82–88, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mills:2021:PPP

[2131] D. Mills and K. Inouye. Problematizing ‘predatory publishing’: a systematic review of factors shaping publishing motives, decisions, and experiences. Learned Publishing, 34(2):89–104, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wang:2021:CSP

[2132] Jing Wang, Willem Halffman, and Hub Zwart. The Chinese scientific publication system: Specific features, specific challenges. Learned Publishing, 34(2):105–115, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2021:CCC

[2133] Zhiqiang Zhang and Lei Zheng. Consumer community cognition, brand loyalty, and behaviour intentions within online publishing communities: an empirical study of Epubit in China. Learned Publishing, 34(2):116– 127, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Coates:2021:HOB

[2134] Adam Coates. How often are basic details of the research process mentioned in social science research papers? Learned Publishing, 34(2):128– 136, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Asai:2021:ARA

[2135] Sumiko Asai. An analysis of revising article processing charges for open access journals between 2018 and 2020. Learned Publishing, 34(2):137– 143, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Severin:2021:PPR

[2136] Anna Severin and Joanna Chataway. Purposes of peer review: a qualitative study of stakeholder expectations and perceptions. Learned Publishing, 34(2):144–155, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Plakhotnik:2021:CCC

[2137] Maria S. Plakhotnik. Communication clarity in calls for papers issued by journal special issues: First impressions matter. Learned Publishing, 34(2):156–163, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ortega:2021:RIT

[2138] José Luis Ortega. The relationship and incidence of three editorial notices in PubPeer: Errata, expressions of concern, and retractions. Learned Publishing, 34(2):164–174, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nguyen:2021:AIP

[2139] Thi Thu Ha Nguyen, Hiep-Hung Pham, Quan-Hoang Vuong, Quoc-Thai Cao, Viet-Hung Dinh, and Dinh Duc Nguyen. The adoption of international publishing within Vietnamese academia from 1986 to 2020: a review. Learned Publishing, 34(2):175–186, April 2021. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Pranic:2021:QRR

[2140] Shelly M. Pranić, Mario Malički, Stjepan Ljudevit Marušić, Bahar Mehmani, and Ana Marušić. Is the quality of reviews reflected in editors’ and authors’ satisfaction with peer review? A cross-sectional study in 12 journals across four research fields. Learned Publishing, 34(2):187– 197, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yoon:2021:EDC

[2141] JungWon Yoon, EunKyung Chung, Janet Schalk, and Jihyun Kim. Examination of data citation guidelines in style manuals and data repositories. Learned Publishing, 34(2):198–215, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Borrego:2021:TAD

[2142] Ángel Borrego, Lluís Anglada, and Ernest Abadal. Transformative agreements: Do they pave the way to open access? Learned Publishing, 34 (2):216–232, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Gazni:2021:JSC

[2143] Ali Gazni and Fereshteh Didegah. Journal self-citation trends in 1975– 2017 and the effect on journal impact and article citations. Learned Publishing, 34(2):233–240, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Nuzzo:2021:PEL

[2144] James L. Nuzzo. Preliminary evidence that letters to the editor are indexed inconsistently in PubMed and in exercise science and physical therapy journals: Implications and resolutions. Learned Publishing, 34 (2):241–252, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Tarkhanov:2021:CAI

[2145] I. A. Tarkhanov, D. V. Fomin-Nilov, and M. V. Fomin. Crypto access: Is it possible to use cryptocurrencies in scholarly periodicals? Learned Publishing, 34(2):253–261, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Mallett:2021:TAJ

[2146] Jonathan J. Mallett, Debra L. Chin, Iris M. Rush, and Justine R. Smith. The transition of ARVO journals to open access. Learned Publishing, 34(2):262–271, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Staines:2021:MAU

[2147] Heather Staines. Mission alignment: University presses versus institutional visions. Learned Publishing, 34(2):272–278, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

BasirianJahromi:2021:SWA

[2148] Reza BasirianJahromi and Amrollah Shamsi. Should we add social media account details to authors’ affiliations? Learned Publishing, 34(2):279– 281, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

McLeod:2021:SAC

[2149] Sam McLeod. Should authors cite sources suggested by peer reviewers? Six antidotes for handling potentially coercive reviewer citation suggestions. Learned Publishing, 34(2):282–286, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hill:2021:PCC

[2150] Tom Hill. The production, circulation, consumption, and ownership of scientific knowledge: Historical perspectives: Review. Learned Publishing, 34(2):287–289, April 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:IIc

[2151] Anonymous. Issue information. Learned Publishing, 34(3):291–293, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:IGD

[2152] Anonymous. Investigating gender differences in journal selection decisions: a survey of academic researchers. Learned Publishing, 34(3):294– 304, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:PIV

[2153] Anonymous. Perspectives on institutional valuing and support for academic and translational outputs in Japan and Australia. Learned Publishing, 34(3):305–314, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:ECB

[2154] Anonymous. Evaluating the correlation between different impact indicators for library and information science journals: Comparing the journal citation reports and Scopus. Learned Publishing, 34(3):315–330, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:PMP

[2155] Anonymous. Politics matters: the power dynamics behind Chinese English-language humanities and social science journals. Learned Publishing, 34(3):331–338, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:ERI

[2156] Anonymous. Ethical review and informed consent guidelines of high impact anthropology, business, and education research journals. Learned Publishing, 34(3):339–346, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:VCS

[2157] Anonymous. Visegrád countries’ scientific productivity in the European context: a 10-year perspective using Web of Science and Scopus. Learned Publishing, 34(3):347–357, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:CIC

[2158] Anonymous. Characteristics of Imperial College London’s COVID-19 research outputs. Learned Publishing, 34(3):358–369, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:STT

[2159] Anonymous. Standardizing terminology for text recycling in research writing. Learned Publishing, 34(3):370–378, July 2021. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:BLB

[2160] Anonymous. Beall’s legacy in the battle against predatory publishers. Learned Publishing, 34(3):379–388, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:APB

[2161] Anonymous. Author preprint behaviour and non-compliance with journal preprint policies: One biomedical journal’s experience. Learned Publishing, 34(3):389–395, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:FAF

[2162] Anonymous. Fear of the academic fake? Journal editorials and the amplification of the ’predatory publishing’ discourse. Learned Publishing, 34(3):396–406, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:KPP

[2163] Anonymous. Knowledge production on predatory publishing: a systematic review. Learned Publishing, 34(3):407–413, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:RIK

[2164] Anonymous. Research on Industry 4.0 and on key related technologies in Vietnam: a bibliometric analysis using Scopus. Learned Publishing, 34(3):414–428, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:ESC

[2165] Anonymous. Evolution of scientific collaboration on COVID-19: a bibliometric analysis. Learned Publishing, 34(3):429–441, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:OOA

[2166] Anonymous. Opportunities for open access: Insights from the Wiley Practitioner Survey. Learned Publishing, 34(3):442–449, July 2021. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:CRR

[2167] Anonymous. COVID-19 Rapid Review cross-publisher initiative: What we have learned and what we are going to do next. Learned Publishing, 34(3):450–453, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:SNG

[2168] Anonymous. scite: the next generation of citations. Learned Publishing, 34(3):454–456, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:DAG

[2169] Anonymous. Did anthropause generate a research pause during the pandemic? The experiences of a non-medical journal. Learned Publishing, 34(3):457–460, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:IPR

[2170] Anonymous. Improving peer-review by developing reviewers’ feedback literacy. Learned Publishing, 34(3):461–467, July 2021. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:RLS

[2171] Anonymous. Research4Life: Landscape and situation analysis: Review. Learned Publishing, 34(3):468–473, July 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:IId

[2172] Anonymous. Issue information. Learned Publishing, 34(4):475–477, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Holt:2021:IDA

[2173] Jade Holt, Andrew Walker, and Phill Jones. Introducing a data availability policy for journals at IOP Publishing: Measuring the impact on authors and editorial teams. Learned Publishing, 34(4):478–486, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Vuong:2021:AOA

[2174] Quan-Hoang Vuong, Huyen Thanh T. Nguyen, Manh-Toan Ho, and Minh-Hoang Nguyen. Adopting open access in an emerging country:

Is gender inequality a barrier in humanities and social sciences? Learned Publishing, 34(4):487–498, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kim:2021:SDP

[2175] Youngseek Kim. A study of the determinants of psychologists’ data sharing and open data badge adoption. Learned Publishing, 34(4):499– 509, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Plakhotnik:2021:IBI

[2176] Maria S. Plakhotnik. Has IMRD become IMRDi in management journals? Implications as an independent structural component of research articles. Learned Publishing, 34(4):510–518, October 2021. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Koch:2021:JEJ

[2177] Tomas Koch and Raf Vanderstraeten. Journal editors and journal indexes: Internationalization pressures in the semi-periphery of the world of science. Learned Publishing, 34(4):519–527, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Chen:2021:PCJ

[2178] Lu Chen, Min Zhang, Wan Xiong, and Qian Liu. Performance of China’s journals indexed in SCIE: an evaluation based on megajournal metrics. Learned Publishing, 34(4):528–536, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Severin:2021:OPR

[2179] Anna Severin and Joanna Chataway. Overburdening of peer reviewers: a multi-stakeholder perspective on causes and effects. Learned Publishing, 34(4):537–546, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Araujo:2021:CEA

[2180] Rafael J. Araújo, Geoffrey S. Shideler, and Marcus B. Reamer. Chief editors in aquatic science and communication are more likely to oversee editorial boards from their own regions. Learned Publishing, 34(4):547– 557, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Urbano:2021:PAC

[2181] Cristóbal Urbano, Sara Tafalla, Ángel Borrego, and Ernest Abadal. Preprints as an alternative to conference proceedings: a hands-on experience at EDICIC Iberian Meeting 2019. Learned Publishing, 34(4): 558–567, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Khatter:2021:RSP

[2182] Amandeep Khatter, Michael Naughton, Hajira Dambha-Miller, and Patrick Redmond. Is rapid scientific publication also high quality? Bibliometric analysis of highly disseminated COVID-19 research papers. Learned Publishing, 34(4):568–577, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhu:2021:HCB

[2183] Haoran Zhu. Home country bias in academic publishing: a case study of the New England Journal of Medicine. Learned Publishing, 34(4): 578–584, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Polonen:2021:WUN

[2184] Janne Pölönen, Sami Syrjämäki, Antti-Jussi Nygård, and Björn Hammarfelt. Who are the users of national open access journals? The case of the Finnish Journal.fi platform. Learned Publishing, 34(4):585–592, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Repiso:2021:PIS

[2185] Rafael Repiso, Jesús Segarra-Saavedra, Tatiana Hidalgo-Marí, and Victoria Tur-Viñes. The prevalence and impact of special issues in communications journals 2015–2019. Learned Publishing, 34(4):593–601, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Hayes:2021:INI

[2186] Victoria Hayes, Emma Williams, Kathleen M. Fairfield, Carolyne Falank, Dina McKelvy, and Robert Bing-You. Impact of a new institutional medical journal on professional identity development and academic cultural change: a qualitative study. Learned Publishing, 34(4):602–611, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ndungu:2021:SJP

[2187] Miriam Wanjiku Ndungu. Scholarly journal publishing standards, policies and guidelines. Learned Publishing, 34(4):612–621, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Bordignon:2021:PCA

[2188] Frederique Bordignon, Liana Ermakova, and Marianne Noel. Overpromotion and caution in abstracts of preprints during the COVID-19 crisis. Learned Publishing, 34(4):622–636, October 2021. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Krauskopf:2021:APC

[2189] Erwin Krauskopf. Article processing charge expenditure in Chile: the current situation. Learned Publishing, 34(4):637–646, October 2021. CO-DEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Ashkenazi:2021:HAA

[2190] Itamar Ashkenazi and Oded Olsha. Honorific authorship and approval of the ICMJE criteria: a survey with a convenience sample. Learned Publishing, 34(4):647–654, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Merga:2021:ALK

[2191] Margaret K. Merga. The academic labour of knowledge mobilization: What scholarly publishers need to know. Learned Publishing, 34(4):655– 665, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Simic:2021:LTO

[2192] Josip Šimić, Matko Marušić, Marijana Gelo, Nikolina Šaravanja, Aleksandra Mišak, and Ana Marušić. Long-term outcomes of 2-day training on planning and writing research on publication output of medical professionals: 11-year cohort study. Learned Publishing, 34(4):666–674, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Marshall:2021:ECR

[2193] Helina Marshall and Maria Fernandes. Early-career researchers shaping publishing strategy. Learned Publishing, 34(4):675–678, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Helgesson:2021:TFC

[2194] Gert Helgesson. The two faces of the corresponding author and the need to separate them. Learned Publishing, 34(4):679–681, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wang:2021:BSJ

[2195] Yu Wang and Liangbin Zhao. Blockchain for scholarly journal evaluation: Potential and prospects. Learned Publishing, 34(4):682–687, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Moussa:2021:JHC

[2196] Salim Moussa. Journal hijacking: Challenges and potential solutions. Learned Publishing, 34(4):688–695, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Baveye:2021:OPR

[2197] Philippe C. Baveye. Objectivity of the peer-review process: Enduring myth, reality, and possible remedies. Learned Publishing, 34(4):696–700, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2021:A

[2198] Anonymous. Award. Learned Publishing, 34(4):701–702, October 2021. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2022:IIa

[2199] Anonymous. Issue information. Learned Publishing, 35(1):1–3, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Conrad:2022:PFR

[2200] Lettie Y. Conrad, Richard Delahunty, and Wendy Ding. The promise and the future of research data sharing. Learned Publishing, 35(1):4–6, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Duan:2022:ROD

[2201] Qingyu Duan, Xiaoguang Wang, and Ningyuan Song. Reuse-oriented data publishing: How to make the shared research data friendlier for researchers. Learned Publishing, 35(1):7–15, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Herbet:2022:DDS

[2202] Marie-Emilia Herbet, Jérémie Leonard, Maria Grazia Santangelo, and Lucie Albaret. Dissimulate or disseminate? A survey on the fate of negative results. Learned Publishing, 35(1):16–29, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Wang:2022:RDP

[2203] Yu Wang, Beibei Chen, Liangbin Zhao, and Yuanxiang Zeng. Research data policies of journals in the Chinese Science Citation Database based on the language, publisher, discipline, access model and metrics. Learned Publishing, 35(1):30–45, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Li:2022:TFO

[2204] Chengzan Li, Yuanchun Zhou, Xiaohuan Zheng, Zeyu Zhang, Lulu Jiang, Zongwen Li, Pengyao Wang, Jianhui Li, Songyuan Xu, and Zhanjie Wang. Tracing the footsteps of open research data in China. Learned Publishing, 35(1):46–55, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cannon:2022:IOF

[2205] Matthew Cannon, Andrew Kelly, and Chris Freeman. Implementing an Open & FAIR data sharing policy — a case study in the earth and environmental sciences. Learned Publishing, 35(1):56–66, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Smit:2022:SRD

[2206] Eefke Smit and Joris van Rossum. STM’s Research Data Year: Results, lessons learnt, and next steps. Learned Publishing, 35(1):67–70, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

MacFarlane:2022:IED

[2207] Aki MacFarlane. The importance of effective data sharing and reuse to funders and others supporting research. Learned Publishing, 35(1):71– 74, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Cousijn:2022:BDS

[2208] Helena Cousijn, Ted Habermann, Elizabeth Krznarich, and Alice Meadows. Beyond data: Sharing related research outputs to make data reusable. Learned Publishing, 35(1):75–80, January 2022. CODEN LEP-UFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Manville:2022:ITO

[2209] Catriona Manville and Grace Melvin. The importance of transparency and openness in research data to drive patient benefit — examples from the United Kingdom. Learned Publishing, 35(1):81–85, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2022:TOR

[2210] Anonymous. Thanking our reviewers 2021. Learned Publishing, 35(1):86– 87, January 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Anonymous:2022:IIb

[2211] Anonymous. Issue information. Learned Publishing, 35(2):89–91, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2022:IAC

[2212] Guijie Zhang, Fangfang Wei, Peixin Wang, and Yikai Liang. Investigating academic conference publications from provincial administrative regions in the mainland of China. Learned Publishing, 35(2):92–104, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Kratochvil:2022:CRT

[2213] Jiří Kratochvíl, Helga Abrahámová, Marta Fialová, and Martina Stodůlková. Citation rules through the eyes of biomedical journal editors. Learned Publishing, 35(2):105–117, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Feenstra:2022:PPP

[2214] Ramón A. Feenstra and Emilio Delgado López-Cózar. Philosophers’ perceptions of pay to publish and open access in Spain: Books versus journals, more than a financial dilemma. Learned Publishing, 35(2):118–129, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Moed:2022:JBL

[2215] Henk F. Moed, Carmen Lopez-Illescas, Vicente P. Guerrero-Bote, and Felix de Moya-Anegon. Journals in Beall’s list perform as a group less well than other open access journals indexed in Scopus but reveal large differences among publishers. Learned Publishing, 35(2):130–139, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Coates:2022:AJU

[2216] Adam Coates. Academic journals’ usernames and the threat of fraudulent accounts on social media. Learned Publishing, 35(2):140–148, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Lei:2022:JRR

[2217] Yan Lei. Is a journal’s ranking related to the reviewer’s academic impact? (An empirical study based on Publons). Learned Publishing, 35(2):149– 162, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2022:CPW

[2218] Bei Zhang, Jing Sun, Huangqun Zhang, and Chengting Xu. Can promotion on WeChat official accounts improve scholarly journals’ academic impact? A micro-level correlation comparison study. Learned Publishing, 35(2):163–174, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Koley:2022:JAE

[2219] Moumita Koley and Kanchan Lala. Are journal archiving and embargo policies impeding the success of India’s open access policy? Learned Publishing, 35(2):175–186, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Khelfaoui:2022:ENP

[2220] Mahdi Khelfaoui and Yves Gingras. Expanding Nature: Product line and brand extensions of a scientific journal. Learned Publishing, 35(2):187– 197, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jamali:2022:SJP

[2221] Hamid R. Jamali, Simon Wakeling, and Alireza Abbasi. Scholarly journal publishing in Australia. Learned Publishing, 35(2):198–208, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Zhang:2022:UPS

[2222] Mei Zhang. University press selection of e-book vendors for US academic libraries: Why work with X but not Y? Learned Publishing, 35(2):209– 218, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jamali:2022:WDJ

[2223] Hamid R. Jamali, Simon Wakeling, and Alireza Abbasi. Why do journals discontinue? A study of Australian ceased journals. Learned Publishing, 35(2):219–228, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Helgesson:2022:EPT

[2224] Gert Helgesson, Igor Radun, Jenni Radun, and Gustav Nilsonne. Editors publishing in their own journals: a systematic review of prevalence and a discussion of normative aspects. Learned Publishing, 35(2):229–240, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Fairclough:2022:QMA

[2225] Ruth Fairclough and Mike Thelwall. Questionnaires mentioned in academic research 1996–2019: Rapid increase but declining citation impact. Learned Publishing, 35(2):241–252, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Yang:2022:DCI

[2226] Lili Yang, Siyun Xu, Jufang Shao, Panzhi Wang, Xianjun Wang, Qi Qi, and Rongwang Yang. Declaration of conflict of interest in medical researchers: a cross-sectional study from China. Learned Publishing, 35 (2):253–260, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Xu:2022:NAE

[2227] Shaoxiong (Brian) Xu and Guangwei Hu. Non-author entities accountable for retractions: a diachronic and cross-disciplinary exploration of reasons for retraction. Learned Publishing, 35(2):261–270, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Galli:2022:LCT

[2228] Carlo Galli, Maria Teresa Colangelo, and Stefano Guizzardi. Linguistic changes in the transition from summaries to abstracts: the case of the Journal of Experimental Medicine. Learned Publishing, 35(2):271– 284, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Jorm:2022:PSH

[2229] Anthony F. Jorm. Publons as a source of high volume, poorly targeted reviewer requests: the need for better standards of practice by publishers.

Learned Publishing, 35(2):285–287, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

Herubel:2022:AMP

[2230] Jean-Pierre V. M. Hérubel. For alma mater: Publishing institutional histories of higher education and university presses: Purposes, genre and scholarly value. Learned Publishing, 35(2):288–291, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).

daSilva:2022:FRN

[2231] Jaime A. Teixeira da Silva and Quan-Hoang Vuong. Fortification of retraction notices to improve their transparency and usefulness. Learned Publishing, 35(2):292–299, April 2022. CODEN LEPUFJ. ISSN 0953-1513 (print), 1741-4857 (electronic).